Fontname

July 2009

Filenames for TEX fonts

Karl Berry

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Filenames for fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Suppliers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Weights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.4 Variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.5 Widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

3 Long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.1 A fontname mapping file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.2 A naming scheme for long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Appendix A Font name lists . . . . . . . . . . . . . . . . . . . 27

A.1 Standard PostScript fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 A.2 Adobe fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 A.3 Apple fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 A.4 Bitstream fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 A.5 DTC fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 A.6 ITC fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 A.7 Linotype fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 A.8 Monotype fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.9 URW fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238

Appendix B Encodings . . . . . . . . . . . . . . . . . . . . . . . . 240

B.1 ‘6w.enc’: Cyrillic CP-1251. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 B.2 ‘8a.enc’: Adobe standard encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 246 B.3 ‘8r.enc’: TEX base encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 B.4 ‘dvips.enc’: Dvips encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 B.5 ‘ec.enc’: Cork encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 B.6 ‘groff.enc’: groff encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 B.7 ‘t5.enc’: T5 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 B.8 ‘qx.enc’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 B.9 ‘texmext.enc’: TEX math extension encoding. . . . . . . . . . . . . . . . 290 B.10 ‘texmsym.enc’: TEX math symbol encoding . . . . . . . . . . . . . . . . . 294 B.11 ‘texmital.enc’: TEX math italic encoding. . . . . . . . . . . . . . . . . . 299 B.12 ‘texnansi.enc’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 B.13 ‘texnansx.enc’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 B.14 ‘xl2.enc’: OT1 + ISO Latin 2 extended. . . . . . . . . . . . . . . . . . . . 316 B.15 ‘xt2.enc’: typewriter OT1 + ISO Latin 2 extended. . . . . . . . . 323

ii

General index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331

1 Introduction

This is Fontname, a naming scheme for (the base part of) external TEX font filenames. This makes at most eight-character names from (almost) arbitrarily complex font names, thus helping portability of TEX documents. We point out some shortcomings of the scheme, and ways of dealing with them.

Please send additions, corrections, or other comments to tex-fonts@tug.org. Email

tex-fonts-request@tug.org to join the list.

Besides this document, the Fontname distribution includes various encoding files, mapping files, and utilities. See the file ‘README’ for details. You can retrieve the most upto-date version of these lists from http://tug.org/fontname. It is also mirrored on all CTAN hosts in the directory ‘tex-archive/doc/fontname’. For a list of CTAN mirrors,

see http://www.ctan.org/tex-archive/CTAN.sites.

This document is in the public domain and may be used freely.

1.1 History

The original inspiration for Fontname was Frank Mittelbach and Rainer Schoepf’s article in TUGboat 11(2) (June 1990), which led to my article in TUGboat 11(4) (November 1990), pages 512–519. Mittelbach wrote another article criticizing the scheme below in TUGboat 13(1) (April 1992), pages 51–53; most of his points are well-taken, but I saw no alternative then, and see no alternative now. Others of his points are addressed in the appropriate sections below.

In August 1995, Fontname version 2 was released, concomitantly with a new release of the PostScript fonts for use with LaTEX and Dvips. After several years of experience, several modifications of the original distribution seemed desirable: more precise specification of encodings (see Section 2.4 [Variants], page 20); machine-readable mapping files (the ‘*.map’ files in the distribution) and standardizing directory names as well as font filenames

(included in ‘supplier.map’ and ‘typeface.map’).

In Fontname version 1, some encoding variant characters (‘0’, ‘2’, etc.) have been reassigned to a ‘7’ or ‘8’ sequence (see Section 2.4 [Variants], page 20). Although this is an incompatible change, the advantages of (1) consistency in naming encoding variants, and (2) having some free variants makes it worthwhile. Also, not so coincidentally, new metrics for PostScript fonts in the Cork encoding were issued—same encoding, different metrics. So a new name served a practical purpose as well.

Many people have contributed to this work. In particular, I gratefully acknowledge Nelson Beebe, Barbara Beeton, Rocky Bernstein, Thierry Bouche, Damian Cugley, Berthold K.P. Horn, Alan Jeffrey, Russell Lang, Pierre MacKay, Sebastian Rahtz, Jean Rivlin, and Tom Rokicki.

1.2 References

Related files and distributions:

• The TEX Directory Structure (TDS) standard, available at (http://tug.org/tds and ‘CTAN:/tds’) specifies a directory tree for all TEX library files. The directory names given in Section 2.1 [Suppliers], page 3 and Section 2.2 [Typefaces], page 4 are intended to mesh with that standard.

• The ‘modes.mf’ file (ftp://tug.org/tex/modes.mf) contains recommended mode

names to use as directory names.

• The Fontinst distribution (http://tug.org/applications/fontinst supports quite general creation of virtual fonts, named according to the scheme given here.

• The Dvips translator (http://tug.org/dvips) supports resident and downloadable PostScript fonts. Dvips was the first place where Fontname became widely known. Dvips also includes Afm2tfm, another program that can create virtual fonts.

• The Dviljk processor (included in TEX Live) contains TEX support for the fonts built into the LaserJet 4.

• ‘Designing New Typefaces with Metafont’, by Richard Southall, Stanford Computer Science Department Technical Report STAN-CS-85-1074, September 1985, for a comprehensive terminology of font design.

2 Filenames for fonts

In this scheme, names may consist of only the letters (monocase a-z), numerals (0-9), and underscore. Having two names which differ only in case is undesirable, since it results in clashes on common Windows filesystems.

Here is the basic scheme (the spaces here are merely for readability):

      S TT W [V...] [N] [E] [DD]

where

S represents the supplier of the font.

TT represents the typeface name.

W represents the weight.

V… represents the variant(s), and is omitted if both it and the width are normal. Many fonts have more than one variant.

N represents the encoding, and is omitted if the encoding is nonstandard. Encodings are subsumed in the section on variants (see Section 2.4 [Variants], page 20).

E represents the width (“expansion”), and is omitted if it is normal.

DD represents the design size (in decimal), and is omitted if the font is linearly scaled. Mittelbach in TUGboat 13(1) proposes using hexadecimal or base-36 notation. In my opinion, the increased range does not make up for the unreadability of the common sizes (e.g., 10 pt fonts would have a suffix ‘a’ (in hex), or ‘j’ (in base 36)).

We used to try to limit font names to eight characters. This was never completely possible, and since DOS filesystems have become rare, it’s no longer an issue.

The weight, variants, and width are probably all best taken from the original name of the font, instead of trying to relate them to some external standard.

2.1 Suppliers

The supplier is the source of a font, typically a (digital) type foundry.

You should use the supplier letter which matches the supplier you obtained the font from, not the original source; for example, Avant Garde was designed by Herb Lubalin for ITC, but Adobe also sells it. The name of the font that you get from Adobe should start with ‘p’. This is because font resellers typically make modifications to the original design.

Notes on specific suppliers:

‘f’ For fonts that are distributed without any specific attribution to the creator, by individuals, or by small foundries. (Unfortunately, we don’t have enough characters to assign one to every font supplier in the world.)

‘r’ obsolete; specifies raw fonts, in the old distribution of Dvips. New fonts should never use ‘r’. (The right thing to do is specify the correct encoding, variant, or whatever the font’s characteristics actually are.)

‘z’ for fonts that just don’t fit well into the naming scheme. The ‘z’ should be followed by the real supplier letter.

Here is the table from the file ‘supplier.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation, directory name, and comment.

0 fontfont  FontFont
2 elsnerflake Elsner & Flake
5 softmake  Softmaker
9 unknown
a autologi  Autologic
b bitstrea  Bitstream

c cg Compugraphic d dtc Digital Typeface Corporation

e apple  Apple

f public small foundries g gnu Free Software Foundation h bh Bigelow & Holmes i itc International Typeface Corporation

j microsft  Microsoft
k softkey  SoftKey
l linotype  Linotype
m monotype  Monotype
n ibm  IBM
o corel  Corel

p adobe Adobe (‘p’ for PostScript) q texgyre TEX Gyre fonts from GUST Typefoundry r - raw (for [obsolete] use with virtual fonts) s sun Storm Type

t paragrap  ParaGraph
u urw  URW
v sil  SIL
w wolfram  Wolfram

y arkandis Arkandis Digital Foundry z - bizarre (fontname is nonstandard)

- autofont Eddie Kohler’s autofont program

- jknappen Joerg Knappen

- mnm Hong Feng, free software in China

- yandy  Y&Y

2.2 Typefaces

For our purposes, a typeface is a collection of related fonts.

We sometimes use the same typeface abbreviation for fonts with different supplier names, where we know (usually by inspection) that the fonts truly are the same. This helps conserve abbreviations.

Notes on specific typefaces:

‘ad’

‘gm’ To explain a common confusion: ‘ad’ is Adobe Garamond (with a ‘FontName’ of ‘AGaramond’). This is quite different from the ‘Garamond’ typeface distributed by Adobe, which is actually ITC Garamond. (Garamond No. 3 (‘f3’), Simoncini Garamond (‘ig’), and Stempel Garamond (‘eg’) are still other variations on the Garamond theme.)

‘lh’
‘ls’
‘lx’

‘r1’ These abbreviations for Lucida typeface subfamilies, et al., solve a problem with names getting too long, or help conserve variant abbreviations. See Section 2.4 [Variants], page 20. Ideally, they would not exist.

‘ns’
‘nt’

‘tm’ Suppliers offer as many variations on the Times New Roman design as they

          do Garamond. Adobe has ‘Times’, ‘TimesNewRoman’, and ‘TimesNewRomanPS’.

Although this last has been used by other foundries (e.g., Monotype) for a Times variation exactly compatible with Adobe’s ‘Times’, Adobe itself also offers it, so we must have abbreviations for all three.

‘pi’ For pi fonts. Since names of pi fonts are rather arbitrary, it is not useful to try to assign typeface abbreviations to all of them. Instead, we use the supplier letter, ‘pi’, an arbitarily-assigned three-digit number, a weight and (if necessary) a variant. For example, ‘mpi001r’. Pi font ‘001’ for one supplier has nothing to do with pi font ‘001’ for another.

Here is the table, from the file ‘typeface.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and directory name, possibly followed by PostScript ‘FontName’s and/or a comment in roman. The ‘FontName’ may be preceded by a letter s and a ‘:’, as in ‘b:Dutch’. In that case, the s is a supplier abbreviation, and we’ve used that abbreviation for the given ‘FontName’ from that supplier.

@c @mapfile{
@c   email = "tex-fonts@tug.org",
@c   date = "18jul10",
@c   url = "http://tug.org/fontname/typeface.map",
@c   supported = "yes",
@c   docstring = "Typeface abbreviations for the TeX Fontname scheme.",
@c   license = "public domain",
@c }
@c
@c
a0 arnopro      ArnoPro
a1 arial        Arial
a2 amasis       Amasis
a3 adlib        AdLib
a4 altgoth2     AlternateGothicNo2
a5 allegro      Allegro

a6 amelia       Amelia
a7 amertext     AmericanText
a8 aurora       Aurora
a9 adsans       Adsans b:Humanist970BT
aa aachen       Aachen
ab aboeckli     ArnoldBoecklin
ac acaslon      ACaslon                                 @r{Adobe Caslon}
ad agaramon     AGaramond                               @r{Adobe Garamond}
ae andalemo     AndaleMono
ag avantgar     AvantGarde m:MAvantGardeGothicPS u:URWGothicL
ah ashleysc     AshleyScript
ai abadi        Abadi
aj ajenson      AJenson
ak arkona       Arkona
al albertus     Albertus b:Flareserif821BT
am amercana     Americana b:Americana
an anna         Anna
ao atqolive     AntiqueOlive b:Incised901BT
ap apollo       Apollo
aq antiqua      Antiqua u:URWAntiqua m:BookAntiqua
ar arcadia      Arcadia
as aldus        Aldus
at atypewri     AmericanTypewriter
au auriol       Auriol b:Freeform721
av avenir       Avenir
ax alexa        Alexa
az azdzgrot     AkzidenzGrotesk b:Gothic725BT
b0 blackoak     Blackoak
b1 blado        Blado
b2 binnyos      BinnyOldStyle
b3 braggado     Braggadocio
b4 bison        Bison b:Brush738BT
b5 binghams     BinghamScript b:Freehand591BT
b6 block        Block b:Gothic821
b7 bernhard     Bernhard BernhardModern
b8 blippo       Blippo
b9 bremen       Bremen
ba barmeno      Barmeno
bb bembo        Bembo b:Aldine401BT
bc banco        Banco
bd bodoni       Bodoni u:URWBodoni
be bell         Bell
bf biffo        Biffo
bg benguiat     Benguiat
bh bauhaus      Bauhaus
bi birch        Birch
bj birka        Birka

bk bookman      Bookman m:MBookmanPS u:URWBookmanL
bl belwe        Belwe
bm bulmer       Bulmer
bn bernard      Bernard
bo balloon      Balloon
bp bodocamp     BodoniCampanile b:Modern735
br berling      Berling b:Revival565BT
bs bakersig     BakerSignet
bt boton        Boton
bu brush        BrushScript
bv baskervi     Baskerville u:URWBasker
bw broadway     Broadway
by berkeley     Berkeley
bz balzano      Balzano
c0 carolina     Carolina
c1 calisto      Calisto
c2 casln224     CaslonTwoTwentyFour
c3 caslon3      CaslonThree
c4 chicago      Chicago
c5 casln540     CaslonFiveForty
c6 clarion      Clarion
c7 compacta     Compacta
c8 calvert      Calvert
c9 codex        Codex b:Calligraphic421BT
ca caslon       Caslon
cb cooper       Cooper
cc cascade      Cascade b:Freehand471BT
cd clarendo     Clarendon u:URWClarendon
ce centenni     Centennial
cf clearfac     Clearface
cg charlema     Charlemagne
ch charter      Charter
ci candida      Candida
cj charis       SILCharis TeXChariot
ck caslonbe     CaslonBookBE
cl cloister     Cloister
cm cm           @r{Computer Modern}
cn century      Century
co cochin       Cochin
cp copperpl     Copperplate
cq cheq         Cheq
cr courier      Courier m:MCourierPS m:CourierTwelveMT u:NimbusMonL
cs cntursbk     CenturySchoolbook
ct cheltenh     Cheltenham
cu centuros     CenturyOldStyle
cv clairvx      Clairvaux
cw cottonwd     Cottonwood

cx caxton       Caxton
cy city         City b:SquareSlabserif711BT
cz choc         Choc b:Staccato555BT
d0 davida       DavidaBold
dc domcasl      Dom DomCasual DomDiagonal
dd ducberry     DucDeBerry
de dellarob     DellaRobbia
dg dogma        Dogma
di diotima      Diotima b:Calligraphic810BT
dj delta        Delta
dk diskus       Diskus
dl dolores      Dolores
dm delima       Delima
do didot        DidotLH
dr doric        Doric
ds dorchesc     DorchesterScript
dt dante        Dante
dy academy      Academy
e0 embassy      Embassy
e1 empire       Empire
e2 metatwo      Meta2
ea electra      Electra b:Transitional521BT
eb baskerbe     BaskervilleBE
ec applecha     e:AppleChancery
ee egyptian     Egyptian EgyptianExtended
ef egyptien     EgyptienneF u:URWEgyptienne
eg stgaramo     StempelGaramond b:OriginalGaramondBT
eh engsft       Engschrift
ei esprit       Esprit
ek eckmann      Eckmann b:Freeform710
el ellingtn     Ellington
en engraver     Engravers
eo amerigo      Amerigo b:AmerigoBT
er eras         Eras
es eurostil     Eurostile b:Square721BT k:Minima y:Electrum
et ehrhardt     Ehrhardt
eu euler        Euler
ev mrseaves     MrsEaves
ex exclsior     Excelsior
f0 flemishs     FlemishScript
f1 fournier     Fournier
f2 facade       Facade
f3 festival     Festival
f4 footligh     Footlight
f5 french       FrenchScript
f9 frutigernext FrutgerNext
fa formata      Formata

fc freestyl     FreeStyleScript
fe forte        Forte
ff fairfld      Fairfield b:Transitional551BT
fg frankgth     FranklinGothic
fi figaro       Figaro
fl flora        Flora
fn fenice       Fenice
fo folio        Folio
fq frizquad     FrizQuadrata
fr frutiger     Frutiger b:Humanist777BT
fs falstaff     Falstaff
ft fettefrk     FetteFraktur Fraktur
fu futura       Futura
fv futurans     FuturaNoStarchPress
fx felixttl     FelixTitling
fy flyer        Flyer
g0 gando        Gando
g1 gothic13     GothicThirteen
g2 giddyup      Giddyup
g3 garamon3     GaramondThree b:AmericanGaramondBT      @r{Garamond No.@:  3}
g4 gymodern     GoudyModern
g5 gycatalg     GoudyCatalogue
g6 gorilla      GorillaITC
g9 ugaramon     GaramondRetrospective
ga gillsana     GillSana
gb gothicbb     GothicBBB
gd garuda       Garuda
ge geneva       e:Geneva
gg garthgra     GarthGraphic
gh galahad      Galahad
gi georgia      Georgia
gj granjon      Granjon b:ElegantGaramondBT
gl galliard     Galliard
gm garamond     Garamond u:URWGaramond
gn gentium      Gentium
go gyoldsty     GoudyOldStyle
gp glypha       Glypha
gq grotesq      Grotesque u:URWGrotesk
gr gloucest     Gloucester
gs gillsans     GillSans m:GillAltOneMT b:Humanist521BT
gt goudytxt     GoudyText
gu guardi       Guardi
gv giovanni     Giovanni
gy goudy        Goudy
gz gazette      GazetteLH
h0 handelgo     HandelGothic
ha hadriano     Hadriano

hb hobo         Hobo Tramp
hc herculnm     Herculanum
hd headline     Headline
hg hiroshig     Hiroshige
hh highland     Highlander
hi hvinsera     HelveticaInserat b:Swiss921BT
hm charme       Charme
hn hvneue       HelveticaNeue
ho horleyos     HorleyOldStyle
hr hvround      HelveticaRounded
hs hnseati      Hanseatic B:Swiss924BT
ht hoeflert     e:HoeflerText
hv helvetic     Helvetica u:NimbusSans u:NimbusSanL b:Swiss721BT b:Monospace821BT b:Swiss911BT d:Sans Newton n:SonoranSansSerif
hw schadow      Schadow
hx huxleyvr     HuxleyVertical
i0 commscr      CommercialScript
i1 impress      Impress
i2 impuls       Impuls
i3 imperial     Imperial u:URWImperial
i4 inconsol     Inconsolata
ia insignia     Insignia
ib ibycus       Ibycus
ic impact       Impact
id isadora      Isadora
if inflex       Inflex
ig scgaramo     SimonciniGaramond b:ItalianGaramond
ii imprint      Imprint
il caecilia     Caecilia
im imago        Imago
in industri     Industria
io ionic        Ionic b:News701BT
ip impressm     Impressum
ir Irianis      IrianisADF
is italnos      ItalianOldStyle
it italia       Italia
iw ironwood     Ironwood
iy iowanold     IowanOldStyle
iz isabella     Isabella
jf jeffersn     Jefferson b:Freehand575BT
jn janson       Janson
jo joanna       Joanna
jp juniper      Juniper
k0 calibri      Calibri
k1 cambria      Cambria
k2 candara      Candara
k3 consolas     Consolas
k4 constant     Constantia

k5 corbel       Corbel
kb kabel        Kabel
kg bankgoth     BankGothic
kl klang        Klang
kn kino         Kino
ko bookmano     BookmanOldStyle
kp kepler       Kepler
kr korinna      Korinna
ks kuenstsc     KuenstlerScript
ku kaufmann     Kaufmann
kw kolo         Kolo
kx kigali       Kigali
l0 lineart      LinearT u:URWLinear
l1 liberty      Liberty
l2 libra        Libra
l3 lucia        Lucia
l4 lucian       Lucian
l5 baltica      Baltica
l6 literatu     Literaturnaya
l7 luxiseri     LuxiSerif
l8 luxisans     LuxiSans
l9 luximono     LuxiMono
lb lubaling     LubalinGraph
lc lucida       Lucida
le legacy       Legacy
lf life         Life
lg lettrgth     LetterGothic ISILGothic
lh lubright     LucidaBright
li latin        Latin
ll linolett     LinoLetter
lm lmodern      LatinModern
ln lino         Lino
lo lithos       Lithos
ls lucsans      LucidaSans
lt lutetia      Lutetia
lw leawood      Leawood
lx lucidfax     LucidaFax
ly lydian       Lydian
lz legacysans   LegacySans
m0 monospac     Monospace
m1 monolisc     MonolineScript
m2 metro        Metro b:Geometric415BT
m3 mirarae      Mirarae
m4 mrearl       MisterEarl
m5 murrayhl     MurrayHill
m6 mittlsft     Mittelschrift
m7 matura       Matura

m8 monotyos     MonotypeOldStyle
ma mandate      Mandate b:Freehand521BT
mc monaco       e:Monaco
md meridien     Meridien
me mercursc     MercuriusScript
mg marigold     Marigold
mh machine      Machine
mi medici       Medici
mj mojo         Mojo
ml melior       Melior b:ZapfElliptical711BT
mn minion       Minion
mo modern       Modern b:Modern20BT
mp memphis      Memphis b:GeometricSlab703BT
mq mesquite     Mesquite
mr madrone      Madrone
ms mistral      Mistral b:Staccato222BT
mt minister     Minister
mv malvern      Malvern
mx maximus      Maximus
my myriad       Myriad
mz amazone      Amazone
n0 normande     Normande
n1 binner       Binner @r{not really a family}
n7 news706      News706
na newaster     NewAster
nb nbaskerv     NewBaskerville
nc ncntrsbk     NewCenturySchlbk u:CenturySchL
nd ncaledon     NewCaledonia
nf nofret       Nofret
ng newsgth      NewsGothic
ni nimrod       Nimrod
nk neograph     Neographik
nl neuland      Neuland b:Informal011BT
nm nimbus       Nimbus
nn nclarend     NewClarendon
no notredam     NotreDame
np nsplanti     NewsPlantin
nr nberolin     NewBerolina
ns timesnps     TimesNewRomanPS
nt timesnew     TimesNewRoman m:TimesNR n:SonoranSerif
nu nuptial      Nuptial
nv novarese     Novarese
nw neueunzi     NeueHammerUnziale
nx nyx          Nyx
ny nyala        Nyala
nz neuzeits     NeuzeitS
o0 oldreadf     OldDreadfulNo7

o1 oranda       Oranda
o2 orbitb       OrbitB
o3 colonna      Colonna
o4 oceansan     OceanSans
o5 comic        Comic
o6 colombine    Colombine l:ColombineLT
o7 oldsty7      OldStyleSeven
o8 officina     Officina
oa ocra         OCR-A
ob ocrb         OCR-B
oc concorde     Concorde
oe oldengli     OldEnglish
of oxford       Oxford
oi ondine       Ondine Mermaid b:FormalScript421BT
oj oldclaud     OldClaude
ol colmcill     Colmcille
om omnia        Omnia
on corona       Corona
op optima       Optima d:Optimum b:ZapfHumanist601BT Omega u:Classico
oq ouch         Ouch
or orator       Orator
os oldstyle     OldStyle
ot coronet      Coronet b:Ribbon131BT
ov octavian     Octavian
ox onyx         Onyx
oy olympian     Olympian l:OlympicLH
oz ozhandcr     OzHandicraft
p0 poplar       Poplar
p1 palacesc     PalaceScript
p2 pepita       Pepita
p3 palette      Palette b:Brush445BT
p4 pioneer      PioneerITC
p5 piranesi     Piranesi
p6 blueprin     Blueprint
p7 petersbu     Petersburg
pa parkave      ParkAvenue
pb ptbarnum     PtBarnum
pc poetica      PoeticaChancery s:Scriptoria
pd placard      Placard
pe prestige     PrestigeElite
pf profile      Profile Profil b:Decorated035BT
pg peignot      Peignot b:Exotic350BT
ph photina      Photina
pi pi           @r{pi fonts---see text and @file{*.map} files}
pj pompeija     Pompeijana
pl palatino     Palatino d:Palton b:ZapfCalligraphic801BT u:URWPalladioL Palmspring k:Palindrome
pm postmed      PostMediaeval

pn plantin      Plantin b:Aldine721BT
po ponderos     Ponderosa
pp perpetua     Perpetua b:Lapidary333BT
pq postatqa     PostAntiqua
pr popplres     PopplResidenz
ps parisian     Parisian
pt present      Present
pu popplaud     PopplLaudatio
pw perrywoo     Perrywood
px popplpfx     PopplPontifex
py playbill     Playbill
pz poliphil     Poliphilus
qr quorum       Quorum
qs bellevue     Bellevue
qt boulevar     Boulevard
r0 romana       Romana
r1 rotis        RotisSemiSerif
r2 cantoria     Cantoria
r3 rotis        RotisSemiSans
r4 rotis        RotisSansSerif
ra raleigh      Raleigh
rc romic        Romic
rd romande      RomandeADF
re requiem      Requiem
rm carmina      Carmina
rn rotation     Rotation
ro rotis        RotisSerif
rp reporter     Reporter
rq russelsq     RussellSquare
rr raphael      Raphael
rs arrus        Arrus
rt carta        Carta
ru runic        Runic
rv revue        Revue
rw rockwell     Rockwell Slate b:GeometricSlab712BT
rc rustican     Rusticana
ry ryumin       Ryumin
rz ruzicka      RuzickaFreehandLH
s0 sanmarco     SanMarco
s1 scotchro     ScotchRoman
s2 script       Script
s3 seagull      Seagull
s4 shotgun      Shotgun ShotgunBlanks
s5 snell        Snell
s6 stuyvesa     Stuyvesant
s7 sassoon      Sassoon
s8 stonesan     StoneSans

s9 segoescr     SegoeScript
sa scala        Scala
sb sabon        Sabon b:ClassicalGaramondBT
sc slimbach     Slimbach
se shelley      Shelley
sf serifa       Serifa
sg serifgth     SerifGothic
sh shannon      Shannon
si stoneinf     StoneInformal
sk skia         e:Skia
sl stencil      Stencil
sm spectrum     Spectrum
sn spartan      Spartan
so sonata       Sonata
sp serpenti     Serpentine
sr snellrnd     SnellRoundhand
ss stschnei     StempelSchneidler
st stone        StoneSerif
sv souvenir     Souvenir
sw swing        Swing
sx syntax       Syntax
sy symbol       Symbol m:MSymbolPS u:StandardSymL
t0 tango        Tango
t1 thunderb     Thunderbird
t2 typo         TypoUpright
tb bebodoni     BodoniBE BBodoni                        @r{Berthold Bodoni}
te timeseur     TimesEuropa
tf tiffany      Tiffany
tg tradegth     TradeGothic
th tahoma       Tahoma
ti techphon     TechPhonetic
tj trajan       Trajan
tk tekton       Tekton
tl castella     Castellar
tm times        Times u:NimbusRoman u:NimbusRomNo9L b:Dutch801BT
tn norasi       Norasi
to torino       Torino b:Industrial736BT
tp tempo        Tempo
tr trebuche     Trebuchet
ts thesansm     TheSansMono
tu doulos       SILDoulos TeXToulouse
tv trumpmed     TrumpMediaeval b:Kuenstler480BT
tw twentyc      TwentiethCentury
ty typewrit     Typewriter u:URWTypewriter
u0 columbus     Columbus
ub babodoni     BauerBodoni
ud studio       Studio

ug bengugth     BenguiatGothic
uh cushing      Cushing
ul universa     Universal
um umbra        Umbra
un univers      Univers b:ZurichBT
ur centaur      Centaur b:Venetian301BT
ut utopia       Utopia
uw usherwoo     Usherwood
uy universi     University
v0 vineta       Vineta
v1 venturissans VenturisSansADF
v2 venturis2    VenturisADFNo2
v3 venturissans2 VenturisSansADFNo2
va activa       Activa
vc vectora      Vectora
vd vandijck     VanDijck
ve veraseri     VeraSerif
vg visigoth     Visigoth
vi victoria     Victoria
vj veljovic     Veljovic
vl versaill     Versailles
vm veramono     VeraSansMono
vn verdana      Verdana
vo venturisold  VenturisOldADF
vr vagrnd       VAGRounded
vs verasans     VeraSans
vt venturis     VenturisADF
vv viva         Viva
w0 weddingt     WeddingT
w1 wiesbswi     WiesbadenSwing
w9 wastrel      Wastrel
wb walbaum      Walbaum
wc snowcap      SnowCap
wd weideman     Weidemann
wg wlklgth      WilhelmKlingsporGotisch
wh wendy        Wendy
wi wingding     Wingdings
wk wilke        Wilke
wn windsor      Windsor
wp warnockp     WarnockPro
wr wittenbe     WitternbergerFrakt
ws weiss        Weiss
wt watertit     WatersTitling
ww willow       Willow
xc eccentri     Eccentric
xl linuxlib     LinuxLibertine
xp exponto      ExPonto

yg centuryg     CenturyGothic
yh strayhor     Strayhorn
ym stymie       Stymie
ys goudysan     GoudySans
za zantiqua     ZAntiqua l:ZapfRenaisSH l:ZapfRenaisSB l:ZapfRenSwaSH
zc zapfchan     ZapfChancery m:ZapfChanceryPS u:URWChanceryL
zd zapfding     ZapfDingbats m:ZapfDingbatsPS u:Dingbats
zf zapfino      Zapfino l:LinotypeZapfino-{One,Two,Three,Four,Ligature,Ornaments}
zg neuzeitg     NeuzeitGrotesk b:Geometric706BT
zp ziptydo      ZiptyDo
zt zeitgeis     Zeitgeist
__agbook       AGBookStencil
__agbookrn     AGBookRounded
__agoldfac     AGOldFace
__agschlbk     AGSchoolbook
__aja          Aja
__albertin     AlbertinaMT
__amigo        Amigo
__andreas      Andreas
__ariadne      Ariadne
__banshee      Banshee
__beesknee     Beesknees
__bellcent     BellCentenial
__berlgrot     BerlinerGrotesk
__bermuda      BermudaLP
__berthold     BertholdScript
__bickham      BickhamScript
__blueisla     BlueIsland
__bodoniof     BodoniOldFaceBE
__bossanov     BossaNovaMVB
__briemaka     BriemAkademi
__briemscr     BriemScript
__buffalog     BuffaloGalMT
__caflisch     CaflischScript
__caliban      Caliban
__catull       Catull
__celestia     CelestiaAntiqua
__cerigo       Cerigo
__chaparra     Chaparral
__christia     Christiana
__colossal     Colossalis
__comenius     ComeniusAntiqua
__concorde     ConcordeNova
__congabra     CongaBrava
__copal        Copal
__coriande     Coriander
__cosmos       Cosmos

__cremona      Cremona
__critter      Critter
__cronos       Cronos
__ctmercur     CTMercurius
__curlz        CurlzMT
__cutout       Cutout
__delphin      Delphin
__elgreco      ElGreco
__else         ElseNPL
__emmascri     Emmascript
__flood        Flood
__florens      FlorensLP
__fusaka       Fusaka
__gallia       MonotypeGallia
__garamobe     GaramondBE
__gillfaci     GillFaciaMT
__gillflor     GillFloriatedCapsMT
__graphite     Graphite
__greymant     GreymantleMVB
__hardwood     HardwoodLP
__immi505      Immi505
__jimbo        Jimbo
__khaki        Khaki
__kidprint     KidprintMT
__kinesis      Kinesis
__kochantq     KochAntiqua
__legault      Legault
__lightlin     LightlineGothicMT
__lotype       LoType
__magnesiu     MagnesiumMVB
__magnolia     MagnoliaMVB
__mahsuri      MahsuriSansMT
__manito       ManitoLP
__mendoza      MendozaRoman
__mezz         Mezz
__micr         MICR
__monalisa     MonaLisa
__motterco     MotterCorpus
__mythos       Mythos
__nueva        Nueva
__oceansan     OceanSans
__origami      Origami
__ozwald       Ozwald
__pastonch     PastonchiMT
__pelican      Pelican
__penumbra     Penumbra
__pepperwo     Pepperwood

__poetica      Poetica
__pompeia      Pompeia
__popplexq     PopplExquisit
__postino      Postino
__quake        Quake
__rad          Rad
__reliq        Reliq
__rosewood     Rosewood
__rulngscr     RulingScript
__russlobl     RusselOblique
__sanvito      Sanvito
__sassafra     Sassafras
__serlio       SerlioLH
__sho          Sho
__shuriken     Shuriken
__smaragd      Smaragd
__spring       SpringLP
__spumoni      SpumoniLP
__studz        Studz
__tiepolo      Tiepolo
__toolbox      Toolbox
__verve        Verve
__volutasc     VolutaScript
__zebrawoo     Zebrawood
__shouyi       Shouyi  @r{for mnm}

2.3 Weights

The weight of a font specifies its boldness. Arranged from lightest to heaviest (more or less):

hairline, extra light, light book, regular, medium demibold, semibold, bold, extra bold, heavy black, ultra, poster

Here is the table, from the file ‘weight.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and any parts of a PostScript ‘FontName’ which use that abbreviation.

     a Thin Hairline
     b Bold
     c Black
     d Demi
     h Heavy Heavyface
     j ExtraLight
     k Book
     l Light
     m Medium
     p Poster

     r Regular Roman
     s Semibold
     u Ultra UltraBlack
     x ExtraBold ExtraBlack

2.4 Variants

Variants include typeface variations (e.g., italic, sans serif), and font encodings (e.g., Adobe standard, TEX text).

A fontname may require multiple variants. To resolve the worst ambiguities, we specify that any encoding variant (‘7’, ‘8’, or ‘9’, see below) come last and any other numeral variant come first (to avoid confusion with a design size). We recommend but do not require that the other variants be given in alphabetical order. (It’s not required because it’s too painful to implement the sorting in TEX, and many existing names already have non-alphabetized variants.)

The letterspacing possibilities introduced by fontinst and virtual fonts have not yet become sufficiently widespread to make standardization beneficial. Likewise for the many possible ways to generate small caps fonts.

Notes on specific variants, both old and new:

‘1’ ‘4’ ‘6’ ‘b’ These variants (‘semisans’, ‘bright’, ‘semiserif’, and ‘fax’) were used only for

single typefaces, so to conserve variant abbreviations, they now have typeface abbreviations instead. Also, for ‘b’, to avoid too-long names; and for the others, to avoid variant vs. design-size ambiguities.

‘g’ ‘h’ Since these variants were not actually used in any font names, they are now

obsolete and may be reassigned in the future.

‘5’ ‘6’ ‘7’ ‘8’ ‘9’ These characters begin a two-character encoding variant. Generally, ‘7’ is for

7-bit encodings, ‘8’ is for 8-bit encodings, and ‘9’ is for expertised encodings with or without oldstyle digits (see the ‘x’ item below), but this is not an absolute rule. Also, ‘5’ is presently used for phonetic encodings and ‘6’ for Cyrillic encodings. The ‘9s’ (‘SuperFont’) variant contain all of Adobe, Latin 1, and Expert glyphs, perhaps with slightly different metrics than the original fonts. For a font to be named with a certain encoding variant, it’s not necessary that all the characters appear in precisely the same positions as in the encoding definition. It’s enough that the usual TEX macros work. In practice, this means that it’s ok for a font to be labelled ‘7t’ if the only difference from Computer Modern is that the ligatures and the ‘lslash’ and ‘Lslash’ characters are different, since it’s impossible to access or change the ligature table from TEX. Standard PostScript fonts don’t have an ‘lslash’ character the way Computer Modern does, but they do have the ‘Lslash’ and ‘lslash’ characters themselves, so the usual TEX \lslash and \Lslash macros can be made to work via ligatures. See the file ‘T1.etx’ file in the fontinst distribution for details.

If a name does not contain a specific encoding variant, its encoding is unspecified. For example, some of the fonts distributed with Dvips(k) have names like ‘ptmr’ for ‘Times-Roman’; they use the Dvips encoding (see Section B.4 [dvips], page 255), which is close to (but not the same as) the TEX text encoding (as in Computer Modern Roman). Similarly, the TFM files distributed with Dvilj(k) for the builtin LaserJet 4 fonts have names like ‘cunm’, since these fonts, while compatible with TEX text, contain many additional characters.

‘9’ These items are needed only because ‘x’ (and possibly ‘j’) followed by a twocharacter regular variant makes some names too long. That is, ‘9t’ is equivalent in meaning to ‘x7t’, and ‘9d’ is equivalent to ‘jx7t’. (This may not be true for all ‘9x’ encodings, though.)

‘k’ ‘z’ These variants (indicating Greek and Cyrillic fonts, respectively) are obsolete;

future fonts in different scripts should be assigned an appropriate encoding abbreviation. These are not currently standardized, due to ignorance of the appropriate encodings. Please send suggestions for abbreviations to the address in Chapter 1 [Introduction], page 1.

‘m’ ‘v’ ‘y’ These math-related variants remain for the sake of typeface-specific math en-

codings, e.g., Lucida. Fonts that use the Computer Modern math encodings should use ‘7m’ (see Section B.11 [texmital], page 299), ‘7v’ (see Section B.9 [texmext], page 290), and ‘7y’ (see Section B.10 [texmsym], page 294).

‘n’ This is used for several different (but very rare) variants: only the Stone typeface has an “informal” variant, and only a few Monotype fonts have a “schoolbook” variant, with different ‘a’, ‘g’, and ‘y’ shapes.

‘r’ Included only if no other variants, including encodings, apply, and either the width (see Section 2.5 [Widths], page 24) is not ‘r’ or a design size is present. I.e., ‘r’ is only used as a placeholder. When the normal version of the typeface is sans serif (e.g., Helvetica), ‘r’ should be used when necessary, not ‘s’. Use ‘s’ only when the typeface family has both serif and sans serif variants.

‘s’ ‘t’ Mittelbach in TUGboat 13(1) suggests that these variants (for ‘sans’ and ‘type-

writer’) should be identified as part of the typeface name, because there are few typeface families with these variants. I feel that since they are logically variants, it’s best to name them that way. But ‘LucidaSans’ (see Section 2.2 [Typefaces], page 4) and a few others are exceptions, to avoid too-long names.

‘x’

‘8x’

‘9?’ ‘8x’ indicates a font in the ‘Expert’ encoding itself. ‘x’ indicates an expertised font, i.e., a composite (virtual) font that includes characters from an ‘8x’ font. And in fact an ‘xee’ sequence is replaced by ‘9e’, to save characters.

Fontname 1 assignments are shown in brackets in the following table, from the file ‘variant.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and either any parts of a PostScript ‘FontName’ which use that abbreviation or the PostScript ‘Encoding’ name.

0 inferior
1 superior

2 proportional digits, not tabular obsolete [3=>7f] Fraction obsolete [4 fax; now typeface ‘lx’, Lucida Fax] 5 escape for (presently) phonetic encodings 6 escape for (mostly) Cyrillic encodings [was ‘SemiSerif’] 7 escape for (mostly) 7-bit encodings 8 escape for (mostly) 8-bit encodings 9 escape for (presently) expert encodings [was oldstyle digits] a Alt Arrows Alternative [was alternate encoding] obsolete [b bright; now typeface lh, Lucida Bright]

c SmallCaps
d Display Titling Caption Headline TallCaps SwashCaps LombardicCaps Festive
e Engraved Copperplate Elite
f Fraktur Gothic OldEnglish Handtooled (‘gothic’ can also be sans)

g SmallText lc only, or designed for small sizes [was grooved, as in the IBM logo]

h Shadow
i Italic Kursiv Ital  text italic

j old-style digits [was invisible] k Greek obsolete

l Outline OpenFace Blanks

m math italic

n Informal Fashion Schlbk  for Stone
o Oblique Obl  slanted
p Ornaments

obsolete [q=>8t Cork (TEX extended) encoding] r roman or sans; often omitted, see text s Gothic sans serif t Monospace fixed-width typewriter u underline or unslanted italic

v MathExtension
w Script Handwritten Swash Calligraphy Cursive Tango Ligature

x built with Adobe expert encoding [was expert-encoded]

y MathSymbol

z Cyrillic font-dependent Cyrillic

5a PhoneticAlternate

5i PhoneticIPA

5s sil-IPA 5t TeX-IPA Fukui Rei, LaTEX T3 5w TeXAfricanLatin wsuipa fonts, LaTEX OT3 5z user 6a T2A 6b Cyrillic part of ISO 8859-5, seven bits 6c T2C 6d Cyrillic CP866 encoding 6g LGR Greek font encoding 6i ISO 8859-5 6k Cyrillic KOI8-R encoding 6m Cyrillic Macintosh encoding 6s Storm extra encoding 6t T2B 6w Cyrillic CP1251 encoding

6x X2

6y LCY

6z user

7a A alternate characters only 7c Dfr Fraktur 7d OsF OSF oldstyle digit encoding

7f Fraction
7k OT2Cyrillic

7m TeXMathItalicEncoding see Section B.11 [texmital], page 299 7t TEX text encoding (as in Computer Modern Roman) 7v TeXMathExtensionEncoding see Section B.9 [texmext], page 290 7y TeXMathSymbolEncoding see Section B.10 [texmsym], page 294

7z user
82 GreekKeys
83 Ibycus1
84 Ibycus2

8a StandardEncoding Adobe standard encoding, see Section B.2 [8a], page 246

8c TeXTextCompanion  LaTEX TS1

8e CE Adobe CE

8f TeXAfricanLatin  LaTEX T4
8g groff

8i TS0 Intersection of TS1/Adobe Standard 8m Macintosh standard encoding 8n LM1 Textures

8q encqxoosix  QX, from GUST

8r TeXBase1Encoding (see Section B.3 [8r], page 252) 8t ECEncoding CorkEncoding (see Section B.5 [ec], page 256), aka tex256.enc

8u XT2Encoding  cmtt + Latin 2, see ‘xl2.enc’
8v TeXVietnamese  T5

8w Windows 3.1 ANSI encoding 8x Expert expert encoding 8y LY1 texnansi

8z XL2Encoding  cmr + Latin 2, see ‘xl2.enc’

9c expert + Text companion 9d expert + oldstyle digits + Cork 9e expert + Cork 9i TS0X Intersection of TS1/Standard/Expert 9o expert + oldstyle digits + TEX text 9s SF SuperFont 9t expert + TEX text 9u Unicode-compatible 9x TeXnANSIEncodingX texnansx, texnansi without repeats

9z user
-  songti  for mnm

2.5 Widths

The width of a font specifies the compression or expansion of the font. Arranged from narrowest to widest (more or less):

ultra compressed, extra condensed, compressed, condensed, narrow regular semiextended, extended, expanded, wide

Expansion or compression of fonts is sometimes done automatically (as by the PostScript scale operator), and sometimes done by humans. In the latter case, the human will also presumably choose a font name which includes ‘Extended’ or ‘Expanded’ or ‘Condensed’ or ‘Narrow’ or whatever according to their own preferences; the abbreviation can follow along. When creating a new synthetically expanded or compressed font for use with TEX, e.g., with Afm2tfm or fontinst, use ‘n’ and ‘e’.

Here is the table, from the file ‘width.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and any parts of a PostScript ‘FontName’ which use that abbreviation.

c Condensed Cond
e Expanded
n Narrow Semicondensed
o UltraCondensed
p Compressed Compact
q ExtraCompressed ExtraCondensed
r Normal Medium Regular (usually omitted)
t Thin
u UltraCompressed
v ExtraExpanded more than Expanded, less than Wide
w Wide
x Extended Elongated
y Semiexpanded

3 Long names

Eight characters is not enough to unambiguously represent all fonts. To do that, we have to allow ourselves arbitrarily long names. Although we certainly cannot use such names for the files themselves, we could still use them in TEX documents given a simple mapping file, as discussed below.

3.1 A fontname mapping file

At the moment, most implementations of TEX look up a TFM file (as part of the \font command), by searching for a file with the name given by the user (possibly in any of series of directories). But if we also looked TFM names up in another file (or set of files), which specifies the actual filename, the fontname given in the TEX source file could be almost anything at all, of any length.

In version 5.851d of Web2c, I implemented this mapping file. Each file ‘texfonts.map’ in a search path is read for abbreviations. The file has a straightforward format: each line specifies the filename and the TEX name for one font, separated by whitespace. Extra information on the line is ignored; then more information could be specified for the benefit of DVI-reading programs in the same file. Comments start with % and continue to the end of the line.

Besides allowing long names, this sort of mapping file has other benefits. TEX source or DVI files can be more easily transported, because the font names in a particular file can be made to work on every system. Also, when combined with a consistent naming scheme, macros could be written to access any of a number of fonts. Right now, each font family has to have specialized macros written to deal with it.

Incidentally, Professor Knuth has approved this as a legitimate “system-dependent” extension; a TEX with such a feature can still be called “TEX”.

3.2 A naming scheme for long names

Once we allow ourselves long names, we can construct a naming scheme to handle arbitrary fonts without much difficulty. Here is one proposal, based on the X Window System font naming conventions.

     supplier-family-weight-variants-width-encoding--size

supplier is the usual Adobe, Autologic, etc., as well as unknown or bizarre—this last meaning the rest of the name is nonstandard. If the supplier is missing, i.e., the name starts with a -, “public domain” is assumed. For fonts made by individuals, the initials of the designer would probably make a good supplier.

family is ComputerModern or Times or whatever.

Everything else is optional. The -- before the size lets one specify a name with, say, a weight and variants, but then skip the width and encoding, but still be able to give a size.

weight and width are as described earlier.

If there is more than one variant, they are separated with some character other than ‘-’, say ‘=’:

     B+H-Lucida-Bold-Sans=Typewriter--10

encoding is what Metafont calls the font_coding_scheme—the layout of the characters in the font. For example, ‘Cork’ or ‘ISOLatin1’ or ‘AdobeAlternate’.

Names are case-sensitive, for consistency with the rest of TEX and with PostScript, etc. Spaces cannot be used in the name, to make it easier for TEX to parse. Likewise, characters with default category codes other than 11 (letter) or 12 (other) should not be used.

Another possibility is to forget all the above, and simply use the vendor’s name (perhaps prefixed by the vendor): ‘Times-Roman’.

Appendix A Font name lists

These sections contain a list of fonts from various suppliers together with the corresponding names.

A.1 Standard PostScript fonts

Abbreviated names for the 35 standard PostScript fonts. An encoding variant, such as ‘8a’ for the Adobe ‘StandardEncoding’, is omitted.

‘pagk’  ‘AvantGarde-Book’
‘pagko’  ‘AvantGarde-BookOblique’
‘pagd’  ‘AvantGarde-Demi’
‘pagdo’  ‘AvantGarde-DemiOblique’
‘pbkd’  ‘Bookman-Demi’
‘pbkdi’  ‘Bookman-DemiItalic’
‘pbkl’  ‘Bookman-Light’
‘pbkli’  ‘Bookman-LightItalic’
‘pcrb’  ‘Courier-Bold’
‘pcrbo’  ‘Courier-BoldOblique’
‘pcrr’  ‘Courier’
‘pcrro’  ‘Courier-Oblique’
‘phvb’  ‘Helvetica-Bold’
‘phvbo’  ‘Helvetica-BoldOblique’
‘phvbrn’ ‘Helvetica-NarrowBold’
‘phvbon’ ‘Helvetica-NarrowBoldOblique’
‘phvr’  ‘Helvetica’
‘phvro’  ‘Helvetica-Oblique’
‘phvrrn’ ‘Helvetica-Narrow’
‘phvron’ ‘Helvetica-NarrowOblique’
‘pncb’  ‘NewCenturySchlbk-Bold’
‘pncbi’  ‘NewCenturySchlbk-BoldItalic’
‘pncri’  ‘NewCenturySchlbk-Italic’
‘pncr’  ‘NewCenturySchlbk-Roman’
‘pplb’  ‘Palatino-Bold’
‘pplbi’  ‘Palatino-BoldItalic’

‘pplri’  ‘Palatino-Italic’
‘pplr’  ‘Palatino-Roman’
‘psyr’  ‘Symbol’
‘ptmb’  ‘Times-Bold’
‘ptmbi’  ‘Times-BoldItalic’
‘ptmri’  ‘Times-Italic’
‘ptmr’  ‘Times-Roman’
‘pzcmi’  ‘ZapfChancery-MediumItalic’
‘pzdr’  ‘ZapfDingbats’

A.2 Adobe fonts

The ‘adobe.map’ file has abbreviations for Adobe fonts.

pcrb8a    Courier-Bold                      A    0.1    cob_____
pcrbo8a   Courier-BoldOblique               A    0.1    cobo____
pcrr8a    Courier                           A    0.1    com_____
pcrro8a   Courier-Oblique                   A    0.1    coo_____
phvr8a    Helvetica                         A    0.2    hv______
phvb8a    Helvetica-Bold                    A    0.2    hvb_____
phvbo8a   Helvetica-BoldOblique             A    0.2    hvbo____
phvro8a   Helvetica-Oblique                 A    0.2    hvo_____
psyr      Symbol                            A    0.3    sy______
ptmb8a    Times-Bold                        A    0.4    tib_____
ptmbi8a   Times-BoldItalic                  A    0.4    tibi____
ptmri8a   Times-Italic                      A    0.4    tii_____
ptmr8a    Times-Roman                       A    0.4    tir_____
phvr8an   Helvetica-Narrow                  A    0.5    hvn_____
phvb8an   Helvetica-Narrow-Bold             A    0.5    hvnb____
phvbo8an  Helvetica-Narrow-BoldOblique      A    0.5    hvnbo___
phvro8an  Helvetica-Narrow-Oblique          A    0.5    hvno____
pplb8a    Palatino-Bold                     A    001    pob_____
pplbi8a   Palatino-BoldItalic               A    001    pobi____
pplri8a   Palatino-Italic                   A    001    poi_____
pplr8a    Palatino-Roman                    A    001    por_____
pbkd8a    Bookman-Demi                      A    002    bkd_____
pbkdi8a   Bookman-DemiItalic                A    002    bkdi____
pbkl8a    Bookman-Light                     A    002    bkl_____
pbkli8a   Bookman-LightItalic               A    002    bkli____
pzcmi8a   ZapfChancery-MediumItalic         A    003    zcmi____
pzdr      ZapfDingbats                      A    003    zd______
pagd8a    AvantGarde-Demi                   A    004    agd_____
pagdo8a   AvantGarde-DemiOblique            A    004    agdo____
pagk8a    AvantGarde-Book                   A    004    agw_____
pagko8a   AvantGarde-BookOblique            A    004    agwo____

pncb8a    NewCenturySchlbk-Bold             A    005    ncb_____
pncbi8a   NewCenturySchlbk-BoldItalic       A    005    ncbi____
pncri8a   NewCenturySchlbk-Italic           A    005    nci_____
pncr8a    NewCenturySchlbk-Roman            A    005    ncr_____
popr8a    Optima                            A    006    op______
popb8a    Optima-Bold                       A    006    opb_____
popbi8a   Optima-BoldItalic                 A    006    opbi____
popbo8a   Optima-BoldOblique                A    006    opbo____
popri8a   Optima-Italic                     A    006    opi_____
popro8a   Optima-Oblique                    A    006    opo_____
psvd8a    Souvenir-Demi                     A    007    sud_____
psvdi8a   Souvenir-DemiItalic               A    007    sudi____
psvl8a    Souvenir-Light                    A    007    sul_____
psvli8a   Souvenir-LightItalic              A    007    suli____
plbd8a    LubalinGraph-Demi                 A    008    lud_____
plbdo8a   LubalinGraph-DemiOblique          A    008    ludo____
plbk8a    LubalinGraph-Book                 A    008    luw_____
plbko8a   LubalinGraph-BookOblique          A    008    luwo____
pgmb8a    Garamond-Bold                     A    009    gab_____
pgmbi8a   Garamond-BoldItalic               A    009    gabi____
pgml8a    Garamond-Light                    A    009    gal_____
pgmli8a   Garamond-LightItalic              A    009    gali____
patb8a    AmericanTypewriter-Bold           A    010    atb_____
patba8a   AmericanTypewriter-BoldA          A    010    atba____
patb8ac   AmericanTypewriter-BoldCond       A    010    atbc____
patba8ac  AmericanTypewriter-BoldCondA      A    010    atbca___
patr8ac   AmericanTypewriter-Cond           A    010    atc_____
patra8ac  AmericanTypewriter-CondA          A    010    atca____
patl8a    AmericanTypewriter-Light          A    010    atl_____
patla8a   AmericanTypewriter-LightA         A    010    atla____
patl8ac   AmericanTypewriter-LightCond      A    010    atlc____
patla8ac  AmericanTypewriter-LightCondA     A    010    atlca___
patm8a    AmericanTypewriter-Medium         A    010    atm_____
patma8a   AmericanTypewriter-MediumA        A    010    atma____
pbgb8a    Benguiat-Bold                     A    011    bgb_____
pbgk8a    Benguiat-Book                     A    011    bgw_____
pfqr8a    FrizQuadrata                      A    011    fq______
pfqb8a    FrizQuadrata-Bold                 A    011    fqb_____
pgpr8a    Glypha                            A    012    gy______
pgpb8a    Glypha-Bold                       A    012    gyb_____
pgpbo8a   Glypha-BoldOblique                A    012    gybo____
pgpro8a   Glypha-Oblique                    A    012    gyo_____
phvc8a    Helvetica-Black                   A    013    hvbl____
phvco8a   Helvetica-BlackOblique            A    013    hvblo___
phvl8a    Helvetica-Light                   A    013    hvl_____
phvlo8a   Helvetica-LightOblique            A    013    hvlo____
phvr8ac   Helvetica-Condensed               A    014    hvc_____

phvb8ac   Helvetica-Condensed-Bold          A    014    hvcb____
phvc8ac   Helvetica-Condensed-Black         A    014    hvcbl___
phvbo8ac  Helvetica-Condensed-BoldObl       A    014    hvcbo___
phvro8ac  Helvetica-Condensed-Oblique       A    014    hvcdo___
phvl8ac   Helvetica-Condensed-Light         A    014    hvcl____
phvlo8ac  Helvetica-Condensed-LightObl      A    014    hvclo___
phvco8ac  Helvetica-Condensed-BlackObl      A    014    hvco____
ptvb8a    TrumpMediaeval-Bold               A    015    tmb_____
ptvbi8a   TrumpMediaeval-BoldItalic         A    015    tmbi____
ptvri8a   TrumpMediaeval-Italic             A    015    tmi_____
ptvr8a    TrumpMediaeval-Roman              A    015    tmr_____
pmlr8a    Melior                            A    016    me______
pmlb8a    Melior-Bold                       A    016    meb_____
pmlbi8a   Melior-BoldItalic                 A    016    mebi____
pmlri8a   Melior-Italic                     A    016    mei_____
pglb8a    Galliard-Bold                     A    017    glb_____
pglbi8a   Galliard-BoldItalic               A    017    glbi____
pglri8a   Galliard-Italic                   A    017    gli_____
pglr8a    Galliard-Roman                    A    017    glr_____
pnbb8a    NewBaskerville-Bold               A    018    nbb_____
pnbbi8a   NewBaskerville-BoldItalic         A    018    nbbi____
pnbri8a   NewBaskerville-Italic             A    018    nbi_____
pnbr8a    NewBaskerville-Roman              A    018    nbr_____
pkrb8a    Korinna-Bold                      A    019    krb_____
pkrbi8a   Korinna-KursivBold                A    019    krkb____
pkrri8a   Korinna-KursivRegular             A    019    krkx____
pkrr8a    Korinna-Regular                   A    019    krrg____
pgyr8a    Goudy                             A    020    go______
pgyb8a    Goudy-Bold                        A    020    gob_____
pgybi8a   Goudy-BoldItalic                  A    020    gobi____
pgyri8a   Goudy-Italic                      A    020    goi_____
psor      Sonata                            A    021    so______
pcub8a    CenturyOldStyle-Bold              A    022    csb_____
pcuri8a   CenturyOldStyle-Italic            A    022    csi_____
pcur8a    CenturyOldStyle-Regular           A    022    csrg____
pfgd8a    FranklinGothic-Demi               A    023    frd_____
pfgdi8a   FranklinGothic-DemiItal           A    023    frdi____
pfgdo8a   FranklinGothic-DemiOblique        A    023    frdo____
pfgh8a    FranklinGothic-Heavy              A    023    frh_____
pfghi8a   FranklinGothic-HeavyItal          A    023    frhi____
pfgho8a   FranklinGothic-HeavyOblique       A    023    frho____
pfgm8a    FranklinGothic-Medium             A    023    frm_____
pfgmi8a   FranklinGothic-MediumItal         A    023    frmi____
pfgk8a    FranklinGothic-Book               A    023    frw_____
pfgki8a   FranklinGothic-BookItal           A    023    frwi____
pfgko8a   FranklinGothic-BookOblique        A    023    frwo____
pctb8a    Cheltenham-Bold                   A    024    chb_____

pctbi8a   Cheltenham-BoldItalic             A    024    chbi____
pctk8a    Cheltenham-Book                   A    024    chw_____
pctki8a   Cheltenham-BookItalic             A    024    chwi____
ppar8a    ParkAvenue                        A    025    pa______
pbdr8a    Bodoni                            A    026    bd______
pbdb8a    Bodoni-Bold                       A    026    bdb_____
pbdbi8a   Bodoni-BoldItalic                 A    026    bdbi____
pbdri8a   Bodoni-Italic                     A    026    bdi_____
pbdp8a    Bodoni-Poster                     A    026    bdps____
plgr8a    LetterGothic                      A    027    lg______
plgb8a    LetterGothic-Bold                 A    027    lgb_____
plgbo8a   LetterGothic-BoldSlanted          A    027    lgbsl___
plgro8a   LetterGothic-Slanted              A    027    lgsl____
pper8a    PrestigeElite                     A    028    pe______
ppeb8a    PrestigeElite-Bold                A    028    peb_____
ppebo8a   PrestigeElite-BoldSlanted         A    028    pebsl___
ppero8a   PrestigeElite-Slanted             A    028    pesl____
porr8a    Orator                            A    029    or______
porro8a   Orator-Slanted                    A    029    orsl____
pngr8a    NewsGothic                        A    030    ng______
pngb8a    NewsGothic-Bold                   A    030    ngb_____
pngbo8a   NewsGothic-BoldOblique            A    030    ngbo____
pngro8a   NewsGothic-Oblique                A    030    ngo_____
ptfr8a    Tiffany                           A    031    tf______
ptfd8a    Tiffany-Demi                      A    031    tfd_____
ptfdi8a   Tiffany-DemiItalic                A    031    tfdi____
ptfh8a    Tiffany-Heavy                     A    031    tfh_____
ptfhi8a   Tiffany-HeavyItalic               A    031    tfhi____
ptfri8a   Tiffany-Italic                    A    031    tfi_____
pcbc8a    CooperBlack                       A    032    cb______
pcbci8a   CooperBlack-Italic                A    032    cbi_____
pburw8a   BrushScript                       A    033    bs______
phbrw8a   Hobo                              A    033    ho______
pslb8a    Stencil                           A    033    st______
paab8a    Aachen-Bold                       A    034    acb_____
pfcrw8a   FreestyleScript                   A    034    fs______
prvb8a    Revue                             A    034    re______
puyr8a    UniversityRoman                   A    034    ur______
prtr      Carta                             A    035    cr______
plcr8a    Lucida                            A    036    lc______
plcb8a    Lucida-Bold                       A    036    lcb_____
plcbi8a   Lucida-BoldItalic                 A    036    lcbi____
plcri8a   Lucida-Italic                     A    036    lci_____
punr8a    Univers                           A    037    uv______
punb8a    Univers-Bold                      A    037    uvb_____
punc8a    Univers-Black                     A    037    uvbl____
punco8a   Univers-BlackOblique              A    037    uvblo___

punbo8a   Univers-BoldOblique               A    037    uvbo____
punl8a    Univers-Light                     A    037    uvl_____
punlo8a   Univers-LightOblique              A    037    uvlo____
punro8a   Univers-Oblique                   A    037    uvo_____
punr8ac   Univers-Condensed                 A    038    uvc_____
punb8ac   Univers-CondensedBold             A    038    uvcb____
punbo8ac  Univers-CondensedBoldOblique      A    038    uvcbo___
punro8ac  Univers-CondensedOblique          A    038    uvcdo___
punl8ac   Univers-CondensedLight            A    038    uvcl____
punlo8ac  Univers-CondensedLightOblique     A    038    uvclo___
pfub8a    Futura-Bold                       A    039    fub_____
pfubo8a   Futura-BoldOblique                A    039    fubo____
pful8a    Futura-Light                      A    039    ful_____
pfulo8a   Futura-LightOblique               A    039    fulo____
pfuk8a    Futura-Book                       A    039    fuw_____
pfuko8a   Futura-BookOblique                A    039    fuwo____
pstr8a    StoneSerif                        A    040    sr______
pstb8a    StoneSerif-Bold                   A    040    srb_____
pstbi8a   StoneSerif-BoldItalic             A    040    srbi____
pstri8a   StoneSerif-Italic                 A    040    sri_____
psts8a    StoneSerif-Semibold               A    040    srsb____
pstsi8a   StoneSerif-SemiboldItalic         A    040    srsbi___
ps8r8a    StoneSans                         A    041    ss______
ps8b8a    StoneSans-Bold                    A    041    ssb_____
ps8bi8a   StoneSans-BoldItalic              A    041    ssbi____
ps8ri8a   StoneSans-Italic                  A    041    ssi_____
ps8s8a    StoneSans-Semibold                A    041    sssb____
ps8si8a   StoneSans-SemiboldItalic          A    041    sssbi___
psir8a    StoneInformal                     A    042    si______
psib8a    StoneInformal-Bold                A    042    sib_____
psibi8a   StoneInformal-BoldItalic          A    042    sibi____
psiri8a   StoneInformal-Italic              A    042    sii_____
psis8a    StoneInformal-Semibold            A    042    sisb____
psisi8a   StoneInformal-SemiboldItalic      A    042    sisbi___
ponr8a    Corona                            A    043    cn______
ponb8a    Corona-Bold                       A    043    cnb_____
ponri8a   Corona-Italic                     A    043    cni_____
pesr8a    Eurostile                         A    044    eu______
pesb8a    Eurostile-Bold                    A    044    eub_____
pesbi8a   Eurostile-BoldOblique             A    044    eubo____
pesd8a    Eurostile-Demi                    A    044    eud_____
pesdi8a   Eurostile-DemiOblique             A    044    eudo____
pesri8a   Eurostile-Oblique                 A    044    euo_____
pexr8a    Excelsior                         A    045    ex______
pexb8a    Excelsior-Bold                    A    045    exb_____
pexri8a   Excelsior-Italic                  A    045    exi_____
pfur8a    Futura                            A    046    fu______

pfux8a    Futura-ExtraBold                  A    046    fueb____
pfuxo8a   Futura-ExtraBoldOblique           A    046    fuebo___
pfuh8a    Futura-Heavy                      A    046    fuh_____
pfuho8a   Futura-HeavyOblique               A    046    fuho____
pfuro8a   Futura-Oblique                    A    046    fuo_____
pfuxo8ac  Futura-CondExtraBoldObl           A    047    fcebo___
pfur8ac   Futura-Condensed                  A    047    fuc_____
pfub8ac   Futura-CondensedBold              A    047    fucb____
pfubo8ac  Futura-CondensedBoldOblique       A    047    fucbo___
pfux8ac   Futura-CondensedExtraBold         A    047    fuceb___
pful8ac   Futura-CondensedLight             A    047    fucl____
pfulo8ac  Futura-CondensedLightOblique      A    047    fuclo___
pfuro8ac  Futura-CondensedOblique           A    047    fuco____
plsr8a    LucidaSans                        A    048    ls______
plsb8a    LucidaSans-Bold                   A    048    lsb_____
plsbi8a   LucidaSans-BoldItalic             A    048    lsbi____
plsri8a   LucidaSans-Italic                 A    048    lsi_____
pmpb8a    Memphis-Bold                      A    049    mpb_____
pmpbi8a   Memphis-BoldItalic                A    049    mpbi____
pmpx8a    Memphis-ExtraBold                 A    049    mpeb____
pmpl8a    Memphis-Light                     A    049    mpl_____
pmpli8a   Memphis-LightItalic               A    049    mpli____
pmpm8a    Memphis-Medium                    A    049    mpm_____
pmpmi8a   Memphis-MediumItalic              A    049    mpmi____
phvx8ap   Helvetica-ExtraCompressed         A    050    hvek____
phvr8ap   Helvetica-Compressed              A    050    hvk_____
phvu8ap   Helvetica-UltraCompressed         A    050    hvuk____
pitb8a    Italia-Bold                       A    051    itb_____
pitm8a    Italia-Medium                     A    051    itm_____
pitk8a    Italia-Book                       A    051    itw_____
pblb8a    Belwe-Bold                        A    052    bwb_____
pblr8ac   Belwe-Condensed                   A    052    bwc_____
pbll8a    Belwe-Light                       A    052    bwl_____
pblm8a    Belwe-Medium                      A    052    bwm_____
pc5ri8a   CaslonFiveForty-Italic            A    053    cfi_____
pc5r8a    CaslonFiveForty-Roman             A    053    cfr_____
pc3ri8a   CaslonThree-Italic                A    053    cti_____
pc3r8a    CaslonThree-Roman                 A    053    ctr_____
pgyx8a    Goudy-ExtraBold                   A    054    goeb____
pgyh8a    Goudy-Heavyface                   A    054    goh_____
pgyhi8a   Goudy-HeavyfaceItalic             A    054    gohi____
pjnb8a    JansonText-Bold                   A    055    jtb_____
pjnbi8a   JansonText-BoldItalic             A    055    jtbi____
pjnri8a   JansonText-Italic                 A    055    jti_____
pjnr8a    JansonText-Roman                  A    055    jtr_____
ierb8a    ItcEras-Bold                      A    056    erb_____
ierk8a    ItcEras-Book                      A    056    erbk____

ierd8a    ItcEras-Demi                      A    056    erd_____
ierl8a    ItcEras-Light                     A    056    erl_____
ierm8a    ItcEras-Medium                    A    056    erm_____
ieru8a    ItcEras-Ultra                     A    056    eru_____
ikbb8a    ItcKabel-Bold                     A    057    kbb_____
ikbd8a    ItcKabel-Demi                     A    057    kbd_____
ikbm8a    ItcKabel-Medium                   A    057    kbm_____
ikbu8a    ItcKabel-Ultra                    A    057    kbu_____
ikbk8a    ItcKabel-Book                     A    057    kbw_____
p_____MICR                              A    058    mi______
poar      OCRA                              A    058    oa______
poara     OCRA-Alternate                    A    058    oaalt___
pobr      OCRB                              A    058    ob______
pobra     OCRB-Alternate                    A    058    obalt___
phnc8a    HelveticaNeue-Black               A    059    hlbl____
phnci8a   HelveticaNeue-BlackItalic         A    059    hlbli___
phnj8a    HelveticaNeue-UltraLight          A    059    hlul____
phnji8a   HelveticaNeue-UltraLightItal      A    059    hluli___
phnb8a    HelveticaNeue-Bold                A    060    hlb_____
phnbi8a   HelveticaNeue-BoldItalic          A    060    hlbi____
phnri8a   HelveticaNeue-Italic              A    060    hli_____
phnr8a    HelveticaNeue-Roman               A    060    hlr_____
phna8a    HelveticaNeue-Thin                A    060    hlt_____
phnai8a   HelveticaNeue-ThinItalic          A    060    hlti____
phnh8a    HelveticaNeue-Heavy               A    061    hlh_____
phnhi8a   HelveticaNeue-HeavyItalic         A    061    hlhi____
phnl8a    HelveticaNeue-Light               A    061    hll_____
phnli8a   HelveticaNeue-LightItalic         A    061    hlli____
phnm8a    HelveticaNeue-Medium              A    061    hlm_____
phnmi8a   HelveticaNeue-MediumItalic        A    061    hlmi____
pttb8a    TimesTen-Bold                     A    062    ttb_____
pttbi8a   TimesTen-BoldItalic               A    062    ttbi____
pttri8a   TimesTen-Italic                   A    062    tti_____
pttr8a    TimesTen-Roman                    A    062    ttr_____
pkur8a    Kaufmann                          A    063    kf______
pkub8a    Kaufmann-Bold                     A    063    kfb_____
pcdr8a    Clarendon                         A    064    cl______
pcdb8a    Clarendon-Bold                    A    064    clb_____
pcdl8a    Clarendon-Light                   A    064    cll_____
ppgb8a    Peignot-Bold                      A    065    pgb_____
ppgd8a    Peignot-Demi                      A    065    pgd_____
ppgl8a    Peignot-Light                     A    065    pgl_____
pndr8a    NewCaledonia                      A    066    nl______
pndb8a    NewCaledonia-Bold                 A    066    nlb_____
pndbi8a   NewCaledonia-BoldItalic           A    066    nlbi____
pndc8a    NewCaledonia-Black                A    066    nlbl____
pndci8a   NewCaledonia-BlackItalic          A    066    nlbli___

pndri8a   NewCaledonia-Italic               A    066    nli_____
pnds8a    NewCaledonia-SemiBold             A    066    nls_____
pndsi8a   NewCaledonia-SemiBoldItalic       A    066    nlsi____
pcfb8a    Clearface-Bold                    A    067    ceb_____
pcfbi8a   Clearface-BoldItalic              A    067    cebi____
pcfc8a    Clearface-Black                   A    067    cebl____
pcfci8a   Clearface-BlackItalic             A    067    cebli___
pcfh8a    Clearface-Heavy                   A    067    ceh_____
pcfhi8a   Clearface-HeavyItalic             A    067    cehi____
pcfr8a    Clearface-Regular                 A    067    cerg____
pcfri8a   Clearface-RegularItalic           A    067    cergi___
pamr8a    Americana                         A    068    am______
pamb8a    Americana-Bold                    A    068    amb_____
pamx8a    Americana-ExtraBold               A    068    ameb____
pamri8a   Americana-Italic                  A    068    ami_____
psgr8a    SerifGothic                       A    069    sg______
psgb8a    SerifGothic-Bold                  A    069    sgb_____
psgc8a    SerifGothic-Black                 A    069    sgbl____
psgx8a    SerifGothic-ExtraBold             A    069    sgeb____
psgh8a    SerifGothic-Heavy                 A    069    sgh_____
psgl8a    SerifGothic-Light                 A    069    sgl_____
pcnr8ax   CenturyExpanded                   A    070    cy______
pcnri8ax  CenturyExpanded-Italic            A    070    cyi_____
psfb8a    Serifa-Bold                       A    071    seb_____
psfc8a    Serifa-Black                      A    071    sebl____
psfri8a   Serifa-Italic                     A    071    sei_____
psfl8a    Serifa-Light                      A    071    sel_____
psfli8a   Serifa-LightItalic                A    071    seli____
psfr8a    Serifa-Roman                      A    071    ser_____
pcarl8a   CaslonOpenFace                    A    072    ca______
pfrb8a    Frutiger-Bold                     A    073    ftb_____
pfrbi8a   Frutiger-BoldItalic               A    073    ftbi____
pfrc8a    Frutiger-Black                    A    073    ftbl____
pfrci8a   Frutiger-BlackItalic              A    073    ftbli___
pfrri8a   Frutiger-Italic                   A    073    fti_____
pfrl8a    Frutiger-Light                    A    073    ftl_____
pfrli8a   Frutiger-LightItalic              A    073    ftli____
pfrr8a    Frutiger-Roman                    A    073    ftr_____
pfru8a    Frutiger-UltraBlack               A    073    ftubl___
pceb8a    Centennial-Bold                   A    074    cib_____
pcebi8a   Centennial-BoldItalic             A    074    cibi____
pcec8a    Centennial-Black                  A    074    cibl____
pceci8a   Centennial-BlackItalic            A    074    cibli___
pceri8a   Centennial-Italic                 A    074    cii_____
pcel8a    Centennial-Light                  A    074    cil_____
pceli8a   Centennial-LightItalic            A    074    cili____
pcer8a    Centennial-Roman                  A    074    cir_____

pegb8a    StempelGaramond-Bold              A    075    gsb_____
pegbi8a   StempelGaramond-BoldItalic        A    075    gsbi____
pegri8a   StempelGaramond-Italic            A    075    gsi_____
pegr8a    StempelGaramond-Roman             A    075    gsr_____
pwsr8a    Weiss                             A    076    we______
pwsb8a    Weiss-Bold                        A    076    web_____
pwsx8a    Weiss-ExtraBold                   A    076    weeb____
pwsri8a   Weiss-Italic                      A    076    wei_____
pg3r8a    GaramondThree                     A    077    gm______
pg3b8a    GaramondThree-Bold                A    077    gmb_____
pg3bi8a   GaramondThree-BoldItalic          A    077    gmbi____
pg3ri8a   GaramondThree-Italic              A    077    gmi_____
pulrk     Universal-GreekwithMathPi         A    078    unvmg___
pulrncp   Universal-NewswithCommPi          A    078    unvnc___
pavh8a    Avenir-Heavy                      A    079    aeh_____
pavho8a   Avenir-HeavyOblique               A    079    aeho____
pavl8a    Avenir-Light                      A    079    ael_____
pavlo8a   Avenir-LightOblique               A    079    aelo____
pavr8a    Avenir-Roman                      A    079    aer_____
pavro8a   Avenir-Oblique                    A    079    aero____
pavc8a    Avenir-Black                      A    080    aebl____
pavco8a   Avenir-BlackOblique               A    080    aeblo___
pavm8a    Avenir-Medium                     A    080    aem_____
pavmo8a   Avenir-MediumOblique              A    080    aemo____
pavk8a    Avenir-Book                       A    080    aew_____
pavko8a   Avenir-BookOblique                A    080    aewo____
pwbb8a    WalbaumBook-Bold                  A    081    wbb_____
pwbbi8a   WalbaumBook-BoldItalic            A    081    wbbi____
pwbri8a   WalbaumBook-Italic                A    081    wbi_____
pwbm8a    WalbaumBook-Medium                A    081    wbm_____
pwbmi8a   WalbaumBook-MediumItalic          A    081    wbmi____
pwbr8a    WalbaumBook-Regular               A    081    wbrg____
paob8a    AntiqueOlive-Bold                 A    082    aqb_____
paoc8a    AntiqueOlive-Black                A    082    aqbl____
paori8a   AntiqueOlive-Italic               A    082    aqi_____
paol8a    AntiqueOlive-Light                A    082    aql_____
paor8a    AntiqueOlive-Roman                A    082    aqr_____
plfb8a    Life-Bold                         A    083    lib_____
plfri8a   Life-Italic                       A    083    lii_____
plfr8a    Life-Roman                        A    083    lir_____
pocri8a   ConcordeBE-Italic                 A    084    qqi_____
pocm8a    ConcordeBE-Medium                 A    084    qqm_____
pocmi8a   ConcordeBE-MediumItalic           A    084    qqmi____
pocr8a    ConcordeBE-Regular                A    084    qqrg____
pg1r8a    Gothic-Thirteen                   A    085    gtt_____
ptph8ac   Tempo-HeavyCondensed              A    085    tehc____
ptphi8ac  Tempo-HeavyCondensedItalic        A    085    tehci___

pcor8a    Cochin                            A    086    cc______
pcob8a    Cochin-Bold                       A    086    ccb_____
pcobi8a   Cochin-BoldItalic                 A    086    ccbi____
pcori8a   Cochin-Italic                     A    086    cci_____
pbhb8a    Bauhaus-Bold                      A    087    bhb_____
pbhd8a    Bauhaus-Demi                      A    087    bhd_____
pbhh8a    Bauhaus-Heavy                     A    087    bhh_____
pbhl8a    Bauhaus-Light                     A    087    bhl_____
pbhm8a    Bauhaus-Medium                    A    087    bhm_____
psbb8a    Sabon-Bold                        A    088    sab_____
psbbi8a   Sabon-BoldItalic                  A    088    sabi____
psbri8a   Sabon-Italic                      A    088    sai_____
psbr8a    Sabon-Roman                       A    088    sar_____
phgb8a    Hiroshige-Bold                    A    089    hrb_____
phgbi8a   Hiroshige-BoldItalic              A    089    hrbi____
phgc8a    Hiroshige-Black                   A    089    hrbl____
phgci8a   Hiroshige-BlackItalic             A    089    hrbli___
phgm8a    Hiroshige-Medium                  A    089    hrm_____
phgmi8a   Hiroshige-MediumItalic            A    089    hrmi____
phgk8a    Hiroshige-Book                    A    089    hrw_____
phgki8a   Hiroshige-BookItalic              A    089    hrwi____
pabr8a    ArnoldBoecklin                    A    090    ab______
pftrf8a   FetteFraktur                      A    090    ff______
pftrf7c   FetteFraktur-Dfr                  A    090    ffdfr___
phir8a    HelveticaInserat-Roman            A    090    hir_____
pptr8a    Present                           A    090    pr______
pdcr8a    DomCasual                         A    091    dc______
pdcb8a    DomCasual-Bold                    A    091    dcb_____
ppqm8a    PostAntiquaBE-Medium              A    092    pbm_____
ppqr8a    PostAntiquaBE-Regular             A    092    pbrg____
pfob8a    Folio-Bold                        A    093    fob_____
pfob8ac   Folio-BoldCondensed               A    093    fobc____
pfox8a    Folio-ExtraBold                   A    093    foeb____
pfol8a    Folio-Light                       A    093    fol_____
pfom8a    Folio-Medium                      A    093    fom_____
plnrw8a   Linoscript                        A    094    lp______
plnrf8a   Linotext                          A    094    lx______
plnrf7c   Linotext-Dfr                      A    094    lxdfr___
pvrb8a    VAGRounded-Bold                   A    095    vrb_____
pvrc8a    VAGRounded-Black                  A    095    vrbl____
pvrl8a    VAGRounded-Light                  A    095    vrl_____
pvra8a    VAGRounded-Thin                   A    095    vrt_____
pazb8a    AkzidenzGroteskBE-Bold            A    096    gfb_____
pazbi8a   AkzidenzGroteskBE-BoldIt          A    096    gfbi____
pazri8a   AkzidenzGroteskBE-It              A    096    gfi_____
pazl8a    AkzidenzGroteskBE-Light           A    096    gfl_____
pazlj8a   AkzidenzGroteskBE-LightOsF        A    096    gflos___

pazm8a    AkzidenzGroteskBE-Md              A    096    gfm_____
pazmi8a   AkzidenzGroteskBE-MdIt            A    096    gfmi____
pazr8a    AkzidenzGroteskBE-Regular         A    096    gfrg____
pazu8a    AkzidenzGroteskBE-Super           A    096    gfsu____
pipb8a    Impressum-Bold                    A    097    imb_____
pipri8a   Impressum-Italic                  A    097    imi_____
pipr8a    Impressum-Roman                   A    097    imr_____
pubb8a    BauerBodoni-Bold                  A    098    bbb_____
pubbi8a   BauerBodoni-BoldItalic            A    098    bbbi____
pubri8a   BauerBodoni-Italic                A    098    bbi_____
pubr8a    BauerBodoni-Roman                 A    098    bbr_____
pnar8a    NewAster                          A    099    na______
pnab8a    NewAster-Bold                     A    099    nab_____
pnabi8a   NewAster-BoldItalic               A    099    nabi____
pnac8a    NewAster-Black                    A    099    nabl____
pnaci8a   NewAster-BlackItalic              A    099    nabli___
pnari8a   NewAster-Italic                   A    099    nai_____
pnas8a    NewAster-SemiBold                 A    099    nas_____
pnasi8a   NewAster-SemiBoldItalic           A    099    nasi____
padb8a    AGaramond-Bold                    A    100    gdb_____
padbi8a   AGaramond-BoldItalic              A    100    gdbi____
padri8a   AGaramond-Italic                  A    100    gdi_____
padr8a    AGaramond-Regular                 A    100    gdrg____
pads8a    AGaramond-Semibold                A    100    gdsb____
padsi8a   AGaramond-SemiboldItalic          A    100    gdsbi___
padbij8a  AGaramond-BoldItalicOsF           A    101    gdbio___
padbj8a   AGaramond-BoldOsF                 A    101    gdbos___
padrij8a  AGaramond-ItalicOsF               A    101    gdio____
padsc8a   AGaramond-SemiboldSC              A    101    gdsbs___
padrc8a   AGaramond-RegularSC               A    101    gdsc____
padsij8a  AGaramond-SemiboldItalicOsF       A    101    gdsis___
padrd8a   AGaramond-Titling                 A    101    gdttl___
padb8x    AGaramondExp-Bold                 A    101    geb_____
padbi8x   AGaramondExp-BoldItalic           A    101    gebi____
padri8x   AGaramondExp-Italic               A    101    gei_____
padr8x    AGaramondExp-Regular              A    101    gerg____
pads8x    AGaramondExp-Semibold             A    101    gesb____
padsi8x   AGaramondExp-SemiboldItalic       A    101    gesbi___
padri7a   AGaramondAlt-Italic               A    101    ggi_____
padr7a    AGaramondAlt-Regular              A    101    ggrg____
pcib8a    Candida-Bold                      A    102    cgb_____
pciri8a   Candida-Italic                    A    102    cgi_____
pcir8a    Candida-Roman                     A    102    cgr_____
pfgr8ac   FranklinGothic-Condensed          A    103    fgc_____
pfgr8aq   FranklinGothic-ExtraCond          A    103    fgec____
pfgr8a    FranklinGothic-Roman              A    103    fgr_____
putb8a    Utopia-Bold                       A    104    utb_____

putbi8a   Utopia-BoldItalic                 A    104    utbi____
putc8a    Utopia-Black                      A    104    utbl____
putri8a   Utopia-Italic                     A    104    uti_____
putr8a    Utopia-Regular                    A    104    utrg____
puts8a    Utopia-Semibold                   A    104    utsb____
putsi8a   Utopia-SemiboldItalic             A    104    utsbi___
putb8x    UtopiaExp-Bold                    A    105    ueb_____
putbi8x   UtopiaExp-BoldItalic              A    105    uebi____
putc8x    UtopiaExp-Black                   A    105    uebl____
putri8x   UtopiaExp-Italic                  A    105    uei_____
putr8x    UtopiaExp-Regular                 A    105    uerg____
puts8x    UtopiaExp-Semibold                A    105    uesb____
putsi8x   UtopiaExp-SemiboldItalic          A    105    uesbi___
putbij8a  Utopia-BoldItalicOsF              A    105    utbio___
putbj8a   Utopia-BoldOsF                    A    105    utbos___
putrij8a  Utopia-ItalicOsF                  A    105    utio____
putrp     Utopia-Ornaments                  A    105    utor____
putrc8a   Utopia-RegularSC                  A    105    utrsc___
putsc8a   Utopia-SemiboldSC                 A    105    utsbs___
putsij8a  Utopia-SemiboldItalicOsF          A    105    utsis___
putrd8a   Utopia-Titling                    A    105    utttl___
putcj8a   Utopia-BlackOsF                   A    105    utzof___
pbyb8a    Berkeley-Bold                     A    106    byb_____
pbybi8a   Berkeley-BoldItalic               A    106    bybi____
pbyc8a    Berkeley-Black                    A    106    bybl____
pbyci8a   Berkeley-BlackItalic              A    106    bybli___
pbymi8a   Berkeley-Italic                   A    106    byi_____
pbym8a    Berkeley-Medium                   A    106    bym_____
pbyk8a    Berkeley-Book                     A    106    byw_____
pbyki8a   Berkeley-BookItalic               A    106    bywi____
paob8ac   AntiqueOlive-BoldCond             A    107    aqbc____
paor8ap   AntiqueOlive-Compact              A    107    aqct____
paohix8a  AntiqueOlive-Nord                 A    107    aqnr____
paohrx8a  AntiqueOlive-NordItalic           A    107    aqnri___
ppsr8a    Parisian                          A    108    pn______
pumr8a    Umbra                             A    108    um______
pmsrw8a   Mistral                           A    109    ms______
prprw8a   Reporter-Two                      A    109    rp______
pccrw8a   CascadeScript                     A    110    ck______
pmirw8a   MediciScript                      A    110    mc______
pnurw8a   NuptialScript                     A    110    nu______
pubb8ac   BauerBodoni-BoldCond              A    111    bbbc____
pubc8a    BauerBodoni-Black                 A    111    bbbl____
pubc8ac   BauerBodoni-BlackCond             A    111    bbblc___
pubci8a   BauerBodoni-BlackItalic           A    111    bbbli___
pmdb8a    Meridien-Bold                     A    112    m0b_____
pmdbi8a   Meridien-BoldItalic               A    112    m0bi____

pmdri8a   Meridien-Italic                   A    112    m0i_____
pmdm8a    Meridien-Medium                   A    112    m0m_____
pmdmi8a   Meridien-MediumItalic             A    112    m0mi____
pmdr8a    Meridien-Roman                    A    112    m0r_____
zpcp29ab  Copperplate-TwentyNineAB          A    113    cp29a___
zpcp29bc  Copperplate-TwentyNineBC          A    113    cp29c___
zpcp30ab  Copperplate-ThirtyAB              A    113    cp30a___
zpcp30bc  Copperplate-ThirtyBC              A    113    cp30c___
zpcp31ab  Copperplate-ThirtyOneAB           A    113    cp31a___
zpcp31bc  Copperplate-ThirtyOneBC           A    113    cp31c___
zpcp32ab  Copperplate-ThirtyTwoAB           A    113    cp32a___
zpcp32bc  Copperplate-ThirtyTwoBC           A    113    cp32c___
zpcp33bc  Copperplate-ThirtyThreeBC         A    113    cp33c___
pnvb8a    Novarese-Bold                     A    114    nob_____
pnvbi8a   Novarese-BoldItalic               A    114    nobi____
pnvm8a    Novarese-Medium                   A    114    nom_____
pnvmi8a   Novarese-MediumItalic             A    114    nomi____
pnvu8a    Novarese-Ultra                    A    114    nou_____
pnvk8a    Novarese-Book                     A    114    now_____
pnvki8a   Novarese-BookItalic               A    114    nowi____
pksbw8a   KuenstlerScript-TwoBold           A    115    kt2b____
pkscw8a   KuenstlerScript-Black             A    115    ktbl____
pksmw8a   KuenstlerScript-Medium            A    115    ktm_____
parr8a    Arcadia                           A    116    a_______
parra8a   Arcadia-A                         A    116    a_aaa___
pinrh8a   Industria-Inline                  A    116    inin____
pinrha8a  Industria-InlineA                 A    116    inina___
piar8a    Insignia                          A    116    ins_____
piara8a   Insignia-A                        A    116    insa____
pinr8a    Industria-Solid                   A    116    insd____
pinra8a   Industria-SolidA                  A    116    insda___
psvb8a    Souvenir-Bold                     A    117    sub_____
psvbi8a   Souvenir-BoldItalic               A    117    subi____
psvm8a    Souvenir-Medium                   A    117    sum_____
psvmi8a   Souvenir-MediumItalic             A    117    sumi____
pbdb8ac   Bodoni-BoldCondensed              A    118    bdbc____
pbdp8ap   Bodoni-PosterCompressed           A    118    bdpsc___
pbdpi8a   Bodoni-PosterItalic               A    118    bdpsi___
pbdk8a    Bodoni-Book                       A    118    bdw_____
pbdki8a   Bodoni-BookItalic                 A    118    bdwi____
prqr8a    RussellSquare                     A    119    ru______
prqro8a   RussellSquare-Oblique             A    119    ruo_____
pcgb8a    Charlemagne-Bold                  A    120    czb_____
pcgr8a    Charlemagne-Regular               A    120    czrg____
ptjb8a    Trajan-Bold                       A    120    tjb_____
ptjr8a    Trajan-Regular                    A    120    tjrg____
plob8a    Lithos-Bold                       A    121    lob_____

ploc8a    Lithos-Black                      A    121    lobl____
ploi8a    Lithos-ExtraLight                 A    121    loel____
plol8a    Lithos-Light                      A    121    lol_____
plor8a    Lithos-Regular                    A    121    lorg____
pcwr8a    Cottonwood                        A    122    c_______
piwr8a    Ironwood                          A    122    i_______
pjpr8a    Juniper                           A    122    j_______
pmqr8a    Mesquite                          A    122    m_______
ppor8a    Ponderosa                         A    122    p_______
ppi026    WoodtypeOrnaments-One             A    122    woor____
ptkb8a    Tekton-Bold                       A    123    tkb_____
ptkbo8a   Tekton-BoldOblique                A    123    tkbo____
ptkro8a   Tekton-Oblique                    A    123    tko_____
ptkr8a    Tekton                            A    123    tkrg____
pagb8ac   AvantGarde-CondBold               A    124    agcb____
pagd8ac   AvantGarde-CondDemi               A    124    agcd____
pagm8ac   AvantGarde-CondMedium             A    124    agcm____
pagk8ac   AvantGarde-CondBook               A    124    agcw____
pssb8a    StempelSchneidler-Bold            A    125    skb_____
pssbi8a   StempelSchneidler-BoldItalic      A    125    skbi____
pssc8a    StempelSchneidler-Black           A    125    skbl____
pssci8a   StempelSchneidler-BlackItalic     A    125    skbli___
pssri8a   StempelSchneidler-Italic          A    125    ski_____
pssl8a    StempelSchneidler-Light           A    125    skl_____
pssli8a   StempelSchneidler-LightItalic     A    125    skli____
pssm8a    StempelSchneidler-Medium          A    125    skm_____
pssmi8a   StempelSchneidler-MedItalic       A    125    skmi____
pssr8a    StempelSchneidler-Roman           A    125    skr_____
ptgr8a    TradeGothic                       A    126    tg______
ptgb8a    TradeGothic-Bold                  A    126    tgb_____
ptgbo8a   TradeGothic-BoldOblique           A    126    tgbo____
ptgc8a    TradeGothic-BoldTwo               A    126    tgb2____
ptgco8a   TradeGothic-BoldTwoOblique        A    126    tgb2o___
ptgl8a    TradeGothic-Light                 A    126    tgl_____
ptglo8a   TradeGothic-LightOblique          A    126    tglo____
ptgro8a   TradeGothic-Oblique               A    126    tgo_____
ptgb8ac   TradeGothic-BoldCondTwenty        A    127    tgbc____
ptgbo8ac  TradeGothic-BoldCondTwentyObl     A    127    tgbco___
ptgr8ac   TradeGothic-CondEighteen          A    127    tgc_____
ptgro8ac  TradeGothic-CondEighteenObl       A    127    tgco____
prar8a    Raleigh                           A    128    ra______
prab8a    Raleigh-Bold                      A    128    rab_____
prad8a    Raleigh-DemiBold                  A    128    rad_____
pram8a    Raleigh-Medium                    A    128    ram_____
pmtb8a    Minister-Bold                     A    129    mzb_____
pmtbi8a   Minister-BoldItalic               A    129    mzbi____
pmtc8a    Minister-Black                    A    129    mzbl____

pmtci8a   Minister-BlackItalic              A    129    mzbli___
pmtl8a    Minister-Light                    A    129    mzl_____
pmtli8a   Minister-LightItalic              A    129    mzli____
pmtk8a    Minister-Book                     A    129    mzw_____
pmtki8a   Minister-BookItalic               A    129    mzwi____
pesb8ac   Eurostile-BoldCondensed           A    130    eubc____
pesb8ax   Eurostile-BoldExtendedTwo         A    130    eubex___
pesr8ac   Eurostile-Condensed               A    130    euc_____
pesr8ax   Eurostile-ExtendedTwo             A    130    euex____
pwdb8a    Weidemann-Bold                    A    131    wdb_____
pwdbi8a   Weidemann-BoldItalic              A    131    wdbi____
pwdc8a    Weidemann-Black                   A    131    wdbl____
pwdci8a   Weidemann-BlackItalic             A    131    wdbli___
pwdm8a    Weidemann-Medium                  A    131    wdm_____
pwdmi8a   Weidemann-MediumItalic            A    131    wdmi____
pwdk8a    Weidemann-Book                    A    131    wdw_____
pwdki8a   Weidemann-BookItalic              A    131    wdwi____
pqrb8a    Quorum-Bold                       A    132    qob_____
pqrc8a    Quorum-Black                      A    132    qobl____
pqrl8a    Quorum-Light                      A    132    qol_____
pqrm8a    Quorum-Medium                     A    132    qom_____
pqrk8a    Quorum-Book                       A    132    qow_____
pcxb8a    Caxton-Bold                       A    133    cxb_____
pcxbi8a   Caxton-BoldItalic                 A    133    cxbi____
pcxl8a    Caxton-Light                      A    133    cxl_____
pcxli8a   Caxton-LightItalic                A    133    cxli____
pcxk8a    Caxton-Book                       A    133    cxw_____
pcxki8a   Caxton-BookItalic                 A    133    cxwi____
psxb8a    Syntax-Bold                       A    134    sxb_____
psxc8a    Syntax-Black                      A    134    sxbl____
psxri8a   Syntax-Italic                     A    134    sxi_____
psxr8a    Syntax-Roman                      A    134    sxr_____
psxu8a    Syntax-UltraBlack                 A    134    sxubl___
pdrb8a    Doric-Bold                        A    135    drb_____
pmxr8a    Maximus                           A    135    mx______
pserd8a   Shelley-AllegroScript             A    136    shal____
pser8a    Shelley-AndanteScript             A    136    shan____
pserw8a   Shelley-VolanteScript             A    136    shvo____
pnzk8a    NeuzeitS-Book                     A    137    new_____
pnzh8a    NeuzeitS-BookHeavy                A    137    newh____
pvlb8a    Versailles-Bold                   A    138    veb_____
pvlbi8a   Versailles-BoldItalic             A    138    vebi____
pvlc8a    Versailles-Black                  A    138    vebl____
pvlci8a   Versailles-BlackItalic            A    138    vebli___
pvlri8a   Versailles-Italic                 A    138    vei_____
pvll8a    Versailles-Light                  A    138    vel_____
pvlli8a   Versailles-LightItalic            A    138    veli____

pvlr8a    Versailles-Roman                  A    138    ver_____
pgmb8ac   Garamond-BoldCondensed            A    139    gabc____
pgmbi8ac  Garamond-BoldCondensedItalic      A    139    gabci___
pgml8ac   Garamond-LightCondensed           A    139    galc____
pgmlu8ac  Garamond-LightCondensedItalic     A    139    galci___
pgmu8ac   Garamond-UltraCondensed           A    139    gauc____
pgmui8ac  Garamond-UltraCondensedItalic     A    139    gauci___
pgmk8ac   Garamond-BookCondensed            A    139    gawc____
pgmki8ac  Garamond-BookCondensedItalic      A    139    gawci___
pctl8a    Cheltenham-Light                  A    140    chl_____
pctli8a   Cheltenham-LightItalic            A    140    chli____
pctu8a    Cheltenham-Ultra                  A    140    chu_____
pctui8a   Cheltenham-UltraItalic            A    140    chui____
pcnb8a    Century-Bold                      A    141    icb_____
pcnbi8a   Century-BoldItalic                A    141    icbi____
pcnl8a    Century-Light                     A    141    icl_____
pcnli8a   Century-LightItalic               A    141    icli____
pcnu8a    Century-Ultra                     A    141    icu_____
pcnui8a   Century-UltraItalic               A    141    icui____
pcnk8a    Century-Book                      A    141    icw_____
pcnki8a   Century-BookItalic                A    141    icwi____
pcnb8ac   Century-BoldCondensed             A    142    icbc____
pcnbi8ac  Century-BoldCondensedItalic       A    142    icbci___
pcnl8ac   Century-LightCondensed            A    142    iclc____
pcnli8ac  Century-LightCondensedItalic      A    142    iclci___
pcnu8ac   Century-UltraCondensed            A    142    icuc____
pcnui8ac  Century-UltraCondensedItalic      A    142    icuci___
pcnk8ac   Century-BookCondensed             A    142    icwc____
pcnki8ac  Century-BookCondensedItalic       A    142    icwci___
pmnb8a    Minion-Bold                       A    143    mob_____
pmnbi8a   Minion-BoldItalic                 A    143    mobi____
pmnc8a    Minion-Black                      A    143    mobl____
pmnrd8a   Minion-DisplayRegular             A    143    mods____
pmnrid8a  Minion-DisplayItalic              A    143    modsi___
pmnri8a   Minion-Italic                     A    143    moi_____
pmnr8a    Minion-Regular                    A    143    morg____
pmns8a    Minion-Semibold                   A    143    mosb____
pmnsi8a   Minion-SemiboldItalic             A    143    mosbi___
pmnb8x    MinionExp-Bold                    A    144    mjb_____
pmnbi8x   MinionExp-BoldItalic              A    144    mjbi____
pmnc8x    MinionExp-Black                   A    144    mjbl____
pmnrd8x   MinionExp-DisplayRegular          A    144    mjds____
pmnrid8x  MinionExp-DisplayItalic           A    144    mjdsi___
pmnri8x   MinionExp-Italic                  A    144    mji_____
pmnr8x    MinionExp-Regular                 A    144    mjrg____
pmns8x    MinionExp-Semibold                A    144    mjsb____
pmnsi8x   MinionExp-SemiboldItalic          A    144    mjsbi___

pmnbij8a  Minion-BoldItalicOsF              A    144    mobio___
pmnbj8a   Minion-BoldOsF                    A    144    mobos___
pmnricd8a Minion-DisplayItalicSC            A    144    modis___
pmnrcd8a  Minion-DisplayRegularSC           A    144    modsc___
pmnric8a  Minion-ItalicSC                   A    144    moisc___
pmnrp     Minion-Ornaments                  A    144    moor____
pmnsc8a   Minion-SemiboldSC                 A    144    mosbs___
pmnrc8a   Minion-RegularSC                  A    144    mosc____
pmnriwd7a Minion-SwashDisplayItalic         A    144    mosdi___
pmnsic8a  Minion-SemiboldItalicSC           A    144    mosic___
pmnsiw7a  Minion-SwashSemiboldItalic        A    144    mossb___
pmnriw7a  Minion-SwashItalic                A    144    moswi___
pmncj8a   Minion-BlackOsF                   A    144    mozof___
mnsb8a    TimesNewRomanPS-Bold              A    145    mtb_____
mnsbi8a   TimesNewRomanPS-BoldItalic        A    145    mtbi____
mnsri8a   TimesNewRomanPS-Italic            A    145    mti_____
pnsr8a    TimesNewRomanPS                   A    145    mtr_____
pbkb8a    Bookman-Bold                      A    146    bkb_____
pbkbi8a   Bookman-BoldItalic                A    146    bkbi____
pbkm8a    Bookman-Medium                    A    146    bkm_____
pbkmi8a   Bookman-MediumItalic              A    146    bkmi____
psrrw8a   SnellRoundhand-Script             A    147    sn______
psrbw8a   SnellRoundhand-BoldScript         A    147    snb_____
psrcw8a   SnellRoundhand-BlackScript        A    147    snbl____
pigb8a    SimonciniGaramond-Bold            A    148    gib_____
pigri8a   SimonciniGaramond-Italic          A    148    gii_____
pigr8a    SimonciniGaramond                 A    148    gir_____
pbcr8a    Banco                             A    149    bc______
phmr8a    Charme                            A    149    cu______
pfyc8ac   Flyer-BlackCondensed              A    149    flblc___
pfyx8ac   Flyer-ExtraBlackCondensed         A    149    flebc___
pwgrf8a   WilhelmKlingsporGotisch           A    149    wk______
pwgrf7c   WilhelmKlingsporGotisch-Dfr       A    149    wkdfr___
pgmu8a    Garamond-Ultra                    A    150    gau_____
pgmui8a   Garamond-UltraItalic              A    150    gaui____
pgmk8a    Garamond-Book                     A    150    gaw_____
pgmki8a   Garamond-BookItalic               A    150    gawi____
prwr8a    Rockwell                          A    151    rk______
prwb8a    Rockwell-Bold                     A    151    rkb_____
prwbi8a   Rockwell-BoldItalic               A    151    rkbi____
prwri8a   Rockwell-Italic                   A    151    rki_____
prwl8a    Rockwell-Light                    A    151    rkl_____
prwli8a   Rockwell-LightItalic              A    151    rkli____
pgsr8a    GillSans                          A    152    gn______
pgsb8a    GillSans-Bold                     A    152    gnb_____
pgsbi8a   GillSans-BoldItalic               A    152    gnbi____
pgsri8a   GillSans-Italic                   A    152    gni_____

pgsl8a    GillSans-Light                    A    152    gnl_____
pgsli8a   GillSans-LightItalic              A    152    gnli____
pspb8a    Serpentine-Bold                   A    153    sfb_____
pspbo8a   Serpentine-BoldOblique            A    153    sfbo____
pspl8a    Serpentine-Light                  A    153    sfl_____
psplo8a   Serpentine-LightOblique           A    153    sflo____
pspm8a    Serpentine-Medium                 A    153    sfm_____
pspmo8a   Serpentine-MediumOblique          A    153    sfmo____
plcrv     LucidaMath-Extension              A    154    lme_____
plcrim    LucidaMath-Italic                 A    154    lmi_____
plcry     LucidaMath-Symbol                 A    154    lms_____
ppi005    EuropeanPi-One                    A    155    euro1___
ppi006    EuropeanPi-Two                    A    155    euro2___
ppi007    EuropeanPi-Three                  A    155    euro3___
ppi008    EuropeanPi-Four                   A    155    euro4___
ppi001    BorderPi-OneFiveOneFiveNine       A    155    j1515___
ppi002    BundesbahnPi-One                  A    156    db1_____
ppi003    BundesbahnPi-Two                  A    156    db2_____
ppi004    BundesbahnPi-Three                A    156    db3_____
pc2b8a    CaslonTwoTwentyFour-Bold          A    157    c2b_____
pc2bi8a   CaslonTwoTwentyFour-BoldIt        A    157    c2bi____
pc2c8a    CaslonTwoTwentyFour-Black         A    157    c2bl____
pc2ci8a   CaslonTwoTwentyFour-BlackIt       A    157    c2bli___
pc2m8a    CaslonTwoTwentyFour-Medium        A    157    c2m_____
pc2mi8a   CaslonTwoTwentyFour-MediumIt      A    157    c2mi____
pc2k8a    CaslonTwoTwentyFour-Book          A    157    c2w_____
pc2ki8a   CaslonTwoTwentyFour-BookIt        A    157    c2wi____
ppi020    MathematicalPi-One                A    158    mh1_____
ppi021    MathematicalPi-Two                A    158    mh2_____
ppi022    MathematicalPi-Three              A    158    mh3_____
ppi023    MathematicalPi-Four               A    158    mh4_____
ppi024    MathematicalPi-Five               A    158    mh5_____
ppi025    MathematicalPi-Six                A    158    mh6_____
plwb8a    Leawood-Bold                      A    159    lwb_____
plwbi8a   Leawood-BoldItalic                A    159    lwbi____
plwc8a    Leawood-Black                     A    159    lwbl____
plwci8a   Leawood-BlackItalic               A    159    lwbli___
plwm8a    Leawood-Medium                    A    159    lwm_____
plwmi8a   Leawood-MediumItalic              A    159    lwmi____
plwk8a    Leawood-Book                      A    159    lww_____
plwki8a   Leawood-BookItalic                A    159    lwwi____
pbir8a    Birch                             A    160    bi______
pb0r8a    Blackoak                          A    160    bo______
pmrr8a    Madrone                           A    160    mq______
pp0r8a    Poplar                            A    160    pp______
pwwr8a    Willow                            A    160    wi______
ppi027    WoodtypeOrnaments-Two             A    160    woor2___

pbbr8a    Bembo                             A    161    bm______
pbbb8a    Bembo-Bold                        A    161    bmb_____
pbbbi8a   Bembo-BoldItalic                  A    161    bmbi____
pbbri8a   Bembo-Italic                      A    161    bmi_____
pgsb8ac   GillSans-BoldCondensed            A    162    gnbc____
pgsr8ac   GillSans-Condensed                A    162    gnc_____
pgsx8a    GillSans-ExtraBold                A    162    gneb____
pgsu8a    GillSans-UltraBold                A    162    gnub____
pgsu8ac   GillSans-UltraBoldCondensed       A    162    gnubc___
prwb8ac   Rockwell-BoldCondensed            A    163    rkbc____
prwr8ac   Rockwell-Condensed                A    163    rkc_____
prwx8a    Rockwell-ExtraBold                A    163    rkeb____
pgpc8a    Glypha-Black                      A    164    gybl____
pgpco8a   Glypha-BlackOblique               A    164    gyblo___
pgpl8a    Glypha-Light                      A    164    gyl_____
pgplo8a   Glypha-LightOblique               A    164    gylo____
pgpa8a    Glypha-Thin                       A    164    gyt_____
pgpao8a   Glypha-ThinOblique                A    164    gyto____
pzcb8a    ZapfChancery-Bold                 A    165    zcb_____
pzcd8a    ZapfChancery-Demi                 A    165    zcd_____
pzcri8a   ZapfChancery-Italic               A    165    zci_____
pzcl8a    ZapfChancery-Light                A    165    zcl_____
pzcli8a   ZapfChancery-LightItalic          A    165    zcli____
pzcr8a    ZapfChancery-Roman                A    165    zcr_____
puwb8a    Usherwood-Bold                    A    166    usb_____
puwbi8a   Usherwood-BoldItalic              A    166    usbi____
puwc8a    Usherwood-Black                   A    166    usbl____
puwci8a   Usherwood-BlackItalic             A    166    usbli___
puwm8a    Usherwood-Medium                  A    166    usm_____
puwmi8a   Usherwood-MediumItalic            A    166    usmi____
puwk8a    Usherwood-Book                    A    166    usw_____
puwki8a   Usherwood-BookItalic              A    166    uswi____
isyb8a    ItcSymbol-Bold                    A    167    isb_____
isybi8a   ItcSymbol-BoldItalic              A    167    isbi____
isyc8a    ItcSymbol-Black                   A    167    isbl____
isyci8a   ItcSymbol-BlackItalic             A    167    isbli___
isym8a    ItcSymbol-Medium                  A    167    ism_____
isymi8a   ItcSymbol-MediumItalic            A    167    ismi____
isyk8a    ItcSymbol-Book                    A    167    isw_____
isyki8a   ItcSymbol-BookItalic              A    167    iswi____
punb8ax   Univers-BoldExt                   A    168    uvbv____
punbo8ax  Univers-BoldExtObl                A    168    uvbvo___
punx8ax   Univers-ExtraBlackExt             A    168    uvfv____
punxo8ax  Univers-ExtraBlackExtObl          A    168    uvfvo___
punr8ax   Univers-Extended                  A    168    uvv_____
punro8ax  Univers-ExtendedObl               A    168    uvvo____
punc8ax   Univers-BlackExt                  A    168    uvzv____

punco8ax  Univers-BlackExtObl               A    168    uvzvo___
poyr8a    Olympian                          A    169    ol______
poyb8a    Olympian-Bold                     A    169    olb_____
poybi8a   Olympian-BoldItalic               A    169    olbi____
poyri8a   Olympian-Italic                   A    169    oli_____
puhb8a    Cushing-Bold                      A    170    iub_____
puhbi8a   Cushing-BoldItalic                A    170    iubi____
puhh8a    Cushing-Heavy                     A    170    iuh_____
puhhi8a   Cushing-HeavyItalic               A    170    iuhi____
puhm8a    Cushing-Medium                    A    170    ium_____
puhmi8a   Cushing-MediumItalic              A    170    iumi____
puhk8a    Cushing-Book                      A    170    iuw_____
puhki8a   Cushing-BookItalic                A    170    iuwi____
pcqr      Cheq                              A    171    cq______
ps7rp     Sassoon-Primary                   A    172    sass____
pfnb8a    Fenice-Bold                       A    173    feb_____
pfnbo8a   Fenice-BoldOblique                A    173    febo____
pfnl8a    Fenice-Light                      A    173    fel_____
pfnlo8a   Fenice-LightOblique               A    173    felo____
pfnr8a    Fenice-Regular                    A    173    ferg____
pfnro8a   Fenice-RegularOblique             A    173    fergo___
pfnu8a    Fenice-Ultra                      A    173    feu_____
pfnuo8a   Fenice-UltraOblique               A    173    feuo____
pctb8ac   Cheltenham-BoldCond               A    174    chbc____
pctbi8ac  Cheltenham-BoldCondItalic         A    174    chbci___
pctl8ac   Cheltenham-LightCond              A    174    chlc____
pctli8ac  Cheltenham-LightCondItalic        A    174    chlci___
pctu8ac   Cheltenham-UltraCond              A    174    chuc____
pctui8ac  Cheltenham-UltraCondItalic        A    174    chuci___
pctk8ac   Cheltenham-BookCond               A    174    chwc____
pctki8ac  Cheltenham-BookCondItalic         A    174    chwci___
phrb8a    HelveticaRounded-Bold             A    175    heb_____
phrb8ac   HelveticaRounded-BoldCond         A    175    hebc____
phrbo8ac  HelveticaRounded-BoldCondObl      A    175    hebco___
phrc8a    HelveticaRounded-Black            A    175    hebl____
phrco8a   HelveticaRounded-BlackObl         A    175    heblo___
phrbo8a   HelveticaRounded-BoldObl          A    175    hebo____
pehra8a   DINEngschrift-Alternate           A    176    dinea___
pehr8a    DINEngschrift                     A    176    dinen___
pm6ra8a   DINMittelschrift-Alternate        A    176    dinma___
pm6r8a    DINMittelschrift                  A    176    dinmi___
pzgb8ac   DINNeuzeitGrotesk-BoldCond        A    176    dngbc___
pzgl8a    DINNeuzeitGrotesk-Light           A    176    dngl____
psnh8a    Spartan-HeavyClassified           A    177    sqhcl___
psnk8a    Spartan-BookClassified            A    177    sqwcl___
pugb8a    BenguiatGothic-Bold               A    178    beb_____
pugbo8a   BenguiatGothic-BoldOblique        A    178    bebo____

pugh8a    BenguiatGothic-Heavy              A    178    beh_____
pugho8a   BenguiatGothic-HeavyOblique       A    178    beho____
pugm8a    BenguiatGothic-Medium             A    178    bem_____
pugmo8a   BenguiatGothic-MediumOblique      A    178    bemo____
pugk8a    BenguiatGothic-Book               A    178    bew_____
pugko8a   BenguiatGothic-BookOblique        A    178    bewo____
ptmx8a    Times-ExtraBold                   A    179    tieb____
ptms8a    Times-Semibold                    A    179    tisb____
ptmsi8a   Times-SemiboldItalic              A    179    tisbi___
pacb8a    ACaslon-Bold                      A    180    awb_____
pacbi8a   ACaslon-BoldItalic                A    180    awbi____
pacri8a   ACaslon-Italic                    A    180    awi_____
pacr8a    ACaslon-Regular                   A    180    awrg____
pacs8a    ACaslon-Semibold                  A    180    awsb____
pacsi8a   ACaslon-SemiboldItalic            A    180    awsbi___
pacb7a    ACaslon-AltBold                   A    181    awab____
pacbi7a   ACaslon-AltBoldItalic             A    181    awabi___
pacri7a   ACaslon-AltItalic                 A    181    awai____
pacr7a    ACaslon-AltRegular                A    181    awarg___
pacs7a    ACaslon-AltSemibold               A    181    awasb___
pacsi7a   ACaslon-AltSemiboldItalic         A    181    awasi___
pacbij8a  ACaslon-BoldItalicOsF             A    181    awbio___
pacbj8a   ACaslon-BoldOsF                   A    181    awbos___
pacrij8a  ACaslon-ItalicOsF                 A    181    awio____
pacrp     ACaslon-Ornaments                 A    181    awor____
pacsc8a   ACaslon-SemiboldSC                A    181    awsbs___
pacrc8a   ACaslon-RegularSC                 A    181    awsc____
pacsic8a  ACaslon-SemiboldItalicOsF         A    181    awsis___
pacsiw7a  ACaslon-SwashSemiboldItalic       A    181    awssb___
pacbiw7a  ACaslon-SwashBoldItalic           A    181    awswb___
pacriw7a  ACaslon-SwashItalic               A    181    awswi___
pacb8x    ACaslonExp-Bold                   A    181    axb_____
pacbi8x   ACaslonExp-BoldItalic             A    181    axbi____
pacri8x   ACaslonExp-Italic                 A    181    axi_____
pacr8x    ACaslonExp-Regular                A    181    axrg____
pacs8x    ACaslonExp-Semibold               A    181    axsb____
pacsi8x   ACaslonExp-SemiboldItalic         A    181    axsbi___
phnj8ax   HelveticaNeue-UltraLigExt         A    182    hlav____
phnjo8ax  HelveticaNeue-UltraLigExtObl      A    182    hlavo___
phnc8ax   HelveticaNeue-BlackExt            A    182    hlzv____
phnco8ax  HelveticaNeue-BlackExtObl         A    182    hlzvo___
phnb8ax   HelveticaNeue-BoldExt             A    183    hlbv____
phnbo8ax  HelveticaNeue-BoldExtObl          A    183    hlbvo___
phna8ax   HelveticaNeue-ThinExt             A    183    hltv____
phnao8ax  HelveticaNeue-ThinExtObl          A    183    hltvo___
phnr8ax   HelveticaNeue-Extended            A    183    hlv_____
phnro8ax  HelveticaNeue-ExtendedObl         A    183    hlvo____

phnh8ax   HelveticaNeue-HeavyExt            A    184    hlhv____
phnho8ax  HelveticaNeue-HeavyExtObl         A    184    hlhvo___
phnl8ax   HelveticaNeue-LightExt            A    184    hllv____
phnlo8ax  HelveticaNeue-LightExtObl         A    184    hllvo___
phnm8ax   HelveticaNeue-MediumExt           A    184    hlmv____
phnmo8ax  HelveticaNeue-MediumExtObl        A    184    hlmvo___
pshb8a    Shannon-Bold                      A    185    sbb_____
pshx8a    Shannon-ExtraBold                 A    185    sbeb____
pshro8a   Shannon-Oblique                   A    185    sbo_____
pshk8a    Shannon-Book                      A    185    sbw_____
pbbx8a    Bembo-ExtraBold                   A    186    bmeb____
pbbxi8a   Bembo-ExtraBoldItalic             A    186    bmebi___
pbbs8a    Bembo-Semibold                    A    186    bmsb____
pbbsi8a   Bembo-SemiboldItalic              A    186    bmsbi___
pggr8a    GarthGraphic                      A    187    gh______
pggb8a    GarthGraphic-Bold                 A    187    ghb_____
pggb8ac   GarthGraphic-BoldCondensed        A    187    ghbc____
pggbi8a   GarthGraphic-BoldItalic           A    187    ghbi____
pggr8ac   GarthGraphic-Condensed            A    187    ghc_____
pggx8a    GarthGraphic-ExtraBold            A    187    gheb____
pggri8a   GarthGraphic-Italic               A    187    ghi_____
pggc8a    GarthGraphic-Black                A    187    ghz_____
phnj8ac   HelveticaNeue-UltraLigCond        A    188    hla_____
phnjo8ac  HelveticaNeue-UltraLigCondObl     A    188    hlao____
phnx8ac   HelveticaNeue-ExtBlackCond        A    188    hlj_____
phnxo8ac  HelveticaNeue-ExtBlackCondObl     A    188    hljo____
phnc8ac   HelveticaNeue-BlackCond           A    188    hlzc____
phnco8ac  HelveticaNeue-BlackCondObl        A    188    hlzco___
phnb8ac   HelveticaNeue-BoldCond            A    189    hlbc____
phnbo8ac  HelveticaNeue-BoldCondObl         A    189    hlbco___
phnr8ac   HelveticaNeue-Condensed           A    189    hlc_____
phnro8ac  HelveticaNeue-CondensedObl        A    189    hlco____
phna8ac   HelveticaNeue-ThinCond            A    189    hltc____
phnao8ac  HelveticaNeue-ThinCondObl         A    189    hltco___
phnh8ac   HelveticaNeue-HeavyCond           A    190    hlhc____
phnho8ac  HelveticaNeue-HeavyCondObl        A    190    hlhco___
phnl8ac   HelveticaNeue-LightCond           A    190    hllc____
phnlo8ac  HelveticaNeue-LightCondObl        A    190    hllco___
phnm8ac   HelveticaNeue-MediumCond          A    190    hlmc____
phnmo8ac  HelveticaNeue-MediumCondObl       A    190    hlmco___
punx8a    Univers-ExtraBlack                A    191    uvf_____
punxo8a   Univers-ExtraBlackObl             A    191    uvfo____
punl8ao   Univers-LightUltraCondensed       A    191    uvluc___
puna8ao   Univers-ThinUltraCondensed        A    191    uvtuc___
punr8ao   Univers-UltraCondensed            A    191    uvuc____
pglc8a    Galliard-Black                    A    192    glbl____
pglci8a   Galliard-BlackItalic              A    192    glbli___

pglu8a    Galliard-Ultra                    A    192    glu_____
pglui8a   Galliard-UltraItalic              A    192    glui____
pg3bij8a  GaramondThree-BoldItalicOsF       A    193    gmbio___
pg3bj8a   GaramondThree-BoldSC              A    193    gmbsc___
pg3rij8a  GaramondThree-ItalicOsF           A    193    gmio____
pg3rc8a   GaramondThree-SC                  A    193    gmsc____
ptmbij8a  Times-BoldItalicOsF               A    194    tibio___
ptmbc8a   Times-BoldSC                      A    194    tibsc___
ptmrij8a  Times-ItalicOsF                   A    194    tiio____
ptmrc8a   Times-RomanSC                     A    194    tirsc___
pddrf8a   DucDeBerry                        A    195    du______
pddrf7c   DucDeBerry-Dfr                    A    195    dudfr___
phcr8a    Herculanum                        A    195    herc____
pomr8a    Omnia                             A    195    omni____
pttbij8a  TimesTen-BoldItalicOsF            A    196    ttbio___
pttbj8a   TimesTen-BoldOsF                  A    196    ttbof___
pttrij8a  TimesTen-ItalicOsF                A    196    ttiof___
pttrc8a   TimesTen-RomanSC                  A    196    ttrsc___
psbbij8a  Sabon-BoldItalicOsF               A    197    sabio___
psbbj8a   Sabon-BoldOsF                     A    197    sabof___
psbrij8a  Sabon-ItalicOsF                   A    197    saiof___
psbrc8a   Sabon-RomanSC                     A    197    sarsc___
pwbbij8a  WalbaumBook-BoldItalicOsF         A    198    wbbio___
pwbbi8x   WalbaumBook-BoldItalicExpert      A    198    wbbix___
pwbbj8a   WalbaumBook-BoldOsF               A    198    wbbos___
pwbb8x    WalbaumBook-BoldExpert            A    198    wbbxp___
pwbrij8a  WalbaumBook-ItalicOsF             A    198    wbios___
pwbri8x   WalbaumBook-ItalicExpert          A    198    wbixp___
pwbmij8a  WalbaumBook-MdItalicOsF           A    198    wbmio___
pwbmi8x   WalbaumBook-MdItalicExpert        A    198    wbmix___
pwbmc8a   WalbaumBook-MediumSC              A    198    wbmsc___
pwbm8x    WalbaumBook-MdExpert              A    198    wbmxp___
pwbr8x    WalbaumBook-RegularExpert         A    198    wbrgx___
pwbrc8a   WalbaumBook-RegularSC             A    198    wbsc____
pubbij8a  BauerBodoni-BoldItalicOsF         A    199    bbbio___
pubbj8a   BauerBodoni-BoldOsF               A    199    bbbof___
pubrij8a  BauerBodoni-ItalicOsF             A    199    bbiof___
pubrc8a   BauerBodoni-RomanSC               A    199    bbrsc___
pmhr8a    Machine                           A    200    ma______
pmhb8a    Machine-Bold                      A    200    mab_____
pr1r8a    RotisSemiSerif                    A    201    rf______
pr1b8a    RotisSemiSerif-Bold               A    201    rfb_____
pror8a    RotisSerif                        A    201    rg______
prob8a    RotisSerif-Bold                   A    201    rgb_____
prori8a   RotisSerif-Italic                 A    201    rgi_____
pr3r8a    RotisSemiSans                     A    202    rc______
pr3b8a    RotisSemiSans-Bold                A    202    rcb_____

pr3x8a    RotisSemiSans-ExtraBold           A    202    rceb____
pr3ri8a   RotisSemiSans-Italic              A    202    rci_____
pr3l8a    RotisSemiSans-Light               A    202    rcl_____
pr3li8a   RotisSemiSans-LightItalic         A    202    rcli____
pr4r8a    RotisSansSerif                    A    203    rb______
pr4b8a    RotisSansSerif-Bold               A    203    rbb_____
pr4x8a    RotisSansSerif-ExtraBold          A    203    rbeb____
pr4ri8a   RotisSansSerif-Italic             A    203    rbi_____
pr4l8a    RotisSansSerif-Light              A    203    rbl_____
pr4li8a   RotisSansSerif-LightItalic        A    203    rbli____
pjnbij8a  JansonText-BoldItalicOsF          A    204    jtbio___
pjnbj8a   JansonText-BoldOsF                A    204    jtbof___
pjnrij8a  JansonText-ItalicOsF              A    204    jtiof___
pjnrc8a   JansonText-RomanSC                A    204    jtrsc___
pgjr8a    Granjon                           A    205    gr______
pgjb8a    Granjon-Bold                      A    205    grb_____
pgjri8a   Granjon-Italic                    A    205    gri_____
pcebij8a  Centennial-BoldItalicOsF          A    206    cibio___
pcebj8a   Centennial-BoldOsF                A    206    cibof___
pcerij8a  Centennial-ItalicOsF              A    206    ciiof___
pcelij8a  Centennial-LightItalicOsF         A    206    cilio___
pcelc8a   Centennial-LightSC                A    206    cilsc___
pcerc8a   Centennial-RomanSC                A    206    cirsc___
pcecij8a  Centennial-BlackItalicOsF         A    206    cizio___
pcecj8a   Centennial-BlackOsF               A    206    cizof___
ppnr8a    Plantin                           A    207    pl______
ppnb8a    Plantin-Bold                      A    207    plb_____
ppnbi8a   Plantin-BoldItalic                A    207    plbi____
ppnri8a   Plantin-Italic                    A    207    pli_____
pteb8a    TimesEuropa-Bold                  A    208    tub_____
ptebi8a   TimesEuropa-BoldItalic            A    208    tubi____
pteri8a   TimesEuropa-Italic                A    208    tui_____
pter8a    TimesEuropa-Roman                 A    208    tur_____
pgjbj8a   Granjon-BoldOsF                   A    209    grbos___
pgjrij8a  Granjon-ItalicOsF                 A    209    grios___
pgjrc8a   Granjon-SC                        A    209    grsc____
ptvbij8a  TrumpMediaeval-BoldItalicOsF      A    210    tmbio___
ptvbj8a   TrumpMediaeval-BoldOsF            A    210    tmbos___
ptvrij8a  TrumpMediaeval-ItalicOsF          A    210    tmios___
ptvrc8a   TrumpMediaeval-SC                 A    210    tmsc____
pbsr8a    BakerSignet                       A    211    ba______
picr8a    Impact                            A    211    impa____
pagb8a    AvantGarde-Bold                   A    212    agb_____
pagbo8a   AvantGarde-BoldObl                A    212    agbo____
pagj8a    AvantGarde-ExtraLight             A    212    agel____
pagjo8a   AvantGarde-ExtraLightObl          A    212    agelo___
pagm8a    AvantGarde-Medium                 A    212    agm_____

pagmo8a   AvantGarde-MediumObl              A    212    agmo____
pplbij8a  Palatino-BoldItalicOsF            A    213    pobio___
pplbj8a   Palatino-BoldOsF                  A    213    pobos___
pplrij8a  Palatino-ItalicOsF                A    213    poios___
pplrc8a   Palatino-SC                       A    213    posc____
pndbij8a  NewCaledonia-BoldItalicOsF        A    214    nlbio___
pndbc8a   NewCaledonia-BoldSC               A    214    nlbsc___
pndrij8a  NewCaledonia-ItalicOsF            A    214    nlios___
pndrc8a   NewCaledonia-SC                   A    214    nlsc____
pppb8a    Perpetua-Bold                     A    215    ptb_____
pppbi8a   Perpetua-BoldItalic               A    215    ptbi____
pppri8a   Perpetua-Italic                   A    215    pti_____
pppr8a    Perpetua                          A    215    ptrg____
pilb8a    Caecilia-Bold                     A    216    ecb_____
pilbi8a   Caecilia-BoldItalic               A    216    ecbi____
pilbij8a  Caecilia-BoldItalicOsF            A    216    ecbio___
pilbj8a   Caecilia-BoldOsF                  A    216    ecbos___
pilh8a    Caecilia-Heavy                    A    216    ech_____
pilhi8a   Caecilia-HeavyItalic              A    216    echi____
pilhij8a  Caecilia-HeavyItalicOsF           A    216    echio___
pilhj8a   Caecilia-HeavyOsF                 A    216    echos___
pilri8a   Caecilia-Italic                   A    216    eci_____
pilrij8a  Caecilia-ItalicOsF                A    216    ecios___
pill8a    Caecilia-Light                    A    216    ecl_____
pilli8a   Caecilia-LightItalic              A    216    ecli____
pillij8a  Caecilia-LightItalicOsF           A    216    eclio___
pillj8a   Caecilia-LightOsF                 A    216    eclos___
pilr8a    Caecilia-Roman                    A    216    ecr_____
pilrj8a   Caecilia-RomanOsF                 A    216    ecros___
pimx8a    Imago-ExtraBold                   A    217    ioeb____
pimxi8a   Imago-ExtraBoldItalic             A    217    ioebi___
piml8a    Imago-Light                       A    217    iol_____
pimli8a   Imago-LightItalic                 A    217    ioli____
pimm8a    Imago-Medium                      A    217    iom_____
pimmi8a   Imago-MediumItalic                A    217    iomi____
pimk8a    Imago-Book                        A    217    iow_____
pimki8a   Imago-BookItalic                  A    217    iowi____
pfab8a    Formata-Bold                      A    218    fmb_____
pfabi8a   Formata-BoldItalic                A    218    fmbi____
pfari8a   Formata-Italic                    A    218    fmi_____
pfal8a    Formata-Light                     A    218    fml_____
pfali8a   Formata-LightItalic               A    218    fmli____
pfam8a    Formata-Medium                    A    218    fmm_____
pfami8a   Formata-MediumItalic              A    218    fmmi____
pfar8a    Formata-Regular                   A    218    fmrg____
pnfb8a    Nofret-Bold                       A    219    nfb_____
pnfbi8a   Nofret-BoldItalic                 A    219    nfbi____

pnfri8a   Nofret-Italic                     A    219    nfi_____
pnfl8a    Nofret-Light                      A    219    nfl_____
pnfli8a   Nofret-LightItalic                A    219    nfli____
pnfm8a    Nofret-Medium                     A    219    nfm_____
pnfmi8a   Nofret-MediumItalic               A    219    nfmi____
pnfr8a    Nofret-Regular                    A    219    nfrg____
ppnb8ac   Plantin-BoldCondensed             A    220    plbc____
ppnl8a    Plantin-Light                     A    220    pll_____
ppnli8a   Plantin-LightItalic               A    220    plli____
ppns8a    Plantin-Semibold                  A    220    plsb____
ppnsi8a   Plantin-SemiboldItalic            A    220    plsbi___
pbrb8a    Berling-Bold                      A    221    bnb_____
pbrbi8a   Berling-BoldItalic                A    221    bnbi____
pbrri8a   Berling-Italic                    A    221    bni_____
pbrr8a    Berling-Roman                     A    221    bnr_____
pcyb8a    City-Bold                         A    222    tzb_____
pcybi8a   City-BoldItalic                   A    222    tzbi____
pcyl8a    City-Light                        A    222    tzl_____
pcyli8a   City-LightItalic                  A    222    tzli____
pcym8a    City-Medium                       A    222    tzm_____
pcymi8a   City-MediumItalic                 A    222    tzmi____
paur8a    Auriol                            A    223    al______
paub8a    Auriol-Bold                       A    223    alb_____
paubi8a   Auriol-BoldItalic                 A    223    albi____
pauc8a    Auriol-Black                      A    223    albl____
pauci8a   Auriol-BlackItalic                A    223    albli___
pauri8a   Auriol-Italic                     A    223    ali_____
ptbb8a    BodoniBE-Bold                     A    224    bpb_____
ptbbi8a   BodoniBE-BoldItalic               A    224    bpbi____
ptbri8a   BodoniBE-Italic                   A    224    bpi_____
ptbl8a    BodoniBE-Light                    A    224    bpl_____
ptbli8a   BodoniBE-LightItalic              A    224    bpli____
ptbm8a    BodoniBE-Medium                   A    224    bpm_____
ptbmi8a   BodoniBE-MediumItalic             A    224    bpmi____
ptbr8a    BodoniBE-Regular                  A    224    bprg____
pegbij8a  StempelGaramond-BoldItalicOsF     A    225    gsbio___
pegbj8a   StempelGaramond-BoldOsF           A    225    gsbof___
pegrij8a  StempelGaramond-ItalicOsF         A    225    gsiof___
pegrc8a   StempelGaramond-RomanSC           A    225    gsrsc___
pnbbij8a  NewBaskerville-BoldItalicOsF      A    226    nbbio___
pnbbj8a   NewBaskerville-BoldSC             A    226    nbbsc___
pnbrij8a  NewBaskerville-ItalicOsF          A    226    nbio____
pnbrc8a   NewBaskerville-SC                 A    226    nbsc____
pgybij8a  Goudy-BoldItalicOsF               A    227    gobio___
pgybj8a   Goudy-BoldOsF                     A    227    gobos___
pgyrij8a  Goudy-ItalicOsF                   A    227    goios___
pgyrc8a   Goudy-SC                          A    227    gosc____

phnbl8a   HelveticaNeue-BoldOutline         A    228    hlbou___
pcvrf7c   Clairvaux-Dfr                     A    229    aodfr___
pcvrf8a   Clairvaux                         A    229    aor_____
pc0rf7c   Carolina-Dfr                      A    229    kldfr___
pc0rf8a   Carolina                          A    229    klr_____
ps0rf7c   SanMarco-Dfr                      A    229    sjdfr___
ps0rf8a   SanMarco                          A    229    sjr_____
pvjb8a    Veljovic-Bold                     A    230    vlb_____
pvjbi8a   Veljovic-BoldItalic               A    230    vlbi____
pvjc8a    Veljovic-Black                    A    230    vlbl____
pvjci8a   Veljovic-BlackItalic              A    230    vlbli___
pvjm8a    Veljovic-Medium                   A    230    vlm_____
pvjmi8a   Veljovic-MediumItalic             A    230    vlmi____
pvjk8a    Veljovic-Book                     A    230    vlw_____
pvjki8a   Veljovic-BookItalic               A    230    vlwi____
pscb8a    Slimbach-Bold                     A    231    spb_____
pscbi8a   Slimbach-BoldItalic               A    231    spbi____
pscc8a    Slimbach-Black                    A    231    spbl____
pscci8a   Slimbach-BlackItalic              A    231    spbli___
pscm8a    Slimbach-Medium                   A    231    spm_____
pscmi8a   Slimbach-MediumItalic             A    231    spmi____
psck8a    Slimbach-Book                     A    231    spw_____
pscki8a   Slimbach-BookItalic               A    231    spwi____
pbvri8a   BaskervilleBook-Italic            A    232    bti_____
pbvm8a    BaskervilleBook-Medium            A    232    btm_____
pbvmi8a   BaskervilleBook-MedItalic         A    232    btmi____
pbvr8a    BaskervilleBook-Regular           A    232    btrg____
murb8a    CentaurMT-Bold                    A    233    nrb_____
murbi8a   CentaurMT-BoldItalic              A    233    nrbi____
murri8a   CentaurMT-Italic                  A    233    nri_____
murria8a  CentaurMT-ItalicA                 A    233    nria____
murr8a    CentaurMT                         A    233    nrr_____
phvr7f    Helvetica-Fraction                A    234    hvfr____
phvb7f    Helvetica-FractionBold            A    234    hvfrb___
pncr7f    NewCenturySchlbk-Fraction         A    234    ncfr____
pncb7f    NewCenturySchlbk-FractionBold     A    234    ncfrb___
pilbic8a  Caecilia-BoldItalicSC             A    235    ecbis___
pilbc8a   Caecilia-BoldSC                   A    235    ecbsc___
pilhic8a  Caecilia-HeavyItalicSC            A    235    echis___
pilhc8a   Caecilia-HeavySC                  A    235    echsc___
pilric8a  Caecilia-ItalicSC                 A    235    ecisc___
pillc8a   Caecilia-LightItalicSC            A    235    eclis___
pillic8a  Caecilia-LightSC                  A    235    eclsc___
pilrc8a   Caecilia-RomanSC                  A    235    ecrsc___
mjob8a    JoannaMT-Bold                     A    236    job_____
mjobi8a   JoannaMT-BoldItalic               A    236    jobi____
mjox8a    JoannaMT-ExtraBold                A    236    joeb____

mjori8a   JoannaMT-Italic                   A    236    joi_____
mjor8a    JoannaMT                          A    236    jor_____
mjos8a    JoannaMT-SemiBold                 A    236    josb____
mjosi8a   JoannaMT-SemiBoldItalic           A    236    josbi___
ppi017    LHolidayPi-One                    A    237    holi1___
ppi018    LHolidayPi-Two                    A    237    holi2___
ppi019    LHolidayPi-Three                  A    237    holi3___
ppi011    LDecorationPi-One                 A    238    deco1___
ppi012    LDecorationPi-Two                 A    238    deco2___
po7r8a    OldStyleSeven                     A    239    os______
po7ri8a   OldStyleSeven-Italic              A    239    osi_____
po7rij8a  OldStyleSeven-ItalicOsF           A    239    osios___
po7rc8a   OldStyleSeven-SC                  A    239    ossc____
ppi013    LGamePi-ChessDraughts             A    240    gamcd___
ppi014    LGamePi-DiceDominoes              A    240    gamdd___
ppi015    LGamePi-EnglishCards              A    240    gamec___
ppi016    LGamePi-FrenchCards               A    240    gamfc___
pc5rij8a  CaslonFiveForty-ItalicOsF         A    241    cfiof___
pc5rc8a   CaslonFiveForty-RomanSC           A    241    cfrsc___
pc3rij8a  CaslonThree-ItalicOsF             A    241    ctiof___
pc3rc8a   CaslonThree-RomanSC               A    241    ctrsc___
mgtrf8a   GoudyTextMT                       A    242    gx______
mgtrf7a   GoudyTextMT-Alternate             A    242    gxalt___
mgtrf7c   GoudyTextMT-Dfr                   A    242    gxdfr___
mgtrfd8a  GoudyTextMT-LombardicCapitals     A    242    gxlom___
pefb8a    EgyptienneF-Bold                  A    243    efb_____
pefc8a    EgyptienneF-Black                 A    243    efbl____
pefri8a   EgyptienneF-Italic                A    243    efi_____
pefr8a    EgyptienneF-Roman                 A    243    efr_____
pwkb8a    Wilke-Bold                        A    244    wlb_____
pwkbi8a   Wilke-BoldItalic                  A    244    wlbi____
pwkc8a    Wilke-Black                       A    244    wlbl____
pwkci8a   Wilke-BlackItalic                 A    244    wlbli___
pwkri8a   Wilke-Italic                      A    244    wli_____
pwkr8a    Wilke-Roman                       A    244    wlr_____
ppxb8a    PopplPontifexBE-Bold              A    245    pjb_____
ppxri8a   PopplPontifexBE-Italic            A    245    pji_____
ppxm8a    PopplPontifexBE-Medium            A    245    pjm_____
ppxm8ac   PopplPontifexBE-MediumCn          A    245    pjmc____
ppxr8a    PopplPontifexBE-Regular           A    245    pjrg____
mtlr8a    CastellarMT                       A    246    el______
malr8a    AlbertusMT                        A    246    lr______
malri8a   AlbertusMT-Italic                 A    246    lri_____
mall8a    AlbertusMT-Light                  A    246    lrl_____
pgvb8a    Giovanni-Bold                     A    247    gvb_____
pgvbi8a   Giovanni-BoldItalic               A    247    gvbi____
pgvc8a    Giovanni-Black                    A    247    gvbl____

pgvci8a   Giovanni-BlackItalic              A    247    gvbli___
pgvk8a    Giovanni-Book                     A    247    gvw_____
pgvki8a   Giovanni-BookItalic               A    247    gvwi____
pazbi8ax  AkzidenzGroteskBE-BoldExIt        A    248    gfbei___
pazb8ax   AkzidenzGroteskBE-BoldEx          A    248    gfbv____
pazr8ax   AkzidenzGroteskBE-Ex              A    248    gfex____
pazl8ax   AkzidenzGroteskBE-LightEx         A    248    gflex___
pazm8ax   AkzidenzGroteskBE-MdEx            A    248    gfmex___
p_____AGOldFace-Bold                    A    249    qwb_____
p_____AGOldFace-BoldOutline             A    249    qwbou___
p_____AGOldFace-Medium                  A    249    qwm_____
p_____AGOldFace-Outline                 A    249    qwou____
p_____AGOldFace-Regular                 A    249    qwrg____
p_____AGOldFace-Shaded                  A    249    qwsh____
p_____GaramondBE-Bold                   A    250    gwb_____
p_____GaramondBE-Condensed              A    250    gwc_____
p_____GaramondBE-Italic                 A    250    gwi_____
p_____GaramondBE-Medium                 A    250    gwm_____
p_____GaramondBE-MediumCn               A    250    gwmc____
p_____GaramondBE-MediumItalic           A    250    gwmi____
p_____GaramondBE-Regular                A    250    gwrg____
pgub8a    Guardi-Bold                       A    251    gub_____
pgubi8a   Guardi-BoldItalic                 A    251    gubi____
pguc8a    Guardi-Black                      A    251    gubl____
pguci8a   Guardi-BlackItalic                A    251    gubli___
pguri8a   Guardi-Italic                     A    251    gui_____
pgur8a    Guardi-Roman                      A    251    gur_____
pkbc8a    Kabel-Black                       A    252    kabl____
pkbh8a    Kabel-Heavy                       A    252    kah_____
pkbl8a    Kabel-Light                       A    252    kal_____
pkbk8a    Kabel-Book                        A    252    kaw_____
ppi009    LAstrologyPi-One                  A    253    astr1___
ppi010    LAstrologyPi-Two                  A    253    astr2___
pprl8a    PopplResidenz-Light               A    254    pzl_____
pprr8a    PopplResidenz-Regular             A    254    pzrg____
pstr5a    StoneSerif-PhoneticAlternate      A    255    srpa____
pstr5i    StoneSerif-PhoneticIPA            A    255    srpi____
ps8r5a    StoneSans-PhoneticAlternate       A    255    sspa____
ps8r5i    StoneSans-PhoneticIPA             A    255    sspi____
prnb8a    Rotation-Bold                     A    256    rnb_____
prnri8a   Rotation-Italic                   A    256    rni_____
prnr8a    Rotation-Roman                    A    256    rnr_____
pasri8a   Aldus-Italic                      A    257    afi_____
pasrij8a  Aldus-ItalicOsF                   A    257    afiof___
pasr8a    Aldus-Roman                       A    257    afr_____
pasrc8a   Aldus-RomanSC                     A    257    afrsc___
mlir8ac   LatinMT-Condensed                 A    258    lnc_____

moxr8a    OnyxMT                            A    258    ox______
mrur8ac   RunicMT-Condensed                 A    258    ric_____
p_____Colossalis-Bold                   A    259    kob_____
p_____Colossalis-Black                  A    259    kobl____
p_____Colossalis-Medium                 A    259    kom_____
p_____Colossalis-Regular                A    259    korg____
p_____BertholdScript-Medium             A    260    qrm_____
p_____BertholdScript-Regular            A    260    qrrg____
pqsrw8a   Bellevue                          A    260    qsrg____
pqtrw8a   Boulevard                         A    260    qtrg____
pfrb8ac   Frutiger-BoldCn                   A    261    ftbc____
pfrc8ac   Frutiger-BlackCn                  A    261    ftblc___
pfrr8ac   Frutiger-Cn                       A    261    ftc_____
pfrx8ac   Frutiger-ExtraBlackCn             A    261    ftebc___
pfrl8ac   Frutiger-LightCn                  A    261    ftlc____
pflb8a    Flora-Bold                        A    262    fab_____
pflm8a    Flora-Medium                      A    262    fam_____
metri8a   EhrhardtMT-Italic                 A    263    ehi_____
metr8a    EhrhardtMT-Regular                A    263    ehrg____
mets8a    EhrhardtMT-Semibold               A    263    ehsb____
metsi8a   EhrhardtMT-SemiboldItalic         A    263    ehsbi___
pppbij8a  Perpetua-BoldItalicOsF            A    264    ptbio___
pppbi8x   Perpetua-BoldItalicExp            A    264    ptbix___
pppbj8a   Perpetua-BoldOsF                  A    264    ptbos___
pppb8x    Perpetua-BoldExp                  A    264    ptbxp___
pppr8x    Perpetua-Exp                      A    264    ptexp___
ppprij8a  Perpetua-ItalicOsF                A    264    ptios___
pppri8x   Perpetua-ItalicExp                A    264    ptixp___
ppprc8a   Perpetua-SC                       A    264    ptsc____
ppxbj8a   PopplPontifexBE-BoldOsF           A    265    pjbos___
ppxb8x    PopplPontifexBE-BoldExp           A    265    pjbxp___
ppxrij8a  PopplPontifexBE-ItalicOsF         A    265    pjios___
ppxri8x   PopplPontifexBE-ItalicExp         A    265    pjixp___
ppxmc8ac  PopplPontifexBE-MediumCnSC        A    265    pjmcs___
ppxm8xc   PopplPontifexBE-MediumCnExp       A    265    pjmcx___
ppxmc8a   PopplPontifexBE-MediumSC          A    265    pjmsc___
ppxm8x    PopplPontifexBE-MediumExp         A    265    pjmxp___
ppxr8x    PopplPontifexBE-RegularExp        A    265    pjrgx___
ppxrc8a   PopplPontifexBE-RegularSC         A    265    pjrsc___
p_____ConcordeNova-Italic               A    266    nvi_____
p_____ConcordeNova-ItalicOsF            A    266    nviof___
p_____ConcordeNova-ItalicExp            A    266    nvixp___
p_____ConcordeNova-Medium               A    266    nvm_____
p_____ConcordeNova-MediumSC             A    266    nvmsc___
p_____ConcordeNova-MediumExp            A    266    nvmxp___
p_____ConcordeNova-Regular              A    266    nvrg____
p_____ConcordeNova-RegularExp           A    266    nvrgx___

p_____ConcordeNova-RegularSC            A    266    nvsc____
peib8a    Esprit-Bold                       A    267    esb_____
peibi8a   Esprit-BoldItalic                 A    267    esbi____
peici8a   Esprit-BlackItalic                A    267    esbli___
peim8a    Esprit-Medium                     A    267    esm_____
peimi8a   Esprit-MediumItalic               A    267    esmi____
peik8a    Esprit-Book                       A    267    esw_____
peiki8a   Esprit-BookItalic                 A    267    eswi____
peic8a    Esprit-Black                      A    267    esz_____
ptbbij8a  BodoniBE-BoldItalicOsF            A    268    bpbio___
ptbbi8x   BodoniBE-BoldItalicExp            A    268    bpbix___
ptbbj8a   BodoniBE-BoldOsF                  A    268    bpbof___
ptbb8x    BodoniBE-BoldExp                  A    268    bpbxp___
ptbrij8a  BodoniBE-ItalicOsF                A    268    bpiof___
ptbri8x   BodoniBE-ItalicExp                A    268    bpixp___
ptblij8a  BodoniBE-LightItalicOsF           A    268    bplio___
ptbli8x   BodoniBE-LightItalicExp           A    268    bplix___
ptblc8a   BodoniBE-LightSC                  A    268    bplsc___
ptbl8x    BodoniBE-LightExp                 A    268    bplxp___
ptbmij8a  BodoniBE-MediumItalicOsF          A    268    bpmio___
ptbmi8x   BodoniBE-MediumItalicExp          A    268    bpmix___
ptbmc8a   BodoniBE-MediumSC                 A    268    bpmsc___
ptbm8x    BodoniBE-MediumExp                A    268    bpmxp___
ptbr8x    BodoniBE-RegularExp               A    268    bprgx___
ptbrc8a   BodoniBE-RegularSC                A    268    bpsc____
melr8a    EllingtonMT                       A    269    ei______
melb8a    EllingtonMT-Bold                  A    269    eib_____
melbi8a   EllingtonMT-BoldItalic            A    269    eibi____
melx8a    EllingtonMT-ExtraBold             A    269    eieb____
melxi8a   EllingtonMT-ExtraBoldItalic       A    269    eiebi___
melri8a   EllingtonMT-Italic                A    269    eii_____
mell8a    EllingtonMT-Light                 A    269    eil_____
melli8a   EllingtonMT-LightItalic           A    269    eili____
pazb8ac   AkzidenzGroteskBE-BoldCn          A    270    gfbc____
pazr8ac   AkzidenzGroteskBE-Cn              A    270    gfc_____
pazx8a    AkzidenzGroteskBE-XBd             A    270    gfeb____
pazl8ac   AkzidenzGroteskBE-LightCn         A    270    gflc____
pazm8ac   AkzidenzGroteskBE-MdCn            A    270    gfmc____
pazmi8ac  AkzidenzGroteskBE-MdCnIt          A    270    gfmci___
pazx8ac   AkzidenzGroteskBE-XBdCn           A    270    gfxc____
pazxi8ac  AkzidenzGroteskBE-XBdCnIt         A    270    gfxci___
pptr8a    Present                           A    271    pr______
pptb8a    Present-Bold                      A    271    prb_____
pptb8ac   Present-BoldCondensed             A    271    prbc____
pptc8a    Present-Black                     A    271    prbl____
pptc8ac   Present-BlackCondensed            A    271    prblc___
pptr8ac   Present-Condensed                 A    271    prc_____

lvcb8a    VectoraLH-Bold                    A    272    vcb_____
lvcbi8a   VectoraLH-BoldItalic              A    272    vcbi____
lvcc8a    VectoraLH-Black                   A    272    vcbl____
lvcci8a   VectoraLH-BlackItalic             A    272    vcbli___
lvcri8a   VectoraLH-Italic                  A    272    vci_____
lvcl8a    VectoraLH-Light                   A    272    vcl_____
lvcli8a   VectoraLH-LightItalic             A    272    vcli____
lvcr8a    VectoraLH-Roman                   A    272    vcr_____
pbab8a    Barmeno-Bold                      A    273    bfb_____
pbax8a    Barmeno-ExtraBold                 A    273    bfeb____
pbam8a    Barmeno-Medium                    A    273    bfm_____
pbar8a    Barmeno-Regular                   A    273    bfrg____
p_____Cosmos-ExtraBold                  A    274    kzeb____
p_____Cosmos-Light                      A    274    kzl_____
p_____Cosmos-LightItalic                A    274    kzli____
p_____Cosmos-Medium                     A    274    kzm_____
pmyMMi8a  MyriadMM-It                       A    275    _mi_____
pmyMM8a   MyriadMM                          A    275    _mrg____
pmnMMi8a  MinionMM-It                       A    276    zgi_____
pmnMM8a   MinionMM                          A    276    zgrg____
pmnMMic8a MinionMM-ItSC                     A    277    zgisc___
pmnMMi8x  MinionMM-ItEp                     A    277    zgix____
pmnMMp    MinionMM-Or                       A    277    zgo_____
pmnMMc8a  MinionMM-SC                       A    277    zgsc____
pmnMMiw7a MinionMM-Sw                       A    277    zgsw____
pmnMM8x   MinionMM-Ep                       A    277    zgx_____
ptmr5a    Times-PhoneticAlternate           A    278    tipa____
ptmr5i    Times-PhoneticIPA                 A    278    tipi____
lpi003    AudioLH-Pi                        A    279    audi1___
lpi002    WarningLH-Pi                      A    279    warn1___
ppmri8a   PostMediaeval-Italic              A    280    pmi_____
ppmm8a    PostMediaeval-Medium              A    280    pmm_____
ppmr8a    PostMediaeval-Regular             A    280    pmrg____
lffb8a    FairfieldLH-Bold                  A    281    fdb_____
lffbi8a   FairfieldLH-BoldItalic            A    281    fdbi____
lffh8a    FairfieldLH-Heavy                 A    281    fdh_____
lffhi8a   FairfieldLH-HeavyItalic           A    281    fdhi____
lffl8a    FairfieldLH-Light                 A    281    fdl_____
lffli8a   FairfieldLH-LightItalic           A    281    fdli____
lffm8a    FairfieldLH-Medium                A    281    fdm_____
lffmi8a   FairfieldLH-MediumItalic          A    281    fdmi____
lffbd8a   FairfieldLH-CaptionBold           A    282    fdbcp___
lffbijw8a FairfieldLH-SwBoldItalicOsF       A    282    fdbio___
lffbc8a   FairfieldLH-BoldSC                A    282    fdbsc___
lffhd8a   FairfieldLH-CaptionHeavy          A    282    fdhcp___
lffhijw8a FairfieldLH-SwHeavyItalicOsF      A    282    fdhio___
lffhc8a   FairfieldLH-HeavySC               A    282    fdhsc___

lffld8a   FairfieldLH-CaptionLight          A    282    fdlcp___
lfflijw8a FairfieldLH-SwLightItalicOsF      A    282    fdlio___
lfflc8a   FairfieldLH-LightSC               A    282    fdlsc___
lffmd8a   FairfieldLH-CaptionMedium         A    282    fdmcp___
lffmijw8a FairfieldLH-SwMediumItalicOsF     A    282    fdmio___
lffmc8a   FairfieldLH-MediumSC              A    282    fdmsc___
lgzb8a    GazetteLH-Bold                    A    283    gzb_____
lgzri8a   GazetteLH-Italic                  A    283    gzi_____
lgzr8a    GazetteLH-Roman                   A    283    gzr_____
lcfbs8a   ClearfaceGothicLH-Bold            A    284    fcb_____
lcfcs8a   ClearfaceGothicLH-Black           A    284    fcbl____
lcfls8a   ClearfaceGothicLH-Light           A    284    fcl_____
lcfms8a   ClearfaceGothicLH-Medium          A    284    fcm_____
lcfrs8a   ClearfaceGothicLH-Roman           A    284    fcr_____
pidb8a    Isadora-Bold                      A    285    irb_____
pidr8a    Isadora-Regular                   A    285    irrg____
pbax8a    Barmeno-ExtraBold                 A    286    bfeb____
pfarl8a   Formata-Outline                   A    286    fmou____
p_____Cosmos-ExtraBold                  A    286    kzeb____
p_____AGOldFace-Shaded                  A    286    qwsh____
p_____AGBook-Stencil                    A    286    tast____
ltgb8ax   TradeGothicLH-BoldExtended        A    287    tgbv____
ltgr8ax   TradeGothicLH-Extended            A    287    tgv_____
mphb8a    PhotinaMT-Bold                    A    288    phb_____
mphbi8a   PhotinaMT-BoldItalic              A    288    phbi____
mphri8a   PhotinaMT-Italic                  A    288    phi_____
mphr8a    PhotinaMT                         A    288    phrg____
mphs8a    PhotinaMT-SemiBold                A    288    phsb____
mphsi8a   PhotinaMT-SemiBoldItalic          A    288    phsbi___
mphu8a    PhotinaMT-UltraBold               A    288    phub____
mphui8a   PhotinaMT-UltraBoldItalic         A    288    phubi___
mdsrw8a   DorchesterScriptMT                A    289    dsscr___
mp2rw8a   PepitaMT                          A    289    epscr___
mbfrw8a   BiffoMT                           A    289    ifscr___
ms2bw8a   ScriptMT-Bold                     A    289    tbscr___
lenb8a    EngraversLH-BoldFace              A    290    ew______
l_____SerlioLH                          A    290    oi______
pplc8a    Palatino-Black                    A    291    pobl____
pplci8a   Palatino-BlackItalic              A    291    pobli___
ppll8a    Palatino-Light                    A    291    pol_____
pplli8a   Palatino-LightItalic              A    291    poli____
pplm8a    Palatino-Medium                   A    291    pom_____
pplmi8a   Palatino-MediumItalic             A    291    pomi____
pocrij8a  ConcordeBE-ItalicOsF              A    292    qqiof___
pocri8x   ConcordeBE-ItalicExp              A    292    qqixp___
pocmij8a  ConcordeBE-MediumItalicOsF        A    292    qqmio___
pocmi8x   ConcordeBE-MediumItalicExp        A    292    qqmix___

pocmc8a   ConcordeBE-MediumSC               A    292    qqmsc___
pocm8x    ConcordeBE-MediumExp              A    292    qqmxp___
pocrc8x   ConcordeBE-RegularExp             A    292    qqrgx___
pocr8a    ConcordeBE-RegularSC              A    292    qqsc____
mc8b8a    CalvertMT-Bold                    A    293    vtb_____
mc8l8a    CalvertMT-Light                   A    293    vtl_____
mc8r8a    CalvertMT                         A    293    vtrg____
mgqr8a    GrotesqueMT                       A    294    gq______
mgqb8a    GrotesqueMT-Bold                  A    294    gqb_____
mgqc8a    GrotesqueMT-Black                 A    294    gqbl____
mgqri8a   GrotesqueMT-Italic                A    294    gqi_____
mgql8a    GrotesqueMT-Light                 A    294    gql_____
mgqli8a   GrotesqueMT-LightItalic           A    294    gqli____
pckb8a    CaslonBookBE-Bold                 A    295    qxb_____
pckri8a   CaslonBookBE-Italic               A    295    qxi_____
pckm8a    CaslonBookBE-Medium               A    295    qxm_____
pckr8a    CaslonBookBE-Regular              A    295    qxrg____
pckbj8a   CaslonBookBE-BoldOsF              A    296    qxbof___
pckb8x    CaslonBookBE-BoldExp              A    296    qxbxp___
pckrij8a  CaslonBookBE-ItalicOsF            A    296    qxiof___
pckri8x   CaslonBookBE-ItalicExp            A    296    qxixp___
pckmc8a   CaslonBookBE-MediumSC             A    296    qxmsc___
pckm8x    CaslonBookBE-MediumExp            A    296    qxmxp___
pckr8x    CaslonBookBE-RegularExp           A    296    qxrgx___
pckrc8a   CaslonBookBE-RegularSC            A    296    qxsc____
mr2b8a    CantoriaMT-Bold                   A    297    eob_____
mr2bi8a   CantoriaMT-BoldItalic             A    297    eobi____
mr2ri8a   CantoriaMT-Italic                 A    297    eoi_____
mr2r8a    CantoriaMT                        A    297    eorg____
mr2x8a    CantoriaMT-ExtraBold              A    298    eoeb____
mr2xi8a   CantoriaMT-ExtraBoldItalic        A    298    eoebi___
mr2l8a    CantoriaMT-Light                  A    298    eol_____
mr2li8a   CantoriaMT-LightItalic            A    298    eoli____
mr2s8a    CantoriaMT-SemiBold               A    298    eosb____
mr2si8a   CantoriaMT-SemiBoldItalic         A    298    eosbi___
mgqb8ax   GrotesqueMT-BoldExtended          A    299    gqbv____
mgqr8ac   GrotesqueMT-Condensed             A    299    gqc_____
mgqr8aq   GrotesqueMT-ExtraCondensed        A    299    gqec____
mgql8ac   GrotesqueMT-LightCondensed        A    299    gqlc____
ppc___Poetica-ChanceryI                 A    300    pvch1___
ppc___Poetica-ChanceryII                A    300    pvch2___
ppc___Poetica-ChanceryIII               A    300    pvch3___
ppc___Poetica-ChanceryIV                A    300    pvch4___
ppc___Poetica-ChanceryExpert            A    300    pvchx___
ppc___Poetica-RomanSmallCapsAlt         A    300    pvrsa___
ppc___Poetica-RomanSmallCaps            A    300    pvrsc___
ppc___Poetica-SuppAmpersands            A    301    pvam____

ppc___Poetica-SuppInitialSwashCaps      A    301    pvisw___
ppc___Poetica-SuppLowercaseAltI         A    301    pvla1___
ppc___Poetica-SuppLowercaseAltII        A    301    pvla2___
ppc___Poetica-SuppLowercaseBegI         A    301    pvlb1___
ppc___Poetica-SuppLowercaseBegII        A    301    pvlb2___
ppc___Poetica-SuppLowercaseEndI         A    301    pvle1___
ppc___Poetica-SuppLowercaseEndII        A    301    pvle2___
ppc___Poetica-SuppLigatures             A    301    pvlg____
ppc___Poetica-SuppOrnaments             A    301    pvor____
ppc___Poetica-SuppSwashCapsI            A    301    pvsw1___
ppc___Poetica-SuppSwashCapsII           A    301    pvsw2___
ppc___Poetica-SuppSwashCapsIII          A    301    pvsw3___
ppc___Poetica-SuppSwashCapsIV           A    301    pvsw4___
msmr8x    SpectrumMT-Expert                 A    302    ekexp___
msmri8a   SpectrumMT-Italic                 A    302    eki_____
msmrij8a  SpectrumMT-ItalicOsF              A    302    ekio____
msmri8x   SpectrumMT-ItalicExpert           A    302    ekixp___
msmr8a    SpectrumMT                        A    302    ekrg____
msms8a    SpectrumMT-SemiBold               A    302    eksb____
msmsc8a   SpectrumMT-SemiBoldOsF            A    302    eksbo___
msms8x    SpectrumMT-SemiBoldExpert         A    302    eksbx___
msmrc8a   SpectrumMT-SC                     A    302    eksc____
p_____AGBookRounded-Bold                A    303    kdb_____
p_____AGBookRounded-BoldCn              A    303    kdbc____
p_____AGBookRounded-BoldCnOutline       A    303    kdbcl___
p_____AGBookRounded-BoldOutline         A    303    kdbou___
p_____AGBookRounded-Medium              A    303    kdm_____
p_____AGBookRounded-MediumOutline       A    303    kdmou___
p_____AGBookRounded-Regular             A    303    kdrg____
p_____BodoniOldFaceBE-Bold              A    304    qvb_____
p_____BodoniOldFaceBE-BoldItalic        A    304    qvbi____
p_____BodoniOldFaceBE-Italic            A    304    qvi_____
p_____BodoniOldFaceBE-Medium            A    304    qvm_____
p_____BodoniOldFaceBE-MediumItalic      A    304    qvmi____
p_____BodoniOldFaceBE-Regular           A    304    qvrg____
mg4ri8a   GoudyModernMT-Italic              A    305    ygi_____
mg4r8a    GoudyModernMT                     A    305    ygrg____
ms1ri8a   ScotchRomanMT-Italic              A    305    ysi_____
ms1r8a    ScotchRomanMT                     A    305    ysrg____
p_____AGSchoolbook-RegularA             A    306    okarg___
p_____AGSchoolbook-Medium               A    306    okm_____
p_____AGSchoolbook-MediumA              A    306    okma____
p_____AGSchoolbook-Regular              A    306    okrg____
pbeb8a    BellGothic-Bold                   A    307    dgb_____
pbec8a    BellGothic-Black                  A    307    dgbl____
pbel8a    BellGothic-Light                  A    307    dgl_____
misb8a    ItalianOldStyleMT-Bold            A    308    ieb_____

misbi8a   ItalianOldStyleMT-BoldItalic      A    308    iebi____
misri8a   ItalianOldStyleMT-Italic          A    308    iei_____
misr8a    ItalianOldStyleMT                 A    308    ierg____
p_____BellCentennial-Address            A    309    bladd___
p_____BellCentennial-BoldListingAlt     A    309    blbla___
p_____BellCentennial-BoldListing        A    309    blbls___
p_____BellCentennial-NameAndNumber      A    309    blnan___
p_____BellCentennial-SubCaption         A    309    blscp___
p_____BodoniOldFaceBE-BoldItOsF         A    310    qvbio___
p_____BodoniOldFaceBE-BoldItExp         A    310    qvbix___
p_____BodoniOldFaceBE-BoldOsF           A    310    qvbos___
p_____BodoniOldFaceBE-BoldExp           A    310    qvbxp___
p_____BodoniOldFaceBE-ItalicSC          A    310    qvisc___
p_____BodoniOldFaceBE-ItalicExp         A    310    qvixp___
p_____BodoniOldFaceBE-MediumItOsF       A    310    qvmio___
p_____BodoniOldFaceBE-MediumItExp       A    310    qvmix___
p_____BodoniOldFaceBE-MediumSC          A    310    qvmsc___
p_____BodoniOldFaceBE-MediumExp         A    310    qvmxp___
p_____BodoniOldFaceBE-RegularExp        A    310    qvrgx___
p_____BodoniOldFaceBE-RegularSC         A    310    qvrsc___
p_____BerlinerGrotesk-Light             A    311    otl_____
p_____BerlinerGrotesk-Medium            A    311    otm_____
mntb8ac   TimesNewRomanMT-BoldCond          A    312    mtbc____
mntr8ac   TimesNewRomanMT-Cond              A    312    mtc_____
mntri8ac  TimesNewRomanMT-CondItalic        A    312    mtci____
movr8x    OctavianMT-Expert                 A    313    ocexp___
movri8a   OctavianMT-Italic                 A    313    oci_____
movrij8a  OctavianMT-ItalicOsF              A    313    ocio____
movri8x   OctavianMT-ItalicExpert           A    313    ocixp___
movr8a    OctavianMT                        A    313    ocrg____
movrc8a   OctavianMT-SC                     A    313    ocsc____
p_____Christiana-Bold                   A    314    hcb_____
p_____Christiana-BoldItalic             A    314    hcbi____
p_____Christiana-Italic                 A    314    hci_____
p_____Christiana-Medium                 A    314    hcm_____
p_____Christiana-MediumItalic           A    314    hcmi____
p_____Christiana-Regular                A    314    hcrg____
p_____Christiana-RegularExpert          A    314    hcrgx___
p_____Christiana-RegularSC              A    314    hcsc____
pmnb6w    MinionCyr-Bold                    A    315    mkb_____
pmnbi6w   MinionCyr-BoldItalic              A    315    mkbi____
pmnri6w   MinionCyr-Italic                  A    315    mki_____
pmnr6w    MinionCyr-Regular                 A    315    mkrg____
pmns6w    MinionCyr-Semibold                A    315    mksb____
pmnsi6w   MinionCyr-SemiboldItalic          A    315    mksbi___
ptbb8ac   BodoniBE-BoldCn                   A    316    bpbc____
ptbbi8ac  BodoniBE-BoldCnItalic             A    316    bpbci___

ptbr8ac   BodoniBE-Condensed                A    316    bpc_____
ptbri8ac  BodoniBE-CnItalic                 A    316    bpci____
ptbm8ac   BodoniBE-MediumCn                 A    316    bpmc____
ptbmi8ac  BodoniBE-MediumCnItalic           A    316    bpmci___
mfsr8a    FalstaffMT                        A    317    fprg____
mifb8a    InflexMT-Bold                     A    317    iqb_____
mm8bl8a   MonotypeOldStyleMT-BoldOut        A    317    mbbou___
pdjb8a    Delta-Bold                        A    318    djb_____
pdjbi8a   Delta-BoldItalic                  A    318    djbi____
pdjl8a    Delta-Light                       A    318    djl_____
pdjli8a   Delta-LightItalic                 A    318    djli____
pdjm8a    Delta-Medium                      A    318    djm_____
pdjmi8a   Delta-MediumItalic                A    318    djmi____
pdjrl8a   Delta-Outline                     A    318    djou____
pdjk8a    Delta-Book                        A    318    djw_____
pdjki8a   Delta-BookItalic                  A    318    djwi____
mnrrw8a   NewBerolinaMT                     A    319    _n______
mp1rw8a   PalaceScriptMT                    A    319    _p______
mp1sw8a   PalaceScriptMT-SemiBold           A    319    _psb____
mahrw8a   AshleyScriptMT                    A    319    le______
mm1rw8a   MonolineScriptMT                  A    319    on______
pbtb8a    Boton-Bold                        A    320    tnb_____
pbtbi8a   Boton-BoldItalic                  A    320    tnbi____
pbtri8a   Boton-Italic                      A    320    tni_____
pbtl8a    Boton-Light                       A    320    tnl_____
pbtli8a   Boton-LightItalic                 A    320    tnli____
pbtm8a    Boton-Medium                      A    320    tnm_____
pbtmi8a   Boton-MediumItalic                A    320    tnmi____
pbtr8a    Boton-Regular                     A    320    tnr_____
mfer8a    ForteMT                           A    321    fz______
mklr8a    KlangMT                           A    321    kn______
mmebw8a   MercuriusMT-BoldScript            A    321    uzbs____
ldob8a    DidotLH-Bold                      A    322    dob_____
ldobj8a   DidotLH-BoldOsF                   A    322    dobos___
ldord8a   DidotLH-Headline                  A    322    dohe____
ldordj8a  DidotLH-HeadlineOsF               A    322    dohes___
ldori8a   DidotLH-Italic                    A    322    doi_____
ldor7a    DidotLH-Initials                  A    322    doini___
ldorij8a  DidotLH-ItalicOsF                 A    322    doios___
ldorp1    DidotLH-OrnamentsOne              A    322    door1___
ldorp2    DidotLH-OrnamentsTwo              A    322    door2___
ldor8a    DidotLH-Roman                     A    322    dor_____
ldorj8a   DidotLH-RomanOsF                  A    322    doros___
ldorc8a   DidotLH-RomanSC                   A    322    dorsc___
pfab8ac   Formata-BoldCondensed             A    323    fmbc____
pfabi8ac  Formata-BoldCondensedItalic       A    323    fmbci___
pfar8ac   Formata-Condensed                 A    323    fmc_____

pfari8ac  Formata-CondensedItalic           A    323    fmci____
pfarl8ac  Formata-CondensedOutline          A    323    fmcou___
pfal8ac   Formata-LightCondensed            A    323    fmlc____
pfali8ac  Formata-LightCondensedItalic      A    323    fmlci___
pfam8ac   Formata-MediumCondensed           A    323    fmmc____
pfami8ac  Formata-MediumCondensedItalic     A    323    fmmci___
pebb8a    BaskervilleBE-Bold                A    324    vib_____
pebri8a   BaskervilleBE-Italic              A    324    vii_____
pebm8a    BaskervilleBE-Medium              A    324    vim_____
pebmi8a   BaskervilleBE-MediumItalic        A    324    vimi____
pebr8a    BaskervilleBE-Regular             A    324    virg____
pb6r8ac   BlockBE-Condensed                 A    325    lkc_____
pb6r8aq   BlockBE-ExtraCn                   A    325    lkec____
pb6ri8aq  BlockBE-ExtraCnIt                 A    325    lkeci___
pb6h8a    BlockBE-Heavy                     A    325    lkh_____
pb6ri8a   BlockBE-Italic                    A    325    lki_____
pb6r8a    BlockBE-Regular                   A    325    lkrg____
pllb8a    LinoLetter-Bold                   A    326    llb_____
pllbi8a   LinoLetter-BoldItalic             A    326    llbi____
pllc8a    LinoLetter-Black                  A    326    llbl____
pllci8a   LinoLetter-BlackItalic            A    326    llbli___
pllri8a   LinoLetter-Italic                 A    326    lli_____
pllm8a    LinoLetter-Medium                 A    326    llm_____
pllmi8a   LinoLetter-MediumItalic           A    326    llmi____
pllr8a    LinoLetter-Roman                  A    326    llr_____
p_____ComeniusAntiqua-Bold              A    327    hzb_____
p_____ComeniusAntiqua-Italic            A    327    hzi_____
p_____ComeniusAntiqua-Medium            A    327    hzm_____
p_____ComeniusAntiqua-Regular           A    327    hzrg____
ppub8a    PopplLaudatio-Bold                A    328    pyb_____
ppubi8a   PopplLaudatio-BoldItalic          A    328    pybi____
ppuri8a   PopplLaudatio-Italic              A    328    pyi_____
ppul8a    PopplLaudatio-Light               A    328    pyl_____
ppuli8a   PopplLaudatio-LightItalic         A    328    pyli____
ppum8a    PopplLaudatio-Medium              A    328    pym_____
ppumi8a   PopplLaudatio-MediumItalic        A    328    pymi____
ppur8a    PopplLaudatio-Regular             A    328    pyrg____
ppub8ac   PopplLaudatio-BoldCondensed       A    329    pybc____
ppur8ac   PopplLaudatio-Condensed           A    329    pyc_____
ppul8ac   PopplLaudatio-LightCondensed      A    329    pylc____
ppum8ac   PopplLaudatio-MediumCondensed     A    329    pymc____
pocb8ac   ConcordeBE-BoldCn                 A    330    qqbc____
pocbl8ac  ConcordeBE-BoldCnOutline          A    330    qqbcl___
pocr8ac   ConcordeBE-Condensed              A    330    qqc_____
pocm8ac   ConcordeBE-MediumCn               A    330    qqmc____
p_____GaramondBE-BoldOsF                A    331    gwbof___
p_____GaramondBE-BoldExpert             A    331    gwbxp___

p_____GaramondBE-CondensedSC            A    331    gwcsc___
p_____GaramondBE-CnExpert               A    331    gwcxp___
p_____GaramondBE-ItalicOsF              A    331    gwiof___
p_____GaramondBE-ItalicExpert           A    331    gwixp___
p_____GaramondBE-MediumCnOsF            A    331    gwmcf___
p_____GaramondBE-MediumCnExpert         A    331    gwmcx___
p_____GaramondBE-MediumItalicOsF        A    331    gwmio___
p_____GaramondBE-MediumItalicExpert     A    331    gwmix___
p_____GaramondBE-MediumSC               A    331    gwmsc___
p_____GaramondBE-MediumExpert           A    331    gwmxp___
p_____GaramondBE-RegularExpert          A    331    gwrgx___
p_____GaramondBE-RegularSC              A    331    gwsc____
p_____GaramondBE-SwashItalic            A    331    gwswi___
pdnb8a    DailyNews-Bold                    A    332    jdb_____
pdnbi8a   DailyNews-BoldItalic              A    332    jdbi____
pdnx8a    DailyNews-ExtraBold               A    332    jdeb____
pdnxi8a   DailyNews-ExtraBoldItalic         A    332    jdebi___
pdnri8a   DailyNews-Italic                  A    332    jdi_____
pdnm8a    DailyNews-Medium                  A    332    jdm_____
pdnmi8a   DailyNews-MediumItalic            A    332    jdmi____
pdnr8a    DailyNews-Regular                 A    332    jdrg____
murbij8a  CentaurMT-BoldItalicOsF           A    333    nrbio___
murbi8x   CentaurMT-BoldItalicExpert        A    333    nrbix___
murbj8a   CentaurMT-BoldOsF                 A    333    nrbof___
murb8x    CentaurMT-BoldExpert              A    333    nrbxp___
murr8x    CentaurMT-Expert                  A    333    nrexp___
murri7a   CentaurMT-ItalicAlternate         A    333    nrial___
murrij8a  CentaurMT-ItalicOsF               A    333    nrio____
murri8x   CentaurMT-ItalicExpert            A    333    nrixp___
murrc8a   CentaurMT-SC                      A    333    nrsc____
murriw8a  CentaurMT-SwashCapitals           A    333    nrswc___
mmob8a    ModernMT-Bold                     A    334    nmb_____
mmobi8a   ModernMT-BoldItalic               A    334    nmbi____
mmor8ac   ModernMT-Condensed                A    334    nmc_____
mmori8ac  ModernMT-CondensedItalic          A    334    nmci____
mmor8ax   ModernMT-Extended                 A    334    nmex____
mmori8ax  ModernMT-ExtendedItalic           A    334    nmexi___
mmor8aw   ModernMT-Wide                     A    334    nmwd____
mmori8aw  ModernMT-WideItalic               A    334    nmwdi___
p_____ElseNPL-Bold                      A    335    enb_____
p_____ElseNPL-Light                     A    335    enl_____
p_____ElseNPL-Medium                    A    335    enm_____
p_____ElseNPL-SemiBold                  A    335    ens_____
p_____Tiepolo-Bold                      A    336    tpb_____
p_____Tiepolo-BoldItalic                A    336    tpbi____
p_____Tiepolo-Black                     A    336    tpbl____
p_____Tiepolo-BlackItalic               A    336    tpbli___

p_____Tiepolo-Book                      A    336    tpw_____
p_____Tiepolo-BookItalic                A    336    tpwi____
pfaric8ac Formata-CnItalicSC                A    337    fmcis___
pfari8xc  Formata-CnItalicExpert            A    337    fmcix___
pfarc8ac  Formata-CondensedSC               A    337    fmcsc___
pfar8xc   Formata-CondensedExpert           A    337    fmcxp___
pfaric8a  Formata-ItalicSC                  A    337    fmisc___
pfari8x   Formata-ItalicExpert              A    337    fmixp___
pfalc8ac  Formata-LightCnSC                 A    337    fmlcs___
pfalic8ac Formata-LightCnItalicSC           A    337    fmlct___
pfal8xc   Formata-LightCnExpert             A    337    fmlcx___
pfali8xc  Formata-LightCnItalicExpert       A    337    fmlcy___
pfalic8a  Formata-LightItalicSC             A    337    fmlis___
pfali8x   Formata-LightItalicExpert         A    337    fmlix___
pfalc8a   Formata-LightSC                   A    337    fmlsc___
pfal8x    Formata-LightExpert               A    337    fmlxp___
pfar8x    Formata-RegularExpert             A    337    fmrgx___
pfarc8a   Formata-RegularSC                 A    337    fmsc____
pnfbij8a  Nofret-BoldItalicOsF              A    338    nfbio___
pnfbi8x   Nofret-BoldItalicExpert           A    338    nfbix___
pnfbj8a   Nofret-BoldOsF                    A    338    nfbof___
pnfb8x    Nofret-BoldExpert                 A    338    nfbxp___
pnfrij8a  Nofret-ItalicOsF                  A    338    nfiof___
pnfri8x   Nofret-ItalicExpert               A    338    nfixp___
pnflij8a  Nofret-LightItalicOsF             A    338    nflio___
pnfli8x   Nofret-LightItalicExpert          A    338    nflix___
pnflc8a   Nofret-LightSC                    A    338    nflsc___
pnfl8x    Nofret-LightExpert                A    338    nflxp___
pnfmij8a  Nofret-MediumItalicOsF            A    338    nfmio___
pnfmi8x   Nofret-MediumItalicExpert         A    338    nfmix___
pnfmc8a   Nofret-MediumSC                   A    338    nfmsc___
pnfm8x    Nofret-MediumExpert               A    338    nfmxp___
pnfr8x    Nofret-RegularExpert              A    338    nfrgx___
pnfrc8a   Nofret-RegularSC                  A    338    nfsc____
p_____Amigo                             A    339    ig______
p_____Pelican                           A    339    lq______
pmgr8a    Marigold                          A    339    my______
pofr8a    Oxford                            A    339    of______
pvgr8a    Visigoth                          A    339    vg______
mapr8x    ApolloMT-Expert                   A    340    ezexp___
mapri8a   ApolloMT-Italic                   A    340    ezi_____
maprij8a  ApolloMT-ItalicOsF                A    340    eziof___
mapri8x   ApolloMT-ItalicExpert             A    340    ezixp___
mapr8a    ApolloMT                          A    340    ezr_____
maprc8a   ApolloMT-SC                       A    340    ezrsc___
maps8a    ApolloMT-SemiBold                 A    340    ezsb____
mapsj8a   ApolloMT-SemiBoldOsF              A    340    ezsbo___

maps8x    ApolloMT-SemiBoldExpert           A    340    ezsbx___
p_____Catull-Bold                       A    341    qlb_____
p_____Catull-Italic                     A    341    qli_____
p_____Catull-Medium                     A    341    qlm_____
p_____Catull-Regular                    A    341    qlrg____
pagMMo8a  ITCAvantGardeMM-Oblique           A    342    zfi_____
pagMM8a   ITCAvantGardeMM                   A    342    zfrg____
mhob8a    HorleyOldStyleMT-Bold             A    343    hsb_____
mhobi8a   HorleyOldStyleMT-BoldItalic       A    343    hsbi____
mhori8a   HorleyOldStyleMT-Italic           A    343    hsi_____
mhol8a    HorleyOldStyleMT-Light            A    343    hsl_____
mholi8a   HorleyOldStyleMT-LightItalic      A    343    hsli____
mhor8a    HorleyOldStyleMT                  A    343    hsrg____
mhos8a    HorleyOldStyleMT-SemiBold         A    343    hssb____
mhosi8a   HorleyOldStyleMT-SbItalic         A    343    hssbi___
p_____Cremona-Bold                      A    344    _cb_____
p_____Cremona-BoldItalic                A    344    _cbi____
p_____Cremona-Italic                    A    344    _ci_____
p_____Cremona-Regular                   A    344    _crg____
pgsb8aq   GillSans-BoldExtraCondensed       A    345    gnbxc___
pgslh8a   GillSans-LightShadowed            A    345    gnls____
pgsrh8a   GillSans-Shadowed                 A    345    gns_____
pgsxd8a   GillSans-ExtraBoldDisplay         A    345    gnxbd___
p_____LoType-Bold                       A    346    ljb_____
p_____LoType-Light                      A    346    ljl_____
p_____LoType-Medium                     A    346    ljm_____
p_____LoType-MediumCondensed            A    346    ljmc____
p_____LoType-MediumItalic               A    346    ljmi____
p_____LoType-Regular                    A    346    ljrg____
mknr8a    KinoMT                            A    347    qi______
mwrrf8a   WittenbergerFrakturMT             A    347    wf______
mwrbf8a   WittenbergerFrakturMT-Bd          A    347    wfb_____
mwdbf7c   WittenbergerFrakturMT-BdDfr       A    347    wfbdf___
mwdrf7c   WittenbergerFrakturMT-Dfr         A    347    wfdfr___
mm7r8a    MaturaMT                          A    347    zz______
mm7rw8a   MaturaMT-ScriptCapitals           A    347    zzscc___
pllbij8a  LinoLetter-BoldItalicOsF          A    348    llbio___
pllbk8a   LinoLetter-BoldOsF                A    348    llbof___
pllbc8a   LinoLetter-BoldSC                 A    348    llbsc___
pllrij8a  LinoLetter-ItalicOsF              A    348    lliof___
pllmij8a  LinoLetter-MediumItalicOsF        A    348    llmio___
pllmj8a   LinoLetter-MediumOsF              A    348    llmof___
pllmc8a   LinoLetter-MediumSC               A    348    llmsc___
pllrj8a   LinoLetter-RomanOsF               A    348    llros___
pllrc8a   LinoLetter-RomanSC                A    348    llrsc___
pllcij8a  LinoLetter-BlackItalicOsF         A    348    llzio___
pllcj8a   LinoLetter-BlackOsF               A    348    llzof___

pllcc8a   LinoLetter-BlackSC                A    348    llzsc___
pbbbij8a  Bembo-BoldItalicOsF               A    349    bmbio___
pbbbi8x   Bembo-BoldItalicExpert            A    349    bmbix___
pbbbj8a   Bembo-BoldOsF                     A    349    bmbos___
pbbb8x    Bembo-BoldExpert                  A    349    bmbxp___
pbbxj8a   Bembo-ExtraBoldOsF                A    349    bmebs___
pbbx8x    Bembo-ExtraBoldExpert             A    349    bmebx___
pbbxij8a  Bembo-ExtraBoldItalicOsF          A    349    bmeis___
pbbxi8x   Bembo-ExtraBoldItalicExpert       A    349    bmeix___
pbbr8x    Bembo-Expert                      A    349    bmexp___
pbbrij8a  Bembo-ItalicOsF                   A    349    bmios___
pbbri8x   Bembo-ItalicExpert                A    349    bmixp___
pbbrc8a   Bembo-SC                          A    349    bmrsc___
pbbsj8a   Bembo-SemiboldOsF                 A    349    bmsbo___
pbbs8x    Bembo-SemiboldExpert              A    349    bmsbx___
pbbsij8a  Bembo-SemiboldItalicOsF           A    349    bmsis___
pbbsi8x   Bembo-SemiboldItalicExpert        A    349    bmsix___
ptkMMo8a  TektonMM-Oblique                  A    350    zji_____
ptkMM8a   TektonMM                          A    350    zjrg____
pvvMM8a   VivaMM                            A    351    zkrg____
pb7b8a    BernhardModern-Bold               A    352    rdb_____
pb7bi8a   BernhardModern-BoldItalic         A    352    rdbi____
pb7ri8a   BernhardModern-Italic             A    352    rdi_____
pb7r8a    BernhardModern-Roman              A    352    rdr_____
p_____Ariadne-Roman                     A    353    aria____
pdiri8a   Diotima-Italic                    A    353    dti_____
pdirij8a  Diotima-ItalicOsF                 A    353    dtiof___
pdir8a    Diotima-Roman                     A    353    dtr_____
pdirj8a   Diotima-RomanOsF                  A    353    dtros___
pdirc8a   Diotima-RomanSC                   A    353    dtrsc___
p_____Smaragd                           A    353    smar____
p_____Smaragd-Alternate                 A    353    smara___
pttb6w    TimesTenCyr-Bold                  A    354    ttyb____
pttbo6w   TimesTenCyr-BoldInclined          A    354    ttybi___
pttro6w   TimesTenCyr-Inclined              A    354    ttyi____
pttr6w    TimesTenCyr-Upright               A    354    ttyu____
panr8a    Anna                              A    355    iarg____
p_____Beesknees                         A    355    ikh_____
p_____MonaLisa-Recut                    A    355    iyin____
p_____MonaLisa-Solid                    A    355    iysd____
p_____Ozwald                            A    355    ozh_____
po8b8a    OfficinaSerif-Bold                A    356    ovb_____
po8bi8a   OfficinaSerif-BoldItalic          A    356    ovbi____
po8k8a    OfficinaSerif-Book                A    356    ovbk____
po8ki8a   OfficinaSerif-BookItalic          A    356    ovwi____
po9b8a    OfficinaSans-Bold                 A    357    owb_____
po9bi8a   OfficinaSans-BoldItalic           A    357    owbi____

po9k8a    OfficinaSans-Book                 A    357    owbk____
po9ki8a   OfficinaSans-BookItalic           A    357    owwi____
p_____SanvitoMM                         A    358    zhrg____
p_____CaflischScriptMM-Alt              A    359    zea_____
p_____CaflischScriptMM                  A    359    zerg____
p_____CaflischScriptMM-Sw               A    359    zesw____
p_____GraphiteMM                        A    360    _grg____
phir6w    HelveticaInseratCyr-Upright       A    361    hiyu____
phvb6w    HelveticaCyr-Bold                 A    361    hvyb____
phvbo6w   HelveticaCyr-BoldInclined         A    361    hvybi___
phvro6w   HelveticaCyr-Inclined             A    361    hvyi____
phvr6w    HelveticaCyr-Upright              A    361    hvyu____
plzb8a    LegacySans-Bold                   A    362    lvb_____
plzbi8a   LegacySans-BoldItalic             A    362    lvbi____
plzm8a    LegacySans-Medium                 A    362    lvm_____
plzmi8a   LegacySans-MediumItalic           A    362    lvmi____
plzu8a    LegacySans-Ultra                  A    362    lvu_____
plzk8a    LegacySans-Book                   A    362    lvw_____
plzki8a   LegacySans-BookItalic             A    362    lvwi____
pleb8a    LegacySerif-Bold                  A    363    lzb_____
plebi8a   LegacySerif-BoldItalic            A    363    lzbi____
plem8a    LegacySerif-Medium                A    363    lzm_____
plemi8a   LegacySerif-MediumItalic          A    363    lzmi____
pleu8a    LegacySerif-Ultra                 A    363    lzu_____
plek8a    LegacySerif-Book                  A    363    lzw_____
pleki8a   LegacySerif-BookItalic            A    363    lzwi____
pexr6w    ExcelsiorCyr-Upright              A    364    exy_____
pexb6w    ExcelsiorCyr-Bold                 A    364    exyb____
pexro6w   ExcelsiorCyr-Inclined             A    364    exyi____
pbvr6w    BaskervilleCyr-Upright            A    365    _by_____
pbvb6w    BaskervilleCyr-Bold               A    365    _byb____
pbvbo6w   BaskervilleCyr-Inclined           A    365    _byi____
p_____Pepperwood-Fill                   A    366    uwfi____
p_____Pepperwood-Outline                A    366    uwou____
p_____Pepperwood-Regular                A    366    uwrg____
p_____Rosewood-Fill                     A    366    uxfi____
p_____Rosewood-Regular                  A    366    uxrg____
p_____Zebrawood-Fill                    A    366    uyfi____
p_____Zebrawood-Regular                 A    366    uyrg____
p_____MendozaRoman-Bold                 A    367    zab_____
p_____MendozaRoman-BoldItalic           A    367    zabi____
p_____MendozaRoman-Medium               A    367    zam_____
p_____MendozaRoman-MediumItalic         A    367    zami____
p_____MendozaRoman-Book                 A    367    zaw_____
p_____MendozaRoman-BookItalic           A    367    zawi____
phhb8a    Highlander-Bold                   A    368    hhb_____
phhbi8a   Highlander-BoldItalic             A    368    hhbi____

phhk8a    Highlander-Book                   A    368    hhbk____
phhm8a    Highlander-Medium                 A    368    hhm_____
phhmi8a   Highlander-MediumItalic           A    368    hhmi____
phhki8a   Highlander-BookItalic             A    368    hhwi____
pbgb8a    Benguiat-Bold                     A    369    bgb_____
pbgbi8a   Benguiat-BoldItalic               A    369    bgbi____
pbgm8a    Benguiat-Medium                   A    369    bgm_____
pbgmi8a   Benguiat-MediumItalic             A    369    bgmi____
pbgk8a    Benguiat-Book                     A    369    bgw_____
pbgki8a   Benguiat-BookItalic               A    369    bgwi____
p_____MezzMM                            A    370    zmrg____
lrzb8a    RuzickaFreehandLH-Bold            A    371    rzb_____
lrzbc8a   RuzickaFreehandLH-BoldSC          A    371    rzbsc___
lrzr8a    RuzickaFreehandLH-Roman           A    371    rzr_____
lrzrc8a   RuzickaFreehandLH-RomanSC         A    371    rzrsc___
pnorf7c   NotreDame-RomanDfr                A    372    nhdfr___
pnorp     NotreDame-Ornaments               A    372    nhorn___
pnorf8a   NotreDame-Roman                   A    372    nhr_____
ppjrp     Pompeijana-Borders                A    372    qbbor___
ppjr8a    Pompeijana-Roman                  A    372    qbr_____
prxrp     Rusticana-Borders                 A    372    rxbor___
prxr8a    Rusticana-Roman                   A    372    rxr_____
lpi005    CaravanLH-One                     A    373    cara1___
lpi006    CaravanLH-Two                     A    373    cara2___
lpi007    CaravanLH-Three                   A    373    cara3___
lpi008    CaravanLH-Four                    A    373    cara4___
leab8a    ElectraLH-Bold                    A    373    etb_____
leabi8a   ElectraLH-BoldCursive             A    373    etbi____
leari8a   ElectraLH-Cursive                 A    373    eti_____
lear8a    ElectraLH-Regular                 A    373    etr_____
leabij8a  ElectraLH-BoldCursiveOsF          A    374    etbio___
leabj8a   ElectraLH-BoldOsF                 A    374    etbos___
leabc8a   ElectraLH-BoldSC                  A    374    etbsc___
leabd8a   ElectraLH-BoldDisplay             A    374    etdb____
leabid8a  ElectraLH-BoldCursiveDisplay      A    374    etdbi___
learid8a  ElectraLH-CursiveDisplay          A    374    etdi____
leard8a   ElectraLH-Display                 A    374    etdr____
learij8a  ElectraLH-CursiveOsF              A    374    etios___
learj8a   ElectraLH-RegularOsF              A    374    etros___
learc8a   ElectraLH-RegularSC               A    374    etrsc___
p_____RulingScript-Two                  A    375    rul2____
p_____Sho-Roman                         A    375    sho_____
p_____WiesbadenSwing-Roman              A    375    wn______
p_____MotterCorpusMM                    A    376    kprg____
p_____PenumbraMM                        A    377    zlrg____
p_____NuevaMM-It                        A    378    twi_____
p_____NuevaMM                           A    378    twr_____

pctbif8a  Cheltenham-HandtooledBdIt         A    379    chiin___
pctbf8a   Cheltenham-HandtooledBold         A    379    chin____
pgmbif8a  Garamond-HandtooledBoldItalic     A    379    gaiin___
pgmbf8a   Garamond-HandtooledBold           A    379    gain____
pcnbif8a  Century-HandtooledBoldItalic      A    379    iciin___
pcnbf8a   Century-HandtooledBold            A    379    icin____
pgmMM8a   ITCGaramondMM                     A    380    gc______
pgmMMi8a  ITCGaramondMM-It                  A    380    gci_____
p_____Cerigo-Bold                       A    381    qcb_____
p_____Cerigo-BoldItalic                 A    381    qcbi____
p_____Cerigo-Medium                     A    381    qcm_____
p_____Cerigo-MediumItalic               A    381    qcmi____
p_____Cerigo-Book                       A    381    qcw_____
p_____Cerigo-BookItalic                 A    381    qcwi____
phab8a    Hadriano-Bold                     A    382    hab_____
phax8a    Hadriano-ExtraBold                A    382    haeb____
phax8ac   Hadriano-ExtraBoldCondensed       A    382    haebc___
phal8a    Hadriano-Light                    A    382    hal_____
prcb8a    Romic-Bold                        A    383    rob_____
prcx8a    Romic-ExtraBold                   A    383    roeb____
prcl8a    Romic-Light                       A    383    rol_____
prcli8a   Romic-LightItalic                 A    383    roli____
prcm8a    Romic-Medium                      A    383    rom_____
p_____Copal-Decorated                   A    384    _ddc____
p_____Copal-Outline                     A    384    _dou____
p_____Copal-Solid                       A    384    _dsd____
pxpMM7a   ExPontoMM-Alternate               A    385    jvalt___
pxpMM7ab  ExPontoMM-Beginning               A    385    jvbeg___
pxpMM7ae  ExPontoMM-Ending                  A    385    jvend___
pxpMM8a   ExPontoMM                         A    385    jvrg____
pxpMMj8a  ExPontoMM-OsF                     A    385    jvros___
pysb8a    GoudySans-Bold                    A    386    hgb_____
pysbi8a   GoudySans-BoldItalic              A    386    hgbi____
pysk8a    GoudySans-Book                    A    386    hgbk____
pyski8a   GoudySans-BookItalic              A    386    hgbki___
pysc8a    GoudySans-Black                   A    386    hgbl____
pysci8a   GoudySans-BlackItalic             A    386    hgbli___
pysm8a    GoudySans-Medium                  A    386    hgm_____
pysmi8a   GoudySans-MediumItalic            A    386    hgmi____
p_____Delphin-I                         A    387    df1_____
p_____Delphin-IA                        A    387    df1a____
p_____Delphin-II                        A    387    df2_____
p_____Delphin-IIA                       A    387    df2a____
pc7b8a    Kompakt                           A    387    kq______
p_____KochAntiqua                       A    387    kv______
pnlb8a    Neuland                           A    387    nq______
p_____Caliban                           A    388    ni______

paxrw8a   Alexa                             A    388    xa______
pbzrw8a   Balzano                           A    388    zn______
p_____JimboMM                           A    389    jprg____
pb7b8ac   Bernhard-BoldCondensed            A    390    dhbc____
pnw1r8a   NeueHammerUnziale-One             A    390    nw1_____
pnw1r7a   NeueHammerUnziale-OneAlt          A    390    nw1al___
pnw2r8a   NeueHammerUnziale-Two             A    390    nw2_____
pnw2r7a   NeueHammerUnziale-TwoAlt          A    390    nw2al___
pclrl8a   Cloister-OpenFace                 A    390    oein____
poir8a    Ondine                            A    390    og______
pghr7a    Galahad-Alternate                 A    391    g_alt___
pghr8a    Galahad-Regular                   A    391    g_rg____
pghrj8a   Galahad-OsF                       A    391    g_ros___
popc8a    Optima-Black                      A    392    opbl____
popci8a   Optima-BlackItalic                A    392    opbli___
popd8a    Optima-DemiBold                   A    392    opd_____
popdi8a   Optima-DemiBoldItalic             A    392    opdi____
popx8a    Optima-ExtraBlack                 A    392    opf_____
popm8a    Optima-Medium                     A    392    opm_____
popmi8a   Optima-MediumItalic               A    392    opmi____
popxi8a   Optima-ExtraBlackItalic           A    392    opxfi___
p_____Aja                               A    393    ja______
p_____PopplExquisit-Medium              A    393    xqm_____
p_____PopplExquisit-Regular             A    393    xqrg____
ppi033    NationalCodes-Africa              A    394    ncaf____
ppi034    NationalCodes-America             A    394    ncam____
ppi035    NationalCodes-AsiaPlus            A    394    ncas____
ppi036    NationalCodes-Europe              A    394    nceu____
ppi037    NationalCodes-Universal           A    394    ncun____
p_____ElGreco                           A    395    eq______
pakm8a    Arkona-Medium                     A    395    xnm_____
pakr8a    Arkona-Regular                    A    395    xnrg____
pfgd8ap   FranklinGothic-DemiCmpr           A    396    frdcp___
pfgdi8ap  FranklinGothic-DemiCmprItal       A    396    frdpi___
pfgd8aq   FranklinGothic-DemiXCmpr          A    396    frdx____
pfgk8ap   FranklinGothic-BookCmpr           A    396    frwcp___
pfgki8ap  FranklinGothic-BookCmprItal       A    396    frwpi___
pfgk8aq   FranklinGothic-BookXCmpr          A    396    frwx____
pfgd8ac   FranklinGothic-DemiCnd            A    397    frdc____
pfgdi8ac  FranklinGothic-DemiCndItal        A    397    frdci___
pfgm8ac   FranklinGothic-MedCnd             A    397    frmc____
pfgmi8ac  FranklinGothic-MedCndItal         A    397    frmci___
pfgk8ac   FranklinGothic-BookCnd            A    397    frwc____
pfgki8ac  FranklinGothic-BookCndItal        A    397    frwci___
p_____Sassafras-Italic                  A    398    xfi_____
p_____Sassafras-LxItalic                A    398    xflgi___
p_____Sassafras-Lx                      A    398    xflgr___

p_____Sassafras-Roman                   A    398    xfr_____
p_____Sassafras-SxItalic                A    398    xfshi___
p_____Sassafras-Sx                      A    398    xfshr___
pkxrd8a   Kigali-Block                      A    399    kxblk___
pkxrw8a   Kigali-ZigZag                     A    399    kxdc____
pkxri8a   Kigali-Italic                     A    399    kxi_____
pkxli8a   Kigali-LxItalic                   A    399    kxlgi___
pkxl8a    Kigali-Lx                         A    399    kxlgr___
pkxr8a    Kigali-Roman                      A    399    kxr_____
pkxsi8a   Kigali-SxItalic                   A    399    kxshi___
pkxs8a    Kigali-Sx                         A    399    kxshr___
pajMMi8a  AJensonMM-It                      A    400    jei_____
pajMM8a   AJensonMM                         A    400    jer_____
pajMM7a   AJensonMM-Alt                     A    401    jealt___
pajMM8x   AJensonMM-Ep                      A    401    jeexp___
pajMMi7a  AJensonMM-ItAlt                   A    401    jeial___
pajMMic8a AJensonMM-ItSC                    A    401    jeisc___
pajMMi8x  AJensonMM-ItEp                    A    401    jeixp___
pajMMc8a  AJensonMM-SC                      A    401    jersc___
pajMMiw8a AJensonMM-Sw                      A    401    jeswc___
mbmb8a    BulmerMT-Bold                     A    402    ybb_____
mbmbi8a   BulmerMT-BoldItalic               A    402    ybbi____
mbmbd8a   BulmerMT-BoldDisplay              A    402    ybdb____
mbmbid8a  BulmerMT-BoldItalicDisplay        A    402    ybdbi___
mbmrd8a   BulmerMT-RegularDisplay           A    402    ybds____
mbmrid8a  BulmerMT-ItalicDisplay            A    402    ybdsi___
mbmri8a   BulmerMT-Italic                   A    402    ybi_____
mbmr8a    BulmerMT-Regular                  A    402    ybrg____
mbms8a    BulmerMT-SemiBold                 A    402    ybsb____
mbmsi8a   BulmerMT-SemiBoldItalic           A    402    ybsi____
mbmb7a    BulmerMT-BoldAlt                  A    403    ybab____
mbmbi7a   BulmerMT-BoldItalicAlt            A    403    ybabi___
mbmr7a    BulmerMT-RegularAlt               A    403    ybarg___
mbms7a    BulmerMT-SemiBoldAlt              A    403    ybasb___
mbmsi7a   BulmerMT-SemiBoldItalicAlt        A    403    ybasi___
mbmbd7a   BulmerMT-BoldDisplayAlt           A    403    ybbda___
mbmbid7a  BulmerMT-BoldItalDisplayAlt       A    403    ybbid___
mbmbi8x   BulmerMT-BoldItalicExpert         A    403    ybbix___
mbmb8x    BulmerMT-BoldExpert               A    403    ybbxp___
mbmri7a   BulmerMT-ItalicAlt                A    403    ybial___
mbmrid7a  BulmerMT-ItalicDisplayAlt         A    403    ybida___
mbmri8x   BulmerMT-ItalicExpert             A    403    ybixp___
mbmrd7a   BulmerMT-RegularDisplayAlt        A    403    ybrda___
mbmr8x    BulmerMT-RegularExpert            A    403    ybrgx___
mbmrc8a   BulmerMT-RegularSC                A    403    ybrsc___
mbmsc8a   BulmerMT-SemiBoldSC               A    403    ybsbs___
mbms8x    BulmerMT-SemiBoldExpert           A    403    ybsbx___

mbmsi8x   BulmerMT-SemiBoldItalExpert       A    403    ybsix___
pizr8a    Isabella                          A    404    iz______
prrr8a    Raphael                           A    404    rr______
pxcr8a    Eccentric                         A    404    xc______
p_____OceanSansMM-It                    A    405    ohi_____
p_____OceanSansMM                       A    405    ohr_____
plcrt8a   LucidaTypewriter                  A    406    tl______
plcbt8a   LucidaTypewriter-Bold             A    406    tlb_____
plcbot8a  LucidaTypewriter-BoldObl          A    406    tlbo____
plcrot8a  LucidaTypewriter-Obl              A    406    tlo_____
plsrt8a   LucidaSansTypewriter              A    407    ty______
plsbt8a   LucidaSansTypewriter-Bd           A    407    tyb_____
plsbot8a  LucidaSansTypewriter-BdObl        A    407    tybo____
plsrot8a  LucidaSansTypewriter-Obl          A    407    tyo_____
p_____BriemScriptMM-Exp                 A    408    iiexp___
p_____BriemScriptMM                     A    408    iirg____
p_____BriemScriptMM-SC                  A    408    iisc____
pkpMM8a   KeplMM                            A    409    k_______
pkpMMi8a  KeplMM-It                         A    409    k_i_____
pkpMM8x   KeplMM-Ep                         A    410    k_exp___
pkpMMic8a KeplMM-ItSC                       A    410    k_isc___
pkpMMi8x  KeplMM-ItEp                       A    410    k_ixp___
pkpMMp1   KeplMM-Or1                        A    410    k_or1___
pkpMMp2   KeplMM-Or2                        A    410    k_or2___
pkpMMp3   KeplMM-Or3                        A    410    k_or3___
pkpMMc8a  KeplMM-SC                         A    410    k_rsc___
pkpMMiw8a KeplMM-Sw                         A    410    k_sw____
pmjr8a    Mojo                              A    411    fbrg____
p_____Shuriken-Boy                      A    412    s_boy___
p_____Andreas                           A    413    hy______
poqr8a    Ouch                              A    414    oq______
p_____CongaBravaMM                      A    415    vbrg____
p_____CongaBravaMM-St                   A    415    vbst____
p_____CronosMM                          A    416    xo______
p_____CronosMM-It                       A    416    xoi_____
p_____CronosMM-Ep                       A    417    xoexp___
p_____CronosMM-ItSC                     A    417    xoisc___
p_____CronosMM-ItEp                     A    417    xoixp___
p_____CronosMM-Or                       A    417    xoor____
p_____CronosMM-SC                       A    417    xorsc___
p_____CronosMM-Sw                       A    417    xosw____
pdkb8a    Diskus-Bold                       A    418    d_b_____
pdkr8a    Diskus                            A    418    d_rg____
potb8a    Coronet-Bold                      A    418    n_b_____
potr8a    Coronet                           A    418    n_rg____
p_____ChaparralMM-It                    A    419    xti_____
p_____ChaparralMM                       A    419    xtr_____

p_____ChaparralMM-Ep                    A    420    xtexp___
p_____ChaparralMM-ItSC                  A    420    xtisc___
p_____ChaparralMM-ItEp                  A    420    xtixp___
p_____ChaparralMM-Or                    A    420    xtor____
p_____ChaparralMM-SC                    A    420    xtrsc___
ppi038    WiesbadenSwing-Dingbats           A    421    wgdb____
p_____BermudaLP-Dots                    A    422    uddc1___
p_____BermudaLP-Squiggle                A    422    uddc2___
p_____BermudaLP-Open                    A    422    udou____
p_____BermudaLP-Solid                   A    422    udrg____
p_____SpumoniLP                         A    423    uu______
pzpr8a    ZiptyDo                           A    424    zp______
pkwr7a    KoloLP-Alternates                 A    425    kwalt___
pkwr8an   KoloLP-Narrow                     A    425    kwn_____
pkwr8a    KoloLP-Regular                    A    425    kwrg____
pkwr8aw   KoloLP-Wide                       A    425    kwwd____
pnxr8a    Nyx                               A    426    nxrg____
p_____Banshee                           A    427    zo______
p_____KinesisMM                         A    428    kj______
p_____KinesisMM-It                      A    428    kji_____
p_____KinesisMM-Ep                      A    429    kjexp___
p_____KinesisMM-ItSC                    A    429    kjisc___
p_____KinesisMM-ItEp                    A    429    kjixp___
p_____KinesisMM-SC                      A    429    kjrsc___
p_____HardwoodLP                        A    430    hw______
p_____ManitoLP                          A    431    mw______
p_____Coriander                         A    432    uc______
p_____SpringLP                          A    433    td______
p_____SpringLP-Light                    A    433    tdl_____
pwhb8a    WendyLP-Bold                      A    434    whb_____
pwhl8a    WendyLP-Light                     A    434    whl_____
pwhm8a    WendyLP-Medium                    A    434    whm_____
p_____BickhamScriptMM-AltI              A    435    b_a1____
p_____BickhamScriptMM-AltII             A    435    b_a2____
p_____BickhamScriptMM-Beg               A    435    b_bg____
p_____BickhamScriptMM-End               A    435    b_en____
p_____BickhamScriptMM-Lig               A    435    b_lig___
p_____BickhamScriptMM-Or                A    435    b_or____
p_____BickhamScriptMM                   A    435    b_rg____
p_____BickhamScriptMM-SwCaps            A    435    b_swc___
pojr8a    OldClaudeLP                       A    436    oj______
pojrj8a   OldClaudeLP-OsF                   A    436    ojosf___
pojrc8a   OldClaudeLP-SC                    A    436    ojsc____
p_____FlorensLP                         A    437    fn______
p_____FlorensLP-Alternates              A    437    fnalt___
p_____BriemAkademiMM                    A    438    dkrg____
p_____Khaki-One                         A    439    hk1_____

p_____Khaki-Two                         A    439    hk2_____
p_____Pompeia-InlineItalic              A    440    vpiit___
p_____Pompeia-Inline                    A    440    vpin____
p_____Flood                             A    441    fyrg____
pwtMM7a   WatersTitlingMM-Alt               A    442    jwalt___
pwtMM7al  WatersTitlingMM-Lig               A    442    jwlig___
pwtMM8a   WatersTitlingMM                   A    442    jwr_____
pwtMMd8a  WatersTitlingMM-TallCaps          A    442    jwtlc___
p_____CelestiaAntiqua-Regular           A    443    zs______
p_____CelestiaAntiqua-Bold              A    443    zsb_____
p_____CelestiaAntiqua-BoldLig           A    443    zsbdl___
p_____CelestiaAntiqua-Italic            A    443    zsi_____
p_____CelestiaAntiqua-ItalicLig         A    443    zsitl___
p_____CelestiaAntiqua-Ligatures         A    443    zslig___
p_____CelestiaAntiqua-Ornaments         A    443    zsor____
p_____CelestiaAntiqua-Semibold          A    443    zssb____
p_____CelestiaAntiqua-SmBdLig           A    443    zssbl___
p_____CelestiaAntiqua-SC                A    443    zssc____
pmyr8ac   Myriad-Condensed                  A    444    wmc_____
pmyb8ac   Myriad-CnBold                     A    444    wmcb____
pmybi8ac  Myriad-CnBoldItalic               A    444    wmcbi___
pmyri8ac  Myriad-CnItalic                   A    444    wmci____
pmys8ac   Myriad-CnSemibold                 A    444    wms_____
pmysi8ac  Myriad-CnSemiboldItalic           A    444    wmsi____
p_____Immi505                           A    445    firg____
p_____MagnoliaMVB                       A    446    xm______
p_____MagnoliaMVB-Ligatures             A    446    xmlig___
p_____BossaNovaMVB                      A    447    xb______
p_____BossaNovaMVB-Alternates           A    447    xbalt___
p_____Emmascript                        A    448    xe______
p_____RusselOblique                     A    449    zr______
p_____RusselOblique-InformalAlt         A    449    zralt___
p_____RusselOblique-Informal            A    449    zrinf___
mdtb8a    DanteMT-Bold                      A    450    dlb_____
mdtbi8a   DanteMT-BoldItalic                A    450    dlbi____
mdtri8a   DanteMT-Italic                    A    450    dli_____
mdtm8a    DanteMT-Medium                    A    450    dlm_____
mdtmi8a   DanteMT-MediumItalic              A    450    dlmi____
mdtr8a    DanteMT-Regular                   A    450    dlrg____
mdtb7a    DanteMT-BoldAlt                   A    451    dlab____
mdtbi7a   DanteMT-BoldItalicAlt             A    451    dlabi___
mdtri7a   DanteMT-ItalicAlt                 A    451    dlai____
mdtr7a    DanteMT-RegularAlt                A    451    dlarg___
mdtm7a    DanteMT-MediumAlt                 A    451    dlasb___
mdtmi7a   DanteMT-MediumItalicAlt           A    451    dlasi___
mdtbij8a  DanteMT-BoldItalicOsF             A    451    dlbio___
mdtbi8x   DanteMT-BoldItalicExpert          A    451    dlbix___

mdtbj8a   DanteMT-BoldOsF                   A    451    dlbos___
mdtb8x    DanteMT-BoldExpert                A    451    dlbxp___
mdtrij8a  DanteMT-ItalicOsF                 A    451    dlios___
mdtri8x   DanteMT-ItalicExpert              A    451    dlixp___
mdtmij8a  DanteMT-MediumItalicOsF           A    451    dlmio___
mdtmi8x   DanteMT-MediumItalicExpert        A    451    dlmix___
mdtmj8a   DanteMT-MediumOsF                 A    451    dlmof___
mdtm8x    DanteMT-MediumExpert              A    451    dlmxp___
mdtr8x    DanteMT-RegularExpert             A    451    dlrgx___
mdtrc8a   DanteMT-RegularSC                 A    451    dlsc____
mdtrd8a   DanteMT-Titling                   A    451    dlttl___
myhb8a    StrayhornMT-Bold                  A    452    yhb_____
myhbi8a   StrayhornMT-BoldItalic            A    452    yhbi____
myhx8a    StrayhornMT-ExtraBold             A    452    yheb____
myhxi8a   StrayhornMT-ExtraBoldItalic       A    452    yhebi___
myhri8a   StrayhornMT-Italic                A    452    yhi_____
myhl8a    StrayhornMT-Light                 A    452    yhl_____
myhli8a   StrayhornMT-LightItalic           A    452    yhli____
myhr8a    StrayhornMT-Regular               A    452    yhrg____
myhbij8a  StrayhornMT-BoldItalicOsF         A    453    yhbio___
myhbj8a   StrayhornMT-BoldOsF               A    453    yhbos___
myhxj8a   StrayhornMT-ExtraBoldOsF          A    453    yhebs___
myhxij8a  StrayhornMT-ExtraBoldItaOsF       A    453    yheis___
myhrij8a  StrayhornMT-ItalicOsF             A    453    yhios___
myhlij8a  StrayhornMT-LightItalicOsF        A    453    yhlio___
myhlc8a   StrayhornMT-LightSC               A    453    yhlsc___
myhrc8a   StrayhornMT-RegularSC             A    453    yhrsc___
p_____Postino-Italic                    A    454    xzi_____
p_____Postino                           A    454    xzr_____
p_____ReliqMM-OsF                       A    455    qeosf___
p_____ReliqMM                           A    455    qer_____
p_____Critter                           A    456    wc______
p_____Cutout                            A    457    wu______
p_____Studz                             A    458    wp______
p_____Rad                               A    459    wr______
p_____Mythos                            A    460    wy______
p_____Toolbox                           A    461    wt______
p_____Quake                             A    462    wq______
pg2r8a    Giddyup-Regular                   A    463    wg______
pg2r      Giddyup-Thangs                    A    463    wgtha___
pmyrl8a   Myriad-Sketch                     A    464    wmske___
pmyrn8a   Myriad-Tilt                       A    464    wmtil___
mbeb8a    BellMT-Bold                       A    465    xlb_____
mbebi8a   BellMT-BoldItalic                 A    465    xlbi____
mberi8a   BellMT-Italic                     A    465    xli_____
mber8a    BellMT-Regular                    A    465    xlrg____
mbes8a    BellMT-SemiBold                   A    465    xlsb____

mbesi8a   BellMT-SemiBoldItalic             A    465    xlsbi___
mber7a    BellMT-RegularAlt                 A    466    xlarg___
mbes7a    BellMT-SemiBoldAlt                A    466    xlasb___
mbesi7a   BellMT-SemiBoldItalicAlt          A    466    xlasi___
mbebi8x   BellMT-BoldItalicExpert           A    466    xlbix___
mbeb8x    BellMT-BoldExpert                 A    466    xlbxp___
mber8x    BellMT-RegularExpert              A    466    xlexp___
mberi7a   BellMT-ItalicAlt                  A    466    xlial___
mberi8x   BellMT-ItalicExpert               A    466    xlixp___
mberx7a   BellMT-RegularExpertAlt           A    466    xlrxa___
mbesc8a   BellMT-SemiBoldSC                 A    466    xlsbs___
mbes8x    BellMT-SemiBoldExpert             A    466    xlsbx___
mberc8a   BellMT-RegularSC                  A    466    xlsc____
mbesi8x   BellMT-SemiBoldItalicExpert       A    466    xlsix___
mbesx7a   BellMT-SemiBoldExpertAlt          A    466    xlsxa___
mf1ri7a   FournierMT-ItalicAlt              A    467    xyali___
mf1r7a    FournierMT-RegularAlt             A    467    xyalt___
mf1ri8a   FournierMT-Italic                 A    467    xyi_____
mf1rij8a  FournierMT-ItalicOsF              A    467    xyios___
mf1ri8x   FournierMT-ItalicExpert           A    467    xyixp___
mf1rp     FournierMT-Ornaments              A    467    xyorn___
mf1r8a    FournierMT-Regular                A    467    xyrg____
mf1r8x    FournierMT-RegularExpert          A    467    xyrgx___
mf1rc8a   FournierMT-RegularSC              A    467    xysc____
mf1rid8a  FournierMT-ItalicTallCaps         A    467    xytti___
mf1rd8a   FournierMT-TallCaps               A    467    xyttl___
p_____CTMercurius-Black                 A    468    uibl____
p_____CTMercurius-BlackItalic           A    468    uibli___
p_____CTMercurius-Light                 A    468    uil_____
p_____CTMercurius-LightItalic           A    468    uili____
p_____CTMercurius-Medium                A    468    uim_____
p_____CTMercurius-MediumItalic          A    468    uimi____
p_____Fusaka-Regular                    A    469    fv______
p_____Fusaka-Alternates                 A    469    fvalt___
p_____MagnesiumMVB                      A    470    uh______
p_____MagnesiumMVB-Grime                A    470    uhdc1___
m_____GillFloriatedCapsMT-Alt           A    471    fwalt___
m_____GillFloriatedCapsMT               A    471    fwcap___
p_____Origami-Bold                      A    472    xrb_____
p_____Origami-BoldItalic                A    472    xrbi____
p_____Origami-Italic                    A    472    xri_____
p_____Origami-Medium                    A    472    xrm_____
p_____Origami-MediumItalic              A    472    xrmi____
p_____Origami-Regular                   A    472    xrrg____
p_____Origami-Semibold                  A    472    xrsb____
p_____Origami-SemiboldItalic            A    472    xrsbi___
p_____GreymantleMVB                     A    473    qg______

p_____GreymantleMVB-Alternates          A    473    qgalt___
p_____GreymantleMVB-Ornaments           A    473    qgor____
p_____VerveMM-A                         A    474    vvaaa___
p_____VerveMM                           A    474    vvr_____
p_____VolutaScript                      A    475    vs______
p_____VolutaScript-Alternates           A    475    vsalt___
p_____VolutaScript-Swash                A    475    vssw____
pcwr8a    Cottonwood                        A    476    c_______
piwr8a    Ironwood                          A    477    i_______
pjpr8a    Juniper                           A    478    j_______
pmqr8a    Mesquite                          A    479    m_______
ppor8a    Ponderosa                         A    480    p_______
pbir8a    Birch                             A    481    bi______
pb0r8a    Blackoak                          A    482    bo______
pmrr8a    Madrone                           A    483    mq______
pp0r8a    Poplar                            A    484    pp______
pwwr8a    Willow                            A    485    wi______
p_____Pepperwood-Fill                   A    486    uwfi____
p_____Pepperwood-Outline                A    486    uwou____
p_____Pepperwood-Regular                A    486    uwrg____
p_____Rosewood-Fill                     A    487    uxfi____
p_____Rosewood-Regular                  A    487    uxrg____
p_____Zebrawood-Fill                    A    488    uyfi____
p_____Zebrawood-Regular                 A    488    uyrg____
ppi026    WoodtypeOrnaments-One             A    489    woor____
ppi027    WoodtypeOrnaments-Two             A    489    woor2___
pg1r8a    Gothic-Thirteen                   A    490    gtt_____
ptph8ac   Tempo-HeavyCondensed              A    491    tehc____
ptphi8ac  Tempo-HeavyCondensedItalic        A    491    tehci___
ppsr8a    Parisian                          A    492    pn______
pumr8a    Umbra                             A    493    um______
mg4ri8a   GoudyModernMT-Italic              A    494    ygi_____
mg4r8a    GoudyModernMT                     A    494    ygrg____
ms1ri8a   ScotchRomanMT-Italic              A    495    ysi_____
ms1r8a    ScotchRomanMT                     A    495    ysrg____
mfsr8a    FalstaffMT                        A    496    fprg____
mifb8a    InflexMT-Bold                     A    497    iqb_____
mm8bl8a   MonotypeOldStyleMT-BoldOut        A    498    mbbou___
mfer8a    ForteMT                           A    499    fz______
mklr8a    KlangMT                           A    500    kn______
mmebw8a   MercuriusMT-BoldScript            A    501    uzbs____
p_____Smaragd                           A    502    smar____
p_____Smaragd-Alternate                 A    502    smara___
p_____Ariadne-Roman                     A    503    aria____
pdiri8a   Diotima-Italic                    A    503    dti_____
pdirij8a  Diotima-ItalicOsF                 A    503    dtiof___
pdir8a    Diotima-Roman                     A    503    dtr_____

pdirj8a   Diotima-RomanOsF                  A    503    dtros___
pdirc8a   Diotima-RomanSC                   A    503    dtrsc___
p_____Legault                           A    504    ul______
p_____Legault-Alternates                A    504    ula_____
p_____Legault-BoldAlternates            A    504    ulab____
p_____Legault-Bold                      A    504    ulb_____
p_____BlueIsland                        A    505    ihrg____
pbdp8ap   Bodoni-PosterCompressed           A    900    bdpsc___
pbax8a    Barmeno-ExtraBold                 A    900    bfeb____
pb0r8a    Blackoak                          A    900    bo______
zpcp33bc  Copperplate-ThirtyThreeBC         A    900    cp33c___
prtr      Carta                             A    900    cr______
pdcr8a    DomCasual                         A    900    dc______
mtlr8a    CastellarMT                       A    900    el______
mp2r8a    PepitaMT                          A    900    epscr___
pgsr8ac   GillSans-Condensed                A    900    gnc_____
pgsu8a    GillSans-UltraBold                A    900    gnub____
mgtrd8a   GoudyTextMT-LombardicCapitals     A    900    gxlom___
piwr8a    Ironwood                          A    900    i_______
mbfrw8a   BiffoMT                           A    900    ifscr___
pimx8a    Imago-ExtraBold                   A    900    ioeb____
mjori8a   JoannaMT-Italic                   A    900    joi_____
mjor8a    JoannaMT                          A    900    jor_____
p_____Colossalis-Bold                   A    900    kob_____
plgr8a    LetterGothic                      A    900    lg______
plor8a    Lithos-Regular                    A    900    lorg____
pmqr8a    Mesquite                          A    900    m_______
pmnrp     Minion-Ornaments                  A    900    moor____
pnurw8a   NuptialScript                     A    900    nu______
pper8a    PrestigeElite                     A    900    pe______
pqsr8a    Bellevue                          A    900    qsrg____
p_____AGOldFace-Outline                 A    900    qwou____
mrur8ac   RunicMT-Condensed                 A    900    ric_____
prwx8a    Rockwell-ExtraBold                A    900    rkeb____
pslr8a    Stencil                           A    900    st______
pwwr8a    Willow                            A    900    wi______
ppi027    WoodtypeOrnaments-Two             A    900    woor2___
pbir8a    Birch                             A    901    bi______
putrd8a   Utopia-Headline                   A    901    uthd____
p_____Critter                           A    901    wc______
pg2r8a    Giddyup-Regular                   A    901    wg______
pg2r      Giddyup-Thangs                    A    901    wgtha___
pmyrd8a   Myriad-Headline                   A    901    wmhd____
pmyrl8a   Myriad-Sketch                     A    901    wmske___
pmyrn8a   Myriad-Tilt                       A    901    wmtil___
p_____Studz                             A    901    wp______
p_____Quake                             A    901    wq______

p_____Rad                               A    901    wr______
p_____Toolbox                           A    901    wt______
p_____Cutout                            A    901    wu______
p_____Mythos                            A    901    wy______
pagd8a    AvantGarde-Demi                   A    902    agd_____
pagdo8a   AvantGarde-DemiOblique            A    902    agdo____
pagk8a    AvantGarde-Book                   A    902    agw_____
pagko8a   AvantGarde-BookOblique            A    902    agwo____
pamr8a    Americana                         A    902    am______
pamx8a    Americana-ExtraBold               A    902    ameb____
pacri8a   ACaslon-Italic                    A    902    awi_____
pacr8a    ACaslon-Regular                   A    902    awrg____
pacs8a    ACaslon-Semibold                  A    902    awsb____
pacsi8a   ACaslon-SemiboldItalic            A    902    awsbi___
pbab8a    Barmeno-Bold                      A    902    bfb_____
pbax8a    Barmeno-ExtraBold                 A    902    bfeb____
pbam8a    Barmeno-Medium                    A    902    bfm_____
pbar8a    Barmeno-Regular                   A    902    bfrg____
pbkd8a    Bookman-Demi                      A    902    bkd_____
pbkdi8a   Bookman-DemiItalic                A    902    bkdi____
pbkl8a    Bookman-Light                     A    902    bkl_____
pbkli8a   Bookman-LightItalic               A    902    bkli____
pb0r8a    Blackoak                          A    902    bo______
pcrb8a    Courier-Bold                      A    902    cob_____
pcrbo8a   Courier-BoldOblique               A    902    cobo____
pcrr8a    Courier                           A    902    com_____
pcrro8a   Courier-Oblique                   A    902    coo_____
prtr      Carta                             A    902    cr______
pfari8a   Formata-Italic                    A    902    fmi_____
pfam8a    Formata-Medium                    A    902    fmm_____
pfami8a   Formata-MediumItalic              A    902    fmmi____
pfar8a    Formata-Regular                   A    902    fmrg____
padb8a    AGaramond-Bold                    A    902    gdb_____
padbi8a   AGaramond-BoldItalic              A    902    gdbi____
padri8a   AGaramond-Italic                  A    902    gdi_____
padr8a    AGaramond-Regular                 A    902    gdrg____
phvr8a    Helvetica                         A    902    hv______
phvb8a    Helvetica-Bold                    A    902    hvb_____
phvbo8a   Helvetica-BoldOblique             A    902    hvbo____
phvr8an   Helvetica-Narrow                  A    902    hvn_____
phvb8an   Helvetica-Narrow-Bold             A    902    hvnb____
phvbo8an  Helvetica-Narrow-BoldOblique      A    902    hvnbo___
phvro8an  Helvetica-Narrow-Oblique          A    902    hvno____
phvro8a   Helvetica-Oblique                 A    902    hvo_____
pkur8a    Kaufmann                          A    902    kf______
ploc8a    Lithos-Black                      A    902    lobl____
plor8a    Lithos-Regular                    A    902    lorg____

pncb8a    NewCenturySchlbk-Bold             A    902    ncb_____
pncbi8a   NewCenturySchlbk-BoldItalic       A    902    ncbi____
pncri8a   NewCenturySchlbk-Italic           A    902    nci_____
pncr8a    NewCenturySchlbk-Roman            A    902    ncr_____
ppar8a    ParkAvenue                        A    902    pa______
ppsr8a    Parisian                          A    902    pn______
pplb8a    Palatino-Bold                     A    902    pob_____
pplbi8a   Palatino-BoldItalic               A    902    pobi____
pplri8a   Palatino-Italic                   A    902    poi_____
pplr8a    Palatino-Roman                    A    902    por_____
p_____Poetica-SuppOrnaments             A    902    pvor____
psyr      Symbol                            A    902    sy______
ptmb8a    Times-Bold                        A    902    tib_____
ptmbi8a   Times-BoldItalic                  A    902    tibi____
ptmri8a   Times-Italic                      A    902    tii_____
ptmr8a    Times-Roman                       A    902    tir_____
ptjb8a    Trajan-Bold                       A    902    tjb_____
ptkb8a    Tekton-Bold                       A    902    tkb_____
ptkr8a    Tekton                            A    902    tkrg____
ppi027    WoodtypeOrnaments-Two             A    902    woor2___
pzcmi8a   ZapfChancery-MediumItalic         A    902    zcmi____
pzdr      ZapfDingbats                      A    902    zd______
mp1rw8a   PalaceScriptMT                    A    903    _p______
pdjk8a    Delta-Book                        A    903    djw_____
pdjki8a   Delta-BookItalic                  A    903    djwi____
lenb8a    EngraversLH-BoldFace              A    903    ew______
pfarl8a   Formata-Outline                   A    903    fmou____
pgsxd8a   GillSans-ExtraBoldDisplay         A    903    gnxbd___
mbfrw8a   BiffoMT                           A    903    ifscr___
p_____MonaLisa-Solid                    A    903    iysd____
pjpr8a    Juniper                           A    903    j_______
p_____LoType-Medium                     A    903    ljm_____
p_____LoType-MediumItalic               A    903    ljmi____
p_____BerlinerGrotesk-Light             A    903    otl_____
prwb8ac   Rockwell-BoldCondensed            A    903    rkbc____
prwr8ac   Rockwell-Condensed                A    903    rkc_____
p_____AGBook-Stencil                    A    903    tast____
ms2bw8a   ScriptMT-Bold                     A    903    tbscr___
p_____Cutout                            A    903    wu______
ms1ri8a   ScotchRomanMT-Italic              A    903    ysi_____
ms1r8a    ScotchRomanMT                     A    903    ysrg____
mm7r8a    MaturaMT                          A    903    zz______
mr2l8a    CantoriaMT-Light                  A    904    eol_____
p_____Beesknees                         A    904    ikh_____
p_____LoType-Medium                     A    904    ljm_____
pm1rw8a   MonolineScriptMT                  A    904    on______
p_____Ozwald                            A    904    ozh_____

mknr8a    KinoMT                            A    904    qi______
pb7r8a    BernhardModern-Bold               A    904    rdb_____
puyr8a    UniversityRoman                   A    904    ur______
pbdp8ap   Bodoni-PosterCompressed           A    905    bdpsc___
pctl8ac   Cheltenham-LightCond              A    905    chlc____
pazl8ac   AkzidenzGroteskBE-LightCn         A    905    gflc____
pgsb8aq   GillSans-BoldExtraCondensed       A    905    gnbxc___
p_____GaramondBE-Condensed              A    905    gwc_____
p_____LoType-MediumCondensed            A    905    ljmc____
mrur8ac   RunicMT-Condensed                 A    905    ric_____
prwr8ac   Rockwell-Condensed                A    905    rkc_____
p_____Mezz-Regular                      A    906    vurg____
p_____Mezz-Black                        A    906    vuz_____
p_____CaflischScript-Bold               A    906    vwb_____
p_____CaflischScript-Regular            A    906    vwrg____
p_____Nueva-BoldExtendedItalic          A    906    vxbei___
p_____Nueva-BoldExtended                A    906    vxbv____
p_____Nueva-Italic                      A    906    vxi_____
p_____Nueva-Roman                       A    906    vxr_____
pmyb8a    Myriad-Bold                       A    906    wmb_____
pmybi8a   Myriad-BoldItalic                 A    906    wmbi____
pmyri8a   Myriad-Italic                     A    906    wmi_____
pmyr8a    Myriad-Roman                      A    906    wmr_____
pmnb8ac   Minion-BoldCondensed              A    907    mobc____
pmnbi8ac  Minion-BoldCondensedItalic        A    907    mobci___
pmnr8ac   Minion-Condensed                  A    907    moc_____
pmnri8ac  Minion-CondensedItalic            A    907    moci____
pvvx8ax   Viva-BoldExtraExtended            A    907    vabxe___
pvvl8ac   Viva-LightCondensed               A    907    valc____
pvvr8a    Viva-Regular                      A    907    varg____
p_____Sanvito-Light                     A    907    vyl_____
p_____Sanvito-Roman                     A    907    vyr_____
p_____Penumbra-SemiboldFlare            A    907    vzsfl___
p_____Penumbra-SemiboldSans             A    907    vzssa___
p_____Penumbra-SemiboldSerif            A    907    vzsse___
ppc___Poetica-ChanceryI                 A    908    pvch1___
ppc___Poetica-SuppLigatures             A    908    pvlg____
ppc___Poetica-SuppOrnaments             A    908    pvor____
ppc___Poetica-SuppSwashCapsI            A    908    pvsw1___
p_____CaflischScript-Bold               A    908    vwb_____
p_____CaflischScript-Regular            A    908    vwrg____
p_____Sanvito-BoldDisplay               A    908    vydb____
p_____Sanvito-Roman                     A    908    vyr_____
p_____Sanvito-Semibold                  A    908    vysb____
pxpr8a    ExPonto-Regular                   A    908    xprg____
padri8a   AGaramond-Italic                  A    912    gdi_____
padr8a    AGaramond-Regular                 A    912    gdrg____

pads8a    AGaramond-Semibold                A    912    gdsb____
padrc8a   AGaramond-RegularSC               A    912    gdsc____
padrd8a   AGaramond-Titling                 A    912    gdttl___
pacri8a   ACaslon-Italic                    A    913    awi_____
pacrp     ACaslon-Ornaments                 A    913    awor____
pacr8a    ACaslon-Regular                   A    913    awrg____
pacs8a    ACaslon-Semibold                  A    913    awsb____
pacrc8a   ACaslon-RegularSC                 A    913    awsc____
pajrd8a   AJenson-RegularDisplay            A    914    jsds____
pajri8a   AJenson-Italic                    A    914    jsi_____
pajr8a    AJenson-Regular                   A    914    jsrg____
pajs8a    AJenson-Semibold                  A    914    jssb____
pajrc8a   AJenson-RegularSC                 A    914    jssc____
p_____Jimbo-Bold                        A    915    jpb_____
pajsd8a   AJenson-SemiboldDisplay           A    915    jssp____
pkpbd8ax  Kepler-BoldExtendedDisplay        A    915    k_bed___
pkpsd8ac  Kepler-SemiboldCondensedDisplay   A    915    k_scd___
p_____Mezz-Bold                         A    915    vub_____
p_____Penumbra-BoldSerif                A    915    vzbsr___
p_____Cronos-Display                    A    915    xodr____
pxpr8a    ExPonto-Regular                   A    915    xprg____
p_____Chaparral-Display                 A    915    xtdr____
zpeurs    EuroSans-Regular                  A    916    _1______
zpeubs    EuroSans-Bold                     A    916    _1b_____
zpeubis   EuroSans-BoldItalic               A    916    _1bi____
zpeuris   EuroSans-Italic                   A    916    _1i_____
zpeurt    EuroMono-Regular                  A    916    _2______
zpeubt    EuroMono-Bold                     A    916    _2b_____
zpeubit   EuroMono-BoldItalic               A    916    _2bi____
zpeurit   EuroMono-Italic                   A    916    _2i_____
zpeur     EuroSerif-Regular                 A    916    _3______
zpeub     EuroSerif-Bold                    A    916    _3b_____
zpeubi    EuroSerif-BoldItalic              A    916    _3bi____
zpeuri    EuroSerif-Italic                  A    916    _3i_____
pmsrw8a   Mistral                           A    917    ms______
ppc___Poetica-ChanceryII                A    917    pvch2___
ppc___Poetica-SuppSwashCapsII           A    917    pvsw2___
pprl8a    PopplResidenz-Light               A    917    pzl_____
pprr8a    PopplResidenz-Regular             A    917    pzrg____
psrrw8a   SnellRoundhand-Script             A    917    sn______
mmebw8a   MercuriusMT-BoldScript            A    917    uzbs____
pwhm8a    WendyLP-Medium                    A    917    whm_____
pbzrw8a   Balzano                           A    917    zn______
ppi028    ScienceTypeCustomPi-No1T               ???    ________
ppi029    ScienceTypeCustomPi-No2T               ???    ________
ppi030    ScienceTypeCustomPi-No3T               ???    ________
ppi031    ScienceTypeCustomPi-No4T               ???    ________

ppi032    ScienceTypeCustomPi-No5T               ???    ________

A.3 Apple fonts

The ‘apple.map’ file has abbreviations for Apple fonts.

eecr8r   AppleChancery
eger8r   Geneva
ehtr8r   HoeflerText
emcr8r   Monaco
epi001   Zeal
eskr8r   Skia

A.4 Bitstream fonts

The ‘bitstrea.map’ file has abbreviations for Bitstream fonts.

bhvl8a    Swiss721BT-Light                  BT  0001    swz721l
bhvli8a   Swiss721BT-LightItalic            BT  0002    swz721li
bhvr8a    Swiss721BT-Roman                  BT  0003    swz721n
bhvri8a   Swiss721BT-Italic                 BT  0004    swz721i
bhvb8a    Swiss721BT-Bold                   BT  0005    swz721b
bhvbi8a   Swiss721BT-BoldItalic             BT  0006    swz721bi
bhvc8a    Swiss721BT-Black                  BT  0007    swz721k
bhvci8a   Swiss721BT-BlackItalic            BT  0008    swz721ki
bhvr8ac   Swiss721BT-RomanCondensed         BT  0009    swz721c
bhvb8ac   Swiss721BT-BoldCondensed          BT  0010    swz721bc
btmr8a    Dutch801BT-Roman                  BT  0011    dut801n
btmri8a   Dutch801BT-Italic                 BT  0012    dut801i
btmb8a    Dutch801BT-Bold                   BT  0013    dut801b
btmbi8a   Dutch801BT-BoldItalic             BT  0014    dut801bi
bopr8a    ZapfHumanist601BT-Roman           BT  0015    zhum601n
bopri8a   ZapfHumanist601BT-Italic          BT  0016    zhum601i
bopb8a    ZapfHumanist601BT-Bold            BT  0017    zhum601b
bopbi8a   ZapfHumanist601BT-BoldItalic      BT  0018    zhu601bi
bsvl8a    SouvenirITCbyBT-Light             BT  0019    souvnrl
bsvli8a   SouvenirITCbyBT-LightItalic       BT  0020    souvnrli
bsvd8a    SouvenirITCbyBT-Demi              BT  0021    souvnrd
bsvdi8a   SouvenirITCbyBT-DemiItalic        BT  0022    souvnrdi
bplr8a    ZapfCalligraphic801BT-Roman       BT  0023    zcal801n
bplri8a   ZapfCalligraphic801BT-Italic      BT  0024    zcal801i
bplb8a    ZapfCalligraphic801BT-Bold        BT  0025    zcal801b
bplbi8a   ZapfCalligraphic801BT-BoldItal    BT  0026    zca801bi
bexr8a    News702BT-Roman                   BT  0027    news702n
bexri8a   News702BT-Italic                  BT  0028    news702i
bexb8a    News702BT-Bold                    BT  0029    news702b
bexbi8a   News702BT-BoldItalic              BT  0030    new702bi
bbdk8a    BodoniBT-Book                     BT  0035    bodonin
bbdki8a   BodoniBT-BookItalic               BT  0036    bodonini

bbdb8a    BodoniBT-Bold                     BT  0037    bodonib
bbdbi8a   BodoniBT-BoldItalic               BT  0038    bodonibi
bcnr8ax   CenturyExpandedBT-Roman           BT  0039    centexn
bcnri8ax  CenturyExpandedBT-Italic          BT  0040    centexi
bcnb8ax   CenturyExpandedBT-Bold            BT  0041    centexb
bcnbi8ax  CenturyExpandedBT-BoldItalic      BT  0042    centexbi
bg3r8a    AmericanGaramondBT-Roman          BT  0043    amrgaran
bg3ri8a   AmericanGaramondBT-Italic         BT  0044    amrgarai
bg3b8a    AmericanGaramondBT-Bold           BT  0045    amrgarab
bg3bi8a   AmericanGaramondBT-BoldItalic     BT  0046    amrgarbi
bngr8a    NewsGothicBT-Roman                BT  0047    newsgotn
bngri8a   NewsGothicBT-Italic               BT  0048    newsgot
bngb8a    NewsGothicBT-Bold                 BT  0049    newsgotb
bngbi8a   NewsGothicBT-BoldItalic           BT  0050    newgotbi
bmlr8a    ZapfElliptical711BT-Roman         BT  0051    zelp711n
bmlri8a   ZapfElliptical711BT-Italic        BT  0052    zelp711i
bmlb8a    ZapfElliptical711BT-Bold          BT  0053    zelp711b
bmlbi8a   ZapfElliptical711BT-BoldItalic    BT  0054    zel711bi
bbdr8a    BodoniBT-Roman                    BT  0055    bodoni
bbdri8a   BodoniBT-Italic                   BT  0056    bodonii
bhbr8a    HoboBT-Regular                    BT  0057    hobon
bonr8a    News705BT-RomanB                  BT  0058    news705n
bonri8a   News705BT-ItalicB                 BT  0059    news705i
bonb8a    News705BT-BoldB                   BT  0060    news705b
b_____BusoramaITCbyBT-Medium            BT  0062    busoramm
bglr8a    GalliardITCbyBT-Roman             BT  0067    galirdn
bglri8a   GalliardITCbyBT-Italic            BT  0068    galirdi
bglb8a    GalliardITCbyBT-Bold              BT  0069    galirdb
bglbi8a   GalliardITCbyBT-BoldItalic        BT  0070    galirdbi
bbkl8a    BookmanITCbyBT-Light              BT  0075    bookmnl
bbkli8a   BookmanITCbyBT-LightItalic        BT  0076    bookmnli
bbkd8a    BookmanITCbyBT-Demi               BT  0079    bookmnd
bbkdi8a   BookmanITCbyBT-DemiItalic         BT  0080    bookmndi
bcsr8a    CenturySchoolbookBT-Roman         BT  0083    censbkn
bcsri8a   CenturySchoolbookBT-Italic        BT  0084    censbki
bcsb8a    CenturySchoolbookBT-Bold          BT  0085    censbkb
bcsbi8a   CenturySchoolbookBT-BoldItalic    BT  0086    censbkbi
bkrr8a    KorinnaITCbyBT-Regular            BT  0087    korinan
bkrri8a   KorinnaITCbyBT-KursivRegular      BT  0088    korinani
bkrb8a    KorinnaITCbyBT-Bold               BT  0089    korinab
bkrbi8a   KorinnaITCbyBT-KursivBold         BT  0090    korinabi
bhva8a    Swiss721BT-Thin                   BT  0100    swz721t
bhvai8a   Swiss721BT-ThinItalic             BT  0101    swz721ti
bhvh8a    Swiss721BT-Heavy                  BT  0102    swz721h
bhvhi8a   Swiss721BT-HeavyItalic            BT  0103    swz721hi
bhvh8aq   Swiss911BT-ExtraCompressed        BT  0105    swz911xc
bhvh8au   Swiss911BT-UltraCompressed        BT  0106    swz911uc

bgor8a    GoudyOldStyleBT-Roman             BT  0107    goudyosn
bgori8a   GoudyOldStyleBT-Italic            BT  0108    goudyosi
bgob8a    GoudyOldStyleBT-Bold              BT  0109    goudyosb
bgobi8a   GoudyOldStyleBT-BoldItalic        BT  0110    goudosbi
bgox8a    GoudyOldStyleBT-ExtraBold         BT  0111    goudosxb
bpnr8a    Aldine721BT-Roman                 BT  0112    ald721n
bpnri8a   Aldine721BT-Italic                BT  0113    ald721i
bpnb8a    Aldine721BT-Bold                  BT  0114    ald721b
bpnbi8a   Aldine721BT-BoldItalic            BT  0115    ald721bi
bjnr8a    KisBT-Roman                       BT  0116    kisn
bjnri8a   KisBT-Italic                      BT  0117    kisi
bc5r8a    Caslon540BT-Roman                 BT  0118    casl540n
bc5ri8a   Caslon540BT-Italic                BT  0119    casl540i
bg1r8a    GothicNo13BT-Regular              BT  0120    gothic13
bbol8a    BalloonBT-Light                   BT  0121    balloonl
bwnr8a    WindsorBT-Roman                   BT  0122    wndsrn
bbgk8a    BenguiatITCbyBT-Book              BT  0123    benguian
bbgki8a   BenguiatITCbyBT-BookItalic        BT  0124    benguini
bbgb8a    BenguiatITCbyBT-Bold              BT  0127    benguiab
bbgbi8a   BenguiatITCbyBT-BoldItalic        BT  0128    benguibi
bbdp8a    PosterBodoniBT-Roman              BT  0129    pstrbodn
bbdpi8a   PosterBodoniBT-Italic             BT  0130    pstrbodi
bbwr8a    BroadwayBT-Regular                BT  0131    brdwayn
bngl8a    NewsGothicBT-Light                BT  0132    newsgotl
bngli8a   NewsGothicBT-LightItalic          BT  0133    newgotli
bopd8a    ZapfHumanist601BT-Demi            BT  0134    zhum601d
bopdi8a   ZapfHumanist601BT-DemiItalic      BT  0135    zhu601di
bopu8a    ZapfHumanist601BT-Ultra           BT  0136    zhum601u
bopui8a   ZapfHumanist601BT-UltraItalic     BT  0137    zhu601ui
bful8a    FuturaBT-Light                    BT  0138    futural
bfuli8a   FuturaBT-LightItalic              BT  0139    futurali
bfuk8a    FuturaBT-Book                     BT  0140    futuran
bfuki8a   FuturaBT-BookItalic               BT  0141    futurani
bfum8a    FuturaBT-Medium                   BT  0142    futuram
bfumi8a   FuturaBT-MediumItalic             BT  0143    futurami
bfub8a    FuturaBT-Bold                     BT  0144    futurab
bfubi8a   FuturaBT-BoldItalic               BT  0145    futurabi
bfuh8a    FuturaBT-Heavy                    BT  0146    futurah
bfuhi8a   FuturaBT-HeavyItalic              BT  0147    futurahi
bfux8a    FuturaBT-ExtraBlack               BT  0148    futuraxk
bfuxi8a   FuturaBT-ExtraBlackItalic         BT  0149    futurxki
bsnk8a    Geometric212BT-Book               BT  0150    geo212n
bsnh8a    Geometric212BT-Heavy              BT  0151    geo212h
bagk8a    AvantGardeITCbyBT-Book            BT  0154    avgardn
bagko8a   AvantGardeITCbyBT-BookOblique     BT  0155    avgardni
bagm8a    AvantGardeITCbyBT-Medium          BT  0156    avgardm
bagmo8a   AvantGardeITCbyBT-MediumOblique   BT  0157    avgardmi

bagd8a    AvantGardeITCbyBT-Demi            BT  0158    avgardd
bagdo8a   AvantGardeITCbyBT-DemiOblique     BT  0159    avgarddo
bkbk8a    KabelITCbyBT-Book                 BT  0162    kabeln
bkbm8a    KabelITCbyBT-Medium               BT  0163    kabelm
bkbd8a    KabelITCbyBT-Demi                 BT  0164    kabeld
bkbb8a    KabelITCbyBT-Bold                 BT  0165    kabelb
bkbu8a    KabelITCbyBT-Ultra                BT  0166    kabelu
bs4r8a    ShotgunBT-Regular                 BT  0167    shotgunn
bhvl8ac   Swiss721BT-LightCondensed         BT  0169    swz721lc
bhvli8ac  Swiss721BT-LightCondensedItalic   BT  0170    sw721lci
bhvri8ac  Swiss721BT-ItalicCondensed        BT  0171    swz721ci
bhvbi8ac  Swiss721BT-BoldCondensedItalic    BT  0172    sw721bci
bhvc8ac   Swiss721BT-BlackCondensed         BT  0173    swz721kc
bhvci8ac  Swiss721BT-BlackCondensedItalic   BT  0174    sw721kci
bhvbl8a   Swiss721BT-BoldOutline            BT  0175    swz721bo
bunl8a    ZurichBT-Light                    BT  0176    zurchl
bunli8a   ZurichBT-LightItalic              BT  0177    zurchli
bunr8a    ZurichBT-Roman                    BT  0178    zurchn
bunri8a   ZurichBT-Italic                   BT  0179    zurchi
bunb8a    ZurichBT-Bold                     BT  0180    zurchb
bunbi8a   ZurichBT-BoldItalic               BT  0181    zurchbi
bunc8a    ZurichBT-Black                    BT  0182    zurchk
bunci8a   ZurichBT-BlackItalic              BT  0183    zurchki
b____BoltBoldITCbyBT-Regular           BT  0184    boltb
bslr8a    StencilBT-Regular                 BT  0191    stenciln
batm8a    AmerTypewriterITCbyBT-Medium      BT  0193    amrtypen
batb8a    AmerTypewriterITCbyBT-Bold        BT  0194    amrtypeb
bd0b8a    DavidaBoldBT-Regular              BT  0195    davidab
bs5r8a    SnellBT-Regular                   BT  0196    snelln
bs5b8a    SnellBT-Bold                      BT  0197    snellb
bs5c8a    SnellBT-Black                     BT  0198    snellk
bburw8a   BrushScriptBT-Regular             BT  0199    brushscn
bful8ac   FuturaBT-LightCondensed           BT  0200    futuralc
bfum8ac   FuturaBT-MediumCondensed          BT  0201    futuramc
bfub8ac   FuturaBT-BoldCondensed            BT  0202    futurabc
bfubi8ac  FuturaBT-BoldCondensedItalic      BT  0203    futurbci
bfux8ac   FuturaBT-ExtraBlackCondensed      BT  0204    futurxkc
bfuxi8ac  FuturaBT-ExtraBlackCondItalic     BT  0205    futuxkci
bfuc8a    FuturaBlackBT-Regular             BT  0206    futurak
botr8a    Ribbon131BT-Regular               BT  0207    ribn131n
botb8a    Ribbon131BT-Bold                  BT  0208    ribn131b
bi1r8a    ImpressBT-Regular                 BT  0209    impressn
ba7r8a    AmericanTextBT-Regular            BT  0211    amrtextn
benrf8a   EngraversOldEnglishBT-Regular     BT  0212    engrolen
bo2r8a    OrbitBbyBT-Regular                BT  0213    orbitbn
bp4r8a    PioneerITCbyBT-Regular            BT  0214    pioneern
bv0r8a    VinetaBT-Regular                  BT  0215    vinetan

byml8a    StymieBT-Light                    BT  0216    stymiel
bymm8a    StymieBT-Medium                   BT  0217    stymiem
bymb8a    StymieBT-Bold                     BT  0218    stymieb
bymx8a    StymieBT-ExtraBold                BT  0219    stymiexb
btfl8a    TiffanyITCbyBT-Light              BT  0220    tifanyl
btfli8a   TiffanyITCbyBT-LightItalic        BT  0221    tifanyli
btfd8a    TiffanyITCbyBT-Demi               BT  0224    tifanyd
btfdi8a   TiffanyITCbyBT-DemiItalic         BT  0225    tifanydi
btfh8a    TiffanyITCbyBT-Heavy              BT  0226    tifanyh
btfhi8a   TiffanyITCbyBT-HeavyItalic        BT  0227    tifanyhi
bcdb8ac   ClarendonBT-BoldCondensed         BT  0228    clrndnbc
bcbcd8a   CooperBT-BlackHeadline            BT  0231    coopbla
bcbcid8a  CooperBT-BlackItalicHeadline      BT  0232    coopblia
bs4rl8a   ShotgunBlanksBT-Regular           BT  0234    shotgunk
bnar8a    Dutch823BT-RomanB                 BT  0235    dut823n
bnari8a   Dutch823BT-ItalicB                BT  0236    dut823i
bnab8a    Dutch823BT-BoldB                  BT  0237    dut823b
berl8a    ErasITCbyBT-Light                 BT  0238    erasl
berk8a    ErasITCbyBT-Book                  BT  0239    erasn
berm8a    ErasITCbyBT-Medium                BT  0240    erasm
berd8a    ErasITCbyBT-Demi                  BT  0241    erasd
berb8a    ErasITCbyBT-Bold                  BT  0242    erasb
beru8a    ErasITCbyBT-Ultra                 BT  0243    erasu
besr8a    Square721BT-Roman                 BT  0246    sqr721n
besb8a    Square721BT-Bold                  BT  0247    sqr721b
besr8ac   Square721BT-RomanCondensed        BT  0248    sqr721c
besb8ac   Square721BT-BoldCondensed         BT  0249    sqr721bc
besr8ax   Square721BT-RomanExtended         BT  0250    sqr721e
besb8ax   Square721BT-BoldExtended          BT  0251    sqr721be
bsfa8a    SerifaBT-Thin                     BT  0252    serifat
bsfai8a   SerifaBT-ThinItalic               BT  0253    serifati
bsfl8a    SerifaBT-Light                    BT  0254    serifal
bsfli8a   SerifaBT-LightItalic              BT  0255    serifali
bsfr8a    SerifaBT-Roman                    BT  0256    serifan
bsfri8a   SerifaBT-Italic                   BT  0257    serifai
bsfb8a    SerifaBT-Bold                     BT  0258    serifab
bsfc8a    SerifaBT-Black                    BT  0259    serifak
bunr8ax   ZurichBT-RomanExtended            BT  0260    zurche
bunb8ax   ZurichBT-BoldExtended             BT  0261    zurchbe
bunc8ax   ZurichBT-BlackExtended            BT  0262    zurchke
bunu8ax   ZurichBT-UltraBlackExtended       BT  0263    zurchuke
ba9r8a    Humanist970BT-RomanC              BT  0264    hum970n
ba9b8a    Humanist970BT-BoldC               BT  0265    hum970b
bndr8a    Transitional511BT-Roman           BT  0266    tran511n
bndri8a   Transitional511BT-Italic          BT  0267    tran511i
bndb8a    Transitional511BT-Bold            BT  0270    tran511b
bndbi8a   Transitional511BT-BoldItalic      BT  0271    tra511bi

bsbr8a    ClassicalGaramondBT-Roman         BT  0277    clsgaran
bsbri8a   ClassicalGaramondBT-Italic        BT  0278    clsgarai
bsbb8a    ClassicalGaramondBT-Bold          BT  0279    clsgarab
bcarf8a   CaslonOldFaceBT-Roman             BT  0280    casolfcn
bcarif8a  CaslonOldFaceBT-Italic            BT  0281    casolfci
bcdl8a    ClarendonBT-Light                 BT  0282    clrndnl
bcdr8a    ClarendonBT-Roman                 BT  0283    clrndnn
bcdb8a    ClarendonBT-Bold                  BT  0284    clrndnb
bcur8a    CenturyOldstyleBT-Roman           BT  0285    cenolstn
bcuri8a   CenturyOldstyleBT-Italic          BT  0286    cenolsti
bcub8a    CenturyOldstyleBT-Bold            BT  0287    cenolstb
bgsl8a    Humanist521BT-Light               BT  0288    hum521l
bgsli8a   Humanist521BT-LightItalic         BT  0289    hum521li
bgsr8a    Humanist521BT-Roman               BT  0290    hum521n
bgsri8a   Humanist521BT-Italic              BT  0291    hum521i
bgsb8a    Humanist521BT-Bold                BT  0292    hum521b
bgsbi8a   Humanist521BT-BoldItalic          BT  0293    hum521bi
bgsx8a    Humanist521BT-ExtraBold           BT  0294    hum521xb
bgsu8a    Humanist521BT-UltraBold           BT  0295    hum521ub
bunl8ac   ZurichBT-LightCondensed           BT  0296    zurchlc
bunli8ac  ZurichBT-LightCondensedItalic     BT  0297    zurchlci
bunr8ac   ZurichBT-RomanCondensed           BT  0298    zurch
bunri8ac  ZurichBT-ItalicCondensed          BT  0299    zurchci
bunb8ac   ZurichBT-BoldCondensed            BT  0300    zurchbc
bunbi8ac  ZurichBT-BoldCondensedItalic      BT  0301    zurchbci
bbbr8a    Aldine401BT-RomanA                BT  0302    ald401n
bbbri8a   Aldine401BT-ItalicA               BT  0303    ald401i
bbbb8a    Aldine401BT-BoldA                 BT  0304    ald401b
bbbbi8a   Aldine401BT-BoldItalicA           BT  0305    ald401bi
bmpl8a    GeometricSlab703BT-Light          BT  0306    geo703l
bmpli8a   GeometricSlab703BT-LightItalic    BT  0307    geo703li
bmpm8a    GeometricSlab703BT-Medium         BT  0308    geo703m
bmpmi8a   GeometricSlab703BT-MediumItalic   BT  0309    geo703mi
bmpb8a    GeometricSlab703BT-Bold           BT  0310    geo703b
bmpbi8a   GeometricSlab703BT-BoldItalic     BT  0311    geo703bi
bmpx8a    GeometricSlab703BT-XtraBold       BT  0312    geo703xb
bmpxi8a   GeometricSlab703BT-XtraBoldItal   BT  0313    ge703xbi
buyr8a    UniversityRomanBT-Regular         BT  0316    univern
buyb8a    UniversityRomanBT-Bold            BT  0317    univernb
bior8a    News701BT-RomanA                  BT  0318    news701n
biori8a   News701BT-ItalicA                 BT  0319    news701i
biob8a    News701BT-BoldA                   BT  0320    news701b
befr8a    HumanistSlabserif712BT-Roman      BT  0321    hums712n
befri8a   HumanistSlabserif712BT-Italic     BT  0322    hums712i
befb8a    HumanistSlabserif712BT-Bold       BT  0323    hums712b
befc8a    HumanistSlabserif712BT-Black      BT  0324    hums712k
bwnl8ac   WindsorBT-LightCondensed          BT  0326    wndsrlc

bwnr8ax   WindsorBT-Elongated               BT  0327    wndsre
bubr8a    BauerBodoniBT-Roman               BT  0328    baubodn
bubri8a   BauerBodoniBT-Italic              BT  0329    baubodi
bubb8a    BauerBodoniBT-Bold                BT  0330    baubodb
bubbi8a   BauerBodoniBT-BoldItalic          BT  0331    baubodbi
bubc8a    BauerBodoniBT-Black               BT  0332    baubodk
bubci8a   BauerBodoniBT-BlackItalic         BT  0333    baubodki
bubb8ac   BauerBodoniBT-BoldCondensed       BT  0334    baubodbc
bubc8ac   BauerBodoniBT-BlackCondensed      BT  0335    baubodkc
biir8a    Dutch766BT-RomanA                 BT  0338    dut766n
biiri8a   Dutch766BT-ItalicA                BT  0339    dut766i
biib8a    Dutch766BT-BoldA                  BT  0340    dut766b
bcahf8a   CaslonOldFaceBT-Heavy             BT  0342    casolfch
bgmk8a    GaramondITCbyBT-Book              BT  0345    garamdn
bgmki8a   GaramondITCbyBT-BookItalic        BT  0346    garamdni
bgmb8a    GaramondITCbyBT-Bold              BT  0347    garamdb
bgmbi8a   GaramondITCbyBT-BoldItalic        BT  0348    garamdbi
bcdc8a    ClarendonBT-Black                 BT  0351    clrndnk
bi3r8a    ImperialBT-Roman                  BT  0352    imperln
bi3ri8a   ImperialBT-Italic                 BT  0353    imperli
bi3b8a    ImperialBT-Bold                   BT  0354    imperlb
b_____BellCentennialBT-NameAndNumber    BT  0355    belcenam
b_____BellCentennialBT-Address          BT  0356    belceadr
b_____BellCentennialBT-SubCaption       BT  0357    belcesub
b_____BellCentennialBT-BoldListing      BT  0358    belcelis
bbers8a   BellGothicBT-Roman                BT  0359    belgothn
bbebs8a   BellGothicBT-Bold                 BT  0360    belgothb
bbecs8a   BellGothicBT-Black                BT  0361    belgothk
bpar8a    ParkAvenueBT-Regular              BT  0362    parkaven
bn7r8a    News706BT-RomanC                  BT  0363    news706n
bn7ri8a   News706BT-ItalicC                 BT  0364    news706i
bn7b8a    News706BT-BoldC                   BT  0365    news706b
blbj8a    LubalinGraphITCbyBT-XtraLight     BT  0366    lubgrxl
blbk8a    LubalinGraphITCbyBT-Book          BT  0367    lubgrhn
blbm8a    LubalinGraphITCbyBT-Medium        BT  0368    lubgrhm
blbd8a    LubalinGraphITCbyBT-Demi          BT  0369    lubgrhd
blbb8a    LubalinGraphITCbyBT-Bold          BT  0370    lubgrhb
bhvl8ax   Swiss721BT-LightExtended          BT  0371    swz721le
bhvr8ax   Swiss721BT-RomanExtended          BT  0372    swz721e
bhvb8ax   Swiss721BT-BoldExtended           BT  0373    swz721be
bhvc8ax   Swiss721BT-BlackExtended          BT  0374    swz721ke
bngr8ac   NewsGothicBT-RomanCondensed       BT  0375    newsgotc
bngri8ac  NewsGothicBT-ItalicCondensed      BT  0376    newgotci
bngb8ac   NewsGothicBT-BoldCondensed        BT  0377    newgotbc
bngbi8ac  NewsGothicBT-BoldCondItalic       BT  0378    newgtbci
bctk8a    CheltenhamITCbyBT-Book            BT  0381    ichlthmn
bctki8a   CheltenhamITCbyBT-BookItalic      BT  0382    ichlthni

bctb8a    CheltenhamITCbyBT-Bold            BT  0383    ichlthmb
bctbi8a   CheltenhamITCbyBT-BoldItalic      BT  0384    ichlthbi
baol8a    Incised901BT-Light                BT  0387    inc901l
baor8a    Incised901BT-Roman                BT  0388    inc901n
baori8a   Incised901BT-Italic               BT  0389    inc901i
baob8a    Incised901BT-Bold                 BT  0390    inc901b
baoc8a    Incised901BT-Black                BT  0391    inc901k
boar      OCRAbyBT-Regular                  BT  0395    ocran
bbhl8a    BauhausITCbyBT-Light              BT  0405    bauhausl
bbhm8a    BauhausITCbyBT-Medium             BT  0406    bauhausm
bbhb8a    BauhausITCbyBT-Bold               BT  0408    bauhausb
bbhh8a    BauhausITCbyBT-Heavy              BT  0409    bauhaush
bcyl8a    SquareSlabserif711BT-Light        BT  0410    sqrs711l
bcym8a    SquareSlabserif711BT-Medium       BT  0411    sqrs711m
bcyb8a    SquareSlabserif711BT-Bold         BT  0412    sqrs711b
bper8a    Prestige12PitchBT-Roman           BT  0413    presti12
blgr8a    LetterGothic12PitchBT-Roman       BT  0414    letgot12
bcdh8a    ClarendonBT-Heavy                 BT  0416    clrndnh
bccr8a    Freehand471BT-Regular             BT  0418    frhnd471
bcrr8a    Courier10PitchBT-Roman            BT  0419    cur10ptn
bcprs8a   CopperplateGothicBT-Roman         BT  0420    copgothn
bcpbs8a   CopperplateGothicBT-Bold          BT  0421    copgothb
bcphs8a   CopperplateGothicBT-Heavy         BT  0422    copgothh
bcprs8ac  CopperplateGothicBT-RomanCond     BT  0423    copgothc
bcpbs8ac  CopperplateGothicBT-BoldCond      BT  0424    copgotbc
bfqr8a    FrizQuadrataITCbyBT-Roman         BT  0428    frzquadn
bfqb8a    FrizQuadrataITCbyBT-Bold          BT  0429    frzquadb
bqrl8a    QuorumITCbyBT-Light               BT  0438    quoruml
bqrm8a    QuorumITCbyBT-Medium              BT  0440    quorumm
bqrc8a    QuorumITCbyBT-Black               BT  0442    quorumk
bcab8a    CaslonBT-Bold                     BT  0443    caslonb
bcabi8a   CaslonBT-BoldItalic               BT  0444    caslonbi
bzcm8a    ZapfChanceryITCbyBT-Medium        BT  0447    zafcham
bzcmi8a   ZapfChanceryITCbyBT-MediumItal    BT  0448    zafchami
bzcd8a    ZapfChanceryITCbyBT-Demi          BT  0449    zafchad
bzcb8a    ZapfChanceryITCbyBT-Bold          BT  0450    zafchab
bgmk8ac   GaramondITCbyBT-BookCondensed     BT  0453    garamdnc
bgmki8ac  GaramondITCbyBT-BookCondItalic    BT  0454    garamnci
bgmb8ac   GaramondITCbyBT-BoldCondensed     BT  0455    garamdbc
bgmbi8ac  GaramondITCbyBT-BoldCondItalic    BT  0456    garambci
bugk8a    BenguiatGothicITCbyBT-Book        BT  0459    bengothn
bugki8a   BenguiatGothicITCbyBT-BookItal    BT  0460    bengotni
bugb8a    BenguiatGothicITCbyBT-Bold        BT  0463    bengothb
bugbi8a   BenguiatGothicITCbyBT-BoldItal    BT  0464    bengotbi
baob8ac   Incised901BT-BoldCondensed        BT  0467    inc901bc
baop8a    Incised901BT-Compact              BT  0468    inc901c
baoh8ax   Incised901BT-Nord                 BT  0469    in901xk

baohi8ax  Incised901BT-NordItalic           BT  0470    in901xki
brwl8a    GeometricSlab712BT-LightA         BT  0483    geo712l
brwli8a   GeometricSlab712BT-LightItalicA   BT  0484    geo712li
brwm8a    GeometricSlab712BT-MediumA        BT  0485    geo712m
brwmi8a   GeometricSlab712BT-MediumItalA    BT  0486    geo712mi
brwb8a    GeometricSlab712BT-BoldA          BT  0487    geo712b
brwx8a    GeometricSlab712BT-ExtraBoldA     BT  0488    geo712xb
bmdr8a    Latin725BT-Roman                  BT  0489    lat725n
bmdri8a   Latin725BT-Italic                 BT  0490    lat725i
bmdm8a    Latin725BT-Medium                 BT  0491    lat725m
bmdmi8a   Latin725BT-MediumItalic           BT  0492    lat725mi
bmdb8a    Latin725BT-Bold                   BT  0493    lat725b
bmdbi8a   Latin725BT-BoldItalic             BT  0494    lat725bi
bgsr8ac   Humanist521BT-RomanCondensed      BT  0495    hum521c
bgsb8ac   Humanist521BT-BoldCondensed       BT  0496    hum521bc
bgsx8ac   Humanist521BT-XtraBoldCondensed   BT  0497    hu521xbc
bamr8a    AmericanaBT-Roman                 BT  0498    amrcanan
bamri8a   AmericanaBT-Italic                BT  0499    amrcanai
bamb8a    AmericanaBT-Bold                  BT  0500    amrcanab
bamx8a    AmericanaBT-ExtraBold             BT  0501    amrcanxb
bamx8ac   AmericanaBT-ExtraBoldCondensed    BT  0502    amrcaxbc
bsnk8ac   Geometric212BT-BookCondensed      BT  0503    geo212nc
bsnh8ac   Geometric212BT-HeavyCondensed     BT  0504    geo212hc
blir8ac   LatinExtraCondensedBT-Regular     BT  0507    latinxc
bctbd8a   CheltenhamBT-BoldHeadline         BT  0508    chltboa
bctbdi8a  CheltenhamBT-BoldItalicHeadline   BT  0509    chltboia
ba6r8a    AmeliaBT-Regular                  BT  0510    amelian
bb8c8a    BlippoBT-Black                    BT  0511    blippok
boxr8a    OnyxBT-Regular                    BT  0519    onyxn
bm2l8a    Geometric415BT-LiteA              BT  0520    geo415l
bm2li8a   Geometric415BT-LiteItalicA        BT  0521    geo415li
bm2m8a    Geometric415BT-MediumA            BT  0522    geo415m
bm2mi8a   Geometric415BT-MediumItalicA      BT  0523    geo415mi
bm2c8a    Geometric415BT-BlackA             BT  0524    geo415k
bm2ci8a   Geometric415BT-BlackItalicA       BT  0525    geo415ki
blgri8a   LetterGothic12PitchBT-Italic      BT  0532    legot12i
blgb8a    LetterGothic12PitchBT-Bold        BT  0533    legot12b
bmhr8a    MachineITCbyBT-Regular            BT  0534    machinen
bhvc8a2   Swiss721BT-BlackNo2               BT  0540    swz721k2
bazb8a    Gothic725BT-BoldA                 BT  0541    goth725b
ba4r8a    AlternateGothicNo2BT-Regular      BT  0542    altgot2n
blgbi8a   LetterGothic12PitchBT-BoldItal    BT  0543    lego12bi
blbjo8a   LubalinGraphITCbyBT-XtraLightOb   BT  0544    lubgrxlo
blbko8a   LubalinGraphITCbyBT-BookOblique   BT  0545    lubgrhno
blbmo8a   LubalinGraphITCbyBT-MediumObl     BT  0546    lubgrhmo
blbdo8a   LubalinGraphITCbyBT-DemiOblique   BT  0547    lubgrhdo
blbbo8a   LubalinGraphITCbyBT-BoldOblique   BT  0548    lubgrhbo

btmx8a    Dutch801BT-ExtraBold              BT  0549    dut801xb
b_____SymbolITCbyBT-Bold                BT  0554    symbolb
b_____SymbolITCbyBT-BoldItalic          BT  0555    symbolbi
b_____Century731BT-RomanA               BT  0558    cen731n
b_____Century731BT-ItalicA              BT  0559    cen731i
b_____Century731BT-BoldA                BT  0560    cen731b
b_____Century731BT-BoldItalicA          BT  0561    cen731bi
bhir8a    Swiss921BT-RegularA               BT  0562    swz921n
bhvcl8a   Swiss721BT-BlackOutline           BT  0563    swz721ko
bcsb8ac   CenturySchoolbookBT-BoldCond      BT  0574    censbkbc
bcbl8a    CooperBT-Light                    BT  0575    cooperl
bcbli8a   CooperBT-LightItalic              BT  0576    cooperli
bcbm8a    CooperBT-Medium                   BT  0577    cooperm
bcbmi8a   CooperBT-MediumItalic             BT  0578    coopermi
bcbb8a    CooperBT-Bold                     BT  0579    cooperb
bcbbi8a   CooperBT-BoldItalic               BT  0580    cooperbi
bcarl8a   CaslonOpenfaceBT-Regular          BT  0581    casopfan
bcrri8a   Courier10PitchBT-Italic           BT  0582    cur10pti
bcrb8a    Courier10PitchBT-Bold             BT  0583    cur10ptb
benrs8a   EngraversGothicBT-Regular         BT  0586    engrgotn
benr8a    EngraversRomanBT-Regular          BT  0587    engrvrn
be0r8a    EmbassyBT-Regular                 BT  0588    embassyn
bkgl8a    BankGothicBT-Light                BT  0589    bnkgothl
bkgm8a    BankGothicBT-Medium               BT  0590    bnkgothm
bf0r8a    FlemishScriptBT-Regular           BT  0591    flemscrn
bp5r8a    PiranesiItalicBT-Regular          BT  0592    piranesi
bser8a    ShelleyAndanteBT-Regular          BT  0593    shlyandn
bserd8a   ShelleyAllegroBT-Regular          BT  0594    shlyalln
bserw8a   ShelleyVolanteBT-Regular          BT  0595    shlyvoln
bhvrt8a   Monospace821BT-Roman              BT  0596    mono821n
bhvrit8a  Monospace821BT-Italic             BT  0597    mono821i
bhvbt8a   Monospace821BT-Bold               BT  0598    mono821b
bhvbit8a  Monospace821BT-BoldItalic         BT  0599    mon821bi
bcsrt8a   CenturySchoolbookBT-Monospace     BT  0600    censbkmn
bdcr8a    DomCasualBT-Regular               BT  0604    domcasn
bpfr8a    Decorated035BT-Regular            BT  0605    deco035n
bt1r8a    ThunderbirdBT-Regular             BT  0606    thundrbn
bb7r8a    BernhardModernBT-Roman            BT  0607    bnhrdmon
boir8a    FormalScript421BT-Regular         BT  0609    frms421n
bmsr8a    Staccato222BT-Regular             BT  0610    stac222n
bcrbi8a   Courier10PitchBT-BoldItalic       BT  0611    cur10pbi
bperi8a   Prestige12PitchBT-Italic          BT  0613    prest12i
bpeb8a    Prestige12PitchBT-Bold            BT  0614    prest12b
bpebi8a   Prestige12PitchBT-BoldItalic      BT  0615    pres12bi
bctr8a    CheltenhamBT-Roman                BT  0620    chlthmn
bctri8a   CheltenhamBT-Italic               BT  0621    chlthmi
bctb8a    CheltenhamBT-Bold                 BT  0622    chlthmb

bctbi8a   CheltenhamBT-BoldItalic           BT  0623    chlthmbi
brmli8a   CarminaBT-LightItalic             BT  0624    carminli
bg0r8a    GandoBT-Regular                   BT  0625    gandon
bpgl8a    Exotic350BT-Light                 BT  0626    exot350l
bpgd8a    Exotic350BT-DemiBold              BT  0627    exo350db
bpgb8a    Exotic350BT-Bold                  BT  0628    exot350b
bcbc8a    CooperBT-Black                    BT  0630    coopbln
bcbci8a   CooperBT-BlackItalic              BT  0631    coopbli
bchb8a    CharterBT-Bold                    BT  0632    chartrb
bchbi8a   CharterBT-BoldItalic              BT  0633    chartrbi
buhh8a    CushingITCbyBT-Heavy              BT  0641    cushngh
buhhi8a   CushingITCbyBT-HeavyItalic        BT  0642    cushnghi
beor8a    AmerigoBT-RomanA                  BT  0643    amrigon
beori8a   AmerigoBT-ItalicA                 BT  0644    amrigoi
beob8a    AmerigoBT-BoldA                   BT  0645    amrigob
bobr      OCRB10PitchBT-Regular             BT  0646    ocrb10n
bchr8a    CharterBT-Roman                   BT  0648    chartrn
bchri8a   CharterBT-Italic                  BT  0649    chartri
bzgm8a    Geometric706BT-MediumB            BT  0660    geo706m
bzgc8a    Geometric706BT-BlackB             BT  0661    geo706k
bzgb8ac   Geometric706BT-BoldCondensedB     BT  0662    geo706bc
bzgc8ac   Geometric706BT-BlackCondensedB    BT  0663    geo706kc
bg6r8a    GorillaITCbyBT-Regular            BT  0664    gorillan
brml8a    CarminaBT-Light                   BT  0667    carminal
brmc8a    CarminaBT-Black                   BT  0668    carminak
bfgk8a    FranklinGothicITCbyBT-Book        BT  0670    frkgotn
bfgki8a   FranklinGothicITCbyBT-BookItal    BT  0671    frkgotni
bfgd8a    FranklinGothicITCbyBT-Demi        BT  0674    frkgotd
bfgdi8a   FranklinGothicITCbyBT-DemiItal    BT  0675    frkgotdi
bfgh8a    FranklinGothicITCbyBT-Heavy       BT  0676    frkgoth
bfghi8a   FranklinGothicITCbyBT-HeavyItal   BT  0677    frkgothi
b_____GrouchITCbyBT-Regular             BT  0678    grouchn
bymli8a   StymieBT-LightItalic              BT  0679    stymieli
bfnr8a    FeniceITCbyBT-Regular             BT  0682    fenicen
bfnri8a   FeniceITCbyBT-RegularItalic       BT  0683    feniceni
bfnb8a    FeniceITCbyBT-Bold                BT  0684    feniceb
bfnbi8a   FeniceITCbyBT-BoldItalic          BT  0685    fenicebi
b_____GrizzlyITCbyBT-Regular            BT  0688    grizzlyn
bymmi8a   StymieBT-MediumItalic             BT  0689    stymiemi
bl2r8a    LibraBT-Regular                   BT  0690    libran
b_____Pica10PitchBT-Roman               BT  0691    pica10n
bymbi8a   StymieBT-BoldItalic               BT  0692    stymiebi
b_____NewtextITCbyBT-Regular            BT  0697    newtxtn
b_____NewtextITCbyBT-RegularItalic      BT  0698    newtxtni
bchc8a    CharterBT-Black                   BT  0709    chartrk
bchci8a   CharterBT-BlackItalic             BT  0710    chartrki
bgjr8a    ElegantGaramondBT-Roman           BT  0711    eleggarn

bgjri8a   ElegantGaramondBT-Italic          BT  0712    eleggari
bgjb8a    ElegantGaramondBT-Bold            BT  0713    eleggarb
bekr8a    Freeform710BT-Regular             BT  0714    frfm710n
bppr8a    Lapidary333BT-Roman               BT  0715    lap333n
bppri8a   Lapidary333BT-Italic              BT  0716    lap333i
bppb8a    Lapidary333BT-Bold                BT  0717    lap333b
bppbi8a   Lapidary333BT-BoldItalic          BT  0718    lap333bi
brmci8a   CarminaBT-BlackItalic             BT  0719    carminki
bpbr8a    PTBarnumBT-Regular                BT  0720    ptbarnmn
brmm8a    CarminaBT-Medium                  BT  0721    carminam
brmmi8a   CarminaBT-MediumItalic            BT  0722    carmini
brmb8a    CarminaBT-Bold                    BT  0723    carminab
brmbi8a   CarminaBT-BoldItalic              BT  0724    carminbi
bi2r8a    ImpulsBT-Regular                  BT  0725    impulsn
bkur8a    KaufmannBT-Regular                BT  0726    kaufmann
bkub8a    KaufmannBT-Bold                   BT  0727    kaufmanb
bbob8a    BalloonBT-Bold                    BT  0728    balonb
bbox8a    BalloonBT-ExtraBold               BT  0729    balonxb
bnvk8a    NovareseITCbyBT-Book              BT  0730    novaren
bnvki8a   NovareseITCbyBT-BookItalic        BT  0731    novareni
bnvb8a    NovareseITCbyBT-Bold              BT  0734    novareb
bnvbi8a   NovareseITCbyBT-BoldItalic        BT  0735    novarebi
bc2k8a    Caslon224ITCbyBT-Book             BT  0745    cas224n
bc2ki8a   Caslon224ITCbyBT-BookItalic       BT  0746    cas224ni
bc2b8a    Caslon224ITCbyBT-Bold             BT  0749    cas224b
bc2bi8a   Caslon224ITCbyBT-BoldItalic       BT  0750    cas224bi
bnabi8a   Dutch823BT-BoldItalicB            BT  0753    dut823bi
borr8a10  OratorBT-TenPitch                 BT  0755    orator10
bvrr8a    VAGRoundedBT-Regular              BT  0756    vagron
bclr8a    CloisterBlackBT-Regular           BT  0757    cloistrk
borr8a15  OratorBT-FifteenPitch             BT  0758    orator15
bhvm8a    Swiss721BT-Medium                 BT  0759    swz721m
bhvmi8a   Swiss721BT-MediumItalic           BT  0760    swz721mi
btvr8a    Kuenstler480BT-Roman              BT  0761    kuen408n
btvri8a   Kuenstler480BT-Italic             BT  0762    kuen408i
btvb8a    Kuenstler480BT-Bold               BT  0763    kuen408b
btvbi8a   Kuenstler480BT-BoldItalic         BT  0764    kue408bi
bhvbl8ac  Swiss721BT-BoldCondensedOutline   BT  0767    sw721bco
bh0r8a    HandelGothicBT-Regular            BT  0768    handgotn
bdcb8a    DomBoldBT-Regular                 BT  0769    domb
beom8a    AmerigoBT-MediumA                 BT  0770    amrigom
beomi8a   AmerigoBT-MediumItalicA           BT  0771    amrigomi
beobi8a   AmerigoBT-BoldItalicA             BT  0772    amrigobi
bc7l8a    CompactaBT-Light                  BT  0783    compctal
bc7r8a    CompactaBT-Roman                  BT  0784    compctan
bc7ri8a   CompactaBT-Italic                 BT  0785    compctai
bc7b8a    CompactaBT-Bold                   BT  0786    compctab

bc7bi8a   CompactaBT-BoldItalic             BT  0787    compctbi
bc7c8a    CompactaBT-Black                  BT  0788    compctak
bbpr8a    Modern735BT-RomanA                BT  0793    mod735n
bbwre8a   BroadwayEngravedBT-Regular        BT  0802    brdwayg
bonbi8a   News705BT-BoldItalicB             BT  0803    new705bi
bpnl8a    Aldine721BT-Light                 BT  0804    ald721l
bpnli8a   Aldine721BT-LightItalic           BT  0805    ald721li
bb4r8a    Brush738BT-RegularA               BT  0806    brus738n
bb6r8ac   Gothic821CondensedBT-Regular      BT  0807    goth821c
bcir8a    CandidaBT-Roman                   BT  0808    candidan
bciri8a   CandidaBT-Italic                  BT  0809    candidai
bcib8a    CandidaBT-Bold                    BT  0810    candidab
bdir8a    Calligraphic810BT-Roman           BT  0811    calg810n
bdiri8a   Calligraphic810BT-Italic          BT  0812    calg810i
bhrb8a    Swiss721BT-BoldRounded            BT  0813    swz721br
bhrc8a    Swiss721BT-BlackRounded           BT  0815    swz721kr
bs3l8a    SeagullBT-Light                   BT  0817    seagulll
bs3m8a    SeagullBT-Medium                  BT  0818    seagullm
bs3b8a    SeagullBT-Bold                    BT  0819    seagullb
bs3h8a    SeagullBT-Heavy                   BT  0820    seagullh
b_____MixageITCbyBT-Bold                BT  0825    mixageb
b_____MixageITCbyBT-BoldItalic          BT  0826    mixagebi
bm3r8a    MiraraeBT-Roman                   BT  0830    miraraen
bm3b8a    MiraraeBT-Bold                    BT  0831    miraraeb
btms8a    Dutch801BT-SemiBold               BT  0834    dut801db
btmsi8a   Dutch801BT-SemiBoldItalic         BT  0835    du801dbi
bsfb8ac   SerifaBT-BoldCondensed            BT  0836    serifabc
bb7ri8a   BernhardModernBT-Italic           BT  0837    bnhrdmoi
bb7b8a    BernhardModernBT-Bold             BT  0838    bnhrdmob
bb7bi8a   BernhardModernBT-BoldItalic       BT  0839    bnhrdmbi
b_____English157BT-Regular              BT  0840    engl157n
blyr8a    LydianBT-Roman                    BT  0841    lydiann
blyri8a   LydianBT-Italic                   BT  0842    lydiani
blyb8a    LydianBT-Bold                     BT  0843    lydianb
blybi8a   LydianBT-BoldItalic               BT  0844    lydianbi
blyrw8a   LydianCursiveBT-Regular           BT  0845    lydianv
bcdr8ac   ClarendonBT-RomanCondensed        BT  0846    clrndnc
bl3r8a    LuciaBT-Regular                   BT  0847    lucian
bfol8a    FolioBT-Light                     BT  0848    foliol
bfoli8a   FolioBT-LightItalic               BT  0849    folioli
bfok8a    FolioBT-Book                      BT  0850    folion
bfom8a    FolioBT-Medium                    BT  0851    foliom
bfob8a    FolioBT-Bold                      BT  0852    foliob
bfox8a    FolioBT-ExtraBold                 BT  0853    folioxb
bfob8ac   FolioBT-BoldCondensed             BT  0854    foliobc
bfrl8a    Humanist777BT-LightB              BT  0855    hum777l
bfrli8a   Humanist777BT-LightItalicB        BT  0856    hum777li

bfrr8a    Humanist777BT-RomanB              BT  0857    hum777n
bfrri8a   Humanist777BT-ItalicB             BT  0858    hum777i
bfrb8a    Humanist777BT-BoldB               BT  0859    hum777b
bfrbi8a   Humanist777BT-BoldItalicB         BT  0860    hum777bi
bfrc8a    Humanist777BT-BlackB              BT  0861    hum777k
bfrci8a   Humanist777BT-BlackItalicB        BT  0862    hum777ki
begr8a    OriginalGaramondBT-Roman          BT  0863    orggaran
begri8a   OriginalGaramondBT-Italic         BT  0864    orggarai
begb8a    OriginalGaramondBT-Bold           BT  0865    orggarab
begbi8a   OriginalGaramondBT-BoldItalic     BT  0866    orggarbi
bpyr8a    PlaybillBT-Regular                BT  0867    playbiln
b_____Century751BT-RomanB               BT  0868    cent751n
b_____Century751BT-ItalicB              BT  0869    cent751i
bngr8aq   NewsGothicBT-ExtraCondensed       BT  0870    newgotxc
bngb8aq   NewsGothicBT-BoldExtraCond        BT  0871    newgtbxc
bsxr8a    Humanist531BT-RomanA              BT  0872    hum531n
bsxb8a    Humanist531BT-BoldA               BT  0874    hum531b
bsxc8a    Humanist531BT-BlackA              BT  0875    hum531k
bsxu8a    Humanist531BT-UltraBlackA         BT  0876    hum531uk
b_____Egyptian710BT-RegularA            BT  0883    egyt710n
bffm8a    Transitional551BT-MediumB         BT  0884    tran551m
bffmi8a   Transitional551BT-MediumItalicB   BT  0885    tra551mi
bm5b8a    MurrayHillBT-Bold                 BT  0886    muryhilb
bcxl8a    CaxtonBT-Light                    BT  0887    caxtonl
bcxli8a   CaxtonBT-LightItalic              BT  0888    caxtonli
bcxk8a    CaxtonBT-Book                     BT  0889    caxtonn
bcxki8a   CaxtonBT-BookItalic               BT  0890    caxtonni
bcxb8a    CaxtonBT-Bold                     BT  0891    caxtonb
bcxbi8a   CaxtonBT-BoldItalic               BT  0892    caxtonbi
brvr8a    RevueBT-Regular                   BT  0895    revuen
br0r8a    RomanaBT-Roman                    BT  0896    romanan
br0b8a    RomanaBT-Bold                     BT  0897    romanab
bssl8a    SchneidlerBT-Light                BT  0899    schndll
bssli8a   SchneidlerBT-LightItalic          BT  0900    schndlli
bssr8a    SchneidlerBT-Roman                BT  0901    schndln
bssri8a   SchneidlerBT-Italic               BT  0902    schndli
bssm8a    SchneidlerBT-Medium               BT  0903    schndlm
bssmi8a   SchneidlerBT-MediumItalic         BT  0904    schndlmi
bssb8a    SchneidlerBT-Bold                 BT  0905    schndlb
bssbi8a   SchneidlerBT-BoldItalic           BT  0906    schndlbi
bssc8a    SchneidlerBT-Black                BT  0907    schndlk
bssci8a   SchneidlerBT-BlackItalic          BT  0908    schndlki
bwnl8a    WindsorBT-Light                   BT  0909    wndsrl
bymx8ac   StymieBT-ExtraBoldCondensed       BT  0910    stymixbc
b_____DeVinneBT-Roman                   BT  0911    devinnen
b_____DeVinneBT-Italic                  BT  0912    devinnei
bear8a    Transitional521BT-RomanA          BT  0913    tran521n

bearw8a   Transitional521BT-CursiveA        BT  0914    tran521i
beab8a    Transitional521BT-BoldA           BT  0915    tran521b
ball8a    Flareserif821BT-Light             BT  0916    flar821l
balr8a    Flareserif821BT-Roman             BT  0917    flar821n
balb8a    Flareserif821BT-Bold              BT  0918    flar821b
bdcri8a   DomDiagonalBT-Regular             BT  0920    domdiagn
bdcbi8a   DomDiagonalBT-Bold                BT  0921    domdiagb
bysl8a    GoudySansITCbyBT-Light            BT  0922    goudsanl
bysli8a   GoudySansITCbyBT-LightItalic      BT  0923    goudsali
bysm8a    GoudySansITCbyBT-Medium           BT  0924    goudsanm
bysmi8a   GoudySansITCbyBT-MediumItalic     BT  0925    goudsami
bysb8a    GoudySansITCbyBT-Bold             BT  0926    goudsanb
bysbi8a   GoudySansITCbyBT-BoldItalic       BT  0927    goudsabi
bysc8a    GoudySansITCbyBT-Black            BT  0928    goudsank
bysci8a   GoudySansITCbyBT-BlackItalic      BT  0929    goudsaki
bnbr8a    NewBaskervilleITCbyBT-Roman       BT  0930    newbaskn
bnbri8a   NewBaskervilleITCbyBT-Italic      BT  0931    newbaski
bnbb8a    NewBaskervilleITCbyBT-Bold        BT  0934    newbaskb
bnbbi8a   NewBaskervilleITCbyBT-BoldItal    BT  0935    newbasbi
bclrl8a   CloisterOpenFaceBT-Regular        BT  0938    cloisofn
bt2r8a    TypoUprightBT-Regular             BT  0939    typouprn
bs2r8a    Script12PitchBT-Roman             BT  0940    scrpt12n
blwk8a    LeawoodITCbyBT-Book               BT  0941    leawodn
blwki8a   LeawoodITCbyBT-BookItalic         BT  0942    leawodni
blwb8a    LeawoodITCbyBT-Bold               BT  0945    leawodb
blwbi8a   LeawoodITCbyBT-BoldItalic         BT  0946    leawodbi
bngd8a    NewsGothicBT-Demi                 BT  0949    newsgotd
bngdi8a   NewsGothicBT-DemiItalic           BT  0950    newsgtdi
b_____CataneoBT-Light                   BT  0951    catanltn
b_____CataneoBT-Regular                 BT  0952    cataneon
b_____CataneoBT-Bold                    BT  0953    catanbon
blfr8a    LifeBT-Roman                      BT  0954    lifen
blfri8a   LifeBT-Italic                     BT  0955    lifei
blfb8a    LifeBT-Bold                       BT  0956    lifeb
blfbi8a   LifeBT-BoldItalic                 BT  0957    lifebi
bunl8aq   ZurichBT-LightExtraCondensed      BT  0958    zurchlxc
bunr8aq   ZurichBT-ExtraCondensed           BT  0959    zurchxc
bunb8aq   ZurichBT-BoldExtraCondensed       BT  0960    zurchbxc
beel8a    Egyptian505BT-Light               BT  0961    egyt505l
beer8a    Egyptian505BT-Roman               BT  0962    egyt505n
beem8a    Egyptian505BT-Medium              BT  0963    egyt505m
beeb8a    Egyptian505BT-Bold                BT  0964    egyt505b
b_____BruceOldStyleBT-Roman             BT  0965    bruolstn
b_____BruceOldStyleBT-Italic            BT  0966    bruolsti
bhsr8a    Swiss924BT-RegularB               BT  0969    swz924n
bunx8a    ZurichBT-ExtraBlack               BT  0970    zurchxk
btmxi8a   Dutch801BT-ExtraBoldItalic        BT  0971    du801xbi

bsbbi8a   ClassicalGaramondBT-BoldItalic    BT  0972    clsgarbi
bbdb8ac   BodoniBT-BoldCondensed            BT  0973    bodonibc
benb8a    EngraversRomanBT-Bold             BT  0974    engrvrb
bi0r8a    CommercialScriptBT-Regular        BT  0976    comscrtn
bpnb8ac   Aldine721BT-BoldCondensed         BT  0977    ald721bc
bs6r8a    StuyvesantBT-Regular              BT  0980    stuyvesn
bnur8a    NuptialBT-Regular                 BT  0981    nuptualn
bftr8a    FrakturBT-Regular                 BT  0983    frakturn
bw0r8a    WeddingTextBT-Regular             BT  0985    wedtxtn
bb7rw8a   BernhardTangoBT-Regular           BT  0986    bnhrdtan
bl1r8a    LibertyBT-Regular                 BT  0987    libertyn
bwnrl8a   WindsorBT-Outline                 BT  0988    wndsro
bp3r8a    Brush445BT-Regular                BT  0990    brus445n
bgql8a    Gothic720BT-LightB                BT  0991    goth720l
bgqli8a   Gothic720BT-LightItalicB          BT  0992    got720li
bgqr8a    Gothic720BT-RomanB                BT  0993    goth720n
bgqri8a   Gothic720BT-ItalicB               BT  0994    goth720i
bgqb8a    Gothic720BT-BoldB                 BT  0995    goth720b
bgqbi8a   Gothic720BT-BoldItalicB           BT  0996    got720bi
baur8a    Freeform721BT-Roman               BT  0997    frfm721n
bauri8a   Freeform721BT-Italic              BT  0998    frfm721i
baub8a    Freeform721BT-Bold                BT  0999    frfm721b
baubi8a   Freeform721BT-BoldItalic          BT  1000    frf721bi
bauc8a    Freeform721BT-Black               BT  1001    frfm721k
bauci8a   Freeform721BT-BlackItalic         BT  1002    frf721ki
bb7b8ac   BernhardBoldCondensedBT-Regular   BT  1003    brhrdbc
bazc8a    Gothic725BT-BlackA                BT  1007    goth720k
bn0r8a    NormandeBT-Roman                  BT  1011    normandn
bn0ri8a   NormandeBT-Italic                 BT  1012    normandi
b_____MattAntiqueBT-Roman               BT  1014    matantqn
b_____MattAntiqueBT-Italic              BT  1015    matantqi
b_____MattAntiqueBT-Bold                BT  1016    matantqb
bmar8a    Freehand521BT-RegularC            BT  1018    frhnd521
bmpm8ac   GeometricSlab703BT-MediumCond     BT  1019    geo703mc
bmpb8ac   GeometricSlab703BT-BoldCond       BT  1020    geo703bc
bmpx8ac   GeometricSlab703BT-XtraBoldCond   BT  1021    ge703xbc
bbll8a    BelweBT-Light                     BT  1022    belwel
bblm8a    BelweBT-Medium                    BT  1023    belwem
bblb8a    BelweBT-Bold                      BT  1024    belweb
bblr8ac   BelweBT-RomanCondensed            BT  1025    belwec
btvc8a    Kuenstler480BT-Black              BT  1026    kuen408k
ba8r8ac   AuroraBT-RomanCondensed           BT  1027    aurorac
ba8b8ac   AuroraBT-BoldCondensed            BT  1028    aurorabc
b_____CataneoBT-RegularSwash            BT  1030    catanswn
b_____CataneoBT-BoldSwash               BT  1031    catanswb
bcbcl8a   CooperBT-BlackOutline             BT  1034    coopblo
btmrd8a   Dutch801BT-RomanHeadline          BT  1035    dut801na

btmrid8a  Dutch801BT-ItalicHeadline         BT  1036    dut801ia
b_____Modern880BT-Roman                 BT  1037    mod880n
b_____Modern880BT-Italic                BT  1038    mod880i
b_____Modern880BT-Bold                  BT  1039    mod880b
bmzr8a    AmazoneBT-Regular                 BT  1040    amazonen
b_____DeVinneBT-Text                    BT  1041    devinnet
b_____DeVinneBT-ItalicText              BT  1042    devinnit
bb5r8a    Freehand591BT-RegularA            BT  1043    frhnd591
b_____Blackletter686BT-Regular          BT  1044    blklt686
benbf8a   EngraversOldEnglishBT-Bold        BT  1045    engroleb
bjfr8a    Freehand575BT-RegularB            BT  1046    frhnd575
btor8a    Industrial736BT-Roman             BT  1047    ind736n
btori8a   Industrial736BT-Italic            BT  1048    ind736i
bgyh8a    GoudyHeavyfaceBT-Regular          BT  1049    goudhfcn
bgyh8ac   GoudyHeavyfaceBT-RegularCond      BT  1050    goudhfcc
bb7rn8a   BernhardFashionBT-Regular         BT  1051    bnhrdfan
bppc8a    Lapidary333BT-Black               BT  1052    lap333k
bgyrf8a   GoudyHandtooledBTsRegular         BT  1053    goudhndn
bg5r8a    GoudyCatalogueBT-Regular          BT  1054    goudcatn
b_____Century725BT-Roman                BT  1055    cen725n
b_____Century725BT-Bold                 BT  1056    cen725b
b_____Century725BT-RomanCondensed       BT  1057    cen725c
b_____Century725BT-BoldCondensed        BT  1058    cen725bc
b_____Century725BT-Italic               BT  1059    cen725i
b_____Century725BT-Black                BT  1060    cen725k
bocr8a    Dutch809BT-RomanC                 BT  1061    dut809n
bocri8a   Dutch809BT-ItalicC                BT  1062    dut809i
bocb8a    Dutch809BT-BoldC                  BT  1063    dut809b
bpsr8a    ParisianBT-Regular                BT  1064    parisnn
bctb8ac   CheltenhamBT-BoldCondensed        BT  1065    chltboc
bctbi8ac  CheltenhamBT-BoldCondItalic       BT  1066    chltboci
bctb8aq   CheltenhamBT-BoldExtraCondensed   BT  1067    chltboxc
bbrr8a    Revival565BT-Roman                BT  1068    revi565n
bbrri8a   Revival565BT-Italic               BT  1069    revi565i
bbrb8a    Revival565BT-Bold                 BT  1070    revi565b
bbrbi8a   Revival565BT-BoldItalic           BT  1071    rev565bi
bubrd8a   BauerBodoniBT-Titling             BT  1072    baubodti
bumr8a    UmbraBT-Regular                   BT  1074    umbran
bbmr8a    BulmerBT-Roman                    BT  1075    bulmern
bbmri8a   BulmerBT-Italic                   BT  1076    bulmeri
bral8a    RaleighBT-Light                   BT  1077    raleighl
brar8a    RaleighBT-Roman                   BT  1078    raleighn
bram8a    RaleighBT-Medium                  BT  1079    raleighm
brad8a    RaleighBT-DemiBold                BT  1080    raleigdb
brab8a    RaleighBT-Bold                    BT  1081    raleighb
brax8a    RaleighBT-ExtraBold               BT  1082    raleigxb
boyr8a    Dutch811BT-RomanD                 BT  1083    dut811n

boyri8a   Dutch811BT-ItalicD                BT  1084    dut811i
boyb8a    Dutch811BT-BoldD                  BT  1085    dut811b
boybi8a   Dutch811BT-BoldItalicD            BT  1086    dut811bi
bwdk8a    WeidemannITCbyBT-Book             BT  1087    weidmnn
bwdki8a   WeidemannITCbyBT-BookItalic       BT  1088    weidmnni
bwdb8a    WeidemannITCbyBT-Bold             BT  1091    weidmnb
bwdbi8a   WeidemannITCbyBT-BoldItalic       BT  1092    weidmnbi
burr8a    Venetian301BT-Roman               BT  1095    ven301n
burri8a   Venetian301BT-Italic              BT  1096    ven301i
burd8a    Venetian301BT-Demi                BT  1097    ven301d
burdi8a   Venetian301BT-DemiItalic          BT  1098    ven301di
burb8a    Venetian301BT-Bold                BT  1099    ven301b
burbi8a   Venetian301BT-BoldItalic          BT  1100    ven301bi
b_____FrysBaskervilleBT-Roman           BT  1101    frybaskn
baar8a    AachenBT-Roman                    BT  1102    aachenn
baab8a    AachenBT-Bold                     BT  1103    aachenb
bhwl8a    SchadowBT-Light                   BT  1104    schadwl
bhwli8a   SchadowBT-LightCursive            BT  1105    schadwli
bhwr8a    SchadowBT-Roman                   BT  1106    schadwn
bhwb8a    SchadowBT-Bold                    BT  1107    schadwb
bhwc8a    SchadowBT-Black                   BT  1108    schadwk
bder8a    DellaRobbiaBT-Roman               BT  1112    delarobn
bdeb8a    DellaRobbiaBT-Bold                BT  1113    delarobb
bhwc8ac   SchadowBT-BlackCondensed          BT  1114    schadwkc
bnlr8a    Informal011BT-Roman               BT  1115    infr011n
bnlc8a    Informal011BT-Black               BT  1116    infr011k
bkbl8a    Geometric231BT-LightC             BT  1126    geo231l
bkbr8a    Geometric231BT-RomanC             BT  1127    geo231n
bkbb8a    Geometric231BT-BoldC              BT  1128    geo231b
bkbh8a    Geometric231BT-HeavyC             BT  1129    geo231h
ba5r8a    AllegroBT-Regular                 BT  1138    allegron
bc9r8a    Calligraphic421BT-RomanB          BT  1139    calg421n
b_____Geometric885BT-RegularD           BT  1140    geo885n
b_____Formal436BT-Regular               BT  1141    frml436n
bl4r8a    LucianBT-Roman                    BT  1142    luciann
bl4b8a    LucianBT-Bold                     BT  1143    lucianb
brsr8a    ArrusBT-Roman                     BT  1145    arrusn
brsri8a   ArrusBT-Italic                    BT  1146    arrusi
brsb8a    ArrusBT-Bold                      BT  1147    arrusb
brsbi8a   ArrusBT-BoldItalic                BT  1148    arrusbi
brsc8a    ArrusBT-Black                     BT  1149    arrusk
brsci8a   ArrusBT-BlackItalic               BT  1150    arruski
bt0r8a    TangoBT-Regular                   BT  1151    tangon
bbsr8a    BakerSignetBT-Roman               BT  1152    bakrsign
bczr8a    Staccato555BT-RegularA            BT  1153    stac555n
ba3r8a    AdLibBT-Regular                   BT  1154    adlibn
bmor8a    Modern20BT-RomanB                 BT  1168    modrn20n

bmori8a   Modern20BT-ItalicB                BT  1169    modrn20i
be1r8a    EmpireBT-Regular                  BT  1176    empiren
bhxr8a    HuxleyVerticalBT-Regular          BT  1177    huxvertn
bozr8a    OzHandicraftBT-Roman              BT  1178    ozhandin
biyr8a    IowanOldStyleBT-Roman             BT  1198    iwolstyn
biyri8a   IowanOldStyleBT-Italic            BT  1199    iwolstyi
biyb8a    IowanOldStyleBT-Bold              BT  1200    iwolstyb
biybi8a   IowanOldStyleBT-BoldItalic        BT  1201    iwolstbi
biyc8a    IowanOldStyleBT-Black             BT  1202    iwolstyk
biyci8a   IowanOldStyleBT-BlackItalic       BT  1203    iwolstki
bo1r8a    OrandaBT-Roman                    BT  1204    orandan
bo1ri8a   OrandaBT-Italic                   BT  1205    orandai
bo1b8a    OrandaBT-Bold                     BT  1206    orandab
bo1bi8a   OrandaBT-BoldItalic               BT  1207    orandabi
bo1r8ac   OrandaBT-RomanCondensed           BT  1208    orandac
bo1b8ac   OrandaBT-BoldCondensed            BT  1209    orandabc
bm4r8a    MisterEarlBT-Regular              BT  1221    mrearln
bb9b8a    BremenBT-Bold                     BT  1231    bremenb
bb9c8a    BremenBT-Black                    BT  1232    bremenk
bwcr8a    SnowCapBT-Regular                 BT  1237    ________
b_____MilleniumBT-BoldExtended          BT  1242    milnumbe
bpi007    SprocketDeluxeBT-Regular          BT  1243    ________
bpi008    SprocketBT-Regular                BT  1244    ________
bpi009    SonicCutThruBT-Heavy              BT  1247    ________
b_____CharterITCbyBT-Roman              BT  1268    ichrtrn
b_____CharterITCbyBT-Italic             BT  1269    ichrtri
b_____CharterITCbyBT-Bold               BT  1270    ichrtrb
b_____CharterITCbyBT-BoldItalic         BT  1271    ichrtrbi
b_____CharterITCbyBT-Black              BT  1272    ichrtrk
b_____CharterITCbyBT-BlackItalic        BT  1273    ichrtrki
bo0r8a    OldDreadfulNo7BT-Regular          BT  7009    ________
bzdr      ZapfDingbatsITCbyBT-Regular       BT  8556    zafdingn
b_____MICR10byBT-Regular                BT  8729    micr010n
b_____MICR12byBT-Regular                BT  8730    micr012n
b_____MICR13byBT-Regular                BT  8731    micr013n
bpi003    UniversalMath1BT-Regular          BT  9014    ________
bpi004    CommercialPiBT-Regular            BT  9025    ________
bsyr      SymbolProportionalBT-Regular      BT  9830    symbpron
bpi010    HolidayPiBT-Regular               BT  9832    ________
bpi001    NewspaperPiBT-Regular             BT  9836    ________

A.5 DTC fonts

The ‘dtc.map’ file has abbreviations for DTC (Digital Typeface Corporation) fonts.

dhvc8rc  SansBlackCondensed-DTC
dhvco8rc SansBlackCondensed-Oblique-DTC
dhvm8rx  SansExtended-Medium-DTC

dhvr8rx  SansExtended-Regular-DTC
dhvro8rc SansExtended-RegularOblique-DTC
dhvx8rc  SansExtraBlackCondensed-DTC
dhvl8rx  SansLightExtended-DTC
dhvh8rc  SansThinCondensed-DTC
dhvho8rc SansThinCondensed-Oblique-DTC
dnbb8r   NewBaskerville-Bold-DTC
dnbbi8r  NewBaskerville-BoldItalic-DTC
dnbri8r  NewBaskerville-RomanItalic-DTC
dnbr8r   NewBaskerville-Roman-DTC
dopr8r   Optimum-Roman-DTC
dopb8r   Optimum-Bold-DTC
dopbo8r  Optimum-BoldItalic-DTC
doro8r   Optimum-RomanItalic-DTC
dpll8r   Palton-Light-DTC

A.6 ITC fonts

The ‘itc.map’ file has abbreviations for ITC (International Typeface Corporation) fonts.

zieuk     EuroSerifITC-Book
zieuki    EuroSerifITC-BookItalic
zieum     EuroSerifITC-Medium
zieumi    EuroSerifITC-MediumItalic
zieub     EuroSerifITC-Bold
zieubi    EuroSerifITC-BoldItalic
zieuc     EuroSerifITC-Black
zieuci    EuroSerifITC-BlackItalic
zieuks    EuroSansITC-Book
zieukis   EuroSansITC-BookItalic
zieums    EuroSansITC-Medium
zieumis   EuroSansITC-MediumItalic
zieubs    EuroSansITC-Bold
zieubis   EuroSansITC-BoldItalic
zieucs    EuroSansITC-Black
zieucis   EuroSansITC-BlackItalic
i_____Aachen-Medium                     ITC 2000
i_____Aachen-Bold                       ITC 2001
i_____Academy-Engraved                  ITC 2002
i_____Aftershock                        ITC 2003
i_____Agincourt                         ITC 2004
i_____Airstream                         ITC 2005
i_____Algerian-Condensed                ITC 2006
i_____Ambrose                           ITC 2007
i_____Angryhog                          ITC 2008
i_____Anna                              ITC 2009

i_____Aquinas                           ITC 2010
i_____Aquitaine-Initials                ITC 2011
i_____Aristocrat                        ITC 2012
i_____Arnova                            ITC 2013
i_____Arriba                            ITC 2014
i_____Arriba-Arriba                     ITC 2015
i_____Artiste                           ITC 2016
i_____Atmosphere                        ITC 2017
i_____AugusteaOpen                      ITC 2018
i_____Avenida                           ITC 2019
i_____BaileyQuad-Bold                   ITC 2020
i_____Balmoral                          ITC 2021
i_____Banco-Light                       ITC 2022
i_____Banco                             ITC 2023
i_____Bang                              ITC 2024
i_____Banner                            ITC 2025
i_____BeckaScript                       ITC 2026
i_____Beesknees                         ITC 2027
i_____Belter                            ITC 2028
i_____Belter-MegaOutline                ITC 2029
i_____Belwe-Mono                        ITC 2030
i_____Belwe-MonoItalic                  ITC 2031
i_____Bendigo                           ITC 2032
i_____Bergell                           ITC 2033
i_____Bertie                            ITC 2034
i_____Bertram                           ITC 2035
i_____BibleScriptFlourishes             ITC 2036
i_____BickleyScript                     ITC 2037
i_____Binary-Light                      ITC 2038
i_____Binary-Bold                       ITC 2039
i_____Bitmax                            ITC 2040
i_____BlackTulip                        ITC 2041
i_____Blackadder                        ITC 2042
i_____Blackmoor                         ITC 2043
i_____Blaze                             ITC 2044
i_____Bluntz                            ITC 2045
i_____Bodoni-Brush                      ITC 2046
i_____BodoniSeventyTwo-Book             ITC 2047
i_____BodoniSeventyTwo-BookItalic       ITC 2048
i_____BodoniSeventyTwo-BookItalicSwash  ITC 2049
i_____BodoniSeventyTwo-Bold             ITC 2050
i_____BodoniSeventyTwo-BoldItalic       ITC 2051
i_____BodoniSeventyTwo-BoldItalicSwash  ITC 2052
i_____Boink                             ITC 2053
i_____Bordeaux-Roman                    ITC 2054
i_____Bordeaux-Italic                   ITC 2055
i_____Bordeaux-RomanBold                ITC 2056

i_____Bordeaux-Display                  ITC 2057
i_____Bordeaux-Script                   ITC 2058
i_____BradleyHand-Regular               ITC 2059
i_____BradleyHand-Bold                  ITC 2060
i_____Braganza-Light                    ITC 2061
i_____Braganza-Regular                  ITC 2062
i_____Brighton-Light                    ITC 2063
i_____Brighton-Medium                   ITC 2064
i_____Brighton-Bold                     ITC 2065
i_____Bronx                             ITC 2066
i_____Buckeroo                          ITC 2067
i_____Burlington                        ITC 2068
i_____BuzzerThree                       ITC 2069
i_____Cabaret                           ITC 2070
i_____CabargaCursiva                    ITC 2071
i_____Campaign                          ITC 2072
i_____CancellarescaScript               ITC 2073
i_____Cancione                          ITC 2074
i_____Caribbean                         ITC 2075
i_____Carlton                           ITC 2076
i_____Carumba                           ITC 2077
i_____Caslon540-ItalicSwashes           ITC 2078
i_____CaxtonRoman-Light                 ITC 2079
i_____CaxtonRoman-LightItalic           ITC 2080
i_____CaxtonRoman-Book                  ITC 2081
i_____CaxtonRoman-Bold                  ITC 2082
i_____CenturyHandtooled-Bold            ITC 2083
i_____CenturyHandtooled-BoldItalic      ITC 2084
i_____Challenge-Bold                    ITC 2085
i_____Challenge-ExtraBold               ITC 2086
i_____Champers                          ITC 2087
i_____CheltenhamHandtooled-Bold         ITC 2088
i_____CheltenhamHandtooled-BoldItalic   ITC 2089
i_____Cherie                            ITC 2090
i_____Chiller                           ITC 2091
i_____Chipper                           ITC 2092
i_____ChocLightbyITC                    ITC 2093
i_____Choc                              ITC 2094
i_____ChromiumOne                       ITC 2095
i_____Citation                          ITC 2096
i_____ClaudeSans-Regular                ITC 2097
i_____ClaudeSans-Italic                 ITC 2098
i_____ClaudeSans-BoldItalic             ITC 2099
i_____Clover                            ITC 2100
i_____CommericalScript                  ITC 2101
i_____Compacta                          ITC 2102
i_____Compacta-Italic                   ITC 2103

i_____Compacta-Bold                     ITC 2104
i_____Coptek                            ITC 2105
i_____Corinthian-Light                  ITC 2106
i_____Corinthian-Medium                 ITC 2107
i_____Corinthian-Bold                   ITC 2108
i_____Corinthian-BoldCondensed          ITC 2109
i_____Crillee-Italic                    ITC 2110
i_____Crillee-BoldItalic                ITC 2111
i_____Crillee-ExtraBoldItalic           ITC 2112
i_____Crillee-ItalicInlineShadow        ITC 2113
i_____Cult                              ITC 2114
i_____Cyberkugel                        ITC 2115
i_____Dancin                            ITC 2116
i_____Data70                            ITC 2117
i_____Demian                            ITC 2118
i_____DemianBold                        ITC 2119
i_____DigitalWoodcuts                   ITC 2120
i_____Digitek                           ITC 2121
i_____Dinitials                         ITC 2122
i_____Dolmen                            ITC 2123
i_____Drycut                            ITC 2124
i_____DynamoShadow                      ITC 2125
i_____Eastwood                          ITC 2126
i_____Edwardian-Medium                  ITC 2127
i_____EdwardianScript-Regular           ITC 2128
i_____EdwardianScript-Bold              ITC 2129
i_____Ellipse-Roman                     ITC 2130
i_____Ellipse-Italic                    ITC 2131
i_____Ellipse-Bold                      ITC 2132
i_____Ellipse-BoldItalic                ITC 2133
i_____Emphasis                          ITC 2134
i_____Enviro                            ITC 2135
i_____Epokha                            ITC 2136
i_____Equinox                           ITC 2137
i_____Etruscan                          ITC 2138
i_____FaithfulFly                       ITC 2139
i_____Farmhaus                          ITC 2140
i_____Fashion-Compressed                ITC 2141
i_____Fashion-Engraved                  ITC 2142
i_____FineHand                          ITC 2143
i_____Flamenco-Inline                   ITC 2144
i_____Flamme                            ITC 2145
i_____Flatiron                          ITC 2146
i_____Flight                            ITC 2147
i_____Fling                             ITC 2148
i_____Florinda                          ITC 2149
i_____Follies                           ITC 2150

i_____Fontoon                           ITC 2151
i_____ForestShaded                      ITC 2152
i_____FrancesUncial                     ITC 2153
i_____Frankfurter-Medium                ITC 2154
i_____Frankfurter                       ITC 2155
i_____Frankfurter-Highlight             ITC 2156
i_____Frankfurter-Inline                ITC 2157
i_____Freddo                            ITC 2158
i_____FreestyleScript                   ITC 2159
i_____FreestyleScript-Bold              ITC 2160
i_____GaramondHandtooled                ITC 2161
i_____GaramondHandtooled-Italic         ITC 2162
i_____Gigi                              ITC 2163
i_____Gill-DisplayCompressed            ITC 2164
i_____Gill-KayoCondensed                ITC 2165
i_____GilliesGothic-ExtraBoldShaded     ITC 2166
i_____Glastonbury                       ITC 2167
i_____GoldenCockerel-InitialsOrnaments  ITC 2168
i_____Grapefruit                        ITC 2169
i_____Gravura                           ITC 2170
i_____Green                             ITC 2171
i_____GreytonScript                     ITC 2172
i_____GrimshawHand                      ITC 2173
i_____Hadfield                          ITC 2174
i_____HandDrawn                         ITC 2175
i_____Harlow                            ITC 2176
i_____HarlowSolid                       ITC 2177
i_____Harvey                            ITC 2178
i_____Hazel                             ITC 2179
i_____Heliotype                         ITC 2180
i_____Helvetica-MediumCondensed         ITC 2181
i_____Helvetica-BoldCondensed           ITC 2182
i_____Highlight                         ITC 2183
i_____Hollyweird                        ITC 2184
i_____Hornpype                          ITC 2185
i_____Ignatius                          ITC 2186
i_____Impakt                            ITC 2187
i_____IndyItalic                        ITC 2188
i_____Informal-Roman                    ITC 2189
i_____Inscription                       ITC 2190
i_____Iris                              ITC 2191
i_____Ironwork                          ITC 2192
i_____Isis                              ITC 2193
i_____Jaft                              ITC 2194
i_____Jambalaya                         ITC 2195
i_____Jazz                              ITC 2196
i_____Jellybaby                         ITC 2197

i_____JohnHandy                         ITC 2198
i_____Jokerman                          ITC 2199
i_____Juanita                           ITC 2200
i_____Juanita-Condensed                 ITC 2201
i_____Juanita-Deco                      ITC 2202
i_____Juanita-Lino                      ITC 2203
i_____Juanita-Xilo                      ITC 2204
i_____Juanita-XiloCondensed             ITC 2205
i_____Juice                             ITC 2206
i_____Kanban                            ITC 2207
i_____Katfish                           ITC 2208
i_____Kendo                             ITC 2209
i_____Kick                              ITC 2210
i_____Klee                              ITC 2211
i_____Klepto                            ITC 2212
i_____Kokoa                             ITC 2213
i_____Kristen                           ITC 2214
i_____Kulukundis                        ITC 2215
i_____LaBamba                           ITC 2216
i_____Lambada                           ITC 2217
i_____Laser                             ITC 2218
i_____LaserChrome                       ITC 2219
i_____Latino-Elongated                  ITC 2220
i_____Laura                             ITC 2221
i_____LCD                               ITC 2222
i_____LeGriffe                          ITC 2223
i_____Lennox-Book                       ITC 2224
i_____Lennox-Medium                     ITC 2225
i_____Lennox-Bold                       ITC 2226
i_____Lexikos                           ITC 2227
i_____Lightnin'                         ITC 2228
i_____LimehouseScript                   ITC 2229
i_____LinoCut                           ITC 2230
i_____Locarno-Light                     ITC 2231
i_____Locarno-Italic                    ITC 2232
i_____RennieMackintosh-Light            ITC 2233
i_____RennieMackintosh-Bold             ITC 2234
i_____Malibu                            ITC 2235
i_____Malstock                          ITC 2236
i_____Marguerita                        ITC 2237
i_____Mastercard                        ITC 2238
i_____Matisse                           ITC 2239
i_____Mekanik-Regular                   ITC 2240
i_____Mekanik-Italic                    ITC 2241
i_____Milano                            ITC 2242
i_____Minska-Light                      ITC 2243
i_____Minska-Medium                     ITC 2244

i_____Minska-Bold                       ITC 2245
i_____Mistral-Light                     ITC 2246
i_____Mistral                           ITC 2247
i_____Mithras                           ITC 2248
i_____MoFunkyFresh                      ITC 2249
i_____MonaLisa-Recut                    ITC 2250
i_____MonaLisa-Solid                    ITC 2251
i_____Montage                           ITC 2252
i_____MotterCorpus-Bold                 ITC 2253
i_____MotterCorpus-CondensedBold        ITC 2254
i_____Musclehead                        ITC 2255
i_____Musica                            ITC 2256
i_____NeoNeo                            ITC 2257
i_____NoovoLight                        ITC 2258
i_____NoovoBold                         ITC 2259
i_____Nora                              ITC 2260
i_____Oberon                            ITC 2261
i_____Odessa                            ITC 2262
i_____OldEnglish                        ITC 2263
i_____OneStrokeScript                   ITC 2264
i_____OneStrokeScript-Bold              ITC 2265
i_____OneStrokeScript-Shaded            ITC 2266
i_____Orange                            ITC 2267
i_____Orbon-Light                       ITC 2268
i_____Orbon-Regular                     ITC 2269
i_____Orbon-Bold                        ITC 2270
i_____Orbon-Black                       ITC 2271
i_____Orlando                           ITC 2272
i_____Outback                           ITC 2273
i_____OutoftheFridge                    ITC 2274
i_____Ozwald                            ITC 2275
i_____Pablo                             ITC 2276
i_____Papyrus                           ITC 2277
i_____Party                             ITC 2278
i_____PendryScript                      ITC 2279
i_____PetersMiro                        ITC 2280
i_____Pink                              ITC 2281
i_____Plaza                             ITC 2282
i_____PleasureBoldShaded                ITC 2283
i_____Pneuma                            ITC 2284
i_____Prague                            ITC 2285
i_____Premier-Lightline                 ITC 2286
i_____Premier-Shaded                    ITC 2287
i_____Princetown                        ITC 2288
i_____Pristina                          ITC 2289
i_____Pritchard                         ITC 2290
i_____Pritchard-LineOut                 ITC 2291

i_____Pump-Regular                      ITC 2292
i_____Pump-DemiBold                     ITC 2293
i_____Quadrus                           ITC 2294
i_____Quixley                           ITC 2295
i_____RageItalic                        ITC 2296
i_____Ragtime                           ITC 2297
i_____Rapier                            ITC 2298
i_____Refracta                          ITC 2299
i_____RegattaCondensed                  ITC 2300
i_____RetailScript                      ITC 2301
i_____Retro-Bold                        ITC 2302
i_____Retro-BoldCondensed               ITC 2303
i_____Riptide                           ITC 2304
i_____Riva                              ITC 2305
i_____Robotik                           ITC 2306
i_____Robotik-Italic                    ITC 2307
i_____Romic-Light                       ITC 2308
i_____Romic-LightItalic                 ITC 2309
i_____Roquette                          ITC 2310
i_____Ru'ach                            ITC 2311
i_____RubberStamp                       ITC 2312
i_____Rundfunk                          ITC 2313
i_____SantaFe                           ITC 2314
i_____Savoye                            ITC 2315
i_____Schizoid                          ITC 2316
i_____Scratch                           ITC 2317
i_____Scriba                            ITC 2318
i_____Scriptease                        ITC 2319
i_____Scriptek                          ITC 2320
i_____Scriptek-Italic                   ITC 2321
i_____Scruff                            ITC 2322
i_____Serengetti                        ITC 2323
i_____Shaman                            ITC 2324
i_____Shatter                           ITC 2325
i_____Sinaloa                           ITC 2326
i_____SkidRow                           ITC 2327
i_____Skylark                           ITC 2328
i_____Slipstream                        ITC 2329
i_____Smack                             ITC 2330
i_____Smudger                           ITC 2331
i_____Snap                              ITC 2332
i_____Spirit                            ITC 2333
i_____Spooky                            ITC 2334
i_____Spotlight                         ITC 2335
i_____Squire                            ITC 2336
i_____Squire-ExtraBold                  ITC 2337
i_____Static                            ITC 2338

i_____Stranger                          ITC 2339
i_____Strobos                           ITC 2340
i_____StudioScript                      ITC 2341
i_____Stylus-Regular                    ITC 2342
i_____Stylus-Bold                       ITC 2343
i_____Superstar                         ITC 2344
i_____Surfboard                         ITC 2345
i_____Synchro                           ITC 2346
i_____Tag                               ITC 2347
i_____Tannhauser                        ITC 2348
i_____Teknik                            ITC 2349
i_____Telegram                          ITC 2350
i_____Temble                            ITC 2351
i_____Tempus-Regular                    ITC 2352
i_____Tempus-Italic                     ITC 2353
i_____TempusSans-Regular                ITC 2354
i_____TempusSans-Italic                 ITC 2355
i_____TigerRag                          ITC 2356
i_____TirantiSolid                      ITC 2357
i_____Trackpad                          ITC 2358
i_____TropicaScript                     ITC 2359
i_____TrueGrit                          ITC 2360
i_____Twang                             ITC 2361
i_____Typados                           ITC 2362
i_____Ulysses                           ITC 2363
i_____UniversityRoman                   ITC 2364
i_____UniversityRoman-Italic            ITC 2365
i_____UniversityRoman-Bold              ITC 2366
i_____VanDijk                           ITC 2367
i_____VanDijkBold                       ITC 2368
i_____Varga                             ITC 2369
i_____Vegas                             ITC 2370
i_____Verkehr                           ITC 2371
i_____Vermont                           ITC 2372
i_____Victorian                         ITC 2373
i_____Victorian-InlineShaded            ITC 2374
i_____ViennaExtended                    ITC 2375
i_____VinerHand                         ITC 2376
i_____Vintage                           ITC 2377
i_____Vinyl-Black                       ITC 2378
i_____Vinyl-Outline                     ITC 2379
i_____Vinyl-SawtoothBlack               ITC 2380
i_____Vinyl-SawtoothOutline             ITC 2381
i_____Vivaldi                           ITC 2382
i_____WadeSans-Light                    ITC 2383
i_____Wanted                            ITC 2384
i_____Waterloo-Bold                     ITC 2385

i_____Westwood                          ITC 2386
i_____WildThing                         ITC 2387
i_____Willow                            ITC 2388
i_____Wisteria                          ITC 2389
i_____Xylo                              ITC 2390
i_____YoungBaroque                      ITC 2391
i_____Zaragoza                          ITC 2392
i_____ZemkeHand                         ITC 2393
i_____Zennor                            ITC 2394
i_____Ziggy                             ITC 2395
i_____Zinjaro                           ITC 2396
i_____MotterSparta                      ITC 2397
i_____Underscript                       ITC 2398
i_____Stenberg                          ITC 2399
i_____Stenberg-Inline                   ITC 2400
i_____Arid                              ITC 2401
i_____Tapioca                           ITC 2402
i_____PiousHenry                        ITC 2403
i_____Korigan-Light                     ITC 2404
i_____Korigan-Bold                      ITC 2405
i_____Abaton                            ITC 2406
i_____Portago                           ITC 2407
i_____Blair-Light                       ITC 2408
i_____Blair-Medium                      ITC 2409
i_____Blair-Bold                        ITC 2410
i_____Scarborough                       ITC 2411
i_____Scarborough-Bold                  ITC 2412
i_____Gema                              ITC 2413
i_____Roswell-Three                     ITC 2414
i_____Roswell-Four                      ITC 2415
i_____Simran                            ITC 2416
i_____Tremor                            ITC 2417
i_____Coconino                          ITC 2418
i_____Stoclet-Light                     ITC 2419
i_____Stoclet-Bold                      ITC 2420
i_____JiggeryPokery                     ITC 2421
i_____Deelirious                        ITC 2422
i_____Flora-Medium                      ITC 2423
i_____Flora-Bold                        ITC 2424
i_____Manhattan                         ITC 2425
i_____Attitudes                         ITC 4000
i_____BackyardBeasties                  ITC 4001
i_____CalligraphicOrnaments             ITC 4002
i_____Celebrations                      ITC 4003
i_____Commercials                       ITC 4004
i_____Connectivities                    ITC 4005
i_____DavesRaves-One                    ITC 4006

i_____DavesRaves-Two                    ITC 4007
i_____DavesRaves-Three                  ITC 4008
i_____Delectables                       ITC 4009
i_____Diversions                        ITC 4010
i_____Diversities                       ITC 4011
i_____Eclectics                         ITC 4012
i_____Energetics                        ITC 4013
i_____Expressions                       ITC 4014
i_____Fontoonies                        ITC 4015
i_____Gargoonies                        ITC 4016
i_____HomeImprovement                   ITC 4017
i_____Incidentals                       ITC 4018
i_____Industrials                       ITC 4019
i_____Inspirations                      ITC 4020
i_____Journeys                          ITC 4021
i_____MoFunkyFresh-Symbols              ITC 4022
i_____Moderns                           ITC 4023
i_____Naturals                          ITC 4024
i_____Organics                          ITC 4025
i_____Organics-Two                      ITC 4026
i_____Primitives                        ITC 4027
i_____Radicals                          ITC 4028
i_____Situations                        ITC 4029
i_____StainedGlass                      ITC 4030
i_____TotSpots                          ITC 4031
i_____Urbans                            ITC 4032
i_____WellBeings                        ITC 4033
i_____Wildlife                          ITC 4034
i_____WildWest                          ITC 4035
i_____Ancestor                          ITC 4036
i_____Shadowettes                       ITC 4037
i_____Beorama                           ITC 4038
i_____Holistics                         ITC 4039
i_____ProfessionalConnections           ITC 4040
i_____Bodoni-Ornaments                  ITC 5000
i_____RennieMackintosh-Ornaments        ITC 5002
i_____TypeEmbellishments-One            ITC 5003
i_____TypeEmbellishments-Two            ITC 5004
i_____TypeEmbellishments-Three          ITC 5005
i_____JapaneseGarden                    ITC 5006
i_____SevenTreasures                    ITC 5007
i_____BodoniSix-Book                    ITC 6000
i_____BodoniSix-BookItalic              ITC 6001
i_____BodoniSix-Bold                    ITC 6002
i_____BodoniSix-BoldItalic              ITC 6003
i_____BodoniTwelve-Book                 ITC 6004
i_____BodoniTwelve-BookItalic           ITC 6005

i_____BodoniTwelve-Bold                 ITC 6006
i_____BodoniTwelve-BoldItalic           ITC 6007
i_____Charlotte-Book                    ITC 6008
i_____Charlotte-BookItalic              ITC 6009
i_____Charlotte-Medium                  ITC 6010
i_____Charlotte-Bold                    ITC 6011
i_____Charter-Regular                   ITC 6012
i_____Charter-RegularItalic             ITC 6013
i_____Charter-Bold                      ITC 6014
i_____Charter-BoldItalic                ITC 6015
i_____Charter-Black                     ITC 6016
i_____Charter-BlackItalic               ITC 6017
i_____Elysium-Book                      ITC 6018
i_____Elysium-BookItalic                ITC 6019
i_____Elysium-Medium                    ITC 6020
i_____Elysium-Bold                      ITC 6021
i_____Figural-Book                      ITC 6022
i_____Figural-BookItalic                ITC 6023
i_____Figural-Medium                    ITC 6024
i_____Figural-Bold                      ITC 6025
i_____FrizQuadrata-Regular              ITC 6026
i_____FrizQuadrata-Italic               ITC 6027
i_____FrizQuadrata-Bold                 ITC 6028
i_____FrizQuadrata-BoldItalic           ITC 6029
i_____Gilgamesh-Book                    ITC 6030
i_____Gilgamesh-BookItalic              ITC 6031
i_____Gilgamesh-Medium                  ITC 6032
i_____Gilgamesh-Bold                    ITC 6033
i_____GoldenCockerel-Roman              ITC 6034
i_____GoldenCockerel-Italic             ITC 6035
i_____GoldenCockerel-Titling            ITC 6036
i_____Humana-Light                      ITC 6037
i_____Humana-LightItalic                ITC 6038
i_____Humana-Medium                     ITC 6039
i_____Humana-MediumItalic               ITC 6040
i_____Humana-Bold                       ITC 6041
i_____Humana-BoldItalic                 ITC 6042
i_____HumanaScript-Light                ITC 6043
i_____HumanaScript-Medium               ITC 6044
i_____HumanaScript-Bold                 ITC 6045
i_____Kallos-Book                       ITC 6046
i_____Kallos-BookItalic                 ITC 6047
i_____Kallos-Medium                     ITC 6048
i_____Kallos-MediumItalic               ITC 6049
i_____Kallos-Bold                       ITC 6050
i_____Kallos-BoldItalic                 ITC 6051
i_____LegacySerif-Book                  ITC 6052

i_____LegacySerif-BookItalic            ITC 6053
i_____LegacySerif-Medium                ITC 6054
i_____LegacySerif-MediumItalic          ITC 6055
i_____LegacySerif-Bold                  ITC 6056
i_____LegacySerif-BoldItalic            ITC 6057
i_____LegacySerif-Ultra                 ITC 6058
i_____MendozaRoman-Book                 ITC 6059
i_____MendozaRoman-BookItalic           ITC 6060
i_____MendozaRoman-Medium               ITC 6061
i_____MendozaRoman-MediumItalic         ITC 6062
i_____MendozaRoman-Bold                 ITC 6063
i_____MendozaRoman-BoldItalic           ITC 6064
i_____Obelisk-Light                     ITC 6065
i_____Obelisk-LightItalic               ITC 6066
i_____Obelisk-Medium                    ITC 6067
i_____Obelisk-Bold                      ITC 6068
i_____OfficinaSerif-Book                ITC 6069
i_____OfficinaSerif-BookItalic          ITC 6070
i_____OfficinaSerif-Bold                ITC 6071
i_____OfficinaSerif-BoldItalic          ITC 6072
i_____StoneInformal-Medium              ITC 6073
i_____StoneInformal-MediumItalic        ITC 6074
i_____StoneInformal-SemiBold            ITC 6075
i_____StoneInformal-SemiBoldItalic      ITC 6076
i_____StoneInformal-Bold                ITC 6077
i_____StoneInformal-BoldItalic          ITC 6078
i_____StoneSerif-Medium                 ITC 6079
i_____StoneSerif-MediumItalic           ITC 6080
i_____StoneSerif-SemiBold               ITC 6081
i_____StoneSerif-SemiBoldItalic         ITC 6082
i_____StoneSerif-Bold                   ITC 6083
i_____StoneSerif-BoldItalic             ITC 6084
i_____Syndor-Book                       ITC 6085
i_____Syndor-BookItalic                 ITC 6086
i_____Syndor-Medium                     ITC 6087
i_____Syndor-MediumItalic               ITC 6088
i_____Syndor-Bold                       ITC 6089
i_____Syndor-BoldItalic                 ITC 6090
i_____Tyfa-Book                         ITC 6091
i_____Tyfa-BookItalic                   ITC 6092
i_____Tyfa-Medium                       ITC 6093
i_____Tyfa-MediumItalic                 ITC 6094
i_____Tyfa-Bold                         ITC 6095
i_____Tyfa-BoldItalic                   ITC 6096
i_____OfficinaSerif-Medium              ITC 6097
i_____OfficinaSerif-MediumItalic        ITC 6098
i_____OfficinaSerif-ExtraBold           ITC 6099

i_____OfficinaSerif-ExtraBoldItalic     ITC 6100
i_____OfficinaSerif-Black               ITC 6101
i_____OfficinaSerif-BlackItalic         ITC 6102
i_____SerifGothic-Light                 ITC 6103
i_____SerifGothic-Regular               ITC 6104
i_____SerifGothic-Bold                  ITC 6105
i_____SerifGothic-ExtraBold             ITC 6106
i_____SerifGothic-Heavy                 ITC 6107
i_____SerifGothic-Black                 ITC 6108
i_____AmericanTypewriter-Light          ITC 6109
i_____AmericanTypewriter-LightItalic    ITC 6110
i_____AmericanTypewriter-Medium         ITC 6111
i_____AmericanTypewriter-MediumItalic   ITC 6112
i_____AmericanTypewriter-Bold           ITC 6113
i_____AmericanTypewriter-BoldItalic     ITC 6114
i_____AmericanTypewriter-CondLight      ITC 6115
i_____AmericanTypewriter-CondMedium     ITC 6116
i_____AmericanTypewriter-CondBold       ITC 6117
i_____LubalinGraph-ExtraLight           ITC 6118
i_____LubalinGraph-ExtraLightOblique    ITC 6119
i_____LubalinGraph-Book                 ITC 6120
i_____LubalinGraph-BookOblique          ITC 6121
i_____LubalinGraph-Medium               ITC 6122
i_____LubalinGraph-MediumOblique        ITC 6123
i_____LubalinGraph-Demi                 ITC 6124
i_____LubalinGraph-DemiOblique          ITC 6125
i_____LubalinGraph-Bold                 ITC 6126
i_____LubalinGraph-BoldOblique          ITC 6127
i_____Benguiat-Book                     ITC 6128
i_____Benguiat-BookItalic               ITC 6129
i_____Benguiat-Medium                   ITC 6130
i_____Benguiat-MediumItalic             ITC 6131
i_____Benguiat-Bold                     ITC 6132
i_____Benguiat-BoldItalic               ITC 6133
i_____Benguiat-CondensedBook            ITC 6134
i_____Benguiat-CondensedBookItalic      ITC 6135
i_____Benguiat-CondensedMedium          ITC 6136
i_____Benguiat-CondensedMediumItalic    ITC 6137
i_____Benguiat-CondensedBold            ITC 6138
i_____Benguiat-CondensedBoldItalic      ITC 6139
i_____Korinna-Regular                   ITC 6140
i_____Korinna-KursivRegular             ITC 6141
i_____Korinna-Bold                      ITC 6142
i_____Korinna-KursivBold                ITC 6143
i_____Korinna-ExtraBold                 ITC 6144
i_____Korinna-KursivExtraBold           ITC 6145
i_____Korinna-Heavy                     ITC 6146

i_____Korinna-KursivHeavy               ITC 6147
i_____Souvenir-Light                    ITC 6148
i_____Souvenir-LightItalic              ITC 6149
i_____Souvenir-Medium                   ITC 6150
i_____Souvenir-MediumItalic             ITC 6151
i_____Souvenir-Demi                     ITC 6152
i_____Souvenir-DemiItalic               ITC 6153
i_____Souvenir-Bold                     ITC 6154
i_____Souvenir-BoldItalic               ITC 6155
i_____Caslon224-Book                    ITC 6156
i_____Caslon224-BookItalic              ITC 6157
i_____Caslon224-Medium                  ITC 6158
i_____Caslon224-MediumItalic            ITC 6159
i_____Caslon224-Bold                    ITC 6160
i_____Caslon224-BoldItalic              ITC 6161
i_____Caslon224-Black                   ITC 6162
i_____Caslon224-BlackItalic             ITC 6163
i_____LubalinGraph-CondBook             ITC 6164
i_____LubalinGraph-CondBookOblique      ITC 6165
i_____LubalinGraph-CondMedium           ITC 6166
i_____LubalinGraph-CondMediumOblique    ITC 6167
i_____LubalinGraph-CondDemi             ITC 6168
i_____LubalinGraph-CondDemiOblique      ITC 6169
i_____LubalinGraph-CondBold             ITC 6170
i_____LubalinGraph-CondBoldOblique      ITC 6171
i_____Galliard-Regular                  ITC 6172
i_____Galliard-Italic                   ITC 6173
i_____Galliard-Bold                     ITC 6174
i_____Galliard-BoldItalic               ITC 6175
i_____Galliard-Black                    ITC 6176
i_____Galliard-BlackItalic              ITC 6177
i_____Galliard-Ultra                    ITC 6178
i_____Galliard-UltraItalic              ITC 6179
i_____Dyadis-Book                       ITC 6180
i_____Dyadis-BookItalic                 ITC 6181
i_____Dyadis-Medium                     ITC 6182
i_____Dyadis-MediumItalic               ITC 6183
i_____Dyadis-Bold                       ITC 6184
i_____Dyadis-BoldItalic                 ITC 6185
i_____BaileySans-Book                   ITC 8000
i_____BaileySans-BookItalic             ITC 8001
i_____BaileySans-Bold                   ITC 8002
i_____BaileySans-BoldItalic             ITC 8003
i_____Charlotte-SansBook                ITC 8004
i_____Charlotte-SansBookItalic          ITC 8005
i_____Charlotte-SansMedium              ITC 8006
i_____Charlotte-SansBold                ITC 8007

i_____Conduit-Light                     ITC 8008
i_____Conduit-LightItalic               ITC 8009
i_____Conduit-Medium                    ITC 8010
i_____Conduit-MediumItalic              ITC 8011
i_____Conduit-Bold                      ITC 8012
i_____Conduit-BoldItalic                ITC 8013
i_____Highlander-Book                   ITC 8014
i_____Highlander-BookItalic             ITC 8015
i_____Highlander-Medium                 ITC 8016
i_____Highlander-MediumItalic           ITC 8017
i_____Highlander-Bold                   ITC 8018
i_____Highlander-BoldItalic             ITC 8019
i_____HumanaSans-Light                  ITC 8020
i_____HumanaSans-LightItalic            ITC 8021
i_____HumanaSans-Medium                 ITC 8022
i_____HumanaSans-MediumItalic           ITC 8023
i_____HumanaSans-Bold                   ITC 8024
i_____HumanaSans-BoldItalic             ITC 8025
i_____LegacySans-Book                   ITC 8026
i_____LegacySans-BookItalic             ITC 8027
i_____LegacySans-Medium                 ITC 8028
i_____LegacySans-MediumItalic           ITC 8029
i_____LegacySans-Bold                   ITC 8030
i_____LegacySans-BoldItalic             ITC 8031
i_____LegacySans-Ultra                  ITC 8032
i_____Odyssee-Light                     ITC 8033
i_____Odyssee-LightItalic               ITC 8034
i_____Odyssee-Medium                    ITC 8035
i_____Odyssee-MediumItalic              ITC 8036
i_____Odyssee-Bold                      ITC 8037
i_____Odyssee-BoldItalic                ITC 8038
i_____Odyssee-Ultra                     ITC 8039
i_____OfficinaSans-Book                 ITC 8040
i_____OfficinaSans-BookItalic           ITC 8041
i_____OfficinaSans-Bold                 ITC 8042
i_____OfficinaSans-BoldItalic           ITC 8043
i_____StoneSans-Medium                  ITC 8044
i_____StoneSans-MediumItalic            ITC 8045
i_____StoneSans-SemiBold                ITC 8046
i_____StoneSans-SemiBoldItalic          ITC 8047
i_____StoneSans-Bold                    ITC 8048
i_____StoneSans-BoldItalic              ITC 8049
i_____Woodland-Light                    ITC 8050
i_____Woodland-Medium                   ITC 8051
i_____Woodland-Demi                     ITC 8052
i_____Woodland-Heavy                    ITC 8053
i_____OfficinaSans-Medium               ITC 8054

i_____OfficinaSans-MediumItalic         ITC 8055
i_____OfficinaSans-ExtraBold            ITC 8056
i_____OfficinaSans-ExtraBoldItalic      ITC 8057
i_____OfficinaSans-Black                ITC 8058
i_____AvantGardeGothic-ExtraLight       ITC 8059
i_____AvantGardeGothic-ExtraLightObl    ITC 8060
i_____AvantGardeGothic-Book             ITC 8061
i_____AvantGardeGothic-BookOblique      ITC 8062
i_____AvantGardeGothic-Medium           ITC 8063
i_____AvantGardeGothic-MediumOblique    ITC 8064
i_____AvantGardeGothic-Demi             ITC 8065
i_____AvantGardeGothic-DemiOblique      ITC 8066
i_____AvantGardeGothic-Bold             ITC 8067
i_____AvantGardeGothic-BoldOblique      ITC 8068
i_____FranklinGothic-Book               ITC 8069
i_____FranklinGothic-BookItalic         ITC 8070
i_____FranklinGothic-Medium             ITC 8071
i_____FranklinGothic-MediumItalic       ITC 8072
i_____FranklinGothic-Demi               ITC 8073
i_____FranklinGothic-DemiItalic         ITC 8074
i_____FranklinGothic-Heavy              ITC 8075
i_____FranklinGothic-HeavyItalic        ITC 8076
i_____FranklinGothic-CondBook           ITC 8077
i_____FranklinGothic-CondBookItal       ITC 8078
i_____FranklinGothic-CondMedium         ITC 8079
i_____FranklinGothic-CondMediumItal     ITC 8080
i_____FranklinGothic-CondDemi           ITC 8081
i_____FranklinGothic-CondDemiItal       ITC 8082
i_____FranklinGothic-ComprBook          ITC 8083
i_____FranklinGothic-ComprBookItal      ITC 8084
i_____FranklinGothic-ComprDemi          ITC 8085
i_____FranklinGothic-ComprDemiItal      ITC 8086
i_____FranklinGothic-ExtraComprBook     ITC 8087
i_____FranklinGothic-ExtraComprDemi     ITC 8088
i_____OfficinaSans-BlackItalic          ITC 8089
i_____BenguiatGothic-Book               ITC 8090
i_____BenguiatGothic-BookItalic         ITC 8091
i_____BenguiatGothic-Medium             ITC 8092
i_____BenguiatGothic-MediumItalic       ITC 8093
i_____BenguiatGothic-Bold               ITC 8094
i_____BenguiatGothic-BoldItalic         ITC 8095
i_____BenguiatGothic-Heavy              ITC 8096
i_____BenguiatGothic-HeavyItalic        ITC 8097
i_____ItcEras-Light                     ITC 8098
i_____ItcEras-Book                      ITC 8099
i_____ItcEras-Medium                    ITC 8100
i_____ItcEras-Demi                      ITC 8101

i_____ItcEras-Bold                      ITC 8102
i_____ItcEras-Ultra                     ITC 8103
i_____AvantGardeGothic-CondBook         ITC 8104
i_____AvantGardeGothic-CondMedium       ITC 8105
i_____AvantGardeGothic-CondDemiBold     ITC 8106
i_____AvantGardeGothic-CondBold         ITC 8107
i_____ItcKabel-Book                     ITC 8108
i_____ItcKabel-Medium                   ITC 8109
i_____ItcKabel-Demi                     ITC 8110
i_____ItcKabel-Bold                     ITC 8111
i_____ItcKabel-Ultra                    ITC 8112

A.7 Linotype fonts

The ‘linotype.map’ file has abbreviations for Linotype fonts.

l_____Aachen-Bold                              1.0
l_____AachenEF-Medium                          1.1
l_____AachenEF-Bold                            1.2
l_____AharoniEF-RegularHd                      2.0
l_____AharoniEF-MediumHd                       2.1
l_____AharoniEF-SemiBoldHd                     2.2
l_____AharoniEF-BoldHd                         2.3
l_____AharoniEF-ExtraBoldHd                    2.4
l_____AharoniEF-OutlineHd                      2.5
l_____AharoniEF-ShadowHD                       2.6
l_____Aja                                      3.0
l_____AkzidenzGrotesk-Light                    4.0
l_____AkzidenzGrotesk-Roman                    4.1
l_____AkzidenzGrotesk-Bold                     4.2
l_____AkzidenzGrotesk-Black                    4.3
l_____AkzidenzGroteskBE-Light                  4.4
l_____AkzidenzGroteskBE-LightOsF               4.5
l_____AkzidenzGroteskBE-Regular                4.6
l_____AkzidenzGroteskBE-Italic                 4.7
l_____AkzidenzGroteskBE-Md                     4.8
l_____AkzidenzGroteskBE-MdIt                   4.9
l_____AkzidenzGroteskBE-Bold                   4.10
l_____AkzidenzGroteskBE-BoldIt                 4.11
l_____AkzidenzGroteskBE-XBd                    4.12
l_____AkzidenzGroteskBE-Super                  4.13
l_____AkzidenzGroteskBE-LightCn                4.14
l_____AkzidenzGroteskBE-Cn                     4.15
l_____AkzidenzGroteskBE-MdCn                   4.16
l_____AkzidenzGroteskBE-MdCnIt                 4.17
l_____AkzidenzGroteskBE-BoldCn                 4.18
l_____AkzidenzGroteskBE-XBdCn                  4.19
l_____AkzidenzGroteskBE-XBdCnIt                4.20

l_____AkzidenzGroteskBE-LightEx                4.21
l_____AkzidenzGroteskBE-Ex                     4.22
l_____AkzidenzGroteskBE-MdEx                   4.23
l_____AkzidenzGroteskBE-BoldEx                 4.24
l_____AkzidenzGroteskBE-BoldExIt               4.25
l_____AGBookRounded-Regular                    4.26
l_____AGBookRounded-Medium                     4.27
l_____AGBookRounded-Bold                       4.28
l_____AGBookRounded-MediumOutline              4.29
l_____AGBookRounded-BoldOutline                4.30
l_____AGBookRounded-BoldCn                     4.31
l_____AGBookRounded-BoldCnOutline              4.32
l_____AGBook-Stencil                           4.33
l_____AGOldFace-Regular                        4.34
l_____AGOldFace-Medium                         4.35
l_____AGOldFace-Bold                           4.36
l_____AGOldFace-Outline                        4.37
l_____AGOldFace-BoldOutline                    4.38
l_____AGOldFace-Shaded                         4.39
l_____AGSchoolbook-Regular                     4.40
l_____AGSchoolbook-RegularA                    4.41
l_____AGSchoolbook-Medium                      4.42
l_____AGSchoolbook-MediumA                     4.43
l_____AlcuinEF-Light                           5.0
l_____AlcuinEF-Regular                         5.1
l_____AlcuinEF-Bold                            5.2
l_____AlcuinEF-ExtraBold                       5.3
l_____Aldus-Roman                              6.0
l_____Aldus-RomanSC                            6.1
l_____Aldus-Italic                             6.2
l_____Aldus-ItalicOsF                          6.3
l_____AleOrnamentsRotatoLL                     7.0
l_____AleOrnamentsSpiratoLL                    7.1
l_____AlexieLL                                 8.0
l_____AlteSchwabacherEF                        9.0
l_____AlternateGothicEF-NoOne                 10.0
l_____AlternateGothicEF-NoTwo                 10.1
l_____AlternateGothicEF-NoThree               10.2
l_____AmericanUncialEF                        11.0
l_____Americana                               12.0
l_____Americana-Italic                        12.1
l_____Americana-Bold                          12.2
l_____Americana-ExtraBold                     12.3
l_____AmericanaEF-Regular                     12.4
l_____AmericanaEF-RegularItalic               12.5
l_____AmericanaEF-Bold                        12.6
l_____AmericanaEF-ExtraBold                   12.7

l_____AnnlieEF-ExtraBold                      14.0
l_____AnnlieEF-ExtraBoldItalic                14.1
l_____AntiqueOlive-Light                      15.0
l_____AntiqueOlive-Roman                      15.1
l_____AntiqueOlive-Italic                     15.2
l_____AntiqueOlive-Bold                       15.3
l_____AntiqueOlive-Black                      15.4
l_____AntiqueOlive-BoldCond                   15.5
l_____AntiqueOlive-Compact                    15.6
l_____AntiqueOlive-Nord                       15.7
l_____AntiqueOlive-NordItalic                 15.8
l_____AntiqueOliveCE-Light                    15.9
l_____AntiqueOliveCE-Roman                    15.10
l_____AntiqueOliveCE-Italic                   15.11
l_____AntiqueOliveCE-Bold                     15.12
l_____AntiqueOliveCE-Black                    15.13
l_____AntiqueOliveEastA-Light                 15.14
l_____AntiqueOliveEastA-Roman                 15.15
l_____AntiqueOliveEastA-Italic                15.16
l_____AntiqueOliveEastA-Bold                  15.17
l_____AntiqueOliveEastA-Black                 15.18
l_____ApolloMT-Roman                          16.0
l_____ApolloMT-RomanSC                        16.1
l_____ApolloMT-RomanExpert                    16.2
l_____ApolloMT-Italic                         16.3
l_____ApolloMT-ItalicOsF                      16.4
l_____ApolloMT-ItalicExpert                   16.5
l_____ApolloMT-SemiBold                       16.6
l_____ApolloMT-SemiBoldOsF                    16.7
l_____ApolloMT-SemiBoldExpert                 16.8
l_____Arcadia                                 17.0
l_____Arcadia-Alternate                       17.1
l_____Ariadne-Roman                           18.0
l_____Arioso                                  19.0
l_____Arkona-Regular                          20.0
l_____Arkona-Medium                           20.1
l_____ArnoldBoecklin                          21.0
l_____AshleyScriptMT                          22.0
l_____NewAster                                23.0
l_____NewAster-Italic                         23.1
l_____NewAster-SemiBold                       23.2
l_____NewAster-SemiBoldItalic                 23.3
l_____NewAster-Bold                           23.4
l_____NewAster-BoldItalic                     23.5
l_____NewAster-Black                          23.6
l_____NewAster-BlackItalic                    23.7
l_____AsterEF-Regular                         23.8

l_____AsterEF-RegularItalic                   23.9
l_____AsterEF-Medium                          23.10
l_____AureliaEF-Light                         24.0
l_____AureliaEF-LightItalic                   24.1
l_____AureliaEF-Book                          24.2
l_____AureliaEF-BookItalic                    24.3
l_____AureliaEF-Bold                          24.4
l_____Auriol                                  25.0
l_____Auriol-Italic                           25.1
l_____Auriol-Bold                             25.2
l_____Auriol-BoldItalic                       25.3
l_____Auriol-Black                            25.4
l_____Auriol-BlackItalic                      25.5
l_____BalderLL                                28.0
l_____BalloonEF-ExtraBold                     29.0
l_____BalloonEF-DropShadow                    29.1
l_____BalmoralEF-Regular                      30.0
l_____BalmoralEF-Alternative                  30.1
l_____Banco                                   31.0
l_____BarbedorEF-Regular                      32.0
l_____BarbedorEF-RegularItalic                32.1
l_____BarbedorEF-Medium                       32.2
l_____BarbedorEF-MediumItalic                 32.3
l_____BarbedorEF-Bold                         32.4
l_____BarbedorEF-BoldItalic                   32.5
l_____BarbedorEF-Black                        32.6
l_____BarbedorEF-BlackItalic                  32.7
l_____BaskervilleCyr-Upright                  33.0
l_____BaskervilleCyr-Inclined                 33.1
l_____BaskervilleCyr-Bold                     33.2
l_____BaskervilleGreek-Upright                33.3
l_____BaskervilleGreek-Inclined               33.4
l_____BaskervilleGreek-Bold                   33.5
l_____BaskervilleGreekP-Upright               33.6
l_____BaskervilleGreekP-Inclined              33.7
l_____BaskervilleGreekP-Bold                  33.8
l_____BaskervilleEF-Regular                   33.9
l_____BaskervilleEF-RegularItalic             33.10
l_____BaskervilleEF-Medium                    33.11
l_____BaskervilleBE-Regular                   33.12
l_____BaskervilleBE-Italic                    33.13
l_____BaskervilleBE-Medium                    33.14
l_____BaskervilleBE-MediumItalic              33.15
l_____BaskervilleBE-Bold                      33.16
l_____NewBaskerville-Roman                    33.17
l_____NewBaskerville-SC                       33.18
l_____NewBaskerville-Italic                   33.19

l_____NewBaskerville-ItalicOsF                33.20
l_____NewBaskerville-Bold                     33.21
l_____NewBaskerville-BoldSC                   33.22
l_____NewBaskerville-BoldItalic               33.23
l_____NewBaskerville-BoldItalicOsF            33.24
l_____NewBaskervilleEF-Roman                  33.25
l_____NewBaskervilleEF-Italic                 33.26
l_____NewBaskervilleEF-SemiBold               33.27
l_____NewBaskervilleEF-SemiBoldItalic         33.28
l_____NewBaskervilleEF-Bold                   33.29
l_____NewBaskervilleEF-BoldItalic             33.30
l_____NewBaskervilleEF-Black                  33.31
l_____NewBaskervilleEF-BlackItalic            33.32
l_____BaskervilleBook-Regular                 33.33
l_____BaskervilleBook-Italic                  33.34
l_____BaskervilleBook-Medium                  33.35
l_____BaskervilleBook-MedItalic               33.36
l_____BaskervilleOldFaceEF                    33.37
l_____AfroditeKATEF-Filly                     34.0
l_____AfroditeKATEF-Ready                     34.1
l_____AfroditeKATEF-Funky                     34.2
l_____BadblocksEF                             34.3
l_____BeatEF-BackBeat                         34.4
l_____BeatEF-HardBeat                         34.5
l_____BloxxEF                                 34.6
l_____BornFreeKATEF-Hungry                    34.7
l_____BornFreeKATEF-Fed                       34.8
l_____BrushableEF                             34.9
l_____CandyEF                                 34.10
l_____ComancheEF                              34.11
l_____CrazeEF                                 34.12
l_____DoublePacEF-Scrambled                   34.13
l_____EastsideEF-Regular                      34.14
l_____EastsideEF-Double                       34.15
l_____FutureWorldEF                           34.16
l_____ItKATEF                                 34.17
l_____JameKATEF-Black                         34.18
l_____JameKATEF-White                         34.19
l_____KiddingKidEF                            34.20
l_____KnockoutEF                              34.21
l_____LittleJoeKATEF-Black                    34.22
l_____LittleJoeKATEF-White                    34.23
l_____MaoMaoEF                                34.24
l_____MaydayEF                                34.25
l_____MiroEF                                  34.26
l_____MrsBeasleyEF                            34.27
l_____NiniveEF                                34.28

l_____RumourEF                                34.29
l_____SiouxEF                                 34.30
l_____StayawhileEF                            34.31
l_____SteelPlateEF                            34.32
l_____TieriliEF                               34.33
l_____TruckEF                                 34.34
l_____TwinPickEF-Hard                         34.35
l_____XabbieEF                                34.36
l_____BellGothic-Light                        36.0
l_____BellGothic-Bold                         36.1
l_____BellGothic-Black                        36.2
l_____BelugaLL                                37.0
l_____Berling-Roman                           38.0
l_____Berling-Italic                          38.1
l_____Berling-Bold                            38.2
l_____Berling-BoldItalic                      38.3
l_____BerlingEF-Regular                       38.4
l_____BerlingEF-RegularItalic                 38.5
l_____BerlingEF-Bold                          38.6
l_____BernhardEF-Antique                      39.1
l_____BernhardEF-Fashion                      39.2
l_____BernhardModern-Roman                    39.3
l_____BernhardModern-Italic                   39.4
l_____BernhardModern-Bold                     39.5
l_____BernhardModern-BoldItalic               39.6
l_____BernhardSchoenschriftEF-Light           39.7
l_____BernhardSchoenschriftEF-Regular         39.8
l_____BertholdScript-Regular                  40.0
l_____BertholdScript-Medium                   40.1
l_____BetonEF-Light                           41.0
l_____BetonEF-DemiBold                        41.1
l_____BetonEF-Bold                            41.2
l_____BetonEF-ExtraBold                       41.3
l_____BetonEF-BoldCondensed                   41.4
l_____BibleScriptEF-Regular                   42.0
l_____BibleScriptEF-Alternates                42.1
l_____BibleScriptEF-Flourishes                42.2
l_____Blackoak                                43.0
l_____BlippoBlackEF                           44.0
l_____BottleneckEF                            45.0
l_____BramleyEF-Light                         46.0
l_____BramleyEF-Medium                        46.1
l_____BramleyEF-Bold                          46.2
l_____BramleyEF-ExtraBold                     46.3
l_____BrightonEF-Light                        47.0
l_____BrightonEF-LightItalic                  47.1
l_____BrightonEF-Medium                       47.2

l_____BrightonEF-Bold                         47.3
lpi___BundesbahnPi-One                        48.0
lpi___BundesbahnPi-Two                        48.1
lpi___BundesbahnPi-Three                      48.2
l_____BusterEF                                49.0
l_____CandiceEF                               50.0
lpi___CaravanLH-One                           51.0
lpi___CaravanLH-Two                           51.1
lpi___CaravanLH-Three                         51.2
lpi___CaravanLH-Four                          51.3
l_____CascadeScript                           52.0
l_____CashEF-Monospace                        53.0
l_____CastellarMT                             54.0
l_____Charlemagne-Regular                     55.0
l_____Charlemagne-Bold                        55.1
l_____Christiana-Regular                      57.0
l_____Christiana-RegularSC                    57.1
l_____Christiana-RegularExpert                57.2
l_____Christiana-Italic                       57.3
l_____Christiana-Medium                       57.4
l_____Christiana-MediumItalic                 57.5
l_____Christiana-Bold                         57.6
l_____Christiana-BoldItalic                   57.7
l_____ChwastBuffalo-BlackCond                 58.0
l_____CirkulusEF                              59.0
l_____Clairvaux                               60.0
l_____Clairvaux-Dfr                           60.1
l_____Clearface-Regular                       61.0
l_____Clearface-RegularItalic                 61.1
l_____Clearface-Bold                          61.2
l_____Clearface-BoldItalic                    61.3
l_____Clearface-Heavy                         61.4
l_____Clearface-HeavyItalic                   61.5
l_____Clearface-Black                         61.6
l_____Clearface-BlackItalic                   61.7
l_____ClearfaceEF-Regular                     61.8
l_____ClearfaceEF-RegularItalic               61.9
l_____ClearfaceEF-Bold                        61.10
l_____ClearfaceEF-BoldItalic                  61.11
l_____ClearfaceEF-Heavy                       61.12
l_____ClearfaceEF-HeavyItalic                 61.13
l_____ClearfaceEF-Black                       61.14
l_____ClearfaceEF-BlackItalic                 61.15
l_____ClearfaceGothicLH-Light                 61.16
l_____ClearfaceGothicLH-Roman                 61.17
l_____ClearfaceGothicLH-Medium                61.18
l_____ClearfaceGothicLH-Bold                  61.19

l_____ClearfaceGothicLH-Black                 61.20
l_____ClearfaceGothicEF-Regular               61.21
l_____ClearfaceGothicEF-Bold                  61.22
l_____ClearfaceGothicEF-ExtraBold             61.23
l_____CloisterOpenFace                        62.0
l_____Cochin                                  63.0
l_____Cochin-Italic                           63.1
l_____Cochin-Bold                             63.2
l_____Cochin-BoldItalic                       63.3
l_____Codex                                   64.0
l_____ColumnaEF-Solid                         65.0
l_____ComeniusAntiqua-Regular                 66.0
l_____ComeniusAntiqua-Italic                  66.1
l_____ComeniusAntiqua-Medium                  66.2
l_____ComeniusAntiqua-Bold                    66.3
l_____CommercialScriptEF                      67.0
l_____Concorde-Roman                          69.0
l_____Concorde-Italic                         69.1
l_____Concorde-Bold                           69.2
l_____Concorde-BoldItalic                     69.3
l_____ConcordeBE-Regular                      69.4
l_____ConcordeBE-RegularSC                    69.5
l_____ConcordeBE-RegularExp                   69.6
l_____ConcordeBE-Italic                       69.7
l_____ConcordeBE-ItalicOsF                    69.8
l_____ConcordeBE-ItalicExp                    69.9
l_____ConcordeBE-Medium                       69.10
l_____ConcordeBE-MediumSC                     69.11
l_____ConcordeBE-MediumEx                     69.12
l_____ConcordeBE-MediumItalic                 69.13
l_____ConcordeBE-MediumItalicOsF              69.14
l_____ConcordeBE-MediumItalicExp              69.15
l_____ConcordeBE-Condensed                    69.16
l_____ConcordeBE-MediumCn                     69.17
l_____ConcordeBE-BoldCn                       69.18
l_____ConcordeBE-BoldCnOutline                69.19
l_____ConcordeNova-Regular                    69.20
l_____ConcordeNova-RegularSC                  69.21
l_____ConcordeNova-RegularExp                 69.22
l_____ConcordeNova-Italic                     69.23
l_____ConcordeNova-ItalicOsF                  69.24
l_____ConcordeNova-ItalicExp                  69.25
l_____ConcordeNova-Medium                     69.26
l_____ConcordeNova-MediumSC                   69.27
l_____ConcordeNova-MediumExp                  69.28
l_____ContactaLL                              70.0
l_____CoolWoolLL                              71.0

l_____CoolWoolCottonClubLL                    71.1
l_____CoolWoolStoneWashedLL                   71.2
l_____CooperBlack                             72.0
l_____CooperBlack-Italic                      72.1
l_____CooperBlackEF-Bold                      72.2
l_____CooperBlackEF-BoldOutline               72.3
l_____CooperBlackEF-BoldCondensed             72.4
l_____CooperBlackEF-BoldCondOutline           72.5
l_____Copal-Solid                             73.0
l_____Copal-Outline                           73.1
l_____Copal-Decorated                         73.2
l_____CorinthianEF-Light                      74.0
l_____CorinthianEF-Medium                     74.1
l_____CorinthianEF-Bold                       74.2
l_____CorinthianEF-ExtraBold                  74.3
l_____Corona-Roman                            75.0
l_____Corona-Italic                           75.1
l_____Corona-Bold                             75.2
l_____CortezEF                                76.0
l_____Cosmos-Light                            77.0
l_____Cosmos-LightItalic                      77.1
l_____Cosmos-Medium                           77.2
l_____Cosmos-ExtraBold                        77.3
l_____CountdownEF                             78.0
l_____CourierCE-Roman                         79.0
l_____CourierCE-Bold                          79.1
l_____Cremona-Regular                         80.0
l_____Cremona-Italic                          80.1
l_____Cremona-Bold                            80.2
l_____Cremona-BoldItalic                      80.3
l_____Critter                                 81.0
l_____CroissantEF                             82.0
l_____Cutout                                  83.0
l_____DailyNews-Regular                       84.0
l_____DailyNews-Italic                        84.1
l_____DailyNews-Medium                        84.2
l_____DailyNews-MediumItalic                  84.3
l_____DailyNews-Bold                          84.4
l_____DailyNews-BoldItalic                    84.5
l_____DailyNews-ExtraBold                     84.6
l_____DailyNews-ExtraBoldItalic               84.7
lpi___DamnedDingbatsEF-Regular                85.0
lpi___DamnedDingbatsEF-Initials               85.1
l_____Data70EF                                86.0
l_____DavidaEF                                87.0
l_____Delphin-I                               88.0
l_____Delphin-II                              88.1

l_____Delphin-IA                              88.2
l_____Delphin-IIA                             88.3
l_____Delta-Light                             89.0
l_____Delta-LightItalic                       89.1
l_____Delta-Book                              89.2
l_____Delta-BookItalic                        89.3
l_____Delta-Medium                            89.4
l_____Delta-MediumItalic                      89.5
l_____Delta-Bold                              89.6
l_____Delta-BoldItalic                        89.7
l_____Delta-Outline                           89.8
l_____DemianEF                                90.0
lpi___DeutscheBahnAG-One                      91.0
lpi___DeutscheBahnAG-Two                      91.1
lpi___DeutscheBahnAG-Three                    91.2
lpi___DeutscheBahnAG-Four                     91.3
lpi___DeutscheBahnAG-Six                      91.4
lpi___DeutscheBahnAG-Five                     91.5
l_____DexGothicEF                             92.0
l_____DigiAntiquaEF-LightCondensed            93.0
l_____Doric-Bold                              95.0
l_____DrugulinEF-RegularHd                    96.0
l_____DrugulinEF-MediumHd                     96.1
l_____DrugulinEF-BoldHd                       96.2
l_____DrugulinEF-ExtraBoldHd                  96.3
l_____EckmannEF                               97.1
l_____EdwardianScriptEF-Regular               98.0
l_____EdwardianScriptEF-RegularA              98.1
l_____EdwardianScriptEF-Bold                  98.2
l_____EdwardianScriptEF-BoldA                 98.3
l_____Egyptian505EF-Light                     99.0
l_____Egyptian505EF-Regular                   99.1
l_____Egyptian505EF-Medium                    99.2
l_____Egyptian505EF-Bold                      99.3
l_____EgyptienneF-Roman                      100.0
l_____EgyptienneF-Italic                     100.1
l_____EgyptienneF-Bold                       100.2
l_____EgyptienneF-Black                      100.3
l_____ElectraLH-Roman                        101.0
l_____ElectraLH-RomanSC                      101.1
l_____ElectraLH-RomanOsF                     101.2
l_____ElectraLH-Cursive                      101.3
l_____ElectraLH-CursiveOsF                   101.4
l_____ElectraLH-Bold                         101.5
l_____ElectraLH-BoldSC                       101.6
l_____ElectraLH-BoldOsF                      101.7
l_____ElectraLH-BoldCursive                  101.8

l_____ElectraLH-BoldCursiveOsF               101.9
l_____ElectraLH-Display                      101.10
l_____ElectraLH-CursiveDisplay               101.11
l_____ElectraLH-BoldDisplay                  101.12
l_____ElectraLH-BoldCursiveDisplay           101.13
l_____ElseNPL-Light                          102.0
l_____ElseNPL-Medium                         102.1
l_____ElseNPL-SemiBold                       102.2
l_____ElseNPL-Bold                           102.3
l_____EnviroEF                               103.0
l_____EscriptLL                              104.0
lpi___EuropeanPi-One                         105.0
lpi___EuropeanPi-Two                         105.1
lpi___EuropeanPi-Three                       105.2
lpi___EuropeanPi-Four                        105.2
l_____ExcaliburEF-Regular                    106.0
l_____ExcaliburEF-Medium                     106.1
l_____FacsimileLL                            107.0
l_____FairfieldLH-Light                      108.0
l_____FairfieldLH-LightSC                    108.1
l_____FairfieldLH-LightItalic                108.2
l_____FairfieldLH-Medium                     108.3
l_____FairfieldLH-MediumSC                   108.4
l_____FairfieldLH-MediumItalic               108.5
l_____FairfieldLH-Bold                       108.6
l_____FairfieldLH-BoldSC                     108.7
l_____FairfieldLH-BoldItalic                 108.8
l_____FairfieldLH-Heavy                      108.9
l_____FairfieldLH-HeavySC                    108.10
l_____FairfieldLH-HeavyItalic                108.11
l_____FairfieldLH-CaptionLight               108.12
l_____FairfieldLH-CaptionMedium              108.13
l_____FairfieldLH-CaptionBold                108.14
l_____FairfieldLH-CaptionHeavy               108.15
l_____FairfieldLH-SwLightItalicOsF           108.16
l_____FairfieldLH-SwMediumItalicOsF          108.17
l_____FairfieldLH-SwBoldItalicOsF            108.18
l_____FairfieldLH-SwHeavyItalicOsF           108.19
l_____FetteGotischEF                         109.0
l_____FlamencoInlineEF                       110.0
l_____Forbes-Bold                            114.0
l_____ForteMT                                115.0
l_____FranklinGothic-Condensed               116.0
l_____FranklinGothic-ExtraCond               116.1
l_____FranklinGothicEF-Regular               116.2
l_____FranklinGothicEF-RegularCond           116.3
l_____FranklinGothicEF-ExtraCond             116.4

l_____FranklinGothic-Roman                   116.5
l_____FranklinGothic-Book                    116.6
l_____FranklinGothic-BookOblique             116.7
l_____FranklinGothic-Demi                    116.8
l_____FranklinGothic-DemiOblique             116.9
l_____FranklinGothic-Heavy                   116.10
l_____FranklinGothic-HeavyOblique            116.11
l_____FranklinGothicEF-Book                  116.12
l_____FranklinGothicEF-BookItalic            116.13
l_____FranklinGothicEF-Medium                116.14
l_____FranklinGothicEF-MediumItalic          116.15
l_____FranklinGothicEF-Demi                  116.16
l_____FranklinGothicEF-DemiItalic            116.17
l_____FranklinGothicEF-Heavy                 116.18
l_____FranklinGothicEF-HeavyItalic           116.19
l_____FranklinGothicEF-BookCnd               116.20
l_____FranklinGothicEF-BookCndSC             116.21
l_____FranklinGothicEF-BookCndItal           116.22
l_____FranklinGothicEF-MedCnd                116.23
l_____FranklinGothicEF-MedCndSC              116.24
l_____FranklinGothicEF-MedCndItal            116.25
l_____FranklinGothicEF-DemiCnd               116.26
l_____FranklinGothicEF-DemiCndItal           116.27
l_____FranklinGothicEF-BookCmpr              116.28
l_____FranklinGothicEF-BookCmprItal          116.29
l_____FranklinGothicEF-DemiCmpr              116.30
l_____FranklinGothicEF-DemiCmprItal          116.31
l_____FranklinGothicEF-BookXCmpr             116.32
l_____FranklinGothicEF-DemiXCmpr             116.33
l_____Frutiger-Light                         117.0
l_____Frutiger-LightItalic                   117.1
l_____Frutiger-Roman                         117.2
l_____Frutiger-Italic                        117.3
l_____Frutiger-Bold                          117.4
l_____Frutiger-BoldItalic                    117.5
l_____Frutiger-Black                         117.6
l_____Frutiger-BlackItalic                   117.7
l_____Frutiger-UltraBlack                    117.8
l_____Frutiger-LightCn                       117.9
l_____Frutiger-Cn                            117.10
l_____Frutiger-BoldCn                        117.11
l_____Frutiger-BlackCn                       117.12
l_____Frutiger ExtraBlackCn                  117.13
l_____Futura-Light                           118.0
l_____Futura-LightOblique                    118.1
l_____Futura-Book                            118.2
l_____Futura-BookOblique                     118.3

l_____Futura                                 118.4
l_____Futura-Oblique                         118.5
l_____Futura-Heavy                           118.6
l_____Futura-HeavyOblique                    118.7
l_____Futura-Bold                            118.8
l_____Futura-BoldOblique                     118.9
l_____Futura-ExtraBold                       118.10
l_____Futura-ExtraBoldOblique                118.11
l_____Futura-CondensedLight                  118.12
l_____Futura-CondensedLightOblique           118.13
l_____Futura-CondensedMedium                 118.14
l_____Futura-CondensedMediumOblique          118.15
l_____Futura-CondensedBold                   118.16
l_____Futura-CondensedBoldOblique            118.17
l_____Futura-CondensedExtraBlack             118.18
l_____Futura-CondExtraBlackOblique           118.19
l_____FuturaEF-Light                         118.22
l_____FuturaEF-LightOblique                  118.23
l_____FuturaEF-Book                          118.24
l_____FuturaEF-BookOblique                   118.25
l_____FuturaEF-Medium                        118.26
l_____FuturaEF-MediumOblique                 118.27
l_____FuturaEF-Bold                          118.28
l_____FuturaEF-DemiBold                      118.29
l_____FuturaEF-BoldOblique                   118.30
l_____FuturaEF-DemiBoldOblique               118.31
l_____FuturaEF-ExtraBold                     118.32
l_____FuturaEF-ExtraBoldOblique              118.33
l_____FuturaEF-Heavy                         118.34
l_____FuturaEF-LightCondensed                118.35
l_____FuturaEF-LightCondensedOblique         118.36
l_____FuturaEF-MediumCondensed               118.37
l_____FuturaEF-MediumCondensedOblique        118.38
l_____FuturaEF-BoldCondensed                 118.39
l_____FuturaEF-BoldCondensedOblique          118.40
l_____FuturaEF-ExtraBoldCondensed            118.41
l_____FuturaEF-ExtraBoldCondOblique          118.42
l_____FuturaBlackEF                          118.43
l_____FuturaDisplayEF                        118.44
l_____FuturaScriptEF                         118.45
l_____FuturaEF-ExtraBoldShaded               118.46
l_____Galahad-Regular                        119.0
l_____Galahad-Alternate                      119.1
l_____Galahad-OsF                            119.2
l_____GarthGraphic                           121.0
l_____GarthGraphic-Italic                    121.1
l_____GarthGraphic-Bold                      121.2

l_____GarthGraphic-BoldItalic                121.3
l_____GarthGraphic-Black                     121.4
l_____GarthGraphic ExtraBlack                121.5
l_____GarthGraphic-Condensed                 121.6
l_____GarthGraphic-BoldCondensed             121.7
l_____GazetteLH-Roman                        122.0
l_____GazetteLH-Italic                       122.1
l_____GazetteLH-Bold                         122.2
lpi___GiacomettiLL                           123.0
l_____GillSans-Light                         124.0
l_____GillSans-LightItalic                   124.1
l_____GillSans                               124.2
l_____GillSans-Italic                        124.3
l_____GillSans-Bold                          124.4
l_____GillSans-BoldItalic                    124.5
l_____GillSans-ExtraBold                     124.6
l_____GillSans-UltraBold                     124.7
l_____GillSans-LightShadowed                 124.8
l_____GillSans-Shadowed                      124.9
l_____GillSans-Condensed                     124.10
l_____GillSans-BoldCondensed                 124.11
l_____GillSans-BoldExtraCondensed            124.12
l_____GillSans-UltraBoldCondensed            124.13
l_____GillSans-ExtraBoldDisplay              124.14
l_____GilliesGothicEF-Light                  125.0
l_____GilliesGothicEF-Bold                   125.1
l_____GilliesGothicEF-ExtraBold              125.2
l_____GilliesGothicEF-XBoShaded              125.3
l_____GlastonburyEF                          126.0
l_____Glypha-Thin                            127.0
l_____Glypha-ThinOblique                     127.1
l_____Glypha-Light                           127.2
l_____Glypha-LightOblique                    127.3
l_____Glypha                                 127.4
l_____Glypha-RomanOblique                    127.5
l_____Glypha-Bold                            127.6
l_____Glypha-BoldOblique                     127.7
l_____Glypha-Black                           127.8
l_____Glypha-BlackOblique                    127.9
l_____GoodiesLL                              128.0
l_____Gothic-Thirteen                        129.0
l_____Goudy-Bold                             130.0
l_____Goudy-BoldOsF                          130.1
l_____Goudy-BoldItalic                       130.2
l_____Goudy-BoldItalicOsF                    130.3
l_____Goudy-ExtraBold                        130.4
l_____Goudy-Heavyface                        130.5

l_____Goudy-HeavyfaceItalic                  130.6
l_____GoudyHandtooledEF                      130.7
l_____GoudyCatalogueEF-Regular               130.8
l_____GoudyCatalogueEF-Italic                130.9
l_____GoudyModernMT-Roman                    130.10
l_____GoudyModernMT-Italic                   130.11
l_____Goudy                                  130.12
l_____Goudy-SC                               130.13
l_____Goudy-Italic                           130.14
l_____Goudy-ItalicOsF                        130.15
l_____GoudySans-Book                         130.16
l_____GoudySans-BookItalic                   130.17
l_____GoudySans-Medium                       130.18
l_____GoudySans-MediumItalic                 130.19
l_____GoudySans-Bold                         130.20
l_____GoudySans-BoldItalic                   130.21
l_____GoudySans-Black                        130.22
l_____GoudySans-BlackItalic                  130.23
l_____GoudySansEF-Book                       130.24
l_____GoudySansEF-BookSC                     130.25
l_____GoudySansEF-BookItalic                 130.26
l_____GoudySansEF-Medium                     130.27
l_____GoudySansEF-MediumSC                   130.28
l_____GoudySansEF-MediumItalic               130.29
l_____GoudySansEF-Bold                       130.30
l_____GoudySansEF-BoldItalic                 130.31
l_____GoudySansEF-Black                      130.32
l_____GoudySansEF-BlackItalic                130.33
l_____GoudyTextMT-Regular                    130.34
l_____GoudyTextMT-Alternate                  130.35
l_____GoudyTextMT-Dfr                        130.36
l_____GoudyTextMT-LombardicCapitals          130.37
l_____Grafiko                                131.0
l_____Granjon                                132.0
l_____Granjon-SC                             132.1
l_____Granjon-Italic                         132.2
l_____Granjon-ItalicOsF                      132.3
l_____Granjon-Bold                           132.4
l_____Granjon-BoldOsF                        132.5
l_____GrotesqueMT-Light                      133.0
l_____GrotesqueMT-LightItalic                133.1
l_____GrotesqueMT                            133.2
l_____GrotesqueMT-Italic                     133.3
l_____GrotesqueMT-Bold                       133.4
l_____GrotesqueMT-Black                      133.5
l_____GrotesqueMT-LightCondensed             133.6
l_____GrotesqueMT-Condensed                  133.7

l_____GrotesqueMT-ExtraCondensed             133.8
l_____GrotesqueMT-BoldExtended               133.9
l_____Guardi-Roman                           134.0
l_____Guardi-Italic                          134.1
l_____Guardi-Bold                            134.2
l_____Guardi-BoldItalic                      134.3
l_____Guardi-Black                           134.4
l_____Guardi-BlackItalic                     134.5
l_____HadfieldEF                             135.0
l_____Hadriano-Light                         136.0
l_____Hadriano-Bold                          136.1
l_____Hadriano-ExtraBold                     136.2
l_____Hadriano-ExtraBoldCondensed            136.3
l_____HanselEF-Regular                       137.0
l_____HanselEF-Alternate                     137.1
l_____HarlowEF-Solid                         138.0
l_____HarlowEF                               138.1
l_____HawthornEF                             139.0
l_____Helvetica-Light                        140.0
l_____Helvetica-LightOblique                 140.1
l_____Helvetica                              140.2
l_____Helvetica-Oblique                      140.3
l_____Helvetica-Bold                         140.4
l_____Helvetica-BoldOblique                  140.5
l_____Helvetica-Black                        140.6
l_____Helvetica-BlackOblique                 140.7
l_____Helvetica-Condensed-Light              140.8
l_____Helvetica-Condensed-LightObl           140.9
l_____Helvetica-Condensed                    140.10
l_____Helvetica-Condensed-Oblique            140.11
l_____Helvetica-Condensed-Bold               140.12
l_____Helvetica-Condensed-BoldObl            140.13
l_____Helvetica-Condensed-Black              140.14
l_____Helvetica-Condensed-BlackObl           140.15
l_____Helvetica-Compressed                   140.16
l_____Helvetica-ExtraCompressed              140.17
l_____Helvetica-UltraCompressed              140.18
l_____HelveticaInserat-Roman                 140.19
l_____HelveticaRounded-Bold                  140.24
l_____HelveticaRounded-BoldObl               140.25
l_____HelveticaRounded-Black                 140.26
l_____HelveticaRounded-BlackObl              140.27
l_____HelveticaRounded-BoldCond              140.28
l_____HelveticaRounded-BoldCondObl           140.29
l_____HelveticaTextbook                      140.30
l_____HelveticaTextbook-Oblique              140.31
l_____HelveticaTextbook-Bold                 140.32

l_____HelveticaTextbook-BoldObl              140.33
l_____HelveticaCE-Roman                      140.36
l_____HelveticaCE-Bold                       140.37
l_____HelveticaCE-NarrowRoman                140.38
l_____HelveticaCE-NarrowBold                 140.39
l_____HelveticaCyr-Upright                   140.40
l_____HelveticaCyr-Inclined                  140.41
l_____HelveticaCyr-Bold                      140.42
l_____HelveticaCyr-BoldInclined              140.43
l_____HelveticaInseratCyr-Upright            140.44
l_____HelveticaGreek-Upright                 140.45
l_____HelveticaGreek-Inclined                140.46
l_____HelveticaGreek-Bold                    140.47
l_____HelveticaGreek-BoldInclined            140.48
l_____HelveticaGreekP-Upright                140.49
l_____HelveticaGreekP-Inclined               140.50
l_____HelveticaGreekP-Bold                   140.51
l_____HelveticaGreekP-BoldInclined           140.52
l_____HelveticaNeue-UltraLight               140.53
l_____HelveticaNeue-UltraLightItalic         140.54
l_____HelveticaNeue-Thin                     140.55
l_____HelveticaNeue-ThinItalic               140.56
l_____HelveticaNeue-Light                    140.57
l_____HelveticaNeue-LightItalic              140.58
l_____HelveticaNeue-Roman                    140.59
l_____HelveticaNeue-Italic                   140.60
l_____HelveticaNeue-Medium                   140.61
l_____HelveticaNeue-MediumItalic             140.62
l_____HelveticaNeue-Bold                     140.63
l_____HelveticaNeue-BoldItalic               140.64
l_____HelveticaNeue-Heavy                    140.65
l_____HelveticaNeue-HeavyItalic              140.66
l_____HelveticaNeue-Black                    140.67
l_____HelveticaNeue-BlackItalic              140.68
l_____HelveticaNeue-BoldOutline              140.69
l_____HelveticaNeue-UltraLigCond             140.70
l_____HelveticaNeue-UltraLigCondObl          140.71
l_____HelveticaNeue-ThinCond                 140.72
l_____HelveticaNeue-ThinCondObl              140.73
l_____HelveticaNeue-LightCond                140.74
l_____HelveticaNeue-LightCondObl             140.75
l_____HelveticaNeue-Condensed                140.76
l_____HelveticaNeue-CondensedObl             140.77
l_____HelveticaNeue-MediumCond               140.78
l_____HelveticaNeue-MediumCondObl            140.79
l_____HelveticaNeue-BoldCond                 140.80
l_____HelveticaNeue-BoldCondObl              140.81

l_____HelveticaNeue-HeavyCond                140.82
l_____HelveticaNeue-HeavyCondObl             140.83
l_____HelveticaNeue-BlackCond                140.84
l_____HelveticaNeue-BlackCondObl             140.85
l_____HelveticaNeue-ExtBlackCond             140.86
l_____HelveticaNeue-ExtBlackCondObl          140.87
l_____HelveticaNeue-UltraLightExt            140.88
l_____HelveticaNeue-UltraLightExtObl         140.89
l_____HelveticaNeue-ThinExt                  140.90
l_____HelveticaNeue-ThinExtObl               140.91
l_____HelveticaNeue-LightExt                 140.92
l_____HelveticaNeue-LightExtObl              140.93
l_____HelveticaNeue-Extended                 140.94
l_____HelveticaNeue-ExtendedObl              140.95
l_____HelveticaNeue-MediumExt                140.96
l_____HelveticaNeue-MediumExtObl             140.97
l_____HelveticaNeue-BoldExt                  140.98
l_____HelveticaNeue-BoldExtObl               140.99
l_____HelveticaNeue-HeavyExt                 140.100
l_____HelveticaNeue-HeavyExtObl              140.101
l_____HelveticaNeue-BlackExt                 140.102
l_____HelveticaNeue-BlackExtObl              140.103
l_____HelveticaNeueCE-Thin                   140.104
l_____HelveticaNeueCE-ThinItalic             140.105
l_____HelveticaNeueCE-Roman                  140.106
l_____HelveticaNeueCE-Italic                 140.107
l_____HelveticaNeueCE-Bold                   140.108
l_____HelveticaNeueCE-BoldItalic             140.109
l_____HelveticaNeueEastA-Thin                140.110
l_____HelveticaNeueEastA-ThinObl             140.111
l_____HelveticaNeueEastA-Roman               140.112
l_____HelveticaNeueEastA-RomanObl            140.113
l_____HelveticaNeueEastA-Bold                140.114
l_____HelveticaNeueEastA-BoldObl             140.115
l_____HeraldusEF-ExtraFett                   141.0
l_____HighlightEF                            143.0
l_____Hiroshige-Book                         144.0
l_____Hiroshige-BookItalic                   144.1
l_____Hiroshige-Medium                       144.2
l_____Hiroshige-MediumItalic                 144.3
l_____Hiroshige-Bold                         144.4
l_____Hiroshige-BoldItalic                   144.5
l_____Hiroshige-Black                        144.6
l_____Hiroshige-BlackItalic                  144.7
l_____Hobo                                   145.0
l_____HoboEF                                 145.1
l_____HogarthScriptEF-Regular                146.0

l_____HondaEF                                147.0
l_____HoratioEF-Light                        148.0
l_____HoratioEF-Medium                       148.1
l_____HoratioEF-Bold                         148.2
l_____HorleyOldStyleMT-Light                 149.0
l_____HorleyOldStyleMT-LightItalic           149.1
l_____HorleyOldStyleMT-Roman                 149.2
l_____HorleyOldStyleMT-Italic                149.3
l_____HorleyOldStyleMT-SemiBold              149.4
l_____HorleyOldStyleMT-SbItalic              149.5
l_____HorleyOldStyleMT-Bold                  149.6
l_____HorleyOldStyleMT-BoldItalic            149.7
lpi___HotmetalBorders-One                    150.0
lpi___HotmetalBorders-Two                    150.1
lpi___HotmetalBorders-Three                  150.2
lpi___HotmetalBorders-Four                   150.3
l_____AmericanTypewriter-Light               153.0
l_____AmericanTypewriter-LightA              153.1
l_____AmericanTypewriter-Medium              153.2
l_____AmericanTypewriter-MediumA             153.3
l_____AmericanTypewriter-Bold                153.4
l_____AmericanTypewriter-BoldA               153.5
l_____AmericanTypewriter-LightCond           153.6
l_____AmericanTypewriter-LightCondA          153.7
l_____AmericanTypewriter-MediumCond          153.8
l_____AmericanTypewriter-MediumCondA         153.9
l_____AmericanTypewriter-BoldCond            153.10
l_____AmericanTypewriter-BoldCondA           153.11
l_____AmericanTypewriterEF-Light             153.14
l_____AmericanTypewriterEF-LightItalic       153.15
l_____AmericanTypewriterEF-Medium            153.16
l_____AmericanTypewriterEF-MediumItalic      153.17
l_____AmericanTypewriterEF-Bold              153.18
l_____AmericanTypewriterEF-BoldItalic        153.19
l_____AmericanTypewriterEF-LightCond         153.20
l_____AmericanTypewriterEF-MediumCond        153.21
l_____AmericanTypewriterEF-BoldCond          153.22
l_____Anna                                   154.0
l_____AnnaEF                                 154.1
l_____AnnaEF-SC                              154.2
l_____AvantGarde-ExtraLight                  155.0
l_____AvantGarde-ExtraLightObl               155.1
l_____AvantGarde-Book                        155.2
l_____AvantGarde-BookOblique                 155.3
l_____AvantGarde-Medium                      155.4
l_____AvantGarde-MediumObl                   155.5
l_____AvantGarde-Demi                        155.6

l_____AvantGarde-DemiOblique                 155.7
l_____AvantGarde-Bold                        155.8
l_____AvantGarde-BoldObl                     155.9
l_____AvantGarde-CondBook                    155.10
l_____AvantGarde-CondMedium                  155.11
l_____AvantGarde-CondDemi                    155.12
l_____AvantGarde-CondBold                    155.13
l_____AvantGardeGothicCE-Book                155.14
l_____AvantGardeGothicCE-Demi                155.15
l_____AvantGardeGothicEF-ExtraLight          155.16
l_____AvantGardeGothicEF-ExtraLightObl       155.17
l_____AvantGardeGothicEF-Book                155.18
l_____AvantGardeGothicEF-BookOblique         155.19
l_____AvantGardeGothicEF-Medium              155.20
l_____AvantGardeGothicEF-MediumOblique       155.21
l_____AvantGardeGothicEF-Demi                155.22
l_____AvantGardeGothicEF-DemiOblique         155.23
l_____AvantGardeGothicEF-Bold                155.24
l_____AvantGardeGothicEF-BoldOblique         155.25
l_____AvantGardeGothicEF-CondBook            155.26
l_____AvantGardeGothicEF-CondMedium          155.27
l_____AvantGardeGothicEF-CondDemi            155.28
l_____AvantGardeGothicEF-CondBold            155.29
l_____AvantGardeMonoEF-Light                 155.30
l_____AvantGardeMonoEF-LightIta              155.31
l_____AvantGardeMonoEF-Bold                  155.32
l_____AvantGardeMonoEF-BoldIta               155.33
l_____Beesknees                              156.0
l_____BeeskneesEF                            156.1
l_____Benguiat-Book                          157.0
l_____Benguiat-BookItalic                    157.1
l_____Benguiat-Medium                        157.2
l_____Benguiat-MediumItalic                  157.3
l_____Benguiat-Bold                          157.4
l_____Benguiat-BoldItalic                    157.5
l_____BenguiatEF-Book                        157.6
l_____BenguiatEF-BookItalic                  157.7
l_____BenguiatEF-Medium                      157.8
l_____BenguiatEF-MediumItalic                157.9
l_____BenguiatEF-Bold                        157.10
l_____BenguiatEF-BoldItalic                  157.11
l_____BenguiatEF-BookCondensed               157.12
l_____BenguiatEF-BookCondensedIta            157.13
l_____BenguiatEF-MediumCondensed             157.14
l_____BenguiatEF-MediumCondensedIta          157.15
l_____BenguiatEF-BoldCondensed               157.16
l_____BenguiatEF-BoldCondensedIta            157.17

l_____Berkeley-Book                          158.0
l_____Berkeley-BookItalic                    158.1
l_____Berkeley-Medium                        158.2
l_____Berkeley-Italic                        158.3
l_____Berkeley-Bold                          158.4
l_____Berkeley-BoldItalic                    158.5
l_____Berkeley-Black                         158.6
l_____Berkeley-BlackItalic                   158.7
l_____BerkeleyOldstyleEF-Book                158.8
l_____BerkeleyOldstyleEF-BookItalic          158.9
l_____BerkeleyOldstyleEF-Medium              158.10
l_____BerkeleyOldstyleEF-MediumItalic        158.11
l_____BerkeleyOldstyleEF-Bold                158.12
l_____BerkeleyOldstyleEF-BoldItalic          158.13
l_____BerkeleyOldstyleEF-Black               158.14
l_____BerkeleyOldstyleEF-BlackItalic         158.15
l_____BoltEF                                 159.0
l_____CharterEF-Regular                      160.0
l_____CharterEF-RegularSC                    160.1
l_____CharterEF-RegularItalic                160.2
l_____CharterEF-RegularItalicOsF             160.3
l_____CharterEF-Bold                         160.4
l_____CharterEF-BoldSC                       160.5
l_____CharterEF-BoldItalic                   160.6
l_____CharterEF-BoldItalicOsF                160.7
l_____CharterEF-Black                        160.8
l_____CharterEF-BlackOsF                     160.9
l_____CharterEF-BlackItalic                  160.10
l_____CharterEF-BlackItalicOsF               160.11
l_____ElanEF-Book                            161.0
l_____ElanEF-BookSC                          161.1
l_____ElanEF-BookItalic                      161.2
l_____ElanEF-Medium                          161.3
l_____ElanEF-MediumSC                        161.4
l_____ElanEF-MediumItalic                    161.5
l_____ElanEF-Bold                            161.6
l_____ElanEF-BoldItalic                      161.7
l_____ElanEF-Black                           161.8
l_____ElanEF-BlackItalic                     161.9
l_____ItcEras-Light                          162.0
l_____ItcEras-Book                           162.1
l_____ItcEras-Medium                         162.2
l_____ItcEras-Demi                           162.3
l_____ItcEras-Bold                           162.4
l_____ItcEras-Ultra                          162.5
l_____ItcErasEF-Light                        162.6
l_____ItcErasEF-Book                         162.7

l_____ItcErasEF-Medium                       162.8
l_____ItcErasEF-Demi                         162.9
l_____ItcErasEF-Bold                         162.10
l_____ItcErasEF-Ultra                        162.11
l_____Fenice-Light                           163.0
l_____Fenice-LightOblique                    163.1
l_____Fenice-Regular                         163.2
l_____Fenice-RegularOblique                  163.3
l_____Fenice-Bold                            163.4
l_____Fenice-BoldOblique                     163.5
l_____Fenice-Ultra                           163.6
l_____Fenice-UltraOblique                    163.7
l_____FeniceEF-Light                         163.8
l_____FeniceEF-LightItalic                   163.9
l_____FeniceEF-Regular                       163.10
l_____FeniceEF-RegularItalic                 163.11
l_____FeniceEF-Bold                          163.12
l_____FeniceEF-BoldItalic                    163.13
l_____FeniceEF-Ultra                         163.14
l_____FeniceEF-UltraItalic                   163.15
l_____FrizQuadrata                           164.0
l_____FrizQuadrata-Bold                      164.1
l_____FrizQuadrataEF-Regular                 164.2
l_____FrizQuadrataEF-RegularSC               164.3
l_____FrizQuadrataEF-RegularItalic           164.4
l_____FrizQuadrataEF-Bold                    164.5
l_____FrizQuadrataEF-BoldItalic              164.6
l_____Giovanni-Book                          165.0
l_____Giovanni-BookItalic                    165.1
l_____Giovanni-Bold                          165.2
l_____Giovanni-BoldItalic                    165.3
l_____Giovanni-Black                         165.4
l_____Giovanni-BlackItalic                   165.5
l_____GiovanniEF-Book                        165.6
l_____GiovanniEF-BookSC                      165.7
l_____GiovanniEF-BookItalic                  165.8
l_____GiovanniEF-Bold                        165.9
l_____GiovanniEF-BoldSC                      165.10
l_____GiovanniEF-BoldItalic                  165.11
l_____GiovanniEF-Black                       165.12
l_____GiovanniEF-BlackItalic                 165.13
l_____Highlander-Book                        166.0
l_____Highlander-BookItalic                  166.1
l_____Highlander-Medium                      166.2
l_____Highlander-MediumItalic                166.3
l_____Highlander-Bold                        166.4
l_____Highlander-BoldItalic                  166.5

l_____HighlanderEF-Book                      166.6
l_____HighlanderEF-BookSC                    166.7
l_____HighlanderEF-BookItalic                166.8
l_____HighlanderEF-Medium                    166.9
l_____HighlanderEF-MediumSC                  166.10
l_____HighlanderEF-MediumItalic              166.11
l_____HighlanderEF-Bold                      166.12
l_____HighlanderEF-BoldItalic                166.13
l_____Isadora-Roman                          167.0
l_____Isadora-Bold                           167.1
l_____IsadoraEF-Regular                      167.2
l_____IsadoraEF-Bold                         167.3
l_____JamilleEF-Book                         168.0
l_____JamilleSCEF-Book                       168.1
l_____JamilleEF-BookItalic                   168.2
l_____JamilleEF-Bold                         168.3
l_____JamilleSCEF-Bold                       168.4
l_____JamilleEF-BoldItalic                   168.5
l_____JamilleEF-Black                        168.6
l_____JamilleEF-BlackItalic                  168.7
l_____Leawood-Book                           169.0
l_____Leawood-BookItalic                     169.1
l_____Leawood-Medium                         169.2
l_____Leawood-MediumItalic                   169.3
l_____Leawood-Bold                           169.4
l_____Leawood-BoldItalic                     169.5
l_____Leawood-Black                          169.6
l_____Leawood-BlackItalic                    169.7
l_____LeawoodEF-Book                         169.8
l_____LeawoodEF-BookSC                       169.9
l_____LeawoodEF-BookItalic                   169.10
l_____LeawoodEF-Medium                       169.11
l_____LeawoodEF-MediumSC                     169.12
l_____LeawoodEF-MediumItalic                 169.13
l_____LeawoodEF-Bold                         169.14
l_____LeawoodEF-BoldItalic                   169.15
l_____LeawoodEF-Black                        169.16
l_____LeawoodEF-BlackItalic                  169.17
l_____LubalinGraph-Book                      170.0
l_____LubalinGraph-BookOblique               170.1
l_____LubalinGraph-Demi                      170.2
l_____LubalinGraph-DemiOblique               170.3
l_____LubalinGraphEF-ExtraLight              170.4
l_____LubalinGraphEF-ExtraLightObl           170.5
l_____LubalinGraphEF-Book                    170.6
l_____LubalinGraphEF-BookOblique             170.7
l_____LubalinGraphEF-Medium                  170.8

l_____LubalinGraphEF-MediumOblique           170.9
l_____LubalinGraphEF-Demi                    170.10
l_____LubalinGraphEF-DemiOblique             170.11
l_____LubalinGraphEF-Bold                    170.12
l_____LubalinGraphEF-BoldOblique             170.13
l_____LubalinGraphEF-BookCond                170.14
l_____LubalinGraphEF-BookCondSC              170.15
l_____LubalinGraphEF-BookCondObl             170.16
l_____LubalinGraphEF-MediumCond              170.17
l_____LubalinGraphEF-MediumCondSC            170.18
l_____LubalinGraphEF-MediumCondObl           170.19
l_____LubalinGraphEF-DemiCond                170.20
l_____LubalinGraphEF-DemiCondObl             170.21
l_____LubalinGraphEF-BoldCond                170.22
l_____LubalinGraphEF-BoldCondObl             170.23
l_____MixageEF-Book                          171.0
l_____MixageSCEF-Book                        171.1
l_____MixageEF-BookItalic                    171.2
l_____MixageEF-Medium                        171.3
l_____MixageSCEF-Medium                      171.4
l_____MixageEF-MediumItalic                  171.5
l_____MixageEF-Bold                          171.6
l_____MixageEF-BoldItalic                    171.7
l_____MixageEF-Black                         171.8
l_____MixageEF-BlackItalic                   171.9
l_____MotterCorpusEF-Regular                 172.0
l_____MotterCorpusEF-RegularCondensed        172.1
l_____NewtextEF-Light                        173.0
l_____NewtextEF-LightItalic                  173.1
l_____NewtextEF-Book                         173.2
l_____NewtextEF-BookItalic                   173.3
l_____NewtextEF-Regular                      173.4
l_____NewtextEF-RegularItalic                173.5
l_____NewtextEF-Demi                         173.6
l_____NewtextEF-DemiItalic                   173.7
l_____Ozwald                                 174.0
l_____OzwaldEF                               174.1
l_____PanacheEF-Book                         175.0
l_____PanacheSCEF-Book                       175.1
l_____PanacheEF-BookItalic                   175.2
l_____PanacheEF-Bold                         175.3
l_____PanacheSCEF-Bold                       175.4
l_____PanacheEF-BoldItalic                   175.5
l_____PanacheEF-Black                        175.6
l_____PanacheEF-BlackItalic                  175.7
l_____Quorum-Light                           176.0
l_____Quorum-Book                            176.1

l_____Quorum-Medium                          176.2
l_____Quorum-Bold                            176.3
l_____Quorum-Black                           176.4
l_____QuorumEF-Light                         176.5
l_____QuorumEF-Book                          176.6
l_____QuorumEF-Medium                        176.7
l_____QuorumEF-Bold                          176.8
l_____QuorumEF-Black                         176.9
l_____RondaEF                                177.0
l_____Slimbach-Book                          178.0
l_____Slimbach-BookItalic                    178.1
l_____Slimbach-Medium                        178.2
l_____Slimbach-MediumItalic                  178.3
l_____Slimbach-Bold                          178.4
l_____Slimbach-BoldItalic                    178.5
l_____Slimbach-Black                         178.6
l_____Slimbach-BlackItalic                   178.7
l_____SlimbachEF-Book                        178.8
l_____SlimbachEF-BookSC                      178.9
l_____SlimbachEF-BookItalic                  178.10
l_____SlimbachEF-Medium                      178.11
l_____SlimbachEF-MediumSC                    178.12
l_____SlimbachEF-MediumItalic                178.13
l_____SlimbachEF-Bold                        178.14
l_____SlimbachEF-BoldItalic                  178.15
l_____SlimbachEF-Black                       178.16
l_____SlimbachEF-BlackItalic                 178.17
l_____StudioScriptEF                         179.0
l_____StudioScriptEF-Alternative             179.1
l_____SyndorEF-Book                          180.0
l_____SyndorEF-BookSC                        180.1
l_____SyndorEF-BookItalic                    180.2
l_____SyndorEF-Medium                        180.3
l_____SyndorEF-MediumSC                      180.4
l_____SyndorEF-MediumItalic                  180.5
l_____SyndorEF-Bold                          180.6
l_____SyndorEF-BoldItalic                    180.7
l_____Tiffany                                181.0
l_____Tiffany-Italic                         181.1
l_____Tiffany-Demi                           181.2
l_____Tiffany-DemiItalic                     181.3
l_____Tiffany-Heavy                          181.4
l_____Tiffany-HeavyItalic                    181.5
l_____TiffanyEF-Light                        181.6
l_____TiffanyEF-LightItalic                  181.7
l_____TiffanyEF-Medium                       181.8
l_____TiffanyEF-MediumItalic                 181.9

l_____TiffanyEF-Demi                         181.10
l_____TiffanyEF-DemiItalic                   181.11
l_____TiffanyEF-Heavy                        181.12
l_____TiffanyEF-HeavyItalic                  181.13
l_____Weidemann-Book                         182.0
l_____Weidemann-BookItalic                   182.1
l_____Weidemann-Medium                       182.2
l_____Weidemann-MediumItalic                 182.3
l_____Weidemann-Bold                         182.4
l_____Weidemann-BoldItalic                   182.5
l_____Weidemann-Black                        182.6
l_____Weidemann-BlackItalic                  182.7
l_____WeidemannEF-Book                       182.8
l_____WeidemannEF-BookSC                     182.9
l_____WeidemannEF-BookItalic                 182.10
l_____WeidemannEF-Medium                     182.11
l_____WeidemannEF-MediumSC                   182.12
l_____WeidemannEF-MediumItalic               182.13
l_____WeidemannEF-Bold                       182.14
l_____WeidemannEF-BoldItalic                 182.15
l_____WeidemannEF-Black                      182.16
l_____WeidemannEF-BlackItalic                182.17
l_____ManhattanEF                            183.0
l_____Imago-Light                            184.0
l_____Imago-LightItalic                      184.1
l_____Imago-Book                             184.2
l_____Imago-BookItalic                       184.3
l_____Imago-Medium                           184.4
l_____Imago-MediumItalic                     184.5
l_____Imago-ExtraBold                        184.6
l_____Imago-ExtraBoldItalic                  184.7
l_____ImpactaLL                              185.0
l_____Impressum-Roman                        186.0
l_____Impressum-Italic                       186.1
l_____Impressum-Bold                         186.2
l_____ImpressumEF-Light                      186.3
l_____ImpressumEF-LightItalic                186.4
l_____ImpressumEF-Medium                     186.5
l_____IndusLL-Normal                         187.0
l_____IndusLL-Bold                           187.1
l_____Industria-Solid                        188.0
l_____Industria-SolidA                       188.1
l_____Industria-Inline                       188.2
l_____Industria-InlineA                      188.3
l_____InflexMT-Bold                          189.0
l_____Ironwood                               190.0
l_____ItalianOldStyleMT-Regular              191.0

l_____ItalianOldStyleMT-Italic               191.1
l_____ItalianOldStyleMT-Bold                 191.2
l_____ItalianOldStyleMT-BoldItalic           191.3
l_____JensonOldStyleEF-BoldCond              192.0
l_____Jiffy                                  193.0
l_____JoannaMT                               194.0
l_____JoannaMT-Italic                        194.1
l_____JoannaMT-SemiBold                      194.2
l_____JoannaMT-SemiBoldItalic                194.3
l_____JoannaMT-Bold                          194.4
l_____JoannaMT-BoldItalic                    194.5
l_____JoannaMT-ExtraBold                     194.6
l_____JuliaScriptEF                          195.0
l_____Juniper                                196.0
l_____KaffeesatzEF-Schwarz                   197.0
l_____KaffeesatzEF-Suess                     197.1
l_____KaffeesatzEF-Weiss                     197.2
l_____KalligraphiaEF                         198.0
l_____Kaufmann                               199.0
l_____Kaufmann-Bold                          199.1
l_____KaufmannEF-Bold                        199.2
l_____Kigali-Roman                           200.0
l_____Kigali-SxItalic                        200.1
l_____Kigali-LxRoman                         200.2
l_____Kigali-LxItalic                        200.3
l_____Kigali-ZigZag                          200.4
l_____Kigali-Block                           200.5
l_____Kigali-SxRoman                         200.6
l_____Kigali-Italic                          200.7
l_____KinoMT                                 201.0
l_____Kismet                                 202.0
l_____KnightsbridgeEF                        203.0
l_____KnightsbridgeEF-Alternate              203.1
l_____KochAntiqua                            204.0
l_____Kompakt                                205.0
l_____KuenstlerScript-Medium                 206.0
l_____KuenstlerScript-TwoBold                206.1
l_____KuenstlerScript-Black                  206.2
l_____KursivschriftStehHaar                  207.0
l_____KursivschriftLiegHaar                  207.1
l_____KursivschriftStehend                   207.2
l_____KursivschriftLiegend                   207.3
l_____KursivschriftRueckLie                  207.4
l_____LakeInformal                           208.0
l_____LakeInformalBold                       208.1
l_____LatienneEF-Roman                       209.0
l_____LatienneEF-RomanSC                     209.1

l_____LatienneEF-Italic                      209.2
l_____LatienneEF-ItalicSC                    209.3
l_____LatienneEF-Medium                      209.4
l_____LatienneEF-MediumSC                    209.5
l_____LatienneEF-MediumItalic                209.6
l_____LatienneEF-Bold                        209.7
l_____LatienneEF-BoldItalic                  209.8
l_____LatienneEF-ItalicSwC                   209.9
l_____LatienneEF-MediumItalicSwC             209.10
l_____LatienneEF-BoldItalicSwC               209.11
l_____LatinMT-Condensed                      210.0
l_____LazybonesEF                            211.0
l_____LeGriffeEF                             212.0
l_____LeGriffeEF-AlternateOne                212.1
l_____LeGriffeEF-AlternateTwo                212.2
l_____LetterGothic                           213.0
l_____LetterGothic-Slanted                   213.1
l_____LetterGothic-Bold                      213.2
l_____LetterGothic-BoldSlanted               213.3
l_____LibertyEF                              214.0
l_____Life-Roman                             215.0
l_____Life-Italic                            215.1
l_____Life-Bold                              215.2
l_____LifeEF-Regular                         215.3
l_____LifeEF-RegularItalic                   215.4
l_____LifeEF-Bold                            215.5
l_____LightlineGothicEF                      216.0
l_____LindsayEF                              217.0
l_____Linoscript                             219.0
l_____Linotext                               220.0
l_____Linotext-Dfr                           220.1
lpi___LAstrologyPi-One                       221.0
lpi___LAstrologyPi-Two                       221.1
l_____Centennial-Light                       222.0
l_____Centennial-LightSC                     222.1
l_____Centennial-LightItalic                 222.2
l_____Centennial-LightItalicOsF              222.3
l_____Centennial-Roman                       222.4
l_____Centennial-RomanSC                     222.5
l_____Centennial-Italic                      222.6
l_____Centennial-ItalicOsF                   222.7
l_____Centennial-Bold                        222.8
l_____Centennial-BoldOsF                     222.9
l_____Centennial-BoldItalic                  222.10
l_____Centennial-BoldItalicOsF               222.11
l_____Centennial-Black                       222.12
l_____Centennial-BlackOsF                    222.13

l_____Centennial-BlackItalic                 222.14
l_____Centennial-BlackItalicOsF              222.15
l_____CentennialCE-Light                     222.16
l_____CentennialCE-LightItalic               222.17
l_____CentennialCE-Roman                     222.18
l_____CentennialCE-Italic                    222.19
l_____CentennialCE-Bold                      222.20
l_____CentennialCE-BoldItalic                222.21
l_____CentennialCE-Black                     222.22
l_____CentennialCE-BlackItalic               222.23
lpi___LDecorationPi-One                      223.0
lpi___LDecorationPi-Two                      223.1
l_____DidotLH-Roman                          224.0
l_____DidotLH-RomanSC                        224.1
l_____DidotLH-RomanOsF                       224.2
l_____DidotLH-Italic                         224.3
l_____DidotLH-ItalicOsF                      224.4
l_____DidotLH-Bold                           224.5
l_____DidotLH-BoldOsF                        224.6
l_____DidotLH-MediumInitials                 224.7
l_____DidotLH-HeadlineRoman                  224.8
l_____DidotLH-HeadlineRomanOsF               224.9
l_____DidotLH-OrnamentsOne                   224.10
l_____DidotLH-OrnamentsTwo                   224.11
lpi___LGamePi-ChessDraughts                  225.0
lpi___LGamePi-DiceDominoes                   225.1
lpi___LGamePi-EnglishCards                   225.2
lpi___LGamePi-FrenchCards                    225.3
lpi___LHolidayPi-One                         227.0
lpi___LHolidayPi-Two                         227.1
lpi___LHolidayPi-Three                       227.2
lpi___TextilLH-PiOne                         228.0
lpi___TextilLH-PiTwo                         228.1
lpi___WarningLH-Pi                           229.0
l_____LiteraEF-Light                         230.0
l_____LiteraEF-Regular                       230.1
l_____LiteraEF-Medium                        230.2
l_____LiteraEF-Heavy                         230.3
l_____Lithos-ExtraLight                      231.0
l_____Lithos-Light                           231.1
l_____Lithos-Regular                         231.2
l_____Lithos-Bold                            231.3
l_____Lithos-Black                           231.4
l_____LuMarcLL-Roman                         232.0
l_____LuMarcLL-Bold                          232.1
l_____Macbeth                                234.0
l_____Madame-Letters                         235.0

l_____Madame-Accents                         235.1
l_____Madame-Numericals                      235.2
l_____MarconiEF-Book                         236.0
l_____MarconiSCEF-Book                       236.1
l_____MarconiEF-BookItalic                   236.2
l_____MarconiEF-SemiBold                     236.3
l_____MarconiEF-SemiBoldItalic               236.4
l_____Maximus                                237.0
l_____Melior                                 238.0
l_____Melior-Italic                          238.1
l_____Melior-Bold                            238.2
l_____Melior-BoldItalic                      238.3
l_____Meridien-Roman                         239.0
l_____Meridien-Italic                        239.1
l_____Meridien-Medium                        239.2
l_____Meridien-MediumItalic                  239.3
l_____Meridien-Bold                          239.4
l_____Meridien-BoldItalic                    239.5
l_____MerlinLL                               240.0
l_____MiamiEF                                241.0
l_____MicrogrammaEF-MediumExtend             242.0
l_____MicrogrammaEF-BoldExtend               242.1
l_____MonantiEF-Regular                      243.0
l_____MonantiEF-SemiBold                     243.1
l_____MonolineScriptMT                       244.0
l_____MurrayHillEF                           246.0
l_____Myriad-Sketch                          247.0
l_____Myriad-Headline                        247.1
l_____Myriad-Tilt                            247.2
l_____Mythos                                 248.0
lpi___NationalCodes-Europe                   249.0
lpi___NationalCodes-Africa                   249.1
lpi___NationalCodes-America                  249.2
lpi___NationalCodes-AsiaPlus                 249.3
lpi___NationalCodes-Universal                249.4
l_____NeueHammerUnziale-One                  250.0
l_____NeueHammerUnziale-OneAlt               250.1
l_____NeueHammerUnziale-TwoAlt               250.2
l_____NeueHammerUnziale-Two                  250.3
l_____NeuzeitSBook                           251.0
l_____NeuzeitSBook-Heavy                     251.1
l_____NevisionCasualEF                       252.0
l_____NewBerolinaMT                          253.0
l_____NewsGothic                             254.0
l_____NewsGothic-Oblique                     254.1
l_____NewsGothic-Bold                        254.2
l_____NewsGothic-BoldOblique                 254.3

l_____NicolasCochinEF-Regular                255.0
l_____NicolasCochinEF-RegularIta             255.1
l_____NicolasCochinEF-Black                  255.2
l_____NikisEF-Light                          256.0
l_____NikisEF-LightItalic                    256.1
l_____NikisEF-SemiBold                       256.2
l_____NikisEF-SemiBoldItalic                 256.3
l_____Nofret-Light                           257.0
l_____Nofret-LightSC                         257.1
l_____Nofret-LightExpert                     257.2
l_____Nofret-LightItalic                     257.3
l_____Nofret-LightItalicOsF                  257.4
l_____Nofret-LightItalicExpert               257.5
l_____Nofret-Regular                         257.6
l_____Nofret-RegularSC                       257.7
l_____Nofret-RegularExpert                   257.8
l_____Nofret-Italic                          257.9
l_____Nofret-ItalicOsF                       257.10
l_____Nofret-ItalicExpert                    257.11
l_____Nofret-Medium                          257.12
l_____Nofret-MediumSC                        257.13
l_____Nofret-MediumExpert                    257.14
l_____Nofret-MediumItalic                    257.15
l_____Nofret-MediumItalicOsF                 257.16
l_____Nofret-MediumItalicExpert              257.17
l_____Nofret-Bold                            257.18
l_____Nofret-BoldOsF                         257.19
l_____Nofret-BoldExpert                      257.20
l_____Nofret-BoldItalic                      257.21
l_____Nofret-BoldItalicOsF                   257.22
l_____Nofret-BoldItalicExpert                257.23
l_____NotreDame-Roman                        258.0
l_____NotreDame-RomanDfr                     258.1
l_____NotreDame-Ornaments                    258.2
l_____NuptialScript                          259.0
l_____OCRA                                   260.0
l_____OCRB                                   260.1
l_____OCRA-A                                 260.2
l_____OCRB-A                                 260.3
l_____OCRBEF-Regular                         260.4
l_____OctavianMT                             261.0
l_____OctavianMT-SC                          261.1
l_____OctavianMT-Expert                      261.2
l_____OctavianMT-Italic                      261.3
l_____OctavianMT-ItalicOsF                   261.4
l_____OctavianMT-ItalicExpert                261.5
l_____OdinEF                                 262.0

l_____OldEnglishEF                           263.0
l_____OldStyleSeven                          264.0
l_____OldStyleSeven-SC                       264.1
l_____OldStyleSeven-Italic                   264.2
l_____OldStyleSeven-ItalicOsF                264.3
l_____MonotypeOldStyleMT-BoldOut             264.4
l_____OldTypewriterEF-ExLight                265.0
l_____OldTypewriterEF-ExLightObl             265.1
l_____OldTypewriterEF-Light                  265.2
l_____OldTypewriterEF-LightObl               265.3
l_____OldTypewriterEF-Regular                265.4
l_____OldTypewriterEF-RegularA               265.5
l_____OldTypewriterEF-RegularObl             265.6
l_____OldTypewriterEF-Bold                   265.7
l_____OldTypewriterEF-BoldObl                265.8
l_____OldTypewriterEF-ExBold                 265.9
l_____OldTypewriterEF-ExBoldObl              265.10
l_____OlympiaEF-Light                        266.1
l_____OlympiaEF-LightOblique                 266.2
l_____OlympiaEF-SemiBold                     266.3
l_____OlympiaEF-SemiBoldOblique              266.4
l_____Olympian                               267.0
l_____Olympian-Italic                        267.1
l_____Olympian-Bold                          267.2
l_____Olympian-BoldItalic                    267.3
l_____Omnia                                  268.0
l_____OndineEF                               269.1
l_____OnyxMT                                 270.0
l_____Orator                                 271.0
l_____Orator-Slanted                         271.1
l_____Oxford                                 272.0
l_____Caecilia-Light                         273.0
l_____Caecilia-LightSC                       273.1
l_____Caecilia-LightOsF                      273.2
l_____Caecilia-LightItalic                   273.3
l_____Caecilia-LightItalicSC                 273.4
l_____Caecilia-LightItalicOsF                273.5
l_____Caecilia-Roman                         273.6
l_____Caecilia-RomanSC                       273.7
l_____Caecilia-RomanOsF                      273.8
l_____Caecilia-Italic                        273.9
l_____Caecilia-ItalicSC                      273.10
l_____Caecilia-ItalicOsF                     273.11
l_____Caecilia-Bold                          273.12
l_____Caecilia-BoldSC                        273.13
l_____Caecilia-BoldOsF                       273.14
l_____Caecilia-BoldItalic                    273.15

l_____Caecilia-BoldItalicSC                  273.16
l_____Caecilia-BoldItalicOsF                 273.17
l_____Caecilia-Heavy                         273.18
l_____Caecilia-HeavySC                       273.19
l_____Caecilia-HeavyOsF                      273.20
l_____Caecilia-HeavyItalic                   273.21
l_____Caecilia-HeavyItalicSC                 273.22
l_____Caecilia-HeavyItalicOsF                273.23
l_____CaeciliaCE-Light                       273.24
l_____CaeciliaCE-LightItalic                 273.25
l_____CaeciliaCE-Roman                       273.26
l_____CaeciliaCE-Italic                      273.27
l_____CaeciliaCE-Bold                        273.28
l_____CaeciliaCE-BoldItalic                  273.29
l_____CaeciliaCE-Heavy                       273.30
l_____CaeciliaCE-HeavyItalic                 273.31
l_____PalaceScriptMT-Regular                 274.0
l_____PalaceScriptMT-Bold                    274.1
l_____Palatino-Light                         275.0
l_____Palatino-LightItalic                   275.1
l_____Palatino-Roman                         275.2
l_____Palatino-SC                            275.3
l_____Palatino-Italic                        275.4
l_____Palatino-ItalicOsF                     275.5
l_____Palatino-Medium                        275.6
l_____Palatino-MediumItalic                  275.7
l_____Palatino-Bold                          275.8
l_____Palatino-BoldOsF                       275.9
l_____Palatino-BoldItalic                    275.10
l_____Palatino-BoldItalicOsF                 275.11
l_____Palatino-Black                         275.12
l_____Palatino-BlackItalic                   275.13
l_____PalatinoCE-Roman                       275.14
l_____PalatinoCE-Italic                      275.15
l_____PalatinoCE-Bold                        275.16
l_____PalatinoCE-BoldItalic                  275.17
l_____PaletteEF                              276.0
l_____Parisian                               277.0
l_____ParkAvenue                             278.0
l_____ParkAvenueEF                           278.1
l_____Pelican                                279.0
l_____PepitaMT                               281.0
l_____Pepperwood-Regular                     282.0
l_____Pepperwood-Fill                        282.1
l_____Pepperwood-Outline                     282.2
l_____Perpetua                               283.0
l_____Perpetua-SC                            283.1

l_____Perpetua-Expert                        283.2
l_____Perpetua-Italic                        283.3
l_____Perpetua-ItalicOsF                     283.4
l_____Perpetua-ItalicExp                     283.5
l_____Perpetua-Bold                          283.6
l_____Perpetua-BoldOsF                       283.7
l_____Perpetua-BoldExp                       283.8
l_____Perpetua-BoldItalic                    283.9
l_____Perpetua-BoldItalicOsF                 283.10
l_____Perpetua-BoldItalicExp                 283.11
l_____PhotinaMT                              284.0
l_____PhotinaMT-Italic                       284.1
l_____PhotinaMT-SemiBold                     284.2
l_____PhotinaMT-SemiBoldItalic               284.3
l_____PhotinaMT-Bold                         284.4
l_____PhotinaMT-BoldItalic                   284.5
l_____PhotinaMT-UltraBold                    284.6
l_____PhotinaMT-UltraBoldItalic              284.7
l_____PhyllisEF                              285.0
l_____PhyllisEF-Initials                     285.1
l_____PioneerEF                              287.0
l_____PlazaEF-Regular                        288.0
l_____PlazaEF-Alternate                      288.1
l_____PlazaEF-Ultra                          288.2
l_____PlazaEF-Inline                         288.3
l_____PlazaEF-Decorative                     288.4
l_____Ponderosa                              289.0
l_____PrestigeElite                          290.0
l_____PrestigeElite-Slanted                  290.1
l_____PrestigeElite-Bold                     290.2
l_____PrestigeElite-BoldSlanted              290.3
l_____PromotorEF-Regular                     291.0
l_____QuentinEF                              292.0
l_____Raleigh                                293.0
l_____Raleigh-Medium                         293.1
l_____Raleigh-DemiBold                       293.2
l_____Raleigh-Bold                           293.3
l_____RenaultEF-Light                        294.0
l_____RenaultEF-LightItalic                  294.1
l_____RenaultEF-Bold                         294.2
l_____RenaultEF-BoldItalic                   294.3
l_____Revue                                  295.0
l_____RevueEF                                295.1
l_____RoemischStehendHar                     296.0
l_____RoemischLiegendHar                     296.1
l_____RoemischRueckLiegHar                   296.2
l_____RoemischStehend                        296.3

l_____RoemischLiegend                        296.4
l_____RoemischRueckLiegend                   296.5
l_____RomanaEF-Light                         297.0
l_____RomanaEF-Book                          297.1
l_____RomanaEF-Medium                        297.2
l_____RomanaEF-Demi                          297.3
l_____RomanaEF-Bold                          297.4
l_____RomanaEF-Ultra                         297.5
l_____Romic-Light                            298.0
l_____Romic-LightItalic                      298.1
l_____Romic-Medium                           298.2
l_____Romic-Bold                             298.3
l_____Romic-ExtraBold                        298.4
l_____RomicEF-Light                          298.5
l_____RomicEF-LightItalic                    298.6
l_____RomicEF-Medium                         298.7
l_____RomicEF-Bold                           298.8
l_____RomicEF-ExtraBold                      298.9
l_____Rundfunk-Antiqua                       299.0
l_____Rundfunk-Grotesk                       299.1
l_____Rusticana-Borders                      300.1
l_____RuzickaFreehandLH-Roman                301.0
l_____RuzickaFreehandLH-RomanSC              301.1
l_____RuzickaFreehandLH-Bold                 301.2
l_____RuzickaFreehandLH-BoldSC               301.3
l_____Salto                                  303.0
l_____SanMarco                               304.0
l_____SanMarco-Dfr                           304.1
l_____SchnitzLL                              305.0
l_____SchwenneNegroLL                        306.0
l_____SchwennelLilaLL                        306.1
l_____Scotch-Roman                           307.0
l_____Scotch-Italic                          307.1
l_____ScriptureEF                            308.0
l_____Serifa-Light                           309.0
l_____Serifa-LightItalic                     309.1
l_____Serifa-Roman                           309.2
l_____Serifa-Italic                          309.3
l_____Serifa-Bold                            309.4
l_____Serifa-Black                           309.5
l_____ShamrockEF                             310.0
l_____ShamrockEF-AlternateOne                310.1
l_____ShamrockEF-AlternateTwo                310.2
l_____Shannon-Book                           311.0
l_____Shannon-Oblique                        311.1
l_____Shannon-Bold                           311.2
l_____Shannon-ExtraBold                      311.3

l_____Shelley-AllegroScript                  312.0
l_____Shelley-AndanteScript                  312.1
l_____Shelley-VolanteScript                  312.2
l_____Sho-Roman                              313.0
l_____SinahRomanLL                           314.0
l_____SinahBoldLL                            314.1
l_____SinahBlackLL                           314.2
l_____SinaloaEF                              315.0
l_____SloganEF                               316.0
l_____Smaragd                                317.0
l_____Smaragd-Alternate                      318.0
l_____SnellRoundhand-Script                  318.1
l_____SnellRoundhand-BoldScript              318.2
l_____SnellRoundhand-BlackScript             318.3
l_____Sonata                                 319.0
l_____SpartanClassified-Book                 320.0
l_____SpartanClassified-Heavy                320.1
l_____SpectrumMT                             321.0
l_____SpectrumMT-SC                          321.1
l_____SpectrumMT-RomanExpert                 321.2
l_____SpectrumMT-Italic                      321.3
l_____SpectrumMT-ItalicOsF                   321.4
l_____SpectrumMT-ItalicExpert                321.5
l_____SpectrumMT-SemiBold                    321.6
l_____SpectrumMT-SemiBoldOsF                 321.7
l_____SpectrumMT-SemiBoldExpert              321.8
l_____SpiritEF-Light                         322.0
l_____SpiritEF-Regular                       322.1
l_____SpiritEF-Bold                          322.2
l_____Sprint                                 323.0
l_____SquireEF-Regular                       324.0
l_____SquireEF-Bold                          324.1
l_____Stencil                                325.0
l_____StencilEF                              325.1
l_____Stop                                   326.0
l_____StopEF                                 326.1
l_____Studz                                  327.0
l_____SuperstarEF-Shadow                     328.0
l_____SuperstarEF-Regular                    328.1
l_____SwiftEF-Light                          329.0
l_____SwiftEF-LightItalic                    329.1
l_____SwiftEF-Regular                        329.2
l_____SwiftSCEF-Regular                      329.3
l_____SwiftEF-RegularItalic                  329.4
l_____SwiftEF-Bold                           329.5
l_____SwiftEF-ExtraBold                      329.6
l_____SwiftEF-BoldCondensed                  329.7

l_____Syntax-Roman                           331.0
l_____Syntax-Italic                          331.1
l_____Syntax-Bold                            331.2
l_____Syntax-Black                           331.3
l_____Syntax-UtraBlack                       331.4
l_____TangoEF                                332.0
l_____TarragonEF                             333.0
l_____Tekton                                 334.0
l_____Tekton-Oblique                         334.1
l_____Tekton-Bold                            334.2
l_____Tekton-BoldOblique                     334.3
l_____Tempo-HeavyCondensed                   335.0
l_____Tempo-HeavyCondensedItalic             335.1
l_____ThorowgoodEF-Regular                   336.0
l_____ThorowgoodEF-RegularItalic             336.1
l_____ThunderbirdEF-ExtraCondensed           337.0
l_____ThunderbirdEF-Regular                  337.1
l_____TiloEF                                 338.0
l_____TimelessEF-Light                       339.0
l_____TimelessEF-LightItalic                 339.1
l_____TimelessEF-Bold                        339.2
l_____TitusEF                                340.0
l_____Toolbox                                341.0
l_____TopographischeZahlentafel              342.0
l_____Trajan-Regular                         343.0
l_____Trajan-Bold                            343.1
l_____TradeGothicLH-BoldExtended             343.2
l_____TrumpMediaeval-Roman                   346.0
l_____TrumpMediaeval-RomanSC                 346.1
l_____TrumpMediaeval-Italic                  346.2
l_____TrumpMediaeval-ItalicOsF               346.3
l_____TrumpMediaeval-Bold                    346.4
l_____TrumpMediaeval-BoldOsF                 346.5
l_____TrumpMediaeval-BoldItalic              346.6
l_____TrumpMediaeval-BoldItalicOsF           346.7
l_____Umbra                                  348.0
l_____Univers-Light                          349.0
l_____Univers-LightOblique                   349.1
l_____Univers                                349.2
l_____Univers-Oblique                        349.3
l_____Univers-Bold                           349.4
l_____Univers-BoldOblique                    349.5
l_____Univers-Black                          349.7
l_____Univers-BlackOblique                   349.8
l_____Univers-ExtraBlack                     349.9
l_____Univers-ExtraBlackOblique              349.10
l_____Univers-CondensedLight                 349.11

l_____Univers-CondensedLightOblique          349.12
l_____Univers-Condensed                      349.13
l_____Univers-CondensedOblique               349.14
l_____Univers-CondensedBold                  349.15
l_____Univers-CondensedBoldOblique           349.16
l_____Univers-ThinUltraCondensed             349.17
l_____Univers-LightUltraCondensedTwo         349.18
l_____Univers-UltraCondensed                 349.19
l_____Univers-Extended                       349.20
l_____Univers-ExtendedObl                    349.21
l_____Univers-BoldExt                        349.22
l_____Univers-BoldExtObl                     349.23
l_____Univers-BlackExt                       349.24
l_____Univers-BlackExtObl                    349.25
l_____Univers-ExtraBlackExt                  349.26
l_____Univers-ExtraBlackExtObl               349.27
l_____UniversCE-Light                        349.28
l_____UniversCE-LightOblique                 349.29
l_____UniversCE-Medium                       349.30
l_____UniversCE-Oblique                      349.31
l_____UniversCE-Bold                         349.32
l_____UniversCE-BoldOblique                  349.33
l_____UniversCE-Black                        349.34
l_____UniversCE-BlackOblique                 349.35
lpi___MathematicalPi-One                     350.0
lpi___MathematicalPi-Two                     350.1
lpi___MathematicalPi-Three                   350.2
lpi___MathematicalPi-Four                    350.3
lpi___MathematicalPi-Five                    350.4
lpi___MathematicalPi-Six                     350.5
lpi___Unversal-GreekwithMathPi               351.0
lpi___MICR                                   351.1
lpi___Unversal-NewswithCommPi                351.2
l_____University-Roman                       352.0
l_____UniversityEF-Roman                     352.1
l_____UniversityEF-RomanItalic               352.2
l_____UniversityEF-Bold                      352.3
l_____VAGRounded-Thin                        353.0
l_____VAGRounded-Light                       353.1
l_____VAGRounded-Bold                        353.2
l_____VAGRounded-Black                       353.3
l_____VAGRundschriftEF                       353.4
l_____VanDijkEF                              354.0
l_____VictorianEF                            355.0
l_____VivaldiEF                              356.0
l_____Walbaum-Roman                          357.0
l_____Walbaum-RomanSC                        357.1

l_____Walbaum-RomanOsF                       357.2
l_____Walbaum-Italic                         357.3
l_____Walbaum-ItalicOsF                      357.4
l_____Walbaum-Bold                           357.5
l_____Walbaum-BoldOsF                        357.6
l_____Walbaum-BoldItalic                     357.7
l_____Walbaum-BoldItalicOsF                  357.8
l_____WalbaumEF-Regular                      357.9
l_____WalbaumEF-RegularItalic                357.10
l_____WalbaumEF-Medium                       357.11
l_____WalbaumEF-MediumItalic                 357.12
l_____WalbaumEF-Bold                         357.13
l_____WalbaumEF-BoldItalic                   357.14
l_____WalbaumBook-Regular                    357.15
l_____WalbaumBook-RegularSC                  357.16
l_____WalbaumBook-RegularExpert              357.17
l_____WalbaumBook-Italic                     357.18
l_____WalbaumBook-ItalicOsF                  357.19
l_____WalbaumBook-ItalicExpert               357.20
l_____WalbaumBook-Medium                     357.21
l_____WalbaumBook-MediumSC                   357.22
l_____WalbaumBook-MdExpert                   357.23
l_____WalbaumBook-MediumItalic               357.24
l_____WalbaumBook-MdItalicOsF                357.25
l_____WalbaumBook-MdItalicExpert             357.26
l_____WalbaumBook-Bold                       357.27
l_____WalbaumBook-BoldOsF                    357.28
l_____WalbaumBook-BoldExpert                 357.29
l_____WalbaumBook-BoldItalic                 357.30
l_____WalbaumBook-BoldItalicOsF              357.31
l_____WalbaumBook-BoldItalicExpert           357.32
l_____WalbaumFrakturEF                       357.33
l_____WhyNotEF                               358.0
l_____WilhelmKlingsporGotisch                359.0
l_____WilhelmKlingsporGotisch-Dfr            359.1
l_____WindsorEF-Light                        360.0
l_____WindsorEF-Bold                         360.1
l_____WindsorEF-UltraHeavy                   360.2
l_____WindsorEF-BoldOutline                  360.3
l_____WindsorEF-LightCondensed               360.4
l_____WindsorEF-ExtraBoldCondensed           360.5
l_____WindsorEF-Elongated                    360.6
ppi___WoodtypeOrnaments-One                  361.0
ppi___WoodtypeOrnaments-Two                  361.1
l_____ZipperEF                               363.0
l_____BelshawEF                              364.0
l_____Belwe-Light                            365.0

l_____Belwe-Medium                           365.1
l_____Belwe-Bold                             365.2
l_____Belwe-Condensed                        365.3
l_____BelweEF-Light                          365.4
l_____BelweEF-Italic                         365.5
l_____BelweEF-Medium                         365.6
l_____BelweEF-Bold                           365.7
l_____BelweEF-BoldCondensed                  365.8
l_____Bembo                                  366.0
l_____Bembo-SC                               366.1
l_____Bembo-Expert                           366.2
l_____Bembo-Italic                           366.3
l_____Bembo-ItalicOsF                        366.4
l_____Bembo-ItalicExpert                     366.5
l_____Bembo-Semibold                         366.6
l_____Bembo-SemiboldOsF                      366.7
l_____Bembo-SemiboldExpert                   366.8
l_____Bembo-SemiboldItalic                   366.9
l_____Bembo-SemiboldItalicOsF                366.10
l_____Bembo-SemiboldItalicExpert             366.11
l_____Bembo-Bold                             366.12
l_____Bembo-BoldOsF                          366.13
l_____Bembo-BoldExpert                       366.14
l_____Bembo-BoldItalic                       366.15
l_____Bembo-BoldItalicOsF                    366.16
l_____Bembo-BoldItalicExpert                 366.17
l_____Bembo-ExtraBold                        366.18
l_____Bembo-ExtraBoldOsF                     366.19
l_____Bembo-ExtraBoldExpert                  366.20
l_____Bembo-ExtraBoldItalic                  366.21
l_____Bembo-ExtraBoldItalicOsF               366.22
l_____Bembo-ExtraBoldItalicExpert            366.23
l_____BerlinerGrotesk-Light                  368.0
l_____BerlinerGrotesk-Medium                 368.1
l_____LoType-Light                           369.0
l_____LoType-Regular                         369.1
l_____LoType-Medium                          369.2
l_____LoType-MediumItalic                    369.3
l_____LoType-Bold                            369.4
l_____LoType-MediumCondensed                 369.5
l_____BiffoMT                                370.0
l_____BigbandTerrazzo                        371.1
l_____Birch                                  372.0
l_____BirdsEF                                373.0
l_____BlockBE-Regular                        375.0
l_____BlockBE-Italic                         375.1
l_____BlockBE-Heavy                          375.2

l_____BlockBE-Condensed                      375.3
l_____BlockBE-ExtraCn                        375.4
l_____BlockBE-ExtraCnItalic                  375.5
l_____BoberiaLL-Light                        376.0
l_____BoberiaLL-LightItalic                  376.1
l_____BoberiaLL-SemiBold                     376.2
l_____Bodoni-Book                            377.0
l_____Bodoni-BookItalic                      377.1
l_____Bodoni                                 377.2
l_____Bodoni-Italic                          377.3
l_____Bodoni-Bold                            377.4
l_____Bodoni-BoldItalic                      377.5
l_____Bodoni-BoldCondensed                   377.6
l_____BodoniEF-Light                         377.7
l_____BodoniEF-LightIta                      377.8
l_____BodoniEF-Regular                       377.9
l_____BodoniEF-RegularIta                    377.10
l_____BodoniEF-DemiBold                      377.11
l_____BodoniEF-Bold                          377.12
l_____BodoniEF-DemiBoldIta                   377.13
l_____BodoniEF-BoldIta                       377.14
l_____BodoniEF-RegularCn                     377.15
l_____BodoniEF-RegularCnIta                  377.16
l_____BodoniEF-BoldCn                        377.17
l_____BodoniEF-DemiBoldCn                    377.18
l_____BodoniEF-BoldCnItalic                  377.19
l_____BodoniEF-DemiBoldCnIta                 377.20
l_____BodoniEF-Black                         377.21
l_____BodoniEF-BlackIta                      377.22
l_____BodoniEF-Ultra                         377.23
l_____BauerBodoni-Roman                      377.24
l_____BauerBodoni-RomanSC                    377.25
l_____BauerBodoni-RomanOsF                   377.26
l_____BauerBodoni-Italic                     377.27
l_____BauerBodoni-ItalicOsF                  377.28
l_____BauerBodoni-Bold                       377.29
l_____BauerBodoni-BoldOsF                    377.30
l_____BauerBodoni-BoldItalic                 377.31
l_____BauerBodoni-BoldItalicOsF              377.32
l_____BauerBodoni-Black                      377.33
l_____BauerBodoni-BlackItalic                377.34
l_____BauerBodoni-BoldCond                   377.35
l_____BauerBodoni-BlackCond                  377.36
l_____BauerBodoniCE-Roman                    377.37
l_____BauerBodoniCE-Italic                   377.38
l_____BauerBodoniCE-Bold                     377.39
l_____BauerBodoniCE-BoldItalic               377.40

l_____BauerBodoniCE-Black                    377.41
l_____BauerBodoniCE-BlackItalic              377.42
l_____BauerBodoniCE-BoldCondensed            377.43
l_____BauerBodoniCE-BlackCondensed           377.44
l_____BauerBodoniEF-Regular                  377.45
l_____BauerBodoniEF-RegularItalic            377.46
l_____BauerBodoniEF-Medium                   377.47
l_____BauerBodoniEF-MediumItalic             377.48
l_____BauerBodoniEF-Bold                     377.49
l_____BauerBodoniEF-BoldItalic               377.50
l_____BodoniBE-Light                         377.51
l_____BodoniBE-LightSC                       377.52
l_____BodoniBE-LightExpert                   377.53
l_____BodoniBE-LightItalic                   377.54
l_____BodoniBE-LightItalicOsF                377.55
l_____BodoniBE-LightItalicExpert             377.56
l_____BodoniBE-Regular                       377.57
l_____BodoniBE-RegularSC                     377.58
l_____BodoniBE-RegularExp                    377.59
l_____BodoniBE-Italic                        377.60
l_____BodoniBE-ItalicOsF                     377.61
l_____BodoniBE-ItalicExp                     377.62
l_____BodoniBE-Medium                        377.63
l_____BodoniBE-MediumSC                      377.64
l_____BodoniBE-MediumExp                     377.65
l_____BodoniBE-MediumItalic                  377.66
l_____BodoniBE-MediumItalicOsF               377.67
l_____BodoniBE-MediumItalicExp               377.68
l_____BodoniBE-Bold                          377.69
l_____BodoniBE-BoldOsF                       377.70
l_____BodoniBE-BoldExp                       377.71
l_____BodoniBE-BoldItalic                    377.72
l_____BodoniBE-BoldItalicOsF                 377.73
l_____BodoniBE-BoldItalicExp                 377.74
l_____BodoniBE-Condensed                     377.75
l_____BodoniBE-CnItalic                      377.76
l_____BodoniBE-MediumCn                      377.77
l_____BodoniBE-MediumCnItalic                377.78
l_____BodoniBE-BoldCn                        377.79
l_____BodoniBE-BoldCnItalic                  377.80
l_____Bodoni-Poster                          377.81
l_____Bodoni-PosterItalic                    377.82
l_____Bodoni-PosterCompressed                377.83
l_____BodoniSixEF-Book                       377.84
l_____BodoniSixEF-BookSC                     377.85
l_____BodoniSixEF-BookItalic                 377.86
l_____BodoniSixEF-Bold                       377.87

l_____BodoniSixEF-BoldItalic                 377.88
l_____BodoniTwelveEF-Book                    377.89
l_____BodoniTwelveEF-BookSC                  377.90
l_____BodoniTwelveEF-BookItalic              377.91
l_____BodoniTwelveEF-Bold                    377.92
l_____BodoniTwelveEF-BoldItalic              377.93
l_____BodoniSeventytwoEF-Book                377.94
l_____BodoniSeventytwoEF-BookSC              377.95
l_____BodoniSeventytwoEF-BookIta             377.96
l_____BodoniSeventytwoEF-Bold                377.97
l_____BodoniSeventytwoEF-BoldIta             377.98
l_____BodoniOldFaceBE-Regular                377.99
l_____BodoniOldFaceBE-RegularSC              377.100
l_____BodoniOldFaceBE-RegularExp             377.101
l_____BodoniOldFaceBE-Italic                 377.102
l_____BodoniOldFaceBE-ItalicSC               377.103
l_____BodoniOldFaceBE-ItalicExp              377.104
l_____BodoniOldFaceBE-Medium                 377.105
l_____BodoniOldFaceBE-MediumSC               377.106
l_____BodoniOldFaceBE-MediumExp              377.107
l_____BodoniOldFaceBE-MediumIt               377.108
l_____BodoniOldFaceBE-MediumItOsF            377.109
l_____BodoniOldFaceBE-MediumItExp            377.110
l_____BodoniOldFaceBE-Bold                   377.111
l_____BodoniOldFaceBE-BoldOsF                377.112
l_____BodoniOldFaceBE-BoldExp                377.113
l_____BodoniOldFaceBE-BoldIt                 377.114
l_____BodoniOldFaceBE-BoldItOsF              377.115
l_____BodoniOldFaceBE-BoldItExp              377.116
l_____BordeauxEF-Regular                     378.0
l_____BordeauxEF-RegularItalic               378.1
l_____BordeauxScriptEF                       378.2
l_____Boulevard                              380.0
l_____BritannicEF-ExtraLight                 382.0
l_____BritannicEF-Light                      382.1
l_____BritannicEF-Medium                     382.2
l_____BritannicEF-Bold                       382.3
l_____BritannicEF-Ultra                      382.4
l_____BrodyEF                                383.0
l_____BronxEF                                384.0
l_____BrushScript                            385.0
l_____BrushScriptEF                          385.1
l_____BuxomEF                                386.0
l_____CabaretEF                              387.0
l_____CalvertMT-Light                        389.0
l_____CalvertMT-Regular                      389.1
l_____CalvertMT-Bold                         389.2

l_____CamelliaEF                             390.0
l_____Candida-Roman                          391.0
l_____Candida-Italic                         391.1
l_____Candida-Bold                           391.2
l_____CandidaEF-Regular                      391.3
l_____CandidaEF-RegularItalic                391.4
l_____CandidaEF-Medium                       391.5
l_____CandidaEF-RegularCondensed             391.6
l_____CandidaEF-MediumCondensed              391.7
l_____CantoriaMT-Light                       392.0
l_____CantoriaMT-LightItalic                 392.1
l_____CantoriaMT-Roman                       392.2
l_____CantoriaMT-Italic                      392.3
l_____CantoriaMT-SemiBold                    392.4
l_____CantoriaMT-SemiBoldItalic              392.5
l_____CantoriaMT-Bold                        392.6
l_____CantoriaMT-BoldItalic                  392.7
l_____CantoriaMT-ExtraBold                   392.8
l_____CantoriaMT-ExtraBoldItalic             392.9
l_____Carolina                               393.0
l_____Carolina-Dfr                           393.1
l_____Carta                                  394.0
l_____CartoonScript-Roman                    395.0
l_____ACaslon-Regular                        396.0
l_____ACaslon-AltRegular                     396.1
l_____ACaslon-RegularSC                      396.2
l_____ACaslonExp-Regular                     396.3
l_____ACaslon-Italic                         396.4
l_____ACaslon-AltItalic                      396.5
l_____ACaslon-ItalicOsF                      396.6
l_____ACaslonExp-Italic                      396.7
l_____ACaslon-Semibold                       396.8
l_____ACaslon-AltSemibold                    396.9
l_____ACaslon-SemiboldSC                     396.10
l_____ACaslonExp-Semibold                    396.11
l_____ACaslon-SemiboldItalic                 396.12
l_____ACaslon-AltSemiboldItalic              396.13
l_____ACaslon-SemiboldItalicOsF              396.14
l_____ACaslonExp-SemiboldItalic              396.15
l_____ACaslon-Bold                           396.16
l_____ACaslon-AltBold                        396.17
l_____ACaslon-BoldOsF                        396.18
l_____ACaslonExp-Bold                        396.19
l_____ACaslon-BoldItalic                     396.20
l_____ACaslon-AltBoldItalic                  396.21
l_____ACaslon-BoldItalicOsF                  396.22
l_____ACaslonExp-BoldItalic                  396.23

l_____ACaslon-SwashItalic                    396.24
l_____ACaslon-SwashSemiboldItalic            396.25
l_____ACaslon-SwashBoldItalic                396.26
l_____ACaslon-Ornaments                      396.27
l_____CaslonThree-Roman                      396.28
l_____CaslonThree-RomanSC                    396.29
l_____CaslonThree-Italic                     396.30
l_____CaslonThree-ItalicOsF                  396.31
l_____CaslonFiveFourty-Roman                 396.32
l_____CaslonFiveFourty-RomanSC               396.33
l_____CaslonFiveFourty-Italic                396.34
l_____CaslonFiveFourty-ItalicOsF             396.35
l_____CaslonOpenFace                         396.36
l_____CaslonFiveFourtyEF-Italic              396.37
l_____CaslonBlackEF                          396.38
l_____CaslonAntiqueEF                        396.39
l_____CaslonGraphiqueEF                      396.40
l_____CaslonTwoTwentyFour-Book               396.41
l_____CaslonTwoTwentyFour-BookIt             396.42
l_____CaslonTwoTwentyFour-Medium             396.43
l_____CaslonTwoTwentyFour-MediumIt           396.44
l_____CaslonTwoTwentyFour-Bold               396.45
l_____CaslonTwoTwentyFour-BoldIt             396.46
l_____CaslonTwoTwentyFour-Black              396.47
l_____CaslonTwoTwentyFour-BlackIt            396.48
l_____CaslonTwoTwentyFourEF-Book             396.49
l_____CaslonTwoTwentyFourEF-BookIt           396.50
l_____CaslonTwoTwentyFourEF-Medium           396.51
l_____CaslonTwoTwentyFourEF-MediumIt         396.52
l_____CaslonTwoTwentyFourEF-Bold             396.53
l_____CaslonTwoTwentyFourEF-BoldI            396.54
l_____CaslonTwoTwentyFourEF-Black            396.55
l_____CaslonTwoTwentyFourEF-BlackIt          396.56
l_____CaslonBookBE-Regular                   396.57
l_____CaslonBookBE-RegularSC                 396.58
l_____CaslonBookBE-RegularExp                396.59
l_____CaslonBookBE-Italic                    396.60
l_____CaslonBookBE-ItalicOsF                 396.61
l_____CaslonBookBE-ItalicExp                 396.62
l_____CaslonBookBE-Medium                    396.63
l_____CaslonBookBE-MediumSC                  396.64
l_____CaslonBookBE-MediumExp                 396.65
l_____CaslonBookBE-Bold                      396.66
l_____CaslonBookBE-BoldOsF                   396.67
l_____CaslonBookBE-BoldExp                   396.68
l_____CastleEF-Light                         397.0
l_____CastleEF-Book                          397.1

l_____CastleEF-Bold                          397.2
l_____CastleEF-Ultra                         397.3
l_____Catull-Regular                         398.0
l_____Catull-Italic                          398.1
l_____Catull-Medium                          398.2
l_____Catull-Bold                            398.3
l_____Caxton-Light                           399.0
l_____Caxton-LightItalic                     399.1
l_____Caxton-BookItalic                      399.2
l_____Caxton-Book                            399.3
l_____Caxton-Bold                            399.4
l_____Caxton-BoldItalic                      399.5
l_____CaxtonEF-Light                         399.6
l_____CaxtonEF-LightItalic                   399.7
l_____CaxtonEF-Book                          399.8
l_____CaxtonEF-BookItalic                    399.9
l_____CaxtonEF-Bold                          399.10
l_____CaxtonEF-BoldItalic                    399.11
l_____CaxtonEF-ExtraBold                     399.12
l_____CaxtonEF-ExtraBoldItalic               399.13
l_____CaxtonEF-BoldCondensed                 399.14
l_____CaxtonEF-ExtraBoldCondensed            399.15
l_____CentaurMT-Regular                      400.0
l_____CentaurMT-RegularSC                    400.1
l_____CentaurMT-RegularExpert                400.2
l_____CentaurMT-Italic                       400.3
l_____CentaurMT-ItalicA                      400.4
l_____CentaurMT-ItalicAlternate              400.5
l_____CentaurMT-ItalicOsF                    400.6
l_____CentaurMT-ItalicExpert                 400.7
l_____CentaurMT-Bold                         400.8
l_____CentaurMT-BoldOsF                      400.9
l_____CentaurMT-BoldExpert                   400.10
l_____CentaurMT-BoldItalic                   400.11
l_____CentaurMT-BoldItalicOsF                400.12
l_____CentaurMT-BoldItalicExpert             400.13
l_____CentaurMT-ItalicSwash                  400.14
l_____ChallengeEF-Bold                       401.0
l_____ChallengeEF-ExtraBold                  401.1
l_____Charme                                 402.0
l_____Cheltenham-Light                       403.0
l_____Cheltenham-LightItalic                 403.1
l_____Cheltenham-Book                        403.2
l_____Cheltenham-BookItalic                  403.3
l_____Cheltenham-Bold                        403.4
l_____Cheltenham-BoldItalic                  403.5
l_____Cheltenham-Ultra                       403.6

l_____Cheltenham-UltraItalic                 403.7
l_____Cheltenham-LightCond                   403.8
l_____Cheltenham-LightCondItalic             403.9
l_____Cheltenham-BookCond                    403.10
l_____Cheltenham-BookCondItalic              403.11
l_____Cheltenham-BoldCond                    403.12
l_____Cheltenham-BoldCondItalic              403.13
l_____Cheltenham-UltraCond                   403.14
l_____Cheltenham-UltraCondItalic             403.15
l_____CheltenhamEF-Light                     403.16
l_____CheltenhamEF-LightItalic               403.17
l_____CheltenhamEF-Book                      403.18
l_____CheltenhamEF-BookItalic                403.19
l_____CheltenhamEF-Bold                      403.20
l_____CheltenhamEF-BoldItalic                403.21
l_____CheltenhamEF-Ultra                     403.22
l_____CheltenhamEF-UltraItalic               403.23
l_____CheltenhamHandtEF-Bold                 403.24
l_____CheltenhamHandtEF-BoldItalic           403.25
l_____CheltenhamEF-LightCond                 403.26
l_____CheltenhamEF-LightCondItalic           403.27
l_____CheltenhamEF-BookCond                  403.28
l_____CheltenhamEF-BookCondItalic            403.29
l_____CheltenhamEF-BoldCond                  403.30
l_____CheltenhamEF-BoldCondItalic            403.31
l_____CheltenhamEF-UltraCond                 403.32
l_____CheltenhamEF-UltraCondItalic           403.33
l_____CheltenhamOldStyleEF                   403.34
l_____Cheq                                   404.0
l_____ChiselEF                               405.0
l_____DigiGroteskNEF-Light                   406.0
l_____DigiGroteskNEF-Bold                    406.1
l_____DigiGroteskNEF-SemiBoldCond            406.2
l_____DigiGroteskNEF-BoldCondens             406.3
l_____DigiGroteskSEF-Light                   406.4
l_____DigiGroteskSEF-SemiBold                406.5
l_____DomCasual                              407.0
l_____DomCasual-Bold                         407.1
l_____DomEF-CasualRegular                    407.2
l_____DomEF-DiagonalRegular                  407.3
l_____DomEF-CasualBold                       407.4
l_____DomEF-DiagonalBold                     407.5
l_____DorchesterScriptMT                     408.0
l_____DynamoEF-Medium                        409.0
l_____DynamoEF-Bold                          409.1
l_____DynamoEF-BoldShadow                    409.2
l_____DynamoEF-BoldCondensed                 409.3

l_____EdisonEF-Book                          410.0
l_____EdisonEF-BookItalic                    410.1
l_____EdisonEF-SemiBold                      410.2
l_____EdisonEF-SemiBoldItalic                410.3
l_____EdisonEF-BoldCondensed                 410.4
l_____EdwardianEF-Light                      411.0
l_____EdwardianEF-LightIta                   411.1
l_____EdwardianEF-Medium                     411.2
l_____EdwardianEF-MediumIta                  411.3
l_____EdwardianEF-Bold                       411.4
l_____EdwardianEF-BoldIta                    411.5
l_____EdwardianEF-ExtraBold                  411.6
l_____EdwardianEF-ExtraBoldIta               411.7
l_____EgizioEF-Condensed                     412.0
l_____EgyptianEF-BoldCondensed               414.0
l_____EgyptianEF-BoldExtended                414.1
l_____EgyptienneEF-MediumCond                415.0
l_____EhrhardtMT-Roman                       416.0
l_____EhrhardtMT-Italic                      416.1
l_____EhrhardtMT-SemiBold                    416.2
l_____EhrhardtMt-SemiBoldItalic              416.3
l_____EirinnAsciiLL                          417.0
l_____EirinnGaelicLL                         417.1
l_____ElGreco                                418.0
l_____ElefontEF                              419.0
l_____EllingtonMT-Light                      420.0
l_____EllingtonMT-LightItalic                420.1
l_____EllingtonMT-Roman                      420.2
l_____EllingtonMT-Italic                     420.3
l_____EllingtonMT-Bold                       420.4
l_____EllingtonMT-BoldItalic                 420.5
l_____EllingtonMT-ExtraBold                  420.6
l_____EllingtonMT-ExtraBoldItalic            420.7
l_____EnglishScriptEF-Regular                421.0
l_____EnglishScriptEF-Bold                   421.1
l_____EnglishScriptEF-DemiBold               421.2
l_____Erbar-LightCondensed                   422.0
l_____Erbar-BoldCondensed                    422.1
l_____EscorialEF-Outline                     423.0
l_____EscorialEF                             423.1
l_____EthnoFontEF                            424.0
l_____Eurostile                              425.0
l_____Eurostile-Oblique                      425.1
l_____Eurostile-Demi                         425.2
l_____Eurostile-DemiOblique                  425.3
l_____Eurostile-Bold                         425.4
l_____Eurostile-BoldOblique                  425.5

l_____Eurostile-Condensed                    425.6
l_____Eurostile-BoldCondensed                425.7
l_____Eurostile-ExtendedTwo                  425.8
l_____Eurostile-BoldExtendedTwo              425.9
l_____EurostileEF-Regular                    425.10
l_____EurostileEF-Black                      425.11
l_____Excelsior                              426.0
l_____Excelsior-Italic                       426.1
l_____Excelsior-Bold                         426.2
l_____ExcelsiorCE-Roman                      426.3
l_____ExcelsiorCE-Italic                     426.4
l_____ExcelsiorCE-Bold                       426.5
l_____ExcelsiorCyr-Upright                   426.6
l_____ExcelsiorCyr-Inclined                  426.7
l_____ExcelsiorCyr-Bold                      426.8
l_____ExcelsiorEastA-Roman                   426.9
l_____ExcelsiorEastA-Italic                  426.10
l_____ExcelsiorEastA-Bold                    426.11
l_____F2FAlLineato                           427.0
l_____F2FCZYKagoTrans                        427.1
l_____F2FEntebbe                             427.2
l_____F2FHogRoach                            427.3
l_____F2FMadZineFear                         427.4
l_____F2FMadZineWip                          427.5
l_____F2FMekkasoTomanik                      427.6
l_____F2FMetamorfosi                         427.7
l_____F2FOCRAlexcyzk                         427.8
l_____F2FPixmix                              427.9
l_____F2FPoisonFlowers                       427.10
l_____F2FProvinciali                         427.11
l_____F2FSimbolico                           427.12
l_____F2FTechLand                            427.13
l_____F2FWhaleTreeRnd                        427.14
l_____F2FZakGlobe                            427.15
l_____Falstaff-Regular                       428.0
l_____FetteFraktur                           429.0
l_____FetteFraktur-Dfr                       429.1
l_____FetteFrakturEF                         429.2
l_____FlashEF-Light                          430.0
l_____FlashEF-Bold                           430.1
l_____Flyer-BlackCondensed                   431.0
l_____Flyer-BlackExtraCondensed              431.1
l_____FlyerEF-Bold                           431.2
l_____FlyerEF-BoldCondensed                  431.3
l_____Folio-Light                            432.0
l_____Folio-Medium                           432.1
l_____Folio-Bold                             432.2

l_____Folio-ExtraBold                        432.3
l_____Folio-BoldCondensed                    432.4
l_____FolioEF-BoldCondensed                  432.5
l_____FolioEF-ExtraCondensed                 432.6
l_____Formata-Light                          435.0
l_____Formata-LightSC                        435.1
l_____Formata-LightExpert                    435.2
l_____Formata-LightItalic                    435.3
l_____Formata-LightItalicSC                  435.4
l_____Formata-LightItalicExpert              435.5
l_____Formata-Regular                        435.6
l_____Formata-RegularSC                      435.7
l_____Formata-RegularExpert                  435.8
l_____Formata-Italic                         435.9
l_____Formata-ItalicSC                       435.10
l_____Formata-ItalicExpert                   435.11
l_____Formata-Medium                         435.12
l_____Formata-MediumItalic                   435.13
l_____Formata-Bold                           435.14
l_____Formata-BoldItalic                     435.15
l_____Formata-Outline                        435.16
l_____Formata-LightCondensed                 435.17
l_____Formata-LightCnSC                      435.18
l_____Formata-LightCnExpert                  435.19
l_____Formata-LightCondensedItalic           435.20
l_____Formata-LightCnItalicSC                435.21
l_____Formata-LightCnItalicExpert            435.22
l_____Formata-Condensed                      435.23
l_____Formata-CondensedOutline               435.24
l_____Formata-CondensedSC                    435.25
l_____Formata-CondensedExpert                435.26
l_____Formata-CondensedItalic                435.27
l_____Formata-CnItalicSC                     435.28
l_____Formata-CnItalicExpert                 435.29
l_____Formata-MediumCondensed                435.30
l_____Formata-MediumCondensedItalic          435.31
l_____Formata-BoldCondensed                  435.32
l_____Formata-BoldCondensedItalic            435.33
l_____FrankfurterEF-Medium                   436.0
l_____FrankfurterEF-Solid                    436.1
l_____FrankfurterEF-Inline                   436.2
l_____FrankfurterEF-Highlight                436.3
l_____FreestyleScript                        437.0
l_____FreestyleScriptEF-Reg                  437.1
l_____FreestyleScriptEF-RegAlt               437.2
l_____FreestyleScriptEF-Bol                  437.3
l_____FreestyleScriptEF-BolAlt               437.4

l_____Bauhaus-Light                          438.0
l_____Bauhaus-Medium                         438.1
l_____Bauhaus-Demi                           438.2
l_____Bauhaus-Bold                           438.3
l_____Bauhaus-Heavy                          438.4
l_____BauhausEF-Light                        438.5
l_____BauhausEF-Medium                       438.6
l_____BauhausEF-DemiBold                     438.7
l_____BauhausEF-Bold                         438.8
l_____BauhausEF-Heavy                        438.9
l_____BenguiatGothic-Book                    439.0
l_____BenguiatGothic-BookOblique             439.1
l_____BenguiatGothic-Medium                  439.2
l_____BenguiatGothic-MediumOblique           439.3
l_____BenguiatGothic-Bold                    439.4
l_____BenguiatGothic-BoldOblique             439.5
l_____BenguiatGothic-Heavy                   439.6
l_____BenguiatGothic-HeavyOblique            439.7
l_____BenguiatGothicEF-Book                  439.8
l_____BenguiatGothicEF-BookItalic            439.9
l_____BenguiatGothicEF-Medium                439.10
l_____BenguiatGothicEF-MediumItalic          439.11
l_____BenguiatGothicEF-Bold                  439.12
l_____BenguiatGothicEF-BoldItalic            439.13
l_____BenguiatGothicEF-Heavy                 439.14
l_____BenguiatGothicEF-HeavyItalic           439.15
l_____Cerigo-Book                            440.0
l_____Cerigo-BookItalic                      440.1
l_____Cerigo-Medium                          440.2
l_____Cerigo-MediumItalic                    440.3
l_____Cerigo-Bold                            440.4
l_____Cerigo-BoldItalic                      440.5
l_____CerigoEF-Book                          440.6
l_____CerigoEF-BookSC                        440.7
l_____CerigoEF-BookItalic                    440.8
l_____CerigoEF-Medium                        440.9
l_____CerigoEF-MediumSC                      440.10
l_____CerigoEF-MediumItalic                  440.11
l_____CerigoEF-Bold                          440.12
l_____CerigoEF-BoldItalic                    440.13
l_____CerigoEF-BookSwashCaps                 440.14
l_____CerigoEF-BookItalicSwashCaps           440.15
l_____Cushing-Book                           441.0
l_____Cushing-BookItalic                     441.1
l_____Cushing-Medium                         441.2
l_____Cushing-MediumItalic                   441.3
l_____Cushing-Bold                           441.4

l_____Cushing-BoldItalic                     441.5
l_____Cushing-Heavy                          441.6
l_____Cushing-HeavyItalic                    441.7
l_____CushingEF-Book                         441.8
l_____CushingEF-BookItalic                   441.9
l_____CushingEF-Medium                       441.10
l_____CushingEF-MediumItalic                 441.11
l_____CushingEF-Bold                         441.12
l_____CushingEF-BoldItalic                   441.13
l_____CushingEF-Heavy                        441.14
l_____CushingEF-HeavyItalic                  441.15
l_____Esprit-Book                            442.0
l_____Esprit-BookItalic                      442.1
l_____Esprit-Medium                          442.2
l_____Esprit-MediumItalic                    442.3
l_____Esprit-Bold                            442.4
l_____Esprit-BoldItalic                      442.5
l_____Esprit-Black                           442.6
l_____Esprit-BlackItalic                     442.7
l_____EspritEF-Book                          442.8
l_____EspritEF-BookSC                        442.9
l_____EspritEF-BookItalic                    442.10
l_____EspritEF-Medium                        442.11
l_____EspritEF-MediumSC                      442.12
l_____EspritEF-MediumItalic                  442.13
l_____EspritEF-Bold                          442.14
l_____EspritEF-BoldItalic                    442.15
l_____EspritEF-Black                         442.16
l_____EspritEF-BlackItalic                   442.17
l_____Flora-Roman                            443.0
l_____Flora-Bold                             443.1
l_____FloraEF-Medium                         443.2
l_____FloraEF-Bold                           443.3
l_____Galliard-Roman                         444.0
l_____Galliard-Italic                        444.1
l_____Galliard-Bold                          444.2
l_____Galliard-BoldItalic                    444.3
l_____Galliard-Black                         444.4
l_____Galliard-BlackItalic                   444.5
l_____Galliard-Ultra                         444.6
l_____Galliard-UltraItalic                   444.7
l_____GalliardEF-Roman                       444.8
l_____GalliardEF-Italic                      444.9
l_____GalliardEF-Bold                        444.10
l_____GalliardEF-BoldItalic                  444.11
l_____GalliardEF-Black                       444.12
l_____GalliardEF-BlackItalic                 444.13

l_____GalliardEF-Ultra                       444.14
l_____GalliardEF-UltraItalic                 444.15
l_____GoldenTypeEF-Original                  445.0
l_____GoldenTypeSCEF-Original                445.1
l_____GoldenTypeEF-Bold                      445.2
l_____GoldenTypeSCEF-Bold                    445.3
l_____GoldenTypeEF-Black                     445.4
l_____IsbellEF-Book                          446.0
l_____IsbellEF-BookItalic                    446.1
l_____IsbellEF-Bold                          446.2
l_____IsbellEF-BoldItalic                    446.3
l_____Korinna-Regular                        447.0
l_____Korinna-KursivRegular                  447.1
l_____Korinna-Bold                           447.2
l_____Korinna-KursivBold                     447.3
l_____KorinnaEF-Regular                      447.4
l_____KorinnaEF-RegularKursiv                447.5
l_____KorinnaEF-Bold                         447.6
l_____KorinnaEF-BoldKursiv                   447.7
l_____KorinnaEF-ExtraBold                    447.8
l_____KorinnaEF-ExtraBoldKursiv              447.9
l_____KorinnaEF-Heavy                        447.10
l_____KorinnaEF-HeavyKursiv                  447.11
l_____LegacySans-Book                        448.0
l_____LegacySans-BookItalic                  448.1
l_____LegacySans-Medium                      448.2
l_____LegacySans-MediumItalic                448.3
l_____LegacySans-Bold                        448.4
l_____LegacySans-BoldItalic                  448.5
l_____LegacySans-Ultra                       448.6
l_____LegacySerif-Book                       448.7
l_____LegacySerif-BookItalic                 448.8
l_____LegacySerif-Medium                     448.9
l_____LegacySerif-MediumItalic               448.10
l_____LegacySerif-Bold                       448.11
l_____LegacySerif-BoldItalic                 448.12
l_____LegacySerif-Ultra                      448.13
l_____LegacySansEF-Book                      448.14
l_____LegacySansEF-BookSC                    448.15
l_____LegacySansEF-BookItalic                448.16
l_____LegacySansEF-Medium                    448.17
l_____LegacySansEF-MediumSC                  448.18
l_____LegacySansEF-MediumItalic              448.19
l_____LegacySansEF-Bold                      448.20
l_____LegacySansEF-BoldItalic                448.21
l_____LegacySansEF-Ultra                     448.22
l_____LegacySerifEF-Book                     448.23

l_____LegacySerifEF-BookSC                   448.24
l_____LegacySerifEF-BookItalic               448.25
l_____LegacySerifEF-Medium                   448.26
l_____LegacySerifEF-MediumSC                 448.27
l_____LegacySerifEF-MediumItalic             448.28
l_____LegacySerifEF-Bold                     448.29
l_____LegacySerifEF-BoldItalic               448.30
l_____LegacySerifEF-Ultra                    448.31
l_____Machine                                449.0
l_____Machine-Bold                           449.1
l_____MachineEF-Regular                      449.2
l_____MachineEF-Bold                         449.3
l_____MonaLisa-Solid                         450.0
l_____MonaLisa-Recut                         450.1
l_____MonaLisaEF-Solid                       450.2
l_____MonaLisaEF-Recut                       450.3
l_____Novarese-Book                          451.0
l_____Novarese-BookItalic                    451.1
l_____Novarese-Medium                        451.2
l_____Novarese-MediumItalic                  451.3
l_____Novarese-Bold                          451.4
l_____Novarese-BoldItalic                    451.5
l_____Novarese-Ultra                         451.6
l_____NovareseEF-Book                        451.7
l_____NovareseEF-BookItalic                  451.8
l_____NovareseEF-Medium                      451.9
l_____NovareseEF-MediumItalic                451.10
l_____NovareseEF-Bold                        451.11
l_____NovareseEF-BoldItalic                  451.12
l_____NovareseEF-Ultra                       451.13
l_____PacellaEF-Book                         452.0
l_____PacellaSCEF-Book                       452.1
l_____PacellaEF-BookItalic                   452.2
l_____PacellaEF-Medium                       452.3
l_____PacellaSCEF-Medium                     452.4
l_____PacellaEF-MediumItalic                 452.5
l_____PacellaEF-Bold                         452.6
l_____PacellaEF-BoldItalic                   452.7
l_____PacellaEF-Black                        452.8
l_____PacellaEF-BlackItalic                  452.9
l_____QuaySansEF-Book                        453.0
l_____QuaySansEF-BookSC                      453.1
l_____QuaySansEF-BookItalic                  453.2
l_____QuaySansEF-Medium                      453.3
l_____QuaySansEF-MediumSC                    453.4
l_____QuaySansEF-MediumItalic                453.5
l_____QuaySansEF-Black                       453.6

l_____QuaySansEF-BlackItalic                 453.7
l_____SerifGothic-Light                      454.0
l_____SerifGothic                            454.1
l_____SerifGothic-Bold                       454.2
l_____SerifGothic-ExtraBold                  454.3
l_____SerifGothic-Heavy                      454.4
l_____SerifGothic-Black                      454.5
l_____SerifGothicEF-Light                    454.6
l_____SerifGothicEF-Regular                  454.7
l_____SerifGothicEF-Bold                     454.8
l_____SerifGothicEF-ExtraBold                454.9
l_____SerifGothicEF-Heavy                    454.10
l_____SerifGothicEF-Black                    454.11
l_____Symbol-Book                            455.0
l_____Symbol-BookItalic                      455.1
l_____Symbol-Medium                          455.2
l_____Symbol-MediumItalic                    455.3
l_____Symbol-Bold                            455.4
l_____Symbol-BoldItalic                      455.5
l_____Symbol-Black                           455.6
l_____Symbol-BlackItalic                     455.7
l_____SymbolEF-Book                          455.8
l_____SymbolEF-BookSC                        455.9
l_____SymbolEF-BookItalic                    455.10
l_____SymbolEF-Medium                        455.11
l_____SymbolEF-MediumSC                      455.12
l_____SymbolEF-MediumItalic                  455.13
l_____SymbolEF-Bold                          455.14
l_____SymbolEF-BoldItalic                    455.15
l_____SymbolEF-Black                         455.16
l_____SymbolEF-BlackItalic                   455.17
l_____Tiepolo-Book                           456.0
l_____Tiepolo-BookItalic                     456.1
l_____Tiepolo-Bold                           456.2
l_____Tiepolo-BoldItalic                     456.3
l_____Tiepolo-Black                          456.4
l_____Tiepolo-BlackItalic                    456.5
l_____TiepoloEF-Book                         456.6
l_____TiepoloEF-BookSC                       456.7
l_____TiepoloEF-BookItalic                   456.8
l_____TiepoloEF-Bold                         456.9
l_____TiepoloEF-BoldSC                       456.10
l_____TiepoloEF-BoldItalic                   456.11
l_____TiepoloEF-Black                        456.12
l_____TiepoloEF-BlackItalic                  456.13
l_____Veljovic-Book                          457.0
l_____Veljovic-BookItalic                    457.1

l_____Veljovic-Medium                        457.2
l_____Veljovic-MediumItalic                  457.3
l_____Veljovic-Bold                          457.4
l_____Veljovic-BoldItalic                    457.5
l_____Veljovic-Black                         457.6
l_____Veljovic-BlackItalic                   457.7
l_____VeljovicEF-Book                        457.8
l_____VeljovicEF-BookOsF                     457.9
l_____VeljovicEF-BookItalic                  457.10
l_____VeljovicEF-BookItalicOsF               457.11
l_____VeljovicEF-Medium                      457.12
l_____VeljovicEF-MediumOsF                   457.13
l_____VeljovicEF-MediumItalic                457.14
l_____VeljovicEF-MediumItalicOsF             457.15
l_____VeljovicEF-Bold                        457.16
l_____VeljovicEF-BoldOsF                     457.17
l_____VeljovicEF-BoldItalic                  457.18
l_____VeljovicEF-BoldItalicOsF               457.19
l_____VeljovicEF-Black                       457.20
l_____VeljovicEF-BlackOsF                    457.21
l_____VeljovicEF-BlackItalic                 457.22
l_____VeljovicEF-BlackItalicOsF              457.23
l_____ZapfBookEF-Light                       458.0
l_____ZapfBookEF-LightItalic                 458.1
l_____ZapfBookEF-Medium                      458.2
l_____ZapfBookEF-MediumItalic                458.3
l_____ZapfBookEF-Demi                        458.4
l_____ZapfBookEF-DemiItalic                  458.5
l_____ZapfBookEF-Heavy                       458.6
l_____ZapfBookEF-HeavyItalic                 458.7
l_____ZapfChancery-Light                     458.8
l_____ZapfChancery-LightItalic               458.9
l_____ZapfChancery-Roman                     458.10
l_____ZapfChancery-Italic                    458.11
l_____ZapfChancery-MediumItalic              458.12
l_____ZapfChancery-Demi                      458.13
l_____ZapfChancery-Bold                      458.14
l_____ZapfChanceryCE-MediumItalic            458.15
l_____ZapfChanceryEF-Light                   458.16
l_____ZapfChanceryEF-LightItalic             458.17
l_____ZapfChanceryEF-Medium                  458.18
l_____ZapfChanceryEF-MediumItalic            458.19
l_____ZapfChanceryEF-Demi                    458.20
l_____ZapfChanceryEF-Bold                    458.21
l_____ZapfInternationalEF-Light              458.23
l_____ZapfInternationalEF-LightIta           458.24
l_____ZapfInternationalEF-Medium             458.25

l_____ZapfInternationalEF-MediumIta          458.26
l_____ZapfInternationalEF-Demi               458.27
l_____ZapfInternationalEF-DemiIta            458.28
l_____ZapfInternationalEF-Heavy              458.29
l_____ZapfInternationalEF-HeavyIta           458.30
l_____AudioLH-Pi                             459.0
l_____EECLH-Pi                               460.0
l_____XmasLH-PiOne                           462.0
l_____XmasLH-PiTwo                           462.1
l_____Lucida                                 463.0
l_____Lucida-Italic                          463.1
l_____Lucida-Bold                            463.2
l_____Lucida-BoldItalic                      463.3
l_____LucidaSans                             463.4
l_____LucidaSans-Italic                      463.5
l_____LucidaSans-Bold                        463.6
l_____LucidaSans-BoldItalic                  463.7
l_____LucidaMath-Symbol                      463.8
l_____LucidaMath-Extension                   463.9
l_____LucidaMath-Italic                      463.10
l_____LucidaEF                               463.11
l_____LucidaEF-Italic                        463.12
l_____LucidaEF-Bold                          463.13
l_____LucidaEF-BoldItalic                    463.14
l_____LucidaBlackletterEF                    463.15
l_____LucidaBrightEF-Roman                   463.16
l_____LucidaBrightEF-RomanIta                463.17
l_____LucidaBrightEF-DemiBold                463.18
l_____LucidaBrightEF-DemiBoldIta             463.19
l_____LucidaBrightNarEF-Roman                463.20
l_____LucidaBrightNarEF-Ita                  463.21
l_____LucidaBrightNarEF-DeBol                463.22
l_____LucidaBrightNarEF-DeBolIta             463.23
l_____LucidaCalligraphyEF                    463.24
l_____LucidaCasualEF-Roman                   463.25
l_____LucidaCasualEF-Italic                  463.26
l_____LucidaFaxEF-Roman                      463.27
l_____LucidaFaxEF-RomanItalic                463.28
l_____LucidaFaxEF-DemiBold                   463.29
l_____LucidaFaxEF-DemiBoldItalic             463.30
l_____LucidaHandwritingEF                    463.31
l_____LucidaTypewriterEF-Regular             463.32
l_____LucidaTypewriterEF-Bold                463.33
l_____LucidaSansEF-Roman                     463.34
l_____LucidaSansEF-RomanIta                  463.35
l_____LucidaSansEF-Bold                      463.36
l_____LucidaSansEF-DemiBold                  463.37

l_____LucidaSansEF-BoldItalic                463.38
l_____LucidaSansEF-DemiBoldIta               463.39
l_____LucidaSansNarEF-Roman                  463.40
l_____LucidaSansNarEF-Ita                    463.41
l_____LucidaSansNarEF-DemBol                 463.42
l_____LucidaSansNarEF-DemBolIta              463.43
l_____Madrone                                465.0
l_____MagnusEF                               466.0
l_____MandarinEF                             467.0
l_____Marigold                               468.0
l_____MatthiaLL-Regular                      469.0
l_____MaturaMT                               470.0
l_____MaturaMT-ScriptCapitals                470.1
l_____MaximaEF-Light                         471.0
l_____MaximaEF-LightOblique                  471.1
l_____MaximaEF-Medium                        471.2
l_____MaximaEF-MediumOblique                 471.3
l_____MaximaEF-Bold                          471.4
l_____MaximaEF-BoldOblique                   471.5
l_____MaximaEF-BoldCondensed                 471.6
l_____MediciScript                           472.0
l_____MercuriusMT-BoldScript                 473.0
l_____Mesquite                               474.0
l_____MetropolitainEF                        475.0
l_____Minion-Regular                         476.0
l_____Minion-RegularSC                       476.1
l_____MinionExp-Regular                      476.2
l_____Minion-Italic                          476.3
l_____Minion-ItalicSC                        476.4
l_____MinionExp-Italic                       476.5
l_____Minion-Semibold                        476.6
l_____Minion-SemiboldSC                      476.7
l_____MinionExp-Semibold                     476.8
l_____Minion-SemiboldItalic                  476.9
l_____Minion-SemiboldItalicSC                476.10
l_____MinionExp-SemiboldItalic               476.11
l_____Minion-Bold                            476.12
l_____Minion-BoldOsF                         476.13
l_____MinionExp-Bold                         476.14
l_____Minion-BoldItalic                      476.15
l_____Minion-BoldItalicOsF                   476.16
l_____MinionExp-BoldItalic                   476.17
l_____Minion-Black                           476.18
l_____Minion-BlackOsF                        476.19
l_____Minion-BlackExpert                     476.20
l_____Minion-Display                         476.21
l_____Minion-DisplaySC                       476.22

l_____MinionExp-Display                      476.23
l_____Minion-DisplayItalic                   476.24
l_____Minion-DisplayItalicSC                 476.25
l_____MinionExp-DisplayItalic                476.26
l_____Minion-SwashItalic                     476.27
l_____Minion-SwashDisplayItalic              476.28
l_____Minion-SwashSemiboldItalic             476.29
l_____Minion-Ornaments                       476.30
l_____MinionCyr-Regular                      476.31
l_____MinionCyr-Italic                       476.32
l_____MinionCyr-Semibold                     476.33
l_____MinionCyr-SemiboldItalic               476.34
l_____MinionCyr-Bold                         476.35
l_____MinionCyr-BoldItalic                   476.36
l_____Minister-Light                         477.0
l_____Minister-LightItalic                   477.1
l_____Minister-Book                          477.2
l_____Minister-BookItalic                    477.3
l_____Minister-Bold                          477.4
l_____Minister-BoldItalic                    477.5
l_____Minister-Black                         477.6
l_____Minister-BlackItalic                   477.7
l_____MinisterEF-Light                       477.8
l_____MinisterEF-LightItalic                 477.9
l_____MinisterEF-Book                        477.10
l_____MinisterEF-BookItalic                  477.11
l_____MinisterEF-Bold                        477.12
l_____ModernNo216EF-Light                    478.0
l_____ModernNo216EF-LightItalic              478.1
l_____ModernNo216EF-Medium                   478.2
l_____ModernNo216EF-MediumItalic             478.3
l_____ModernNo216EF-Bold                     478.4
l_____ModernNo216EF-BoldItalic               478.5
l_____ModernNo216EF-Heavy                    478.6
l_____ModernNo216EF-HeavyItalic              478.7
l_____ModernMT-Bold                          478.8
l_____ModernMT-BoldItalic                    478.9
l_____ModernMT-Condensed                     478.10
l_____ModernMT-CondensedItalic               478.11
l_____ModernMT-Extended                      478.12
l_____ModernMT-ExtendedItalic                478.13
l_____ModernMT-Wide                          478.14
l_____ModernMT-WideItalic                    478.15
l_____ModernExtendedEF-Book                  478.16
l_____ModernExtendedEF-BookIta               478.17
l_____ModernExtendedEF-Bold                  478.18
l_____ModifiedGothic                         479.0

l_____MonterreyEF                            480.0
l_____MotterFeminaEF                         481.0
l_____MysteryEF-White                        482.0
l_____MysteryEF-Black                        482.1
l_____MysteryEF-Mixed                        482.2
l_____PinxitAstroLL                          485.0
l_____PinxitOfficeLL                         485.1
l_____PinxitPrivateLL                        485.2
l_____Plak-Black                             486.0
l_____Plak-BlackCondensed                    486.1
l_____Plak-BlackExtraCondensed               486.2
l_____Plantin-Light                          487.0
l_____Plantin-LightItalic                    487.1
l_____Plantin                                487.2
l_____Plantin-Italic                         487.3
l_____Plantin-Semibold                       487.4
l_____Plantin-SemiboldItalic                 487.5
l_____Plantin-Bold                           487.6
l_____Plantin-BoldItalic                     487.7
l_____Plantin-BoldCondensed                  487.8
l_____PlaybillEF                             488.0
l_____Poetica-SuppLowercaseAltI              489.0
l_____Poetica-SuppLowercaseAltII             489.1
l_____Poetica-RomanSmallCaps                 489.2
l_____Poetica-RomanSmallCapsAlt              489.3
l_____Poetica-SuppLigatures                  489.4
l_____Poetica-SuppLowercaseBegI              489.5
l_____Poetica-SuppLowercaseBegII             489.6
l_____Poetica-SuppLowercaseEndI              489.7
l_____Poetica-SuppLowercaseEndII             489.8
l_____Poetica-SuppAmpersands                 489.9
l_____Poetica-ChanceryExpert                 489.10
l_____Poetica-ChanceryI                      489.11
l_____Poetica-ChanceryII                     489.12
l_____Poetica-ChanceryIII                    489.13
l_____Poetica-ChanceryIV                     489.14
l_____Poetica-SuppInitialsSwashCaps          489.15
l_____Poetica-SwashCapsI                     489.16
l_____Poetica-SwashCapsII                    489.17
l_____Poetica-SwashCapsIII                   489.18
l_____Poetica-SwashCapsIV                    489.19
l_____Poetica-SuppOrnaments                  489.20
l_____Pompeijana-Roman                       490.0
l_____Pompeijana-Borders                     490.1
l_____PopplExquisit-Regular                  491.0
l_____PopplExquisit-Medium                   491.1
l_____PopplLaudatio-Light                    491.2

l_____PopplLaudatio-LightItalic              491.3
l_____PopplLaudatio-Regular                  491.4
l_____PopplLaudatio-Italic                   491.5
l_____PopplLaudatio-Medium                   491.6
l_____PopplLaudatio-MediumItalic             491.7
l_____PopplLaudatio-Bold                     491.8
l_____PopplLaudatio-BoldItalic               491.9
l_____PopplLaudatio-LightCondensed           491.10
l_____PopplLaudatio-Condensed                491.11
l_____PopplLaudatio-MediumCondensed          491.12
l_____PopplLaudatio-BoldCondensed            491.13
l_____PopplPontifex-Roman                    491.14
l_____PopplPontifex-RomanSC                  491.15
l_____PopplPontifex-Italic                   491.16
l_____PopplPontifex-Bold                     491.17
l_____PopplPontifex-Black                    491.18
l_____PopplPontifex-BoldCondensed            491.19
l_____PopplPontifexBE-Regular                491.20
l_____PopplPontifexBE-RegularSC              491.21
l_____PopplPontifexBE-RegularExp             491.22
l_____PopplPontifexBE-Italic                 491.23
l_____PopplPontifexBE-ItalicOsF              491.24
l_____PopplPontifexBE-ItalicExp              491.25
l_____PopplPontifexBE-Medium                 491.26
l_____PopplPontifexBE-MediumSC               491.27
l_____PopplPontifexBE-MediumExp              491.28
l_____PopplPontifexBE-Bold                   491.29
l_____PopplPontifexBE-BoldOsF                491.30
l_____PopplPontifexBE-BoldExp                491.31
l_____PopplPontifexBE-MediumCn               491.32
l_____PopplPontifexBE-MediumCnSC             491.33
l_____PopplPontifexBE-MediumCnExp            491.34
l_____PopplResidenz-Light                    491.35
l_____PopplResidenz-Regular                  491.36
l_____PostAntiqua-Roman                      492.0
l_____PostAntiqua-Bold                       492.1
l_____PostAntiquaBE-Regular                  492.2
l_____PostAntiquaBE-Medium                   492.3
l_____PostMediaeval-Regular                  492.4
l_____PostMediaeval-Italic                   492.5
l_____PostMediaeval-Medium                   492.6
l_____PraxisEF-Light                         493.0
l_____PraxisEF-LightSC                       493.1
l_____PraxisEF-Regular                       493.2
l_____PraxisEF-RegularSC                     493.3
l_____PraxisEF-SemiBold                      493.4
l_____PraxisEF-Bold                          493.5

l_____PraxisEF-Heavy                         493.6
l_____PremierEF                              494.0
l_____PremierEF-Alternate                    494.1
l_____PremierEF-Shaded                       494.2
l_____Present                                495.0
l_____Present-Bold                           495.1
l_____Present-Black                          495.2
l_____Present-Condensed                      495.3
l_____Present-BoldCondensed                  495.4
l_____Present-BlackCondensed                 495.5
l_____PrincetownEF                           496.0
l_____ProteusEF-Light                        497.0
l_____ProteusEF-Book                         497.1
l_____ProteusEF-Medium                       497.2
l_____ProteusEF-Bold                         497.3
l_____Quake                                  498.0
l_____QuartzEF                               499.0
l_____QuasariaLL-Regular                     500.0
l_____QuasariaLL-Oblique                     500.1
l_____QuasariaLL-Bold                        500.2
l_____QuasariaLL-BoldOblique                 500.3
l_____Rad                                    501.0
l_____RadiantEFText-Light                    502.0
l_____RadiantEFText-Book                     502.1
l_____RadiantEFText-Medium                   502.2
l_____RadiantEFText-DemiBold                 502.3
l_____RadiantEFText-Bold                     502.4
l_____RadiantEFText-Black                    502.5
l_____RadiantEFDisplay-CondNo2               502.6
l_____RadiantEFDisplay-CondNo1               502.7
l_____RadiantEFDisplay-CondNo4               502.8
l_____RadiantEFDisplay-CondNo3               502.9
l_____RadiantEFDisplay-CondNo6               502.10
l_____RadiantEFDisplay-CondNo5               502.11
l_____RageItalicEF                           503.0
l_____ReadMyHandLL                           504.0
l_____ReadMyHandLL-Bold                      504.1
l_____Reporter-Two                           505.0
l_____RialtoEF                               506.0
l_____Rockwell-Light                         507.0
l_____Rockwell-LightItalic                   507.1
l_____Rockwell                               507.2
l_____Rockwell-Italic                        507.3
l_____Rockwell-Bold                          507.4
l_____Rockwell-BoldItalic                    507.5
l_____Rockwell-ExtraBold                     507.6
l_____Rockwell-Condensed                     507.7

l_____Rockwell-BoldCondensed                 507.8
l_____RomanScriptEF                          508.0
l_____RomanScriptEF-Alternate                508.1
l_____Rosewood-Fill                          509.0
l_____Rosewood-Regular                       509.1
l_____Rotation-Roman                         510.0
l_____Rotation-Italic                        510.1
l_____Rotation-Bold                          510.2
l_____Roundy-Roman                           511.0
l_____Roundy-Swash                           511.1
l_____RulingScript                           512.0
l_____RulingScript-Addition                  512.1
l_____RussellSquare                          513.0
l_____RussellSquare-Oblique                  513.1
l_____Sabon-Roman                            514.0
l_____Sabon-RomanSC                          514.1
l_____Sabon-RomanOsF                         514.2
l_____Sabon-Italic                           514.3
l_____Sabon-ItalicOsF                        514.4
l_____Sabon-Bold                             514.5
l_____Sabon-BoldOsF                          514.6
l_____Sabon-BoldItalic                       514.7
l_____Sabon-BoldItalicOsF                    514.8
l_____Sassafras-Roman                        517.0
l_____Sassafras-SxRoman                      517.1
l_____Sassafras-LxRoman                      517.2
l_____Sassafras-LxItalic                     517.3
l_____Sassafras-SxItalic                     517.4
l_____Sassafras-Italic                       517.5
l_____StempelSchneidler-Light                518.0
l_____StempelSchneidler-LightItalic          518.1
l_____StempelSchneidler-Roman                518.2
l_____StempelSchneidler-Italic               518.3
l_____StempelSchneidler-Medium               518.4
l_____StempelSchneidler-MedItalic            518.5
l_____StempelSchneidler-Bold                 518.6
l_____StempelSchneidler-BoldItalic           518.7
l_____StempelSchneidler-Black                518.8
l_____StempelSchneidler-BlackItalic          518.9
l_____SchneidlerEF-Light                     518.10
l_____SchneidlerEF-AmaltheaLight             518.11
l_____SchneidlerEF-Roman                     518.12
l_____SchneidlerEF-AmaltheaRoman             518.13
l_____SchneidlerEF-Medium                    518.14
l_____SchneidlerEF-AmaltheaMedium            518.15
l_____SchneidlerEF-Bold                      518.16
l_____SchneidlerEF-AmaltheaBold              518.17

l_____SchneidlerEF-Black                     518.18
l_____SchneidlerEF-AmaltheaBlack             518.19
l_____SchreibmaschineEF-Regular              519.0
l_____Serpentine-Light                       520.0
l_____Serpentine-LightOblique                520.1
l_____Serpentine-Medium                      520.2
l_____Serpentine-MediumOblique               520.3
l_____Serpentine-Bold                        520.4
l_____Serpentine-BoldOblique                 520.5
l_____SerpentineEF-Bold                      520.6
l_____SerpentineEF-BoldItalic                520.7
l_____VarioEF-Regular                        522.0
l_____VarioEF-RegularItalic                  522.1
l_____VectoraLH-Light                        523.0
l_____VectoraLH-LightItalic                  523.1
l_____VectoraLH-Roman                        523.2
l_____VectoraLH-Italic                       523.3
l_____VectoraLH-Bold                         523.4
l_____VectoraLH-BoldItalic                   523.5
l_____VectoraLH-Black                        523.6
l_____VectoraLH-BlackItalic                  523.7
l_____VegasEF                                524.0
l_____Versailles-Light                       526.0
l_____Versailles-LightItalic                 526.1
l_____Versailles-Roman                       526.2
l_____Versailles-Italic                      526.3
l_____Versailles-Bold                        526.4
l_____Versailles-BoldItalic                  526.5
l_____Versailles-Black                       526.6
l_____Versailles-BlackItalic                 526.7
l_____Visigoth                               527.0
l_____VladimirScriptEF                       528.0
l_____Westside                               530.0
l_____WiesbadenSwing-Roman                   531.0
l_____Willow                                 532.0
l_____WittenbergerFrakturMT                  533.0
l_____WittenbergerFrakturMT-Dfr              533.1
l_____WittenbergerFrakturMT-Bd               533.2
l_____WittenbergerFrakturMT-BdDfr            533.3
l_____Zebrawood-Fill                         534.0
l_____Zebrawood-Regular                      534.1
l_____AgrafieLL                              535.0
l_____AlbertusMT-Light                       536.0
l_____AlbertusMT-Roman                       536.1
l_____AlbertusMT-Italic                      536.2
l_____Alexa                                  537.0
l_____AlgerianEF                             538.0

l_____Amigo                                  540.0
l_____AngroEF-Light                          541.0
l_____AngroEF-Bold                           541.1
l_____AnzeigenGroteskEF                      542.0
l_____ArsisEF-Regular                        543.0
l_____ArsisEF-RegularItalic                  543.1
l_____AugusteaOpenEF                         544.0
l_____Avenir-Light                           545.0
l_____Avenir-LightOblique                    545.1
l_____Avenir-Book                            545.2
l_____Avenir-BookOblique                     545.3
l_____Avenir-Roman                           545.4
l_____Avenir-Oblique                         545.5
l_____Avenir-Medium                          545.6
l_____Avenir-MediumOblique                   545.7
l_____Avenir-Heavy                           545.8
l_____Avenir-HeavyOblique                    545.9
l_____Avenir-Black                           545.10
l_____Avenir-BlackOblique                    545.11
l_____BakerSignet                            546.0
l_____Balzano                                547.0
l_____Barmeno-Regular                        548.0
l_____Barmeno-Medium                         548.1
l_____Barmeno-Bold                           548.2
l_____Barmeno-ExtraBold                      548.3
l_____BellCentennial-Address                 549.0
l_____BellCentennial-SubCaption              549.1
l_____BellCentennial-NameAndNumber           549.2
l_____BellCentennial-BoldListing             549.3
l_____BellCentennial-BoldListingAlt          549.4
l_____City-Light                             550.0
l_____City-LightItalic                       550.1
l_____City-Medium                            550.2
l_____City-MediumItalic                      550.3
l_____City-Bold                              550.4
l_____City-BoldItalic                        550.5
lpi___BorderPi-OneFiveOneFiveNine            551.0
l_____BroadwayEF                             552.2
l_____BroadwayEF-Engraved                    552.3
l_____Caliban                                553.0
l_____CarouselEF                             554.0
l_____Century-Light                          555.0
l_____Century-LightItalic                    555.1
l_____Century-Book                           555.2
l_____Century-BookItalic                     555.3
l_____Century-Bold                           555.4
l_____Century-BoldItalic                     555.5

l_____Century-Ultra                          555.6
l_____Century-UltraItalic                    555.7
l_____Century-LightCondensed                 555.8
l_____Century-LightCondensedItalic           555.9
l_____Century-BookCondensed                  555.10
l_____Century-BookCondensedItalic            555.11
l_____Century-BoldCondensed                  555.12
l_____Century-BoldCondensedItalic            555.13
l_____Century-UltraCondensed                 555.14
l_____Century-UltraCondensedItalic           555.15
l_____CenturyEF-Light                        555.16
l_____CenturyEF-LightItalic                  555.17
l_____CenturyEF-Book                         555.18
l_____CenturyEF-BookItalic                   555.19
l_____CenturyEF-Bold                         555.20
l_____CenturyEF-BoldItalic                   555.21
l_____CenturyEF-Ultra                        555.22
l_____CenturyEF-UltraItalic                  555.23
l_____CenturyHandtEF-Bold                    555.24
l_____CenturyHandtEF-BoldItalic              555.25
l_____CenturyEF-LightCondensed               555.26
l_____CenturyEF-LightCondensedItalic         555.27
l_____CenturyEF-BookCondensed                555.28
l_____CenturyEF-BookCondensedItalic          555.29
l_____CenturyEF-BoldCondensed                555.30
l_____CenturyEF-BoldCondensedItalic          555.31
l_____CenturyEF-UltraCondensed               555.32
l_____CenturyEF-UltraCondensedItalic         555.33
l_____CenturyExpanded                        555.34
l_____CenturyExpanded-Italic                 555.35
l_____CenturyExpandedEF-Regular              555.36
l_____CenturyExpandedEF-RegularItalic        555.37
l_____CenturyOldStyle-Regular                555.38
l_____CenturyOldStyle-Italic                 555.39
l_____CenturyOldStyle-Bold                   555.40
l_____CenturyOldStyleEF-Regular              555.41
l_____CenturyOldStyleEF-RegularItalic        555.42
l_____CenturyOldStyleEF-Bold                 555.43
l_____NewCenturySchlbk-Roman                 555.44
l_____NewCenturySchlbk-Italic                555.45
l_____NewCenturySchlbk-Bold                  555.46
l_____NewCenturySchlbk-BoldItalic            555.47
l_____NewCenturySchlbkCE-Roman               555.50
l_____NewCenturySchlbkCE-Italic              555.51
l_____NewCenturySchlbkCE-Bold                555.52
l_____NewCenturySchlbkCE-BoldItalic          555.53
l_____NewCenturySchlbkGreek-Upright          555.54

l_____NewCenturySchlbkGreek-Inclined         555.55
l_____NewCenturySchlbkGreek-Bold             555.56
l_____NewCenturySchlbkGreekP-Upright         555.57
l_____NewCenturySchlbkGreekP-Inclined        555.58
l_____NewCenturySchlbkGreekP-Bold            555.59
l_____CenturySchoolbookEF-Reg                555.60
l_____CenturySchoolbookEF-RegIta             555.61
l_____CenturySchoolbookEF-Bol                555.62
l_____CenturySchoolbookEF-BolIta             555.63
l_____ChesterfieldEF                         556.0
l_____Choc-Roman                             557.0
l_____ChromiumOneEF                          558.0
l_____Clarendon-Light                        559.0
l_____Clarendon                              559.1
l_____Clarendon-Bold                         559.2
l_____ClarendonEF-Light                      559.3
l_____ClarendonEF-Medium                     559.4
l_____ClarendonEF-Bold                       559.5
l_____ClarendonEF-ExtraBold                  559.6
l_____ClarendonEF-BoldExpanded               559.7
l_____Colossalis-Regular                     560.0
l_____Colossalis-Medium                      560.1
l_____Colossalis-Bold                        560.2
l_____Colossalis-Black                       560.3
l_____CompactaEF-Light                       561.0
l_____CompactaEF-Regular                     561.1
l_____CompactaEF-RegularItalic               561.2
l_____CompactaEF-Bold                        561.3
l_____CompactaEF-BoldItalic                  561.4
l_____CompactaEF-Black                       561.5
l_____CompactaEF-RegularOutline              561.6
l_____CompactaEF-BoldOutline                 561.7
l_____ConferenceEF                           562.0
l_____Copperplate-TwentyNineBC               563.0
l_____Copperplate-ThirtyBC                   563.1
l_____Copperplate-ThirtyTwoBC                563.2
l_____Copperplate-ThirtyOneBC                563.3
l_____Copperplate-ThrityThreeBC              563.4
l_____Copperplate-TwentyNineAB               563.5
l_____Copperplate-ThirtyAB                   563.6
l_____Copperplate-ThirtyTwoAB                563.7
l_____Copperplate-ThirtyOneAB                563.8
l_____CopperplateEF-Medium                   563.10
l_____CopperplateEF-Bold                     563.11
l_____CopperplateEF-LightCondensed           563.12
l_____CopperplateEF-MediumCondensed          563.13
l_____CopperplateEF-BoldCondensed            563.14

l_____CopperplateEF-Light                    563.9
l_____Cottonwood                             564.0
l_____CrilleeEF-LightItalic                  565.0
l_____CrilleeEF-Italic                       565.1
l_____CrilleeEF-BoldItalic                   565.2
l_____CrilleeEF-ExtraBoldItalic              565.3
l_____DINMittelschrift                       566.0
l_____DINMittelschrift-Alternate             566.1
l_____DINEngschrift                          566.2
l_____DINEngschrift-Alternate                566.3
l_____DIN1451EF-MittelNeu                    566.4
l_____DIN1451EF-MittelAlt                    566.5
l_____DIN1451EF-EngNeu                       566.6
l_____DIN1451EF-EngAlt                       566.7
l_____DINNeuzeitGrotesk-Light                566.8
l_____DINNeuzeitGrotesk-BoldCond             566.9
l_____DalcoraHE                              567.0
lpi___DecoNumbersLH-Circle                   568.0
lpi___DecoNumbersLH-OCRA                     568.1
lpi___DecoNumbersLH-Serlio                   568.2
lpi___DecoNumbersLH-Square                   568.3
lpi___DecoNumbersLH-Triangle                 568.4
l_____DemosEF-Medium                         569.0
l_____DemosEF-MediumSC                       569.1
l_____DemosEF-MediumItalic                   569.2
l_____DemosEF-SemiBold                       569.3
l_____DextorEF                               570.0
l_____Diotima-Roman                          571.0
l_____Diotima-RomanSC                        571.1
l_____Diotima-RomanOsF                       571.2
l_____Diotima-Italic                         571.3
l_____Diotima-ItalicOsF                      571.4
l_____DucDeBerry                             572.0
l_____DucDeBerry-Dfr                         572.1
l_____EinhornEF                              573.0
l_____EngraversLH-BoldFace                   574.0
l_____EngraversEF-Roman                      574.1
l_____EngraversEF-Bold                       574.2
l_____GaladrielEF                            578.0
l_____StempelGaramond-Roman                  579.0
l_____StempelGaramond-RomanSC                579.1
l_____StempelGaramond-RomanOsF               579.2
l_____StempelGaramond-Italic                 579.3
l_____StempelGaramond-ItalicOsF              579.4
l_____StempelGaramond-Bold                   579.5
l_____StempelGaramond-BoldOsF                579.6
l_____StempelGaramond-BoldItalic             579.7

l_____StempelGaramond-BoldItalicOsF          579.8
l_____StempelGaramondCE-Roman                579.9
l_____StempelGaramondCE-Italic               579.10
l_____StempelGaramondCE-Bold                 579.11
l_____StempelGaramondCE-BoldItalic           579.12
l_____StempelGaramondEastA-Roman             579.13
l_____StempelGaramondEastA-Italic            579.14
l_____StempelGaramondEastA-Bold              579.15
l_____StempelGaramondEastA-BoldItalic        579.16
l_____AGaramond-Regular                      579.17
l_____AGaramondAlt-Regular                   579.18
l_____AGaramond-RegularSC                    579.19
l_____AGaramondExp-Regular                   579.20
l_____AGaramond-Italic                       579.21
l_____AGaramondAlt-Italic                    579.22
l_____AGaramond-ItalicOsF                    579.23
l_____AGaramondExp-Italic                    579.24
l_____AGaramond-Semibold                     579.25
l_____AGaramond-SemiboldSC                   579.26
l_____AGaramondExp-Semibold                  579.27
l_____AGaramond-SemiboldItalic               579.28
l_____AGaramond-SemiboldItalicOsF            579.29
l_____AGaramondExp-SemiboldItalic            579.30
l_____AGaramond-Bold                         579.31
l_____AGaramond-BoldOsF                      579.32
l_____AGaramondExp-Bold                      579.33
l_____AGaramond-BoldItalic                   579.34
l_____AGaramond-BoldItalicOsF                579.35
l_____AGaramondExp-BoldItalic                579.36
l_____AGaramond-Titling                      579.37
l_____GaramondBE-Regular                     579.38
l_____GaramondBE-RegularSC                   579.39
l_____GaramondBE-RegularExpert               579.40
l_____GaramondBE-Italic                      579.41
l_____GaramondBE-ItalicOsF                   579.42
l_____GaramondBE-ItalicExpert                579.43
l_____GaramondBE-Medium                      579.44
l_____GaramondBE-MediumSC                    579.45
l_____GaramondBE-MediumExpert                579.46
l_____GaramondBE-MediumItalic                579.47
l_____GaramondBE-MediumItalicOsF             579.48
l_____GaramondBE-MediumItalicExpert          579.49
l_____GaramondBE-Bold                        579.50
l_____GaramondBE-BoldOsF                     579.51
l_____GaramondBE-BoldExpert                  579.52
l_____GaramondBE-Condensed                   579.53
l_____GaramondBE-CondensedSC                 579.54

l_____GaramondBE-CnExpert                    579.55
l_____GaramondBE-MediumCn                    579.56
l_____GaramondBE-MediumCnOsF                 579.57
l_____GaramondBE-MediumCnExpert              579.58
l_____GaramondBE-SwashItalic                 579.59
l_____Garamond-Light                         579.60
l_____Garamond-LightItalic                   579.61
l_____Garamond-Book                          579.62
l_____Garamond-BookItalic                    579.63
l_____Garamond-Bold                          579.64
l_____Garamond-BoldItalic                    579.65
l_____Garamond-Ultra                         579.66
l_____Garamond-UltraItalic                   579.67
l_____Garamond-LightCondensed                579.68
l_____Garamond-LightCondensedItalic          579.69
l_____Garamond-BookCondensed                 579.70
l_____Garamond-BookCondensedItalic           579.71
l_____Garamond-BoldCondensed                 579.72
l_____Garamond-BoldCondensedItalic           579.73
l_____Garamond-UltraCondensed                579.74
l_____Garamond-UltraCondensedItalic          579.75
l_____GaramondEF-Light                       579.76
l_____GaramondEF-LightItalic                 579.77
l_____GaramondEF-Book                        579.78
l_____GaramondEF-BookItalic                  579.79
l_____GaramondEF-Bold                        579.80
l_____GaramondEF-BoldItalic                  579.81
l_____GaramondEF-Ultra                       579.82
l_____GaramondEF-UltraItalic                 579.83
l_____GaramondHandtEF-Bold                   579.84
l_____GaramondHandtEF-BoldItalic             579.85
l_____GaramondEF-LightCondensed              579.86
l_____GaramondEF-LightCondensedItalic        579.87
l_____GaramondEF-BookCondensed               579.88
l_____GaramondEF-BookCondensedItalic         579.89
l_____GaramondEF-BoldCondensed               579.90
l_____GaramondEF-BoldCondensedItalic         579.91
l_____GaramondEF-UltraCondensed              579.92
l_____GaramondEF-UltraCondensedItalic        579.93
l_____SimonciniGaramond                      579.94
l_____SimonciniGaramond-Italic               579.95
l_____SimonciniGaramond-Bold                 579.96
l_____GaramondThree                          579.97
l_____GaramondThree-SC                       579.98
l_____GaramondThree-Italic                   579.99
l_____GaramondThree-ItalicOsF                579.100
l_____GaramondThree-Bold                     579.101

l_____GaramondThree-BoldSC                   579.102
l_____GaramondThree-BoldItalic               579.103
l_____GaramondThree-BoldItalicOsF            579.104
l_____GaramondNo5EF-Light                    579.105
l_____GaramondNo5EF-LightItalic              579.106
l_____GaramondNo5EF-Bold                     579.107
l_____Giddyup                                580.0
l_____GlaserStencilEF                        581.0
l_____Grace                                  582.0
l_____GrafiloneLL-Light                      583.0
l_____GrafiloneLL-SemiBold                   583.1
l_____GrafiloneLL-Bold                       583.2
l_____HaMangaIrregularLL-Regular             584.0
l_____HaMangaIrregularLL-Italic              584.1
l_____HaMangaIrregularPicturesLL             584.2
l_____HandelGothicEF-Light                   585.0
l_____HandelGothicEF-Medium                  585.1
l_____HandelGothicEF-Bold                    585.2
l_____Herculanum                             586.0
l_____HoTomLL-Regular                        587.0
l_____HollanderEF-Regular                    588.0
l_____HollanderSCEF-Regular                  588.1
l_____HollanderEF-RegularItalic              588.2
l_____HollanderEF-Bold                       588.3
l_____HorndonEF                              589.0
l_____BarcelonaEF-Book                       591.0
l_____BarcelonaEF-BookItalic                 591.1
l_____BarcelonaEF-Medium                     591.2
l_____BarcelonaEF-MediumItalic               591.3
l_____BarcelonaEF-Bold                       591.4
l_____BarcelonaEF-BoldItalic                 591.5
l_____BarcelonaEF-Heavy                      591.6
l_____BarcelonaEF-HeavyItalic                591.7
l_____Bookman-Light                          592.0
l_____Bookman-LightItalic                    592.1
l_____Bookman-Medium                         592.2
l_____Bookman-MediumItalic                   592.3
l_____Bookman-Demi                           592.4
l_____Bookman-DemiItalic                     592.5
l_____Bookman-Bold                           592.6
l_____Bookman-BoldItalic                     592.7
l_____BookmanCE-Light                        592.8
l_____BookmanCE-LightItalic                  592.9
l_____BookmanCE-Demi                         592.10
l_____BookmanCE-DemiItalic                   592.11
l_____BookmanEF-Light                        592.12
l_____BookmanEF-LightItalic                  592.13

l_____BookmanEF-Medium                       592.14
l_____BookmanEF-MediumItalic                 592.15
l_____BookmanEF-Demi                         592.16
l_____BookmanEF-DemiItalic                   592.17
l_____BookmanEF-Bold                         592.18
l_____BookmanEF-BoldItalic                   592.19
l_____GammaEF-Book                           593.0
l_____GammaSCEF-Book                         593.1
l_____GammaEF-BookItalic                     593.2
l_____GammaEF-Medium                         593.3
l_____GammaSCEF-Medium                       593.4
l_____GammaEF-MediumItalic                   593.5
l_____GammaEF-Bold                           593.6
l_____GammaEF-BoldItalic                     593.7
l_____GammaEF-Black                          593.8
l_____GammaEF-BlackItalic                    593.9
l_____Italia-Book                            594.0
l_____Italia-Medium                          594.1
l_____Italia-Bold                            594.2
l_____ItaliaEF-Book                          594.3
l_____ItaliaEF-Medium                        594.4
l_____ItaliaEF-Bold                          594.5
l_____ItaliaEF-Condense                      594.6
l_____MendozaRoman-Book                      595.0
l_____MendozaRoman-BookItalic                595.1
l_____MendozaRoman-Medium                    595.2
l_____MendozaRoman-MediumItalic              595.3
l_____MendozaRoman-Bold                      595.4
l_____MendozaRoman-BoldItalic                595.5
l_____MendozaRomanEF-Book                    595.6
l_____MendozaRomanEF-BookSC                  595.7
l_____MendozaRomanEF-BookItalic              595.8
l_____MendozaRomanEF-Medium                  595.9
l_____MendozaRomanEF-MediumSC                595.10
l_____MendozaRomanEF-MediumItalic            595.11
l_____MendozaRomanEF-Bold                    595.12
l_____MendozaRomanEF-BoldItalic              595.13
l_____OfficinaSans-Book                      596.0
l_____OfficinaSans-BookItalic                596.1
l_____OfficinaSans-Bold                      596.2
l_____OfficinaSans-BoldItalic                596.3
l_____OfficinaSerif-Book                     596.4
l_____OfficinaSerif-BookItalic               596.5
l_____OfficinaSerif-Bold                     596.6
l_____OfficinaSerif-BoldItalic               596.7
l_____OfficinaSansEF-Book                    596.8
l_____OfficinaSansEF-BookItalic              596.9

l_____OfficinaSansEF-Bold                    596.10
l_____OfficinaSansEF-BoldItalic              596.11
l_____OfficinaSerifEF-Book                   596.12
l_____OfficinaSerifEF-BookItalic             596.13
l_____OfficinaSerifEF-Bold                   596.14
l_____OfficinaSerifEF-BoldItalic             596.15
l_____StoneInformal                          597.0
l_____StoneInformal-Italic                   597.1
l_____StoneInformal-SemiBold                 597.2
l_____StoneInformal-SemiBoldItalic           597.3
l_____StoneInformal-Bold                     597.4
l_____StoneInformal-BoldItalic               597.5
l_____StoneSans                              597.6
l_____StoneSans-Italic                       597.7
l_____StoneSans-SemiBold                     597.8
l_____StoneSans-SemiBoldItalic               597.9
l_____StoneSans-Bold                         597.10
l_____StoneSans-BoldItalic                   597.11
l_____StoneSerif                             597.12
l_____StoneSerif-Italic                      597.13
l_____StoneSerif-SemiBold                    597.14
l_____StoneSerif-SemiBoldItalic              597.15
l_____StoneSerif-Bold                        597.16
l_____StoneSerif-BoldItalic                  597.17
l_____StoneSans-PhoneticIPA                  597.18
l_____StoneSans-PhoneticAlternate            597.19
l_____StoneSerif-PhoneticIPA                 597.20
l_____StoneSerif-PhoneticAlternate           597.21
l_____StoneInformalEF-Medium                 597.22
l_____StoneInformalEF-MediumSC               597.23
l_____StoneInformalEF-MediumItalic           597.24
l_____StoneInformalEF-SemiBold               597.25
l_____StoneInformalEF-SemiBoldSC             597.26
l_____StoneInformalEF-SemiBoldItalic         597.27
l_____StoneInformalEF-Bold                   597.28
l_____StoneInformalEF-BoldItalic             597.29
l_____StoneSansEF-Medium                     597.30
l_____StoneSansEF-MediumSC                   597.31
l_____StoneSansEF-MediumItalic               597.32
l_____StoneSansEF-SemiBold                   597.33
l_____StoneSansEF-SemiBoldSC                 597.34
l_____StoneSansEF-SemiBoldItalic             597.35
l_____StoneSansEF-Bold                       597.36
l_____StoneSansEF-BoldItalic                 597.37
l_____StoneSerifEF-Medium                    597.38
l_____StoneSerifEF-MediumSC                  597.39
l_____StoneSerifEF-MediumItalic              597.40

l_____StoneSerifEF-SemiBold                  597.41
l_____StoneSerifEF-SemiBoldSC                597.42
l_____StoneSerifEF-SemiBoldItalic            597.43
l_____StoneSerifEF-Bold                      597.44
l_____StoneSerifEF-BoldItalic                597.45
l_____StoneSansEF-MediumPhonetic             597.46
l_____StoneSansEF-MediumPhoneticPlus         597.47
l_____StoneSerifEF-MediumPhonetic            597.48
l_____StoneSerifEF-MediumPhoneticPlus        597.49
l_____Usherwood-Book                         598.0
l_____Usherwood-BookItalic                   598.1
l_____Usherwood-Medium                       598.2
l_____Usherwood-MediumItalic                 598.3
l_____Usherwood-Bold                         598.4
l_____Usherwood-BoldItalic                   598.5
l_____Usherwood-Black                        598.6
l_____Usherwood-BlackItalic                  598.7
l_____UsherwoodEF-Book                       598.8
l_____UsherwoodEF-BookItalic                 598.9
l_____UsherwoodEF-Medium                     598.10
l_____UsherwoodEF-MediumItalic               598.11
l_____UsherwoodEF-Bold                       598.12
l_____UsherwoodEF-BoldItalic                 598.13
l_____UsherwoodEF-Black                      598.14
l_____UsherwoodEF-BlackItalic                598.15
l_____Impact                                 600.0
l_____ImprimeurClassiqueEF-Reg               601.0
l_____ImprimeurClassiqueEF-Obl               601.1
l_____ImprimeurNouveauEF-Reg                 601.2
l_____ImprimeurNouveauEF-Obl                 601.3
l_____Insignia                               602.0
l_____Insignia-A                             602.1
l_____JansonText-Roman                       603.0
l_____JansonText-RomanSC                     603.1
l_____JansonText-Italic                      603.2
l_____JansonText-ItalicOsF                   603.3
l_____JansonText-Bold                        603.4
l_____JansonText-BoldOsF                     603.5
l_____JansonText-BoldItalic                  603.6
l_____JansonText-BoldItalicOsF               603.7
l_____Kabel-Light                            604.0
l_____Kabel-Book                             604.1
l_____Kabel-Heavy                            604.2
l_____Kabel-Black                            604.3
l_____ItcKabel-Book                          604.4
l_____ItcKabel-Medium                        604.5
l_____ItcKabel-Demi                          604.6

l_____ItcKabel-Bold                          604.7
l_____ItcKabel-Ultra                         604.8
l_____ItcKabelEF-Book                        604.9
l_____ItcKabelEF-Medium                      604.10
l_____ItcKabelEF-Demi                        604.11
l_____ItcKabelEF-Bold                        604.12
l_____ItcKabelEF-Ultra                       604.13
l_____KapitelliaEF-Bold                      605.0
l_____KlangMT                                606.0
l_____KolossEF                               607.0
l_____LCDEF                                  608.0
l_____LinoLetter-Roman                       610.0
l_____LinoLetter-RomanSC                     610.1
l_____LinoLetter-RomanOsF                    610.2
l_____LinoLetter-Italic                      610.3
l_____LinoLetter-ItalicOsF                   610.4
l_____LinoLetter-Medium                      610.5
l_____LinoLetter-MediumSC                    610.6
l_____LinoLetter-MediumOsF                   610.7
l_____LinoLetter-MediumItalic                610.8
l_____LinoLetter-MediumItalicOsF             610.9
l_____LinoLetter-Bold                        610.10
l_____LinoLetter-BoldSC                      610.11
l_____LinoLetter-BoldOsF                     610.12
l_____LinoLetter-BoldItalic                  610.13
l_____LinoLetter-BoldItalicOsF               610.14
l_____LinoLetter-Black                       610.15
l_____LinoLetter-BlackSC                     610.16
l_____LinoLetter-BlackOsF                    610.17
l_____LinoLetter-BlackItalic                 610.18
l_____LinoLetter-BlackItalicOsF              610.19
l_____TechnicalLH-PiOne                      612.0
l_____TechnicalLH-PiTwo                      612.1
l_____MagnaEF-Light                          613.0
l_____MagnaEF-Italic                         613.1
l_____MagnaEF-Bold                           613.2
l_____MateoRomanLL                           614.0
l_____MateoBoldLL                            614.1
l_____MateoBoldOutlineLL                     614.2
l_____Memphis-Light                          615.0
l_____Memphis-LightItalic                    615.1
l_____Memphis-Medium                         615.2
l_____Memphis-MediumItalic                   615.3
l_____Memphis-Bold                           615.4
l_____Memphis-BoldItalic                     615.5
l_____Memphis-ExtraBold                      615.6
l_____MetroLite-Two                          616.0

l_____MetroMedium-Two                        616.1
l_____MetroBlack-Two                         616.2
l_____Mistral                                617.0
l_____Monotype Script-Bold                   618.0
l_____Napoleon-BoldCondensed                 619.0
l_____NeulandStar                            620.1
l_____NewCaledonia                           621.0
l_____NewCaledonia-RomanSC                   621.1
l_____NewCaledonia-Italic                    621.2
l_____NewCaledonia-ItalicOsF                 621.3
l_____NewCaledonia-SemiBold                  621.4
l_____NewCaledonia-SemiBoldItalic            621.5
l_____NewCaledonia-Bold                      621.6
l_____NewCaledonia-BoldSC                    621.7
l_____NewCaledonia-BoldItalic                621.8
l_____NewCaledonia-BoldItalicOsF             621.9
l_____NewCaledonia-Black                     621.10
l_____NewCaledonia-BlackItalic               621.11
l_____NoweAtenyLL-Two                        622.0
l_____OctopussEF                             623.0
l_____OctopussEF-Shaded                      623.1
l_____OldTown536EF                           624.0
l_____Optima                                 625.0
l_____Optima-Oblique                         625.1
l_____Optima-Medium                          625.2
l_____Optima-MediumOblique                   625.3
l_____Optima-Demi                            625.4
l_____Optima-DemiOblique                     625.5
l_____Optima-Bold                            625.6
l_____Optima-BoldOblique                     625.7
l_____Optima-Black                           625.8
l_____Optima-BlackOblique                    625.9
l_____Optima-ExtraBlack                      625.10
l_____Optima-ExtraBlackOblique               625.11
l_____OptimaCE-Roman                         625.12
l_____OptimaCE-Oblique                       625.13
l_____OptimaCE-Bold                          625.14
l_____OptimaCE-BoldOblique                   625.15
l_____OptimaEastA-Roman                      625.16
l_____OptimaEastA-RomanOblique               625.17
l_____OptimaEastA-Bold                       625.18
l_____OptimaEastA-BoldOblique                625.19
l_____PaddingtonEF                           626.0
l_____PaddingtonEF-Alternate                 626.1
l_____PapyrusEF-Regular                      627.0
l_____PapyrusEF-Alternates                   627.1
l_____Peignot-Light                          628.0

l_____Peignot-Demi                           628.1
l_____Peignot-Bold                           628.2
l_____PendryScriptEF-Regular                 629.0
l_____PendryScriptEF-Alternates              629.1
l_____PiccadillyEF                           630.0
l_____Poplar                                 631.0
l_____PumpEF-Light                           632.0
l_____PumpEF-Medium                          632.1
l_____PumpEF-Demi                            632.2
l_____PumpEF-Bold                            632.3
l_____PumpEF-Triline                         632.4
l_____RailroadGothic                         633.0
l_____RotisSansSerif-Light                   634.0
l_____RotisSansSerif-LightItalic             634.1
l_____RotisSansSerif                         634.2
l_____RotisSansSerif-Italic                  634.3
l_____RotisSansSerif-Bold                    634.4
l_____RotisSansSerif-ExtraBold               634.5
l_____RotisSerif                             634.6
l_____RotisSerif-Italic                      634.7
l_____RotisSerif-Bold                        634.8
l_____RotisSemiSans-Light                    634.9
l_____RotisSemiSans-LightItalic              634.10
l_____RotisSemiSans                          634.11
l_____RotisSemiSans-Italic                   634.12
l_____RotisSemiSans-Bold                     634.13
l_____RotisSemiSans-ExtraBold                634.14
l_____RotisSemiSerif                         634.15
l_____RotisSemiSerif-Bold                    634.16
l_____RunicMT-Condensed                      635.0
l_____SassoonPrimary                         636.0
l_____SerlioLH                               637.0
l_____ShelleyEF                              638.0
l_____SierraEF-Regular                       639.0
l_____SierraEF-RegularItalic                 639.1
l_____SierraEF-Bold                          639.2
l_____SierraEF-BoldItalic                    639.3
l_____SlipstreamEF                           640.0
l_____Souvenir-Light                         641.0
l_____Souvenir-LightItalic                   641.1
l_____Souvenir-Medium                        641.2
l_____Souvenir-MediumItalic                  641.3
l_____Souvenir-Demi                          641.4
l_____Souvenir-DemiItalic                    641.5
l_____Souvenir-Bold                          641.6
l_____Souvenir-BoldItalic                    641.7
l_____SouvenirGreek-Light                    641.8

l_____SouvenirGreek-Demi                     641.9
l_____SouvenirGreekP-Light                   641.10
l_____SouvenirGreekP-Demi                    641.11
l_____SouvenirEF-Light                       641.12
l_____SouvenirEF-LightItalic                 641.13
l_____SouvenirEF-Medium                      641.14
l_____SouvenirEF-MediumItalic                641.15
l_____SouvenirEF-Demi                        641.16
l_____SouvenirEF-DemiItalic                  641.17
l_____SouvenirEF-Bold                        641.18
l_____SouvenirEF-BoldItalic                  641.19
l_____SouvenirMonoEF-Light                   641.20
l_____SouvenirMonoEF-LightItalic             641.21
l_____SouvenirMonoEF-Bold                    641.22
l_____SpringfieldEF                          642.0
l_____StentorEF                              643.0
l_____StymieEF-Light                         644.0
l_____StymieEF-Medium                        644.1
l_____StymieEF-Bold                          644.2
l_____StymieEF-Black                         644.3
l_____StymieEF-MediumCondensed               644.4
l_____StymieEF-BoldCondensed                 644.5
l_____SynchroEF                              645.0
l_____SynchroEF-Alternate                    645.1
l_____TeaChest                               646.0
l_____Times-Roman                            648.0
l_____Times-RomanSC                          648.1
l_____Times-Italic                           648.2
l_____Times-ItalicOsF                        648.3
l_____Times-SemiBold                         648.4
l_____Times-SemiBoldItalic                   648.5
l_____Times-Bold                             648.6
l_____Times-BoldSC                           648.7
l_____Times-BoldItalic                       648.8
l_____Times-BoldItalicOsF                    648.9
l_____TimesTen-Roman                         648.10
l_____TimesTen-RomanSC                       648.11
l_____TimesTen-RomanOsF                      648.12
l_____TimesTen-Italic                        648.13
l_____TimesTen-ItalicOsF                     648.14
l_____TimesTen-Bold                          648.15
l_____TimesTen-BoldOsF                       648.16
l_____TimesTen-BoldItalic                    648.17
l_____TimesTen-BoldItalicOsF                 648.18
l_____TimesEighteen-Roman                    648.19
l_____TimesEighteen-Bold                     648.20
l_____Times-ExtraBold                        648.21

l_____TimesEuropa-Roman                      648.22
l_____TimesEuropa-Italic                     648.23
l_____TimesEuropa-Bold                       648.24
l_____TimesEuropa-BoldItalic                 648.25
l_____Times-PhoneticIPA                      648.26
l_____Times-PhoneticAlternate                648.27
l_____TimesCE-Roman                          648.28
l_____TimesCE-Italic                         648.29
l_____TimesCE-Bold                           648.30
l_____TimesCE-BoldItalic                     648.31
l_____TimesTenCE-Roman                       648.32
l_____TimesTenCE-Italic                      648.33
l_____TimesTenCE-Bold                        648.34
l_____TimesTenCE-BoldItalic                  648.35
l_____TimesTenCyr-Upright                    648.36
l_____TimesTenCyr-Inclined                   648.37
l_____TimesTenCyr-Bold                       648.38
l_____TimesTenCyr-BoldInclined               648.39
l_____TimesTenEastA-Roman                    648.40
l_____TimesTenEastA-Italic                   648.41
l_____TimesTenEastA-Bold                     648.42
l_____TimesTenEastA-BoldItalic               648.43
l_____TimesTenGreek-Upright                  648.44
l_____TimesTenGreek-Inclined                 648.45
l_____TimesTenGreek-Bold                     648.46
l_____TimesTenGreek-BdInclined               648.47
l_____TimesTenGreekP-Upright                 648.48
l_____TimesTenGreekP-Inclined                648.49
l_____TimesTenGreekP-Bold                    648.50
l_____TimesTenGreekP-BoldInclined            648.51
l_____TimesModernEF                          648.52
l_____TimesNewRomanPS-Roman                  648.53
l_____TimesNewRomanPS-Italic                 648.54
l_____TimesNewRomanPS-Bold                   648.55
l_____TimesNewRomanPS-BoldItalic             648.56
l_____TimesNewRomanMT-Cond                   648.57
l_____TimesNewRomanMT-CondItalic             648.58
l_____TimesNewRomanMT-BoldCond               648.59
l_____TradeGothic-Light                      649.0
l_____TradeGothic-LightOblique               649.1
l_____TradeGothic                            649.2
l_____TradeGothic-Oblique                    649.3
l_____TradeGothic-Bold                       649.4
l_____TradeGothic-BoldTwo                    649.5
l_____TradeGothic-BoldOblique                649.6
l_____TradeGothic-BoldTwoOblique             649.7
l_____TradeGothic-CondEighteen               649.8

l_____TradeGothic-CondEighteenObl            649.9
l_____TradeGothic-BoldCondTwenty             649.10
l_____TradeGothic-BoldCondTwentyObl          649.11
l_____TradeGothicLH-Extended                 649.12
l_____Utopia-Regular                         651.0
l_____Utopia-RegularSC                       651.1
l_____UtopiaExp-Regular                      651.2
l_____Utopia-Italic                          651.3
l_____Utopia-ItalicOsF                       651.4
l_____UtopiaExp-Italic                       651.5
l_____Utopia-Semibold                        651.6
l_____Utopia-SemiboldSC                      651.7
l_____UtopiaExp-Semibold                     651.8
l_____Utopia-SemiboldItalic                  651.9
l_____Utopia-SemiboldItalicOsF               651.10
l_____UtopiaExp-SemiboldItalic               651.11
l_____Utopia-Bold                            651.12
l_____Utopia-BoldOsF                         651.13
l_____UtopiaExp-Bold                         651.14
l_____Utopia-BoldItalic                      651.15
l_____Utopia-BoldItalicOsF                   651.16
l_____UtopiaExp-BoldItalic                   651.17
l_____Utopia-Black                           651.18
l_____Utopia-BlackOsF                        651.19
l_____Utopia-BlackExpert                     651.20
l_____Utopia-Headline                        651.21
l_____Utopia-TitlingCapitals                 651.22
l_____Utopia-Ornaments                       651.23
l_____VenusMager                             652.0
l_____VenusKursivMager                       652.1
l_____VenusLinkskursiv                       652.2
l_____VenusHalbfett                          652.3
l_____VenusKursivHalbfett                    652.4
l_____VenusEgypHalbfett                      652.5
l_____VenusEgypHalbfettKurs                  652.6
l_____VoltaEF-Regular                        653.0
l_____VoltaEF-Medium                         653.1
l_____VoltaEF-MediumItalic                   653.2
l_____VoltaEF-Bold                           653.3
l_____Weiss                                  654.0
l_____Weiss-Italic                           654.1
l_____Weiss-Bold                             654.2
l_____Weiss-ExtraBold                        654.3
l_____WeissAntiquaEF-Regular                 654.4
l_____WeissAntiquaEF-RegularItalic           654.5
l_____WeissAntiquaEF-Medium                  654.6
l_____WeissAntiquaEF-Bold                    654.7

l_____WeissAntiquaEF-ExtraBold               654.8
l_____WeissAntiquaEF-UltraBold               654.9
l_____WeissAntiquaEF-BoldCondensed           654.10
l_____WeissAntiquaEF-ExtraBoldCondensed      654.11
l_____Wilke-Roman                            655.0
l_____Wilke-Italic                           655.1
l_____Wilke-Bold                             655.2
l_____Wilke-BoldItalic                       655.3
l_____Wilke-Black                            655.4
l_____Wilke-BlackItalic                      655.5
l_____Woodstock                              656.0
l_____Bellevue                               657.0
l_____BinnerEF                               658.0
l_____Boton-Light                            659.0
l_____Boton-LightItalic                      659.1
l_____Boton-Regular                          659.2
l_____Boton-Italic                           659.3
l_____Boton-Medium                           659.4
l_____Boton-MediumItalic                     659.5
l_____Boton-Bold                             659.6
l_____Boton-BoldItalic                       659.7
l_____AachenCE-Bold                         9501.0
l_____AdHoc                                 9501.1
l_____AdHoc-Bis                             9501.2
l_____AdHoc-SC                              9501.3
l_____AdHoc-BisItalic                       9501.4
l_____AdHoc-Italic                          9501.5
l_____LinotypeAmelia                        9501.8
l_____Andreas                               9501.9
l_____BaskervilleClassico                   9501.13
l_____BaskervilleClassico-Italic            9501.14
l_____BaskervilleClassico-BoldItalic        9501.15
l_____BaskervilleClassico-SC                9501.16
l_____BaskervilleClassico-Bold              9501.17
l_____LinotypeBecket                        9501.18
l_____Bernhard-BoldCondensed                9501.19
l_____Birka                                 9501.20
l_____Birka-SC                              9501.21
l_____Birka-SemiBoldItalic                  9501.22
l_____Birka-SemiBold                        9501.23
l_____Birka-BoldItalic                      9501.24
l_____Birka-Italic                          9501.25
l_____Birka-Bold                            9501.26
l_____LinotypeBlackWhiteHeadline            9501.27
l_____LinotypeBlackWhiteOutline             9501.28
l_____LinotypeBlackWhiteLaserOutline        9501.29
l_____LinotypeBlackWhite                    9501.30

l_____LinotypeBlackWhiteLaser               9501.31
l_____BodoniHE-Light                        9501.32
l_____BodoniHE-LightItalic                  9501.33
l_____BodoniHE-SemiBold                     9501.34
l_____BodoniHE-SemiBoldItalic               9501.35
l_____BodoniHE-BookItalic                   9501.36
l_____BodoniHE-Book                         9501.37
l_____BodoniHE-Bold                         9501.38
l_____BodoniHE-ExtraBold                    9501.39
l_____BodoniHE-BoldItalic                   9501.40
l_____BodoniHE-SemiBoldCondensedItalic      9501.41
l_____BodoniHE-SemiBoldCondensed            9501.42
l_____BodoniHE-BoldCondensedItalic          9501.43
l_____BodoniHE-BoldCondensed                9501.44
l_____BodoniHE-ExtraBoldItalic              9501.45
l_____BodoniClassico                        9501.46
l_____BodoniClassico-BoldItalic             9501.47
l_____BodoniClassico-SC                     9501.48
l_____BodoniClassico-Bold                   9501.49
l_____BodoniClassico-Italic                 9501.50
l_____Carniola                              9501.52
l_____Carniola-Bold                         9501.53
l_____Carniola-SemiBold                     9501.54
l_____Carniola-SC                           9501.55
l_____Carniola-SemiBoldItalic               9501.56
l_____Carniola-BoldItalic                   9501.57
l_____Carniola-Italic                       9501.58
l_____CaslonHE-Light                        9501.59
l_____CaslonHE-LightItalic                  9501.60
l_____CaslonHE-Italic                       9501.61
l_____CaslonHE-Roman                        9501.62
l_____CaslonClassico                        9501.63
l_____CaslonClassico-Bold                   9501.64
l_____CaslonClassico-SC                     9501.65
l_____CaslonClassico-BoldItalic             9501.66
l_____CaslonClassico-Italic                 9501.67
l_____Century-HandtooledBold                9501.68
l_____Century-HandtooledBoldItalic          9501.69
l_____NewCenturySchlbkCyr-Roman             9501.70
l_____NewCenturySchlbkCyr-Bold              9501.71
l_____NewCenturySchlbkCyr-Italic            9501.72
l_____NewCenturySchlbkCyr-BoldItalic        9501.73
l_____Cheltenham-HandtooledBd               9501.74
l_____Cheltenham-HandtooledBdIt             9501.75
l_____CloisterOpenFace                      9501.76
l_____Linotype CMC-Seven                    9501.77
l_____CooperBlackCE-Regualr                 9501.78

l_____CooperBlackCE-Italic                  9501.79
l_____CourierCyr                            9501.80
l_____CourierCyr-Oblique                    9501.81
l_____CourierCyr-BoldOblique                9501.82
l_____CourierCyr-Bold                       9501.83
l_____LinotypeDavida                        9501.84
l_____Devin                                 9501.95
l_____Devin-Bold                            9501.96
l_____Devin-Italic                          9501.97
l_____Devin-SemiBold                        9501.98
l_____Devin-SC                              9501.99
l_____Devin-SemiBoldItalic                  9501.100
l_____Devin-BoldItalic                      9501.101
l_____Dialog-Condensed                      9501.102
l_____Dialog-SemiBoldSC                     9501.103
l_____Dialog-LightSC                        9501.104
l_____Dialog-ExtraBoldItalic                9501.105
l_____Dialog-ExtraBold                      9501.106
l_____Dialog-SemiBoldItalic                 9501.107
l_____Dialog-SemiBold                       9501.108
l_____Dialog-LightItalic                    9501.109
l_____Dialog-Light                          9501.110
l_____Dialog-SC                             9501.111
l_____Dialog-BoldItalic                     9501.112
l_____Dialog-Bold                           9501.113
l_____Dialog-Italic                         9501.114
l_____Dialog                                9501.115
l_____DialogCond-Bold                       9501.116
l_____DialogCond-Italic                     9501.117
l_____DialogCond-BoldItalic                 9501.118
l_____DialogCond-SemiBoldItalic             9501.119
l_____DialogCond-SemiBold                   9501.120
l_____DigiAntiqua-LightCondensed            9501.121
l_____DigiGrotesk-BoldCondensed             9501.122
l_____EmonaCond                             9501.124
l_____EmonaCond-Bold                        9501.125
l_____EmonaCond-BoldItalic                  9501.126
l_____Emona-Bold                            9501.127
l_____Emona-SemiBold                        9501.128
l_____Emona-SC                              9501.129
l_____Emona-SemiBoldItalic                  9501.130
l_____Emona-Italic                          9501.131
l_____Emona                                 9501.132
l_____EmonaCond-Italic                      9501.133
l_____Emona-BoldItalic                      9501.134
l_____EsperantoCond                         9501.135
l_____Esperanto                             9501.136

l_____Esperanto-Bold                        9501.137
l_____Esperanto-SemiBold                    9501.138
l_____Esperanto-SC                          9501.139
l_____Esperanto-SemiBoldItalic              9501.140
l_____Esperanto-BoldItalic                  9501.141
l_____Esperanto-Italic                      9501.142
l_____EsperantoCond-SC                      9501.143
l_____EsperantoCond-Bold                    9501.144
l_____EsperantoCond-Italic                  9501.145
l_____EsperantoCond-BoldItalic              9501.146
l_____ExcelsiorInsertio                     9501.147
l_____F2FAlRetto                            9501.148
l_____F2FBoneR-Book                         9501.149
l_____F2FCzykagoSemiSerif                   9501.150
l_____F2FLego-Stoned                        9501.151
l_____F2FLovegrid                           9501.152
l_____F2FMadame-Butterfly                   9501.153
l_____F2FMadZine-Dirt                       9501.154
l_____F2FMadZine-Script                     9501.155
l_____F2FOCR-AlexczykShake                  9501.156
l_____OCR-BczykNorm                         9501.157
l_____OCR-BczykBold                         9501.158
l_____F2FPrototipa-Multipla                 9501.159
l_____F2FScreen-Scream                      9501.160
l_____F2FShakkarakk                         9501.161
l_____F2FShpeetz                            9501.162
l_____F2FTagliatelle-Sugo                   9501.163
l_____F2FTwins                              9501.164
l_____F2FTyrell-Corp                        9501.165
l_____Linotype FehrleDisplay                9501.166
l_____Fortuna                               9501.167
l_____Fortuna-Bold                          9501.168
l_____FranklinGothic-BookCmpr               9501.169
l_____FranklinGothic-DemiCmprItal           9501.170
l_____FranklinGothic-DemiXCmpr              9501.171
l_____FranklinGothic-BookXCmpr              9501.172
l_____FranklinGothic-DemiCndItal            9501.173
l_____FranklinGothic-DemiCnd                9501.174
l_____FranklinGothic-MediumCndItal          9501.175
l_____FranklinGothic-MediumCnd              9501.176
l_____FranklinGothic-BookCndItal            9501.177
l_____FranklinGothic-BookCnd                9501.178
l_____FranklinGothic-DemiCmpr               9501.179
l_____FranklinGothic-BookCmprItal           9501.180
l_____FreestyleScriptCE                     9501.181
l_____FrutigerCE-LightItalic                9501.182
l_____FrutigerCE-Roman                      9501.183

l_____FrutigerCE-Black                      9501.184
l_____FrutigerCE-Italic                     9501.185
l_____FrutigerCE-Light                      9501.186
l_____FrutigerCE-UltraBlack                 9501.187
l_____FrutigerCE-BlackItalic                9501.188
l_____FrutigerCE-BoldItalic                 9501.189
l_____FrutigerCE-Bold                       9501.190
l_____FuturaCE-Light                        9501.191
l_____FuturaCE-Bold                         9501.192
l_____FuturaCE-Medium                       9501.193
l_____FuturaCE-BoldOblique                  9501.194
l_____FuturaCE-LightOblique                 9501.195
l_____FuturaCE-ExtraBoldOblique             9501.196
l_____FuturaCE-ExtraBold                    9501.197
l_____FuturaCE-HeavyOblique                 9501.198
l_____FuturaCE-Heavy                        9501.199
l_____FuturaCE-MediumOblique                9501.200
l_____FuturaCE-BookOblique                  9501.201
l_____FuturaCE-Book                         9501.202
l_____Garamond-HandtooledBold               9501.206
l_____Garamond-HandtooledBoldItalic         9501.207
l_____GaramondClassico                      9501.208
l_____GaramondClassico-BoldItalic           9501.209
l_____GaramondClassico-SC                   9501.210
l_____GaramondClassico-Italic               9501.211
l_____GaramondClassico-Bold                 9501.212
l_____GillSansCE-Light                      9501.213
l_____GillSansCE-Roman                      9501.214
l_____GillSansCE-Bold                       9501.215
l_____GillSansCE-BoldItalic                 9501.216
l_____GillSansCE-Italic                     9501.217
l_____GillSansCE-LightItalic                9501.218
l_____GoudyModern94                         9501.219
l_____GoudyModern94-Italic                  9501.220
l_____GoudyModern94-Bold                    9501.221
l_____GoudyModern94-BoldItalic              9501.222
l_____GoudyModern94-SC                      9501.223
l_____GoudyVillage                          9501.224
l_____GoudyVillage-SC                       9501.225
l_____GoudyVillage-BoldItalic               9501.226
l_____GoudyVillage-Italic                   9501.227
l_____GoudyVillage-Bold                     9501.228
l_____GriffoClassico                        9501.229
l_____GriffoClassico-Bold                   9501.230
l_____GriffoClassico-BoldItalic             9501.231
l_____GriffoClassico-SC                     9501.232
l_____GriffoClassico-Italic                 9501.233

l_____HelveticaNeueCE-Light                 9501.238
l_____HelveticaNeueCE-Black                 9501.239
l_____HelveticaNeueCE-UltraLightIt          9501.240
l_____HelveticaNeueCE-BlackItalic           9501.241
l_____HelveticaNeueCE-Medium                9501.242
l_____HelveticaNeueCE-Heavy                 9501.243
l_____HelveticaNeueCE-HeavyItalic           9501.244
l_____HelveticaNeueCE-UltraLight            9501.245
l_____HelveticaNeueCE-MediumItalic          9501.246
l_____HelveticaNeueCE-LightItalic           9501.247
lpi___Linotype HieroglyphsPi-One            9501.248
lpi___Linotype HieroglyphsPi-Two            9501.249
l_____Humana                                9501.250
l_____Humana-SC                             9501.251
l_____Isolde                                9501.252
l_____Isolde-Italic                         9501.253
l_____Isolde-BoldItalic                     9501.254
l_____Isolde-SC                             9501.255
l_____Isolde-Bold                           9501.256
l_____JensonClassico                        9501.257
l_____JensonClassico-Italic                 9501.258
l_____JensonClassico-BoldItalic             9501.259
l_____JensonClassico-Bold                   9501.260
l_____JensonClassico-SC                     9501.261
l_____Jesper                                9501.262
l_____Jonatan                               9501.263
l_____Kalix                                 9501.264
l_____Kalix-SC                              9501.265
l_____Kalix-Italic                          9501.266
l_____Kalix-Bold                            9501.267
l_____Kalix-BoldItalic                      9501.268
l_____Kalix-SemiBoldItalic                  9501.269
l_____Kalix-SemiBold                        9501.270
l_____Kasper                                9501.271
lpi___MacKeyCaps                            9501.272
l_____KisClassico                           9501.281
l_____KisClassico-BoldItalic                9501.282
l_____KisClassico-SC                        9501.283
l_____KisClassico-Italic                    9501.284
l_____KisClassico-Bold                      9501.285
lpi___Linobats                              9501.288
l_____LucidaTypewriter                      9501.289
l_____LucidaSansTypewriter                  9501.290
l_____LucidaSansTypewriter-Bd               9501.291
l_____LucidaSansTypewriter-BdObl            9501.292
l_____LucidaTypewriter-BoldObl              9501.293
l_____LucidaTypewriter-Bold                 9501.294

l_____LucidaTypewriter-Obl                  9501.295
l_____LucidaSansTypewriter-Obl              9501.296
l_____LinotypeLutherscheFraktur             9501.297
l_____LinotypeLutherscheFraktur-Dfr         9501.298
l_____Manuskript                            9501.299
l_____Manuskript-BoldItalic                 9501.300
l_____Manuskript-SC                         9501.301
l_____Manuskript-Bold                       9501.302
l_____Manuskript-Italic                     9501.303
l_____MarcoPolo                             9501.304
l_____MarcoPolo-SC                          9501.305
l_____MarcoPolo-Italic                      9501.306
l_____Memento                               9501.307
l_____Memento-Italic                        9501.308
l_____Memento-SemiBold                      9501.309
l_____Memento-SemiBoldItalic                9501.310
l_____Memento-BoldItalic                    9501.311
l_____Memento-Bold                          9501.312
l_____Memento-SC                            9501.313
l_____Miramar                               9501.314
l_____Miramar-SC                            9501.315
l_____Miramar-BoldItalic                    9501.316
l_____Miramar-Bold                          9501.317
l_____Miramar-Italic                        9501.318
l_____Mojo                                  9501.319
l_____Monti-Roman                           9501.320
l_____Monti-SemiBold                        9501.321
l_____Napoleon-Light                        9501.322
l_____Napoleon-Roman                        9501.323
l_____Napoleon-Italic                       9501.324
l_____Napoleon-Bold                         9501.325
l_____Napoleon-BoldCondensed                9501.326
l_____Napoleon-SemiBold                     9501.327
l_____Neuland                               9501.328
l_____NormaCompr                            9501.329
l_____NormaCompr-Bold                       9501.330
l_____NormaCompr-Italic                     9501.331
l_____NormaCompr-SC                         9501.332
l_____NormaCond-Bold                        9501.333
l_____NormaCond                             9501.334
l_____Norma-SemiBoldItalic                  9501.335
l_____Norma-SemiBold                        9501.336
l_____Norma-BoldItalic                      9501.337
l_____Norma-Bold                            9501.338
l_____Norma-Italic                          9501.339
l_____Norma                                 9501.340
l_____NormaCond-SC                          9501.341

l_____NormaCond-Italic                      9501.342
l_____Norma-SC                              9501.343
l_____NormaCond-BoldItalic                  9501.344
l_____NormaCompr-BoldItalic                 9501.345
l_____Nyfors                                9501.346
l_____OdenseCompr                           9501.347
l_____OdenseCompr-Italic                    9501.348
l_____OdenseCond-Bold                       9501.349
l_____OdenseCond-Italic                     9501.350
l_____OdenseCond-BoldItalic                 9501.351
l_____Odense-Bold                           9501.352
l_____Odense-Italic                         9501.353
l_____Odense                                9501.354
l_____Odense-SemiBoldSC                     9501.355
l_____Odense-ExtraBoldItalic                9501.356
l_____Odense-ExtraBold                      9501.357
l_____Odense-SemiBoldItalic                 9501.358
l_____Odense-SemiBold                       9501.359
l_____Odense-SC                             9501.360
l_____Odense-BoldItalic                     9501.361
l_____Odense-Neon                           9501.362
l_____OdenseCond                            9501.363
l_____Omnibus                               9501.364
l_____Omnibus-Italic                        9501.365
l_____Omnibus-BoldItali                     9501.366
l_____Omnibus-SemiBoldItalic                9501.367
l_____Omnibus-SemiBold                      9501.368
l_____Omnibus-SC                            9501.369
l_____Omnibus-Bold                          9501.370
l_____Ondine                                9501.371
l_____Optima-Medium                         9501.372
l_____Optima-DemiBoldItalic                 9501.373
l_____Optima-BlackItalic                    9501.374
l_____Optima-ExtraBlackItalic               9501.375
l_____Optima-ExtraBlack                     9501.376
l_____Optima-Black                          9501.377
l_____Optima-DemiBold                       9501.378
l_____Optima-MediumItalic                   9501.379
l_____Ouch                                  9501.380
l_____LinotypePaintIt                       9501.381
l_____LinotypePaintIt-Black                 9501.382
l_____Pax                                   9501.383
l_____Pax-SemiBoldItalic                    9501.384
l_____Pax-SemiBold                          9501.385
l_____PaxCond-SC                            9501.386
l_____PaxCond-BoldItalic                    9501.387
l_____PaxCond-Bold                          9501.388

l_____PaxCond-Italic                        9501.389
l_____PaxCond                               9501.390
l_____Pax-SC                                9501.391
l_____Pax-Italic                            9501.392
l_____Pax-BoldItalic                        9501.393
l_____Pax-Bold                              9501.394
l_____PaxTwo                                9501.395
l_____PaxTwo-Italic                         9501.396
l_____PaxTwo-Bold                           9501.397
l_____PaxTwo-SemiBold                       9501.398
l_____PaxTwo-SC                             9501.399
l_____PaxTwo-SemiBoldItalic                 9501.400
l_____PaxTwo-BoldItalic                     9501.401
l_____Persona                               9501.402
l_____Pierrot                               9501.403
l_____Ragnar                                9501.406
l_____Ragnar-SC                             9501.407
l_____Ragnar-SemiBoldItalic                 9501.408
l_____Ragnar-Bold                           9501.409
l_____Ragnar-BoldItalic                     9501.410
l_____Ragnar-SemiBold                       9501.411
l_____Ragnar-Italic                         9501.412
l_____ResPublicaCond                        9501.413
l_____ResPublica-Italic                     9501.414
l_____ResPublica-BoldItalic                 9501.415
l_____ResPublica-SC                         9501.416
l_____ResPublica-Bold                       9501.417
l_____ResPublica                            9501.418
l_____ResPublicaCond-Italic                 9501.419
l_____ResPublica-SemiBoldItalic             9501.420
l_____ResPublica-SemiBold                   9501.421
l_____RevueCE                               9501.422
l_____Rustika                               9501.423
l_____Rustika-Bold                          9501.424
l_____Rustika-BoldItalic                    9501.425
l_____Rustika-Italic                        9501.426
l_____Rustika-SC                            9501.427
l_____SabonCE-Roman                         9501.428
l_____SabonCE-Italic                        9501.429
l_____SabonCE-BoldItalic                    9501.430
l_____SabonCyr-Bold                         9501.431
l_____SabonGreek-Roman                      9501.432
l_____SabonGreek-Bold                       9501.433
l_____SabonGreek-BoldItalic                 9501.434
l_____SabonGreek-Italic                     9501.435
l_____SabonCyr-Italic                       9501.436
l_____SabonCyr-Roman                        9501.437

l_____SabonCE-Bold                          9501.438
l_____SabonCyr-BoldItalic                   9501.439
l_____Saga                                  9501.440
l_____Saga-BoldItalic                       9501.441
l_____Saga-SemiBold                         9501.442
l_____Saga-SemiBoldItalic                   9501.443
l_____Saga-SC                               9501.444
l_____Saga-Bold                             9501.445
l_____Saga-Italic                           9501.446
l_____Linotype SallweyScript                9501.447
l_____SanMarcoCyr                           9501.448
l_____Saphir                                9501.449
l_____Semper                                9501.456
l_____Semper-Italic                         9501.457
l_____Semper-BoldItalic                     9501.458
l_____Semper-SC                             9501.459
l_____Semper-Bold                           9501.460
l_____Shuriken-Boy                          9501.461
l_____LinotypeSpitz-Light                   9501.462
l_____LinotypeSpitz-Book                    9501.463
l_____LinotypeSpitz-Black                   9501.464
l_____LinotypeSpitz-Bold                    9501.465
l_____LinotypeSpitz-Medium                  9501.466
l_____StencilCyr                            9501.467
l_____Tiemann-Light                         9501.468
l_____Tiemann-Roman                         9501.469
l_____Trajanus-Roman                        9501.470
l_____Trajanus-Black                        9501.471
l_____Trajanus-BlackItalic                  9501.472
l_____Trajanus-BoldItalic                   9501.473
l_____Trajanus-Italic                       9501.474
l_____Trajanus-Bold                         9501.475
l_____Transport                             9501.476
l_____Transport-BoldItalic                  9501.477
l_____Transport-Light                       9501.478
l_____Transport-LightItalic                 9501.479
l_____Transport-SC                          9501.480
l_____Transport-LightSC                     9501.481
l_____Transport-Bold                        9501.482
l_____Transport-Italic                      9501.483
l_____UniversCyr-Light                      9501.484
l_____UniversCyr-LightOblique               9501.485
l_____UniversCyr                            9501.486
l_____UniversCyr-Oblique                    9501.487
l_____UniversCyr-Bold                       9501.488
l_____UniversCyr-BoldOblique                9501.489
l_____UniversCyr-Black                      9501.490

l_____UniversCyr-BlackOblique               9501.491
l_____UniversityRomanCE                     9501.492
l_____UniversityCyr-Roman                   9501.493
l_____Vega                                  9501.494
l_____Vega-Bold                             9501.495
l_____Vega-SemiBold                         9501.496
l_____Vega-SC                               9501.497
l_____Vega-SemiBoldItalic                   9501.498
l_____Vega-BoldItalic                       9501.499
l_____Vega-Italic                           9501.500
l_____WiesbadenSwingCyr-Roman               9501.501

A.8 Monotype fonts

The ‘monotype.map’ file has abbreviations for Monotype fonts.

mapr8x    ApolloMT-Expert                   MT  1001    ezexp___
maprc8a   ApolloMT-SC                       MT  1001    ezi_____
mapri8x   ApolloMT-ItalicExpert             MT  1001    eziof___
mapri8a   ApolloMT-Italic                   MT  1001    ezixp___
mapr8a    ApolloMT                          MT  1001    ezr_____
maprij8a  ApolloMT-ItalicOsF                MT  1001    ezrsc___
maps8a    ApolloMT-SemiBold                 MT  1001    ezsb____
mapsj8a   ApolloMT-SemiBoldOsF              MT  1001    ezsbo___
maps8x    ApolloMT-SemiBoldExpert           MT  1001    ezsbx___
mnrrw8a   NewBerolinaMT                     MT  1002    _n______
mp1rw8a   PalaceScriptMT                    MT  1002    _p______
mp1sw8a   PalaceScriptMT-SemiBold           MT  1002    _psb____
mahrw8a   AshleyScriptMT                    MT  1002    le______
mm1rw8a   MonolineScriptMT                  MT  1002    on______
mbvr8a    BaskervilleMT                     MT  1003    bas_____
mbvb8a    BaskervilleMT-Bold                MT  1003    basb____
mbvbi8a   BaskervilleMT-BoldItalic          MT  1003    basbi___
mbvri8a   BaskervilleMT-Italic              MT  1003    basi____
mbvs8a    BaskervilleMT-SemiBold            MT  1003    bass____
mbvsi8a   BaskervilleMT-SemiBoldItalic      MT  1003    bassi___
mbbr8a    Bembo                             MT  1004    bm______
mbbb8a    Bembo-Bold                        MT  1004    bmb_____
mbbbi8a   Bembo-BoldItalic                  MT  1004    bmbi____
mbbri8a   Bembo-Italic                      MT  1004    bmi_____
mbbx8a    Bembo-ExtraBold                   MT  1005    bmeb____
mbbxi8a   Bembo-ExtraBoldItalic             MT  1005    bmebi___
mbbs8a    Bembo-SemiBold                    MT  1005    bmsb____
mbbsi8a   Bembo-SemiBoldItalic              MT  1005    bmsbi___
mdsrw8a   DorchesterScriptMT                MT  1006    dsscr___
mp2rw8a   PepitaMT                          MT  1006    epscr___

mbfrw8a   BiffoMT                           MT  1006    ifscr___
mswbw8a   SwingMT-Bold                      MT  1006    swib____
ms2bw8a   ScriptMT-Bold                     MT  1006    tbscr___
mb1ri8a   BladoMT-Italic                    MT  1007    blai____
mpzr8a    PoliphilusMT                      MT  1007    pol_____
mvdr8a    VanDijckMT                        MT  1007    van_____
mvdri8a   VanDijckMT-Italic                 MT  1007    vani____
mbdr8a    BodoniMT                          MT  1008    bod_____
mbdb8a    BodoniMT-Bold                     MT  1008    bodb____
mbdbi8a   BodoniMT-BoldItalic               MT  1008    bodbi___
mbdri8a   BodoniMT-Italic                   MT  1008    bodi____
mbdu8a    BodoniMT-UltraBold                MT  1009    bodu____
mbdui8a   BodoniMT-UltraBoldItalic          MT  1009    bodui___
mbdk8a    BodoniMT-Book                     MT  1009    bodw____
mbdki8a   BodoniMT-BookItalic               MT  1009    bodwi___
moxr8a    OnyxMT                            MT  1009    ox______
mbdb8ac   BodoniMT-CondensedBold            MT  1010    bodc____
mbdbi8ac  BodoniMT-CondensedBoldItalic      MT  1010    bodci___
mbdc8a    BodoniMT-Black                    MT  1010    bodk____
mbdci8a   BodoniMT-BlackItalic              MT  1010    bodki___
mb3r8a    BraggadocioMT                     MT  1011    bra_____
mfir8a    FigaroMT                          MT  1011    fig_____
mfer8a    ForteMT                           MT  1011    for_____
mklr8a    KlangMT                           MT  1011    kla_____
mc8b8a    CalvertMT-Bold                    MT  1012    vtb_____
mc8l8a    CalvertMT-Light                   MT  1012    vtl_____
mc8r8a    CalvertMT                         MT  1012    vtrg____
mcnb8a    CenturyMT-Bold                    MT  1013    cenb____
mcnbi8a   CenturyMT-BoldItalic              MT  1013    cenbi___
mcnb8ax   CenturyMT-Expanded                MT  1013    cenx____
mcnbi8ax  CenturyMT-ExpandedItalic          MT  1013    cenxi___
mcubi8a   CenturyOSMT-BoldItalic            MT  1015    cobi____
mcur8a    CenturyOSMT                       MT  1015    cos_____
mcub8a    CenturyOSMT-Bold                  MT  1015    cosb____
mcuri8a   CenturyOSMT-Italic                MT  1015    cosi____
mcdr8a    ClarendonMT                       MT  1016    cld_____
meer8a    EgyptianMT-Extended               MT  1016    egyex___
mnnr8a    NewClarendonMT                    MT  1016    ncl_____
mnnb8a    NewClarendonMT-Bold               MT  1016    nclb____
mc6r8a    ClarionMT                         MT  1017    cla_____
mc6b8a    ClarionMT-Bold                    MT  1017    clab____
mc6ri8a   ClarionMT-Italic                  MT  1017    clai____
mcfb8a    ClearfaceMT-Bold                  MT  1018    cleb____
mcfbs8a   ClearfaceGothicMT-Bold            MT  1018    clgb____
mcfds8a   ClearfaceGothicMT-DemiBold        MT  1018    clgdb___
mcrr8a12  CourierTwelveMT                   MT  1019    cot_____
mtyre8a   TypewriterEliteMT                 MT  1019    tye_____

mtyrg8a   TypewriterGothicMT                MT  1019    tyg_____
mtyr8a    TypewriterMT                      MT  1019    typ_____
metr8a    EhrhardtMT                        MT  1020    ehr_____
metri8a   EhrhardtMT-Italic                 MT  1020    ehri____
mets8a    EhrhardtMT-SemiBold               MT  1020    ehrs____
metsi8a   EhrhardtMT-SemiBoldItalic         MT  1020    ehrsi___
menr8a    EngraversMT                       MT  1021    eng_____
menb8a    EngraversMT-Bold                  MT  1021    engb____
menrf8a   EngraversOldEnglishMT             MT  1021    engoe___
mfsra8a   FalstaffMT                        MT  1022    fprg____
mhdr8a    HeadlineMT-Bold                   MT  1022    head____
mpdr8ac   PlacardMT-Condensed               MT  1022    plcc____
mpdb8ac   PlacardMT-CondensedBold           MT  1022    plccb___
mgmr8a    GaramondMT                        MT  1023    gar_____
mgmb8a    GaramondMT-Bold                   MT  1023    garb____
mgmbi8a   GaramondMT-BoldItalic             MT  1023    garbi___
mgmri8a   GaramondMT-Italic                 MT  1023    gari____
mgmria8a  GaramondMT-AlternativeItalic      MT  1023    garit___
mgsr8a    GillSans                          MT  1024    gn______
mgsb8a    GillSans-Bold                     MT  1024    gnb_____
mgsbi8a   GillSans-BoldItalic               MT  1024    gnbi____
mgsri8a   GillSans-Italic                   MT  1024    gni_____
mgsl8a    GillSans-Light                    MT  1024    gnl_____
mgsli8a   GillSans-LightItalic              MT  1024    gnli____
mgsb8ac   GillSans-BoldCondensed            MT  1025    gnbc____
mgsr8ac   GillSans-Condensed                MT  1025    gnc_____
mgsx8a    GillSans-ExtraBold                MT  1025    gneb____
mgsu8a    GillSans-UltraBold                MT  1025    gnub____
mgsu8ac   GillSans-UltraBoldCondensed       MT  1025    gnubc___
mgrb8a    GloucesMT-Bold                    MT  1026    glb_____
mgrb8ac   GloucesMT-CondBold                MT  1026    glcb____
mgrb8ax   GloucesMT-ExtendedBold            MT  1026    gleb____
mgrb8aq   GloucesMT-ExtraCondensedBold      MT  1026    glecb___
mgrr8a    GloucesOldStyleMT                 MT  1026    glos____
mg5r8a    GoudyCatalogueMT                  MT  1027    goc_____
mg5ri8a   GoudyCatalogueMT-Italic           MT  1027    goci____
mgor8a    GoudyOldStyleMT                   MT  1027    gos_____
mgori8a   GoudyOldStyleMT-Italic            MT  1027    gosi____
mgyb8a    GoudyMT-Bold                      MT  1027    goub____
mgybi8a   GoudyMT-BoldItalic                MT  1027    goubi___
mg4r8a    GoudyModernMT                     MT  1028    gom_____
mg4ri8a   GoudyModernMT-Italic              MT  1028    gomi____
mgyx8a    GoudyMT-ExtraBold                 MT  1028    goueb___
mgtr8a    GoudyTextMT                       MT  1028    goutx___
mgqr8a    GrotesqueMT                       MT  1029    gq______
mgqc8a    GrotesqueMT-Black                 MT  1029    gqb_____
mgqb8a    GrotesqueMT-Bold                  MT  1029    gqbl____

mgqri8a   GrotesqueMT-Italic                MT  1029    gqi_____
mgql8a    GrotesqueMT-Light                 MT  1029    gql_____
mgqli8a   GrotesqueMT-LightItalic           MT  1029    gqli____
mgqc8a    GrotesqueMT-Black                 MT  1030    gqbl____
mgqb8ax   GrotesqueMT-BoldExtended          MT  1030    gqbv____
mgqr8ac   GrotesqueMT-Condensed             MT  1030    gqc_____
mgqr8aq   GrotesqueMT-ExtraCondensed        MT  1030    gqec____
mgql8ac   GrotesqueMT-LightCondensed        MT  1030    gqlc____
miir8a    ImprintMT                         MT  1032    im______
miib8a    ImprintMT-Bold                    MT  1032    imb_____
miibi8a   ImprintMT-BoldItalic              MT  1032    imbi____
miiri8a   ImprintMT-Italic                  MT  1032    imi_____
mifr8a    InflexMT                          MT  1033    iqb_____
moer8a    OldEnglishTextMT                  MT  1033    mbbou___
mosbl8a   MonotOldStyleMT-BoldOutline       MT  1033    old_____
mior8a    IonicMT                           MT  1034    ion_____
miob8a    IonicMT-Bold                      MT  1034    ionb____
miori8a   IonicMT-Italic                    MT  1034    ioni____
misb8a    ItalianOldStyleMT-Bold            MT  1035    ieb_____
misbi8a   ItalianOldStyleMT-BoldItalic      MT  1035    iebi____
misri8a   ItalianOldStyleMT-Italic          MT  1035    iei_____
misr8a    ItalianOldStyleMT                 MT  1035    ierg____
mmob8a    ModernMT-Bold                     MT  1036    nmb_____
mmobi8a   ModernMT-BoldItalic               MT  1036    nmbi____
mmor8ax   ModernMT-Extended                 MT  1036    nmex____
mmori8ax  ModernMT-ExtendedItalic           MT  1036    nmexi___
mnpr8a    NewsPlantinMT                     MT  1038    npl_____
mnpb8a    NewsPlantinMT-Bold                MT  1038    nplb____
mnpbi8a   NewsPlantinMT-BoldItalic          MT  1038    nplbi___
mnpri8a   NewsPlantinMT-Italic              MT  1038    npli____
mosr8a    OldStyleMT                        MT  1039    ost_____
mosb8a    OldStyleMT-Bold                   MT  1039    ostb____
mosbi8a   OldStyleMT-BoldItalic             MT  1039    ostbi___
mosri8a   OldStyleMT-Italic                 MT  1039    osti____
mpnr8a    Plantin                           MT  1040    pl______
mpnb8a    Plantin-Bold                      MT  1040    plb_____
mpnbi8a   Plantin-BoldItalic                MT  1040    plbi____
mpnri8a   Plantin-Italic                    MT  1040    pli_____
mpnb8ac   Plantin-BoldCondensed             MT  1041    plbc____
mpnl8a    Plantin-Light                     MT  1041    pll_____
mpnli8a   Plantin-LightItalic               MT  1041    plli____
mpns8a    Plantin-SemiBold                  MT  1041    plsb____
mpnsi8a   Plantin-SemiBoldItalic            MT  1041    plsbi___
mrwr8a    Rockwell                          MT  1042    rk______
mrwb8a    Rockwell-Bold                     MT  1042    rkb_____
mrwbi8a   Rockwell-BoldItalic               MT  1042    rkbi____
mrwri8a   Rockwell-Italic                   MT  1042    rki_____

mrwl8a    Rockwell-Light                    MT  1042    rkl_____
mrwli8a   Rockwell-LightItalic              MT  1042    rkli____
mrwb8ac   Rockwell-BoldCondensed            MT  1043    rkbc____
mrwr8ac   Rockwell-Condensed                MT  1043    rkc_____
mrwx8a    Rockwell-ExtraBold                MT  1043    rkeb____
msbr8a    SabonMT                           MT  1044    sab_____
msbri8a   SabonMT-Italic                    MT  1044    sabi____
msbs8a    SabonMT-SemiBold                  MT  1044    sabs____
msbsi8a   SabonMT-SemiBoldItalic            MT  1044    sabsi___
mntr8a    TimesNRMT                         MT  1046    tim_____
mntb8a    TimesNRMT-Bold                    MT  1046    timb____
mntbi8a   TimesNRMT-BoldItalic              MT  1046    timbi___
mntri8a   TimesNRMT-Italic                  MT  1046    timi____
mntx8a    TimesNRMT-ExtraBold               MT  1047    timeb___
mnts8a    TimesNRMT-SemiBold                MT  1047    tims____
mntsi8a   TimesNRMT-SemiBoldItalic          MT  1047    timsi___
mntb8ac   TimesNewRomanMT-BoldCond          MT  1048    mtbc____
mntr8ac   TimesNewRomanMT-Cond              MT  1048    mtc_____
mntri8ac  TimesNewRomanMT-CondItalic        MT  1048    mtci____
mtwb8a    TwCenMT-Bold                      MT  1049    tcb_____
mtwbi8a   TwCenMT-BoldItalic                MT  1049    tcbi____
mtwl8a    TwCenMT-Light                     MT  1049    tcl_____
mtwli8a   TwCenMT-LightItalic               MT  1049    tcli____
mtwm8a    TwCenMT-Medium                    MT  1049    tcm_____
mtwmi8a   TwCenMT-MediumItalic              MT  1049    tcmi____
mtwb8ac   TwCenMT-CondensedBold             MT  1050    tccb____
mtwm8ac   TwCenMT-CondensedMedium           MT  1050    tccm____
mtws8a    TwCenMT-SemiBold                  MT  1050    tcsb____
mtwsm8a   TwCenMT-SemiMedium                MT  1050    tcsm____
mtwx8ac   TwCenMT-CondensedExtraBold        MT  1051    tcceb___
mtwu8ac   TwCenMT-CondensedUltraBold        MT  1051    tccub___
mtwxi8a   TwCenMT-ExtraBoldItalic           MT  1051    tcebi___
mtwu8a    TwCenMT-UltraBold                 MT  1051    tcub____
mtwui8a   TwCenMT-UltraBoldItalic           MT  1051    tcubi___
mtwx8a    TwCenMT-ExtraBold                 MT  1051    tweb____
mwbr8a    WalbaumMT                         MT  1052    wl______
mwbri8a   WalbaumMT-Italic                  MT  1052    wli_____
mwbm8a    WalbaumMT-Medium                  MT  1052    wlm_____
mwbmi8a   WalbaumMT-MediumItalic            MT  1052    wlmi____
mnsb8a    TimesNewRomanPS-Bold              MT  1053    mtb_____
mnsbi8a   TimesNewRomanPS-BoldItalic        MT  1053    mtbi____
mnsri8a   TimesNewRomanPS-Italic            MT  1053    mti_____
mnsr8a    TimesNewRomanPS                   MT  1053    mtr_____
ma1r8a    ArialMT                           MT  1054    ari_____
ma1b8a    ArialMT-Bold                      MT  1054    ariab___
ma1bi8a   ArialMT-BoldItalic                MT  1054    aribi___
ma1ri8a   ArialMT-Italic                    MT  1054    arii____

ma1r8an   ArialNarrowMT                     MT  1055    an______
ma1b8an   ArialNarrowMT-Bold                MT  1055    anb_____
ma1ri8an  ArialNarrowMT-BoldItalic          MT  1055    anbi____
ma1bi8an  ArialNarrowMT-Italic              MT  1055    ani_____
mcrr8a    CourierPS                         MT  1058    cr______
mcrb8a    CourierPS-Bold                    MT  1058    crb_____
mcrbi8a   CourierPS-BoldItalic              MT  1058    crbi____
mcrri8a   CourierPS-Italic                  MT  1058    cri_____
msyr      SymbolPS                          MT  1058    syps____
mnir8a    NimrodMT                          MT  1061    nim_____
mnib8a    NimrodMT-Bold                     MT  1061    nimb____
mnibi8a   NimrodMT-BoldItalic               MT  1061    nimbi___
mniri8a   NimrodMT-Italic                   MT  1061    nimi____
ma1ci8a   ArialMT-BlackItalic               MT  1064    arbli___
ma1c8a    ArialMT-Black                     MT  1064    aribl___
ma1l8a    ArialMT-Light                     MT  1064    aril____
ma1li8a   ArialMT-LightItalic               MT  1064    arili___
mburw8a   BrushScriptMT                     MT  1071    ________
mi0rw8a   CommercialScriptMT                MT  1071    ________
mr5rw8a   FrenchScriptMT                    MT  1071    ________
motb8a    CoronetMT-Bold                    MT  1071    corb____
mmebw8a   MercuriusMT-BoldScript            MT  1071    uzbs____
mppb8a    Perpetua-Bold                     MT  1072    ptb_____
mppbi8a   Perpetua-BoldItalic               MT  1072    ptbi____
mppri8a   Perpetua-Italic                   MT  1072    pti_____
mppr8a    Perpetua                          MT  1072    ptrg____
mair8ac   AbadiMT-Condensed                 MT  1074    abac____
maib8ac   AbadiMT-CondensedBold             MT  1074    abacb___
maix8ac   AbadiMT-CondensedExtraBold        MT  1074    abace___
mail8ac   AbadiMT-CondensedLight            MT  1074    abacl___
mtlr8a    CastellerMT                       MT  1075    el______
malr8a    AlbertusMT                        MT  1075    lr______
malri8a   AlbertusMT-Italic                 MT  1075    lri_____
mall8a    AlbertusMT-Light                  MT  1075    lrl_____
ma1x8a    ArialMT-ExtraBold                 MT  1076    arie____
ma1xi8a   ArialMT-ExtraBoldItalic           MT  1076    ariei___
ma1m8a    ArialMT-Medium                    MT  1076    arim____
ma1mi8a   ArialMT-MediumItalic              MT  1076    arimi___
ma1r8ac   ArialMT-Condensed                 MT  1077    aric____
ma1b8ac   ArialMT-CondensedBold             MT  1077    aricb___
ma1x8ac   ArialMT-CondensedExtraBold        MT  1077    arice___
ma1l8ac   ArialMT-CondensedLight            MT  1077    aricl___
mbvr8x    BskvillExpMT                      MT  1078    bay_____
mbvb8x    BskvillExpMT-Bold                 MT  1078    bayb____
mbvbi8x   BskvillExpMT-BoldItalic           MT  1078    baybi___
mbvri8x   BskvillExpMT-Italic               MT  1078    bayi____
mbvs8x    BskvillExpMT-SemiBold             MT  1078    bays____

mbvsi8x   BskvillExpMT-SemiBoldItal         MT  1078    baysi___
mber8a    BellMT                            MT  1079    bl______
mbeb8a    BellMT-Bold                       MT  1079    blb_____
mbebi8a   BellMT-BoldItalic                 MT  1079    blbi____
mberi8a   BellMT-Italic                     MT  1079    bli_____
mbes8a    BellMT-SemiBold                   MT  1079    bls_____
mbesi8a   BellMT-SemiBoldItalic             MT  1079    blsi____
mbbr7a    Bembo-Alt                         MT  1080    bea_____
mbbb7a    Bembo-AltBold                     MT  1080    beab____
mbbbi7a   Bembo-AltBoldItalic               MT  1080    beabi___
mbbri7a   Bembo-AltItalic                   MT  1080    beai____
mbbb8x    BemboExpert-Bold                  MT  1080    bey_____
mbbbi8x   BemboExpert-BoldItalic            MT  1080    beyb____
mbbri8x   BemboExpert-Italic                MT  1080    beybi___
mbbr8x    BemboExpert-Roman                 MT  1080    beyi____
mbbx7a    Bembo-AltExtraBold                MT  1081    beae____
mbbs7a    Bembo-AltSemiBold                 MT  1081    beas____
mbbsi7a   Bembo-AltSemiBoldItalic           MT  1081    beasi___
mbbx8x    BemboExpert-ExtraBold             MT  1081    beye____
mbbxi8x   BemboExpert-ExtraBoldItalic       MT  1081    beyei___
mbbs8x    BemboExpert-SemiBold              MT  1081    beys____
mbbsi8x   BemboExpert-SemiBoldItalic        MT  1081    beysi___
mbnr8ac   BernardMT-Condensed               MT  1082    berc____
mc7b8a    CompactaMT-Bold                   MT  1082    comb____
mnkr8a    NeographikMT                      MT  1082    neo_____
mrur8ac   RunicMT-Condensed                 MT  1082    runc____
metr8x    EhrhardtExptMT                    MT  1083    ehy_____
metri8x   EhrhardtExptMT-Italic             MT  1083    ehyi____
mets8x    EhrhardtExptMT-SemiBold           MT  1083    ehys____
metsi8x   EhrhardtExptMT-SemiBoldItal       MT  1083    ehysi___
mjnr8a    JansonMT                          MT  1085    jan_____
mjnb8a    JansonMT-Bold                     MT  1085    janb____
mjnbi8a   JansonMT-BoldItalic               MT  1085    janbi___
mjnri8a   JansonMT-Italic                   MT  1085    jani____
mppr8x    Perpetua-Expert                   MT  1087    pey_____
mpprc8a   Perpetua-SC                       MT  1087    ptsc____
mpnbi8x   PlantinExp-BoldItalic             MT  1088    plbiy___
mpnb8x    PlantinExp-Bold                   MT  1088    plby____
mpnri8x   PlantinExp-Italic                 MT  1088    pliy____
mpnr8x    PlantinExp                        MT  1088    ply_____
mpnra8x   PlantinExp-AltFigures             MT  1089    plya____
mpnl8x    PlantinExp-Light                  MT  1089    plyl____
mpnli8x   PlantinExp-LightItalic            MT  1089    plyli___
mpns8x    PlantinExp-SemiBold               MT  1089    plys____
mpnsi8x   PlantinExp-SemiBoldItalic         MT  1089    plysi___
zmsnr34   Spartan-ThreeFourMT               MT  1091    spf_____
zmsnr12   Spartan-OneTwoMT                  MT  1091    spo_____

zmsnr4    Spartan-FourMT                    MT  1091    spr_____
zmsnr13   Spartan-OneThreeMT                MT  1091    spt_____
zmsnr24   Spartan-TwoFourMT                 MT  1091    spw_____
zmsnb34   Spartan-ThreeFourMT-Bold          MT  1092    spfb____
zmsnb12   Spartan-OneTwoMT-Bold             MT  1092    spob____
zmsnb4    Spartan-FourMT-Bold               MT  1092    sprb____
zmsnb13   Spartan-OneThreeMT-Bold           MT  1092    sptb____
zmsnb24   Spartan-TwoFourMT-Bold            MT  1092    spwb____
zmsnri34  Spartan-ThreeFourMT-Italic        MT  1093    spfi____
zmsnri12  Spartan-OneTwoMT-Italic           MT  1093    spoi____
zmsnri4   Spartan-FourMT-Italic             MT  1093    spri____
zmsnri13  Spartan-OneThreeMT-Italic         MT  1093    spti____
zmsnri24  Spartan-TwoFourMT-Italic          MT  1093    spwi____
mntm8a    TimesNRMT-Medium                  MT  1094    timm____
mntmi8a   TimesNRMT-MediumItalic            MT  1094    timmi___
mtmrg8a   TimesSmallTextMT                  MT  1094    tst_____
mtmbg8a   TimesSmallTextMT-Bold             MT  1094    tstb____
mtmrig8a  TimesSmallTextMT-Italic           MT  1094    tsti____
melr8a    EllingtonMT                       MT  1095    ei______
melb8a    EllingtonMT-Bold                  MT  1095    eib_____
melbi8a   EllingtonMT-BoldItalic            MT  1095    eibi____
melx8a    EllingtonMT-ExtraBold             MT  1095    eieb____
melxi8a   EllingtonMT-ExtraBoldItalic       MT  1095    eiebi___
melri8a   EllingtonMT-Italic                MT  1095    eii_____
mell8a    EllingtonMT-Light                 MT  1095    eil_____
melli8a   EllingtonMT-LightItalic           MT  1095    eili____
mntr8x    TimesNRExpertMT                   MT  1096    tiy_____
mntb8x    TimesNRExpertMT-Bold              MT  1096    tiyb____
mntbi8x   TimesNRExpertMT-BoldItalic        MT  1096    tiybi___
mntri8x   TimesNRExpertMT-Italic            MT  1096    tiyi____
mair8a    AbadiMT                           MT  1098    aba_____
maib8a    AbadiMT-Bold                      MT  1098    abab____
maibi8a   AbadiMT-BoldItalic                MT  1098    ababi___
mairi8a   AbadiMT-Italic                    MT  1098    abai____
maij8a    AbadiMT-ExtraLight                MT  1098    abel____
maiji8a   AbadiMT-ExtraLightItalic          MT  1098    abeli___
mail8a    AbadiMT-Light                     MT  1099    abal____
maili8a   AbadiMT-LightItalic               MT  1099    abali___
maix8a    AbadiMT-ExtraBold                 MT  1099    abeb____
maixi8a   AbadiMT-ExtraBoldItalic           MT  1099    abebi___
mntrk     TimesNRDualGreekMT                MT  1100    td______
mntbk     TimesNRDualGreekMT-Bold           MT  1100    tdb_____
mntbik    TimesNRDualGreekMT-BoldIncl       MT  1100    tdbi____
mntrik    TimesNRDualGreekMT-Inclined       MT  1100    tdi_____
mgmrwd8a  GaramondMT-SwashCaps              MT  1101    garsc___
mgmrw8a   GaramondMT-Swash                  MT  1101    gasw____
mgmr8x    GaramondExpertMT                  MT  1101    gay_____

mgmb8x    GaramondExpertMT-Bold             MT  1101    gayb____
mgmbi8x   GaramondExpertMT-BoldItalic       MT  1101    gaybi___
mgmri8x   GaramondExpertMT-Italic           MT  1101    gayi____
mb2r8a    BinnyOldStyleMT                   MT  1102    bin_____
mb2ri8a   BinnyOldStyleMT-Italic            MT  1102    bini____
murb8a    CentaurMT-Bold                    MT  1103    nrb_____
murbi8a   CentaurMT-BoldItalic              MT  1103    nrbi____
murri8a   CentaurMT-Italic                  MT  1103    nri_____
murr8a    CentaurMT                         MT  1103    nrr_____
murd8a    CentaurFestiveMT-Italic
mc1r8a    CalistoMT                         MT  1104    cto_____
mc1b8a    CalistoMT-Bold                    MT  1104    ctob____
mc1bi8a   CalistoMT-BoldItalic              MT  1104    ctobi___
mc1ri8a   CalistoMT-Italic                  MT  1104    ctoi____
mr2b8a    CantoriaMT-Bold                   MT  1105    eob_____
mr2bi8a   CantoriaMT-BoldItalic             MT  1105    eobi____
mr2ri8a   CantoriaMT-Italic                 MT  1105    eoi_____
mr2r8a    CantoriaMT                        MT  1105    eorg____
mr2x8a    CantoriaMT-ExtraBold              MT  1106    eoeb____
mr2xi8a   CantoriaMT-ExtraBoldItalic        MT  1106    eoebi___
mr2l8a    CantoriaMT-Light                  MT  1106    eol_____
mr2li8a   CantoriaMT-LightItalic            MT  1106    eoli____
mr2s8a    CantoriaMT-SemiBold               MT  1106    eosb____
mr2si8a   CantoriaMT-SemiBoldItalic         MT  1106    eosbi___
mfxr8a    FelixTitlingMT                    MT  1107    felt____
mpprd8a   PerpetuaTitlingMT                 MT  1107    pert____
mppbd8a   PerpetuaTitlingMT-Bold            MT  1107    pertb___
mppld8a   PerpetuaTitlingMT-Light           MT  1107    pertl___
mf4r8a    FootlightMT                       MT  1108    ft______
mf4b8a    FootlightMT-Bold                  MT  1108    ftb_____
mf4bi8a   FootlightMT-BoldItalic            MT  1108    ftbi____
mf4ri8a   FootlightMT-Italic                MT  1108    fti_____
mf4x8a    FootlightMT-ExtraBold             MT  1109    fteb____
mf4xi8a   FootlightMT-ExtraBoldItalic       MT  1109    ftebi___
mf4l8a    FootlightMT-Light                 MT  1109    ftl_____
mf4li8a   FootlightMT-LightItalic           MT  1109    ftli____
mjnr8x    JansonExpertMT                    MT  1111    jny_____
mjnb8x    JansonExpertMT-Bold               MT  1111    jnyb____
mjnbi8x   JansonExpertMT-BoldItalic         MT  1111    jnybi___
mjnri8x   JansonExpertMT-Italic             MT  1111    jnyi____
mf3rd8a   FestivalTitlingMT                 MT  1112    fst_____
mknr8a    KinoMT                            MT  1112    kin_____
mm7r8a    MaturaMT                          MT  1112    mat_____
mm7rw7a   MaturaMT-ScriptCapitals           MT  1112    matsc___
mvird8ac  VictoriaTitlingMT-Condensed       MT  1112    vict____
mztr8a    ZeitgeistMT                       MT  1115    zei_____
mztr7a    ZeitgeistMT-Alternative           MT  1115    zeial___

zmztcm8x  ZeitgeistExpertMT-Cameo           MT  1115    zeiay___
mztb8a    ZeitgeistMT-Bold                  MT  1115    zeib____
mztb8x    ZeitgeistExpertMT-Bold            MT  1115    zeiby___
mztr8ac   ZeitgeistMT-Condensed             MT  1115    zeic____
zmztcm8a  ZeitgeistMT-Cameo                 MT  1115    zeica___
zmztcz8a  ZeitgeistMT-CrazyPaving           MT  1115    zeicp___
mztr8xc   ZeitgeistExpertMT-Condensed       MT  1115    zeicy___
mztri8a   ZeitgeistMT-Italic                MT  1115    zeii____
mztri8x   ZeitgeistExpertMT-Italic          MT  1115    zeiiy___
mztr8x    ZeitgeistExpertMT                 MT  1115    zeiy____
murriw7a  CentaurSwashMT                    MT  1116    censw___
murr7a    CentaurMT-Alt                     MT  1116    cta_____
murb7a    CentaurMT-AltBold                 MT  1116    ctab____
murbi7a   CentaurMT-AltBoldItalic           MT  1116    ctabi___
murri7a   CentaurMT-AltItalic               MT  1116    ctai____
murb8x    CentaurExpertMT-Bold              MT  1116    ctbiy___
murbi8x   CentaurExpertMT-BoldItalic        MT  1116    ctrby___
murri8x   CentaurExpertMT-Italic            MT  1116    ctriy___
murr8x    CentaurExpertMT                   MT  1116    ctry____
molr8a    ColmcilleMT                       MT  1117    cm______
molr7a    ColmcilleAltMT                    MT  1117    cma_____
molb7a    ColmcilleAltMT-Bold               MT  1117    cmab____
molbi7a   ColmcilleAltMT-BoldItalic         MT  1117    cmabi___
molri7a   ColmcilleAltMT-Italic             MT  1117    cmai____
molb8a    ColmcilleMT-Bold                  MT  1117    cmb_____
molbi8a   ColmcilleMT-BoldItalic            MT  1117    cmbi____
molri8a   ColmcilleMT-Italic                MT  1117    cmi_____
molrp     ColmcilleOrnamentsMT              MT  1117    cmo_____
mntr8a7   TimesNRSevenMT                    MT  1118    sv______
mntb8a7   TimesNRSevenMT-Bold               MT  1118    svb_____
mntbi8a7  TimesNRSevenMT-BoldItalic         MT  1118    svbi____
mntri8a7  TimesNRSevenMT-Italic             MT  1118    svi_____
mo3r8a    ColonnaMT                         MT  1119    colo____
miirh8a   ImprintMT-Shadow                  MT  1119    is______
miirih8a  ImprintMT-ShadowItalic            MT  1119    isi_____
ma2r8a    AmasisMT                          MT  1120    ama_____
ma2b8a    AmasisMT-Bold                     MT  1120    amab____
ma2bi8a   AmasisMT-BoldItalic               MT  1120    amabi___
ma2ri8a   AmasisMT-Italic                   MT  1120    amai____
ma2c8a    AmasisMT-Black                    MT  1121    amabl___
ma2l8a    AmasisMT-Light                    MT  1121    amal____
ma2li8a   AmasisMT-LightItalic              MT  1121    amali___
ma2m8a    AmasisMT-Medium                   MT  1121    amam____
ma2mi8a   AmasisMT-MediumItalic             MT  1121    amami___
ma2ci8a   AmasisMT-BlackItalic              MT  1121    ambli___
msbrij8a  SabonMT-ItalicOSF                 MT  1124    saboi___
msbsj8a   SabonMT-SemiBoldOSF               MT  1124    sabos___

msbrc8a   SabonMT-SCOSF                     MT  1124    sabsc___
msbsij8a  SabonMT-SemiBoldItalicOSF         MT  1124    sasio___
mb1ri8x   BladoExpertMT-Italic              MT  1125    blaiy___
mpzr8x    PoliphilusExpertMT                MT  1125    poly____
mvdra8x   VanDijckExpertMT-AltFigs          MT  1125    vana____
mvdria8x  VanDijckExpertMT-AltFigsIta       MT  1125    vanai___
mvdr8x    VanDijckExpertMT                  MT  1125    vany____
mvdri8x   VanDijckExpertMT-Italic           MT  1125    vanyi___
mber7a    BellMT-Alt                        MT  1126    blt_____
mberi7a   BellMT-AltItalic                  MT  1126    blti____
mbes7a    BellMT-AltSemiBold                MT  1126    blts____
mbesi7a   BellMT-AltSemiBoldItalic          MT  1126    bltsi___
mber8x    BellExpertMT                      MT  1126    bly_____
mbera8x   BellExpertMT-Alt                  MT  1126    blya____
mbesa8x   BellExpertMT-AltSemiBold          MT  1126    blyas___
mbeb8x    BellExpertMT-Bold                 MT  1126    blyb____
mbebi8x   BellExpertMT-BoldItalic           MT  1126    blybi___
mberi8x   BellExpertMT-Italic               MT  1126    blyi____
mbes8x    BellExpertMT-SemiBold             MT  1126    blys____
mbesi8x   BellExpertMT-SemiBoldItalic       MT  1126    blysi___
mgsx8ac   GillSansMT-ExCondensedBold        MT  1133    gsecb___
mgslh8a   GillSansShadowedMT-Light          MT  1133    gsls____
mgsrh8a   GillSansMT-Shadow                 MT  1133    gss_____
mgsra8a   GillAltOneMT                      MT  1134    gla_____
mgsba8a   GillAltOneMT-Bold                 MT  1134    glab____
mgsbia8a  GillAltOneMT-BoldItalic           MT  1134    glabi___
mgsria8a  GillAltOneMT-Italic               MT  1134    glai____
mgsla8a   GillAltOneMT-Light                MT  1134    glal____
mgslia8a  GillAltOneMT-LightItalic          MT  1134    glali___
mdtr8a    DanteMT                           MT  1149    dn______
mdtb8a    DanteMT-Bold                      MT  1149    dnb_____
mdtbi8a   DanteMT-BoldItalic                MT  1149    dnbi____
mdtri8a   DanteMT-Italic                    MT  1149    dni_____
mdtm8a    DanteMT-Medium                    MT  1149    dnm_____
mdtmi8a   DanteMT-MediumItalic              MT  1149    dnmi____
mdtr7a    DanteMT-Alt                       MT  1150    dna_____
mdtb7a    DanteMT-AltBold                   MT  1150    dnab____
mdtbi7a   DanteMT-AltBoldItalic             MT  1150    dnabi___
mdtri7a   DanteMT-AltItalic                 MT  1150    dnai____
mdtm7a    DanteMT-AltMedium                 MT  1150    dnam____
mdtmi7a   DanteMT-AltMediumItalic           MT  1150    dnami___
mdtr8x    DanteExpertMT                     MT  1150    dne_____
mdtb8x    DanteExpertMT-Bold                MT  1150    dneb____
mdtbi8x   DanteExpertMT-BoldItalic          MT  1150    dnebi___
mdtri8x   DanteExpertMT-Italic              MT  1150    dnei____
mdtm8x    DanteExpertMT-Medium              MT  1150    dnem____
mdtmi8x   DanteExpertMT-MediumItalic        MT  1150    dnemi___

mdtrd8a   DanteTitlingMT                    MT  1150    dnt_____
maqr8a    BookAntiqua                       MT  1152    bk______
maqb8a    BookAntiqua-Bold                  MT  1152    bkb_____
maqbi8a   BookAntiqua-BoldItalic            MT  1152    bkbi____
maqri8a   BookAntiqua-Italic                MT  1152    bki_____
mkor8a    BookmanOldStyle                   MT  1153    bd______
mkob8a    BookmanOldStyle-Bold              MT  1153    bdb_____
mkobi8a   BookmanOldStyle-BoldItalic        MT  1153    bdbi____
mkori8a   BookmanOldStyle-Italic            MT  1153    bdi_____
mygr8a    CenturyGothic                     MT  1154    cg______
mygb8a    CenturyGothic-Bold                MT  1154    cgb_____
mygbi8a   CenturyGothic-BoldItalic          MT  1154    cgbi____
mygri8a   CenturyGothic-Italic              MT  1154    cgi_____
mcsr8a    CenturySchoolbook                 MT  1155    cb______
mcsb8a    CenturySchoolbook-Bold            MT  1155    cbb_____
mcsbi8a   CenturySchoolbook-BoldItalic      MT  1155    cbbi____
mcsri8a   CenturySchoolbook-Italic          MT  1155    cbi_____
mzcmi8a   MonotypeCorsiva                   MT  1156    mci_____
mzdr      MonotypeSorts                     MT  1156    ms______
mpi001    ArabesqueOrnamentsOneMT           MT  1157    araba___
mpi002    ArabesqueOrnamentsTwoMT           MT  1157    arabb___
mpi003    ArabesqueOrnamentsThreeMT         MT  1157    arabc___
mpi004    RococoOrnamentsOneMT              MT  1157    rcoa____
mpi005    RococoOrnamentsTwoMT              MT  1157    rcob____
mpi006    RococoOrnamentsThreeMT            MT  1157    rcoc____
mgtr7a    GoudyTextMT-Alternative           MT  1181    gouta___
mgtr7c    GoudyTextMT-Dfr                   MT  1181    goutd___
mgtrd8a   GoudyTextMT-LombardicCapitals     MT  1181    goutl___
mgtr8a    GoudyTextMT                       MT  1181    goutx___
mlir8ac   LatinCondensedMT                  MT  1182    lnc_____
moxr8a    OnyxMT                            MT  1182    ox______
mrur8ac   RunicMT-Condensed                 MT  1182    ric_____
mpi007    ChristmasOrnamentsMT-One          MT  1183    chra____
mpi008    ChristmasOrnamentsMT-Two          MT  1183    chrb____
mpi009    ChristmasOrnamentsMT-Three        MT  1183    chrc____
mpi010    ChristmasOrnamentsMT-Four         MT  1183    chrd____
mpi011    ChristmasOrnamentsMT-Five         MT  1183    chre____
mpi012    ChristmasOrnamentsMT-Six          MT  1183    chrf____
mpi013    ContemporaryOrnMT-One             MT  1183    cora____
mpi014    ContemporaryOrnMT-Two             MT  1183    corb____
mpi015    ContemporaryOrnMT-Three           MT  1183    corc____
mpi016    ContemporaryOrnMT-Four            MT  1183    cord____
mpi017    ContemporaryOrnMT-Five            MT  1183    core____
mntr9s    TimesNewRomanSF                   MT  1194    tsf_____
mntb9s    TimesNewRomanSF-Bold              MT  1194    tsfb____
mntbi9s   TimesNewRomanSF-BoldItalic        MT  1194    tsfbi___
mntri9s   TimesNewRomanSF-Italic            MT  1194    tsfi____

mu0rp1    ColumbusOrnamentsOneMT            MT  1196    cuora___
mu0rp2    ColumbusOrnamentsTwoMT            MT  1196    cuorb___
mu0r8x    ColumbusExpMT                     MT  1196    cuy_____
mu0b8x    ColumbusExpMT-Bold                MT  1196    cuyb____
mu0bi8x   ColumbusExpMT-BoldItalic          MT  1196    cuybi___
mu0ri8x   ColumbusExpMT-Italic              MT  1196    cuyi____
mu0s8x    ColumbusExpMT-SemiBold            MT  1196    cuys____
mu0si8x   ColumbusExpMT-SemiBoldItal        MT  1196    cuysi___
ma1rz     ArialCyrMT                        MT  1205    ac______
ma1bz     ArialCyrMT-Bold                   MT  1205    acb_____
ma1biz    ArialCyrMT-BoldInclined           MT  1205    acbi____
ma1riz    ArialCyrMT-Inclined               MT  1205    aci_____
mcrrz     CourierCyr                        MT  1206    cc______
mcrbz     CourierCyr-Bold                   MT  1206    ccb_____
mcrbiz    CourierCyr-BoldInclined           MT  1206    ccbi____
mcrriz    CourierCyr-Inclined               MT  1206    cci_____
mntrz     TimesNRomCyrMT                    MT  1207    trc_____
mntbz     TimesNRomCyrMT-Bold               MT  1207    trcb____
mntbiz    TimesNRomCyrMT-BoldInclined       MT  1207    trcbi___
mntriz    TimesNRomCyrMT-Inclined           MT  1207    trci____
mgsbd8a   GillSansDisplayMT-Bold            MT  1209    gdb_____
mgsbd8ac  GillSansDisplayMT-CondBold        MT  1209    gdcb____
mgsbd8ax  GillSansDisplayMT-ExtraBold       MT  1209    gdeb____
mg4ri8a   GoudyModernMT-Italic              MT  1210    ygi_____
mg4r8a    GoudyModernMT                     MT  1210    ygrg____
ms1ri8a   ScotchRomanMT-Italic              MT  1210    ysi_____
ms1r8a    ScotchRomanMT                     MT  1210    ysrg____
mfsr8a    FalstaffMT                        MT  1211    fprg____
mifb8a    InflexMT-Bold                     MT  1211    iqb_____
mm8bl8a   MonotypeOldStyleMT-BoldOut        MT  1211    mbbou___
mfer8a    ForteMT                           MT  1212    fz______
mklr8a    KlangMT                           MT  1212    kn______
mmebw8a   MercuriusMT-BoldScript            MT  1212    uzbs____
mjox8a    JoannaMT-ExtraBold                MT  1213    joae____
mjob8a    JoannaMT-Bold                     MT  1213    job_____
mjobi8a   JoannaMT-BoldItalic               MT  1213    jobi____
mjori8a   JoannaMT-Italic                   MT  1213    joi_____
mjor8a    JoannaMT                          MT  1213    jor_____
mjos8a    JoannaMT-SemiBold                 MT  1213    josb____
mjosi8a   JoannaMT-SemiBoldItalic           MT  1213    josbi___
mphb8a    PhotinaMT-Bold                    MT  1214    phb_____
mphbi8a   PhotinaMT-BoldItalic              MT  1214    phbi____
mphri8a   PhotinaMT-Italic                  MT  1214    phi_____
mphr8a    PhotinaMT                         MT  1214    phrg____
mphs8a    PhotinaMT-SemiBold                MT  1214    phsb____
mphsi8a   PhotinaMT-SemiBoldItalic          MT  1214    phsbi___
mphu8a    PhotinaMT-UltraBold               MT  1214    phub____

mphui8a   PhotinaMT-UltraBoldItalic         MT  1214    phubi___
msmr8x    SpectrumMT-Expert                 MT  1215    ekexp___
msmri8a   SpectrumMT-Italic                 MT  1215    eki_____
msmrij8a  SpectrumMT-ItalicOsF              MT  1215    ekio____
msmri8x   SpectrumMT-ItalicExpert           MT  1215    ekixp___
msmr8a    SpectrumMT                        MT  1215    ekrg____
msms8a    SpectrumMT-SemiBold               MT  1215    eksb____
msmsj8a   SpectrumMT-SemiBoldOsF            MT  1215    eksbo___
msms8x    SpectrumMT-SemiBoldExpert         MT  1215    eksbx___
msmrc8a   SpectrumMT-SC                     MT  1215    eksc____
mhrr8a    ArialRoundedMT                    MT  1216    arr_____
mhrb8a    ArialRoundedMT-Bold               MT  1216    arrb____
mhrx8a    ArialRoundedMT-ExtraBold          MT  1216    arreb___
mhrl8a    ArialRoundedMT-Light              MT  1216    arrl____
mpi101    AlmanacMT                         MT  1222    pia_____
mpi102    BotanicalMT                       MT  1222    pib_____
mpi103    DirectionsMT                      MT  1222    pie_____
mpi104    SportsOneMT                       MT  1222    pif_____
mpi105    SportsTwoMT                       MT  1222    pig_____
mpi106    TransportMT                       MT  1222    pij_____
mpi107    VacationMT                        MT  1222    pik_____
mpi201    CircleFrameMT                     MT  1222    pic_____
mpi202    CircleFrameNegMT                  MT  1222    pid_____
mpi203    SquareFrameMT                     MT  1222    pih_____
mpi204    SquareFrameNegMT                  MT  1222    pii_____
m_____PastonchiMT-RegularAlt            MT  1223    pna_____
m_____PastonchiMT-Italic                MT  1223    pni_____
m_____PastonchiMT-ItalicOsF             MT  1223    pnosi___
m_____PastonchiMT-Regular               MT  1223    pnr_____
m_____PastonchiMT-RegularSC             MT  1223    pnsc____
m_____PastonchiMT-Swash                 MT  1223    pnsw____
m_____PastonchiMT-SwashCaps             MT  1223    pnswc___
m_____PastonchiMT-Titling               MT  1223    pnt_____
m_____PastonchiMT-TitlingAlt            MT  1223    pnta____
m_____PastonchiMT-RegularExpert         MT  1223    pny_____
m_____PastonchiMT-ItalicExpert          MT  1223    pnyi____
movr8x    OctavianMT-Expert                 MT  1224    ocexp___
movri8a   OctavianMT-Italic                 MT  1224    oci_____
movrij8a  OctavianMT-ItalicOsF              MT  1224    ocio____
movri8x   OctavianMT-ItalicExpert           MT  1224    ocixp___
movr8a    OctavianMT                        MT  1224    ocrg____
movrc8a   OctavianMT-SC                     MT  1224    ocsc____
mnsr8x    TimesNRPSExpert                   MT  1225    tpy_____
mnsb8x    TimesNRPSExpert-Bold              MT  1225    tpyb____
mnsbi8x   TimesNRPSExpert-BoldItalic        MT  1225    tpybi___
mnsri8x   TimesNRPSExpert-Italic            MT  1225    tpyi____
ma1r9s    ArialSFMT                         MT  1226    asf_____

ma1b9s    ArialSFMT-Bold                    MT  1226    asfb____
ma1bi9s   ArialSFMT-BoldItalic              MT  1226    asfbi___
ma1ri9s   ArialSFMT-Italic                  MT  1226    asfi____
mdmb8a    DelimaMT-Bold                     MT  1234    dlb_____
mdmbi8a   DelimaMT-BoldItalic               MT  1234    dlbi____
mdmri8a   DelimaMT-Italic                   MT  1234    dli_____
mdml8a    DelimaMT-Light                    MT  1234    dll_____
mdmli8a   DelimaMT-LightItalic              MT  1234    dlli____
mdmr8a    DelimaMT                          MT  1234    dlr_____
mdmx8a    DelimaMT-ExtraBold                MT  1235    dleb____
mdmxi_DelimaMT-ExtraBoldItalic          MT  1235    dlebi___
mdms8a    DelimaMT-SemiBold                 MT  1235    dlsb____
mdmsi8a   DelimaMT-SemiBoldItalic           MT  1235    dlsbi___
mo4rMM8a  OceanSansMM                       MT  1236    ________
mo4riMM8a OceanSansMM-It                    MT  1236    ________
mlyr8a    LydianMT                          MT  1238    ly______
mlyrw8a   LydianCursiveMT                   MT  1238    lyc_____
mwrrf8a   WittenbergerFraktMT               MT  1239    wf______
mwrbf8a   WittenbergerFraktMT-Bd            MT  1239    wfb_____
mwrrf7c   WittenbergerFraktMT-BdDfr         MT  1239    wfbdf___
mwrbf7c   WittenbergerFraktMT-Dfr           MT  1239    wfdfr___
mgsh8a    GillSans-Heavy                    MT  1240    gilh____
mgshi8a   GillSans-HeavyItalic              MT  1240    gilhi___
mgsk8a    GillSans-Book                     MT  1240    gilk____
mgski8a   GillSans-BookItalic               MT  1240    gilki___
mgsha8a   GillAltOneMT-Heavy                MT  1241    glah____
mgshia8a  GillAltOneMT-HeavyItalic          MT  1241    glahi___
mgska8a   GillAltOneMT-Book                 MT  1241    glak____
mgskia8a  GillAltOneMT-BookItalic           MT  1241    glaki___
mf1r8a    FournierMT                        MT  1242    fn______
mf1r7a    FournierMT-Alt                    MT  1242    fna_____
mf1ri7a   FournierMT-AltItalic              MT  1242    fnai____
mf1ri8a   FournierMT-Italic                 MT  1242    fni_____
mf1rp     FournierOrnamentsMT               MT  1242    fno_____
mf1rd8a   FournierMT-TallCaps               MT  1242    fnt_____
mf1rid8a  FournierMT-TallCapsItalic         MT  1242    fnti____
mf1r8x    FournierExpertMT                  MT  1242    fny_____
mf1ri8x   FournierExpertMT-Italic           MT  1242    fnyi____
miir8a    ImprintExpertMT                   MT  1243    imy_____
miib8a    ImprintExpertMT-Bold              MT  1243    imyb____
miibi8a   ImprintExpertMT-BoldItalic        MT  1243    imybi___
miiri8a   ImprintExpertMT-Italic            MT  1243    imyi____
mpwr8a    PerrywoodMT                       MT  1244    py______
mpwb8a    PerrywoodMT-Bold                  MT  1244    pyb_____
mpwbi8a   PerrywoodMT-BoldItalic            MT  1244    pybi____
mpwri8a   PerrywoodMT-Italic                MT  1244    pyi_____
mpwx8a    PerrywoodMT-ExtraBold             MT  1245    pye_____

mpwxi8a   PerrywoodMT-ExtraBoldItalc        MT  1245    pyei____
mpwl8a    PerrywoodMT-Light                 MT  1245    pyl_____
mpwli8a   PerrywoodMT-LightItalic           MT  1245    pyli____
mpws8a    PerrywoodMT-SemiBold              MT  1245    pys_____
mpwsi8a   PerrywoodMT-SemiBoldItalic        MT  1245    pysi____
mphbij8a  PhotinaMT-BoldItalicOsF           MT  1246    phbio___
mphbj8a   PhotinaMT-BoldOsF                 MT  1246    phbo____
mphrij8a  PhotinaMT-ItalicOsF               MT  1246    phio____
mphrc8a   PhotinaMT-RegularSC               MT  1246    phsc____
mphsij8a  PhotinaMT-SemiBoldItalicOsF       MT  1246    phsio___
mphsj8a   PhotinaMT-SemiBoldOsF             MT  1246    phso____
mphuij8a  PhotinaMT-UltraBoldItOsF          MT  1246    phuio___
mphuj8a   PhotinaMT-UltraBoldOsF            MT  1246    phuo____
mphr8x    PhotinaMT-RegularExpert           MT  1246    phy_____
mphb8x    PhotinaMT-BoldExpert              MT  1246    phyb____
mphbi8x   PhotinaMT-BoldItalicExpert        MT  1246    phybi___
mphri8x   PhotinaMT-ItalicExpert            MT  1246    phyi____
mphs8x    PhotinaMT-SemiBoldExpert          MT  1246    phys____
mphsi8x   PhotinaMT-SemiBoldItExpert        MT  1246    physi___
mphu8x    PhotinaMT-UltraBoldExpert         MT  1246    phyu____
mphui8x   PhotinaMT-UltraBoldItExpert       MT  1246    phyui___
mbmb8a    BulmerMT-Bold                     MT  1247    ybb_____
mbmbi8a   BulmerMT-BoldItalic               MT  1247    ybbi____
mbmbd8a   BulmerMT-BoldDisplay              MT  1247    ybdb____
mbmbid8a  BulmerMT-BoldItalicDisplay        MT  1247    ybdbi___
mbmrd8a   BulmerMT-RegularDisplay           MT  1247    ybds____
mbmrid8a  BulmerMT-ItalicDisplay            MT  1247    ybdsi___
mbmri8a   BulmerMT-Italic                   MT  1247    ybi_____
mbmr8a    BulmerMT-Regular                  MT  1247    ybrg____
mbms8a    BulmerMT-SemiBold                 MT  1247    ybsb____
mbmsi8a   BulmerMT-SemiBoldItalic           MT  1247    ybsi____
mbmb7a    BulmerMT-BoldAlt                  MT  1248    ybab____
mbmbi7a   BulmerMT-BoldItalicAlt            MT  1248    ybabi___
mbmr7a    BulmerMT-RegularAlt               MT  1248    ybarg___
mbms7a    BulmerMT-SemiBoldAlt              MT  1248    ybasb___
mbmsi7a   BulmerMT-SemiBoldItalicAlt        MT  1248    ybasi___
mbmbd7a   BulmerMT-BoldDisplayAlt           MT  1248    ybbda___
mbmbid7a  BulmerMT-BoldItalDisplayAlt       MT  1248    ybbid___
mbmbi8x   BulmerMT-BoldItalicExpert         MT  1248    ybbix___
mbmb8x    BulmerMT-BoldExpert               MT  1248    ybbxp___
mbmri7a   BulmerMT-ItalicAlt                MT  1248    ybial___
mbmrid7a  BulmerMT-ItalicDisplayAlt         MT  1248    ybida___
mbmri8x   BulmerMT-ItalicExpert             MT  1248    ybixp___
mbmrd7a   BulmerMT-RegularDisplayAlt        MT  1248    ybrda___
mbmr8x    BulmerMT-RegularExpert            MT  1248    ybrgx___
mbmrc8a   BulmerMT-RegularSC                MT  1248    ybrsc___
mbmsc8a   BulmerMT-SemiBoldSC               MT  1248    ybsbs___

mbms8x    BulmerMT-SemiBoldExpert           MT  1248    ybsbx___
mbmsi8x   BulmerMT-SemiBoldItalExpert       MT  1248    ybsix___
mo4b8a    OceanSansMT-Bold                  MT  1249    oeb_____
mo4bi8a   OceanSansMT-BoldItalic            MT  1249    oebi____
mo4k8a    OceanSansMT-Book                  MT  1249    oek_____
mo4ki8a   OceanSansMT-BookItalic            MT  1249    oeki____
mo4x8a    OceanSansMT-ExtraBold             MT  1250    oeeb____
mo4xi8a   OceanSansMT-ExtraBoldItalic       MT  1250    oeebi___
mo4l8a    OceanSansMT-Light                 MT  1250    oel_____
mo4li8a   OceanSansMT-LightItalic           MT  1250    oeli____
mo4s8a    OceanSansMT-SemiBold              MT  1250    oesb____
mo4si8a   OceanSansMT-SemiBoldItalic        MT  1250    oesbi___
mmob8a    ModernMT-Bold                     MT  1259    nmb_____
mmobi8a   ModernMT-BoldItalic               MT  1259    nmbi____
mmor8ac   ModernMT-Condensed                MT  1259    nmc_____
mmori8ac  ModernMT-CondensedItalic          MT  1259    nmci____
mmor8ax   ModernMT-Extended                 MT  1259    nmex____
mmori8ax  ModernMT-ExtendedItalic           MT  1259    nmexi___
mmor8aw   ModernMT-Wide                     MT  1259    nmwd____
mmori8aw  ModernMT-WideItalic               MT  1259    nmwdi___
mwbr8x    WalbaumExpMT                      MT  1268    wly_____
mwbri8x   WalbaumExpMT-Italic               MT  1268    wlyi____
mwbm8x    WalbaumExpMT-Medium               MT  1268    wlym____
mwbmi8x   WalbaumExpMT-MediumItalic         MT  1268    wlymi___
mlgr8a    LetterGothicMT                    MT  1269    lc______
mlgb8a    LetterGothicMT-Bold               MT  1269    lcb_____
mlgbo8a   LetterGothicMT-BoldOblique        MT  1269    lcbi____
mlgro8a   LetterGothicMT-Oblique            MT  1269    lci_____
ma1rzn    ArialNarrowCyrMT                  MT  1270    anc_____
ma1bzn    ArialNarrowCyrMT-Bold             MT  1270    ancb____
ma1bizn   ArialNarrowCyrMT-BoldInc          MT  1270    ancbi___
ma1rizn   ArialNarrowCyrMT-Inclined         MT  1270    anci____
mygrz     CenturyGothicCyr                  MT  1271    cgc_____
mygbz     CenturyGothicCyr-Bold             MT  1271    cgcb____
mygbiz    CenturyGothicCyr-BoldInc          MT  1271    cgcbi___
mygriz    CenturyGothicCyr-Inclined         MT  1271    cgci____
mkorz     BookmanOldStyleCyr                MT  1272    bdc_____
mkobz     BookmanOldStyleCyr-Bold           MT  1272    bdcb____
mkobiz    BookmanOldStyleCyr-BoldInc        MT  1272    bdcbi___
mkoriz    BookmanOldStyleCyr-Inc            MT  1272    bdci____
mnirz     NimrodCyrMT                       MT  1273    nic_____
mnibz     NimrodCyrMT-Bold                  MT  1273    nicb____
mnibiz    NimrodCyrMT-BoldInclined          MT  1273    nicbi___
mniriz    NimrodCyrMT-Inclined              MT  1273    nici____
mptrz     PlantinCyrMT                      MT  1274    plc_____
mptbz     PlantinCyrMT-Bold                 MT  1274    plcb____
mptbiz    PlantinCyrMT-BoldInclined         MT  1274    plcbi___

mptriz    PlantinCyrMT-Inclined             MT  1274    plci____
maqrz     BookAntiquaCyr                    MT  1275    bkc_____
maqbz     BookAntiquaCyr-Bold               MT  1275    bkcb____
maqbiz    BookAntiquaCyr-BoldInclined       MT  1275    bkcbi___
maqriz    BookAntiquaCyr-Inclined           MT  1275    bkci____
mzcmiz    MonotypeCorsivaCyr                MT  1276    mcc_____
mzcmizaa  MonotypeCorsivaCyr-AltOne         MT  1276    mccaa___
mzcmizab  MonotypeCorsivaCyr-AltTwo         MT  1276    mccab___
mzcmizac  MonotypeCorsivaCyr-AltThree       MT  1276    mccac___
mcsrz     CenturySchoolbookCyr              MT  1277    cbc_____
mcsbz     CenturySchoolbookCyr-Bold         MT  1277    cbcb____
mcsbiz    CenturySchoolbookCyr-BolInc       MT  1277    cbcbi___
mcsriz    CenturySchoolbookCyr-Inc          MT  1277    cbci____
mp6r8a    BlueprintMT                       MT  1283    bp______
mp6b8a    BlueprintMT-Bold                  MT  1283    bpb_____
mp6bi8a   BlueprintMT-BoldItalic            MT  1283    bpbi____
mp6ri8a   BlueprintMT-Italic                MT  1283    bpi_____
mn1r8a    BinnerGothicMT                    MT  1289    bg______
mfgr8aq   FrankGothMT-ExtraCondensed        MT  1289    fgec____
m_____LightlineGothicMT                 MT  1289    lg______
mwnr8a    WindsorMT                         MT  1289    wds_____
mygr__CenturyGothicTur                  MT  1293    cgt_____
mygb__CenturyGothicTur-Bold             MT  1293    cgtb____
mygbi__CenturyGothicTur-BoldItalic       MT  1293    cgtbi___
mygri__CenturyGothicTur-Italic           MT  1293    cgti____
mygr__CenturyGothicCE                   MT  1293    cgu_____
mygb__CenturyGothicCE-Bold              MT  1293    cgub____
mygbi__CenturyGothicCE-BoldItalic        MT  1293    cgubi___
mygri__CenturyGothicCE-Italic            MT  1293    cgui____
ma1r__n   ArialNarrowTur                    MT  1294    ant_____
ma1b__n   ArialNarrowTur-Bold               MT  1294    antb____
ma1bi__n  ArialNarrowTur-BoldItalic         MT  1294    antbi___
ma1ri__n  ArialNarrowTur-Italic             MT  1294    anti____
ma1r__n   ArialNarrowCE                     MT  1294    anu_____
ma1b__n   ArialNarrowCE-Bold                MT  1294    anub____
ma1bi__n  ArialNarrowCE-BoldItalic          MT  1294    anubi___
ma1ri__n  ArialNarrowCE-Italic              MT  1294    anui____
maqr__BookAntiquaTur                    MT  1295    bkt_____
maqb__BookAntiquaTur-Bold               MT  1295    bktb____
maqbi__BookAntiquaTur-BoldItalic         MT  1295    bktbi___
maqri__BookAntiquaTur-Italic             MT  1295    bkti____
maqr__BookAntiquaCE                     MT  1295    bku_____
maqb__BookAntiquaCE-Bold                MT  1295    bkub____
maqbi__BookAntiquaCE-BoldItalic          MT  1295    bkubi___
maqri__BookAntiquaCE-Italic              MT  1295    bkui____
mkor__BookmanOldStyleTur                MT  1296    bdt_____
mkob__BookmanOldStyleTur-Bold           MT  1296    bdtb____

mkobi__BookmanOldStyleTur-BoldIta        MT  1296    bdtbi___
mkori__BookmanOldStyleTur-Italic         MT  1296    bdti____
mkor__BookmanOldStylCE                  MT  1296    bdu_____
mkob__BookmanOldStylCE-Bold             MT  1296    bdub____
mkobi__BookmanOldStylCE-BoldItalic       MT  1296    bdubi___
mkori__BookmanOldStylCE-Italic           MT  1296    bdui____
mcsr__CenturySchoolbookTur              MT  1297    cst_____
mcsb__CenturySchoolbookTur-Bold         MT  1297    cstb____
mcsbi__CenturySchoolbookTur-BolIta       MT  1297    cstbi___
mcsri__CenturySchoolbookTur-Italic       MT  1297    csti____
mcsr__CenturySchoolbookCE               MT  1297    csu_____
mcsb__CenturySchoolbookCE-Bold          MT  1297    csub____
mcsbi__CenturySchoolbookCE-BoldIta       MT  1297    csubi___
mcsri__CenturySchoolbookCE-Italic        MT  1297    csui____
mzcmi__MonotypeCorsivaTur                MT  1298    mcit____
mzcmi__MonotypeCorsivaCE                 MT  1298    mciu____
ma1rk     ArialDualGreekMT                  MT  1299    ard_____
ma1bk     ArialDualGreekMT-Bold             MT  1299    ardb____
ma1bik    ArialDualGreekMT-BoldIncl         MT  1299    ardbi___
ma1rik    ArialDualGreekMT-Inclined         MT  1299    ardi____
ma1r__ArialTur                          MT  1306    art_____
ma1b__ArialTur-Bold                     MT  1306    artb____
ma1bi__ArialTur-BoldItalic               MT  1306    artbi___
ma1ri__ArialTur-Italic                   MT  1306    arti____
ma1r__ArialCE                           MT  1306    aru_____
ma1b__ArialCE-Bold                      MT  1306    arub____
ma1bi__ArialCE-BoldItalic                MT  1306    arubi___
ma1ri__ArialCE-Italic                    MT  1306    arui____
mntr__TimesNewRomanTur                  MT  1307    tit_____
mntb__TimesNewRomanTur-Bold             MT  1307    titb____
mntbi__TimesNewRomanTur-BoldItalic       MT  1307    titbi___
mntri__TimesNewRomanTur-Italic           MT  1307    titi____
mntr__TimesNewRomanCE                   MT  1307    tiu_____
mntb__TimesNewRomanCE-Bold              MT  1307    tiub____
mntbi__TimesNewRomanCE-BoldItalic        MT  1307    tiubi___
mntri__TimesNewRomanCE-Italic            MT  1307    tiui____
mcrr__CourierTur                        MT  1308    crt_____
mcrb__CourierTur-Bold                   MT  1308    crtb____
mcrbi__CourierTur-BoldItalic             MT  1308    crtbi___
mcrri__CourierTur-Italic                 MT  1308    crti____
mcrr__CourierCE                         MT  1308    cru_____
mcrb__CourierCE-Bold                    MT  1308    crub____
mcrbi__CourierCE-BoldItalic              MT  1308    crubi___
mcrri__CourierCE-Italic                  MT  1308    crui____
mpnrn8a   PlantinMT-Schlbk                  MT  1309    psb_____
mpnbn8a   PlantinMT-SchlbkBold              MT  1309    psbb____
mpnbin8a  PlantinMT-SchlbkBoldItalic        MT  1309    psbbi___

mpnrin8a  PlantinMT-SchlbkItalic            MT  1309    psbi____
mn1p8a    BinnerPosterMT                    MT  1310    binp____
mgqbd8ac  GrotesqueDisplayMT-BoldCond       MT  1310    gdbc____
mgqbd8ax  GrotesqueDisplayMT-BoldExt        MT  1310    gdbe____
micr8a    ImpactMT                          MT  1310    imp_____
mjobj8a   JoannaMT-BoldOSF                  MT  1315    joob____
mjobij8a  JoannaMT-BoldItalicOSF            MT  1315    joobi___
mjoxj8a   JoannaMT-ExtraBoldOSF             MT  1315    jooeb___
mjorij8a  JoannaMT-ItalicOSF                MT  1315    jooi____
mjosc8a   JoannaMT-SemiBoldSCOSF            MT  1315    joos____
mjosij8a  JoannaMT-SemiBoldItalicOSF        MT  1315    joosi___
mjorc8a   JoannaMT-SCOSF                    MT  1315    josc____
mjor8x    JoannaMT-Expert                   MT  1315    joy_____
mjob8x    JoannaMT-BoldExpert               MT  1315    joyb____
mjobi8x   JoannaMT-BoldItalicExpert         MT  1315    joybi___
mjox8x    JoannaMT-ExtraBoldExpert          MT  1315    joyeb___
mjori8x   JoannaMT-ItalicExpert             MT  1315    joyi____
mjos8x    JoannaMT-SemiBoldExpert           MT  1315    joys____
mjosi8a   JoannaMT-SemiBoldItaExpert        MT  1315    joysi___
mgsrk     GillSansDualGreekMT               MT  1329    gsd_____
mgsbk     GillSansDualGreekMT-Bold          MT  1329    gsdb____
mgsbik    GillSansDualGreekMT-BoldInc       MT  1329    gsdbi___
mgsrik    GillSansDualGreekMT-Incl          MT  1329    gsdi____
mnir__NimrodTur                         MT  1330    nit_____
mnib__NimrodTur-Bold                    MT  1330    nitb____
mnibi__NimrodTur-BoldItalic              MT  1330    nitbi___
mniri__NimrodTur-Italic                  MT  1330    niti____
mnir__NimrodCE                          MT  1330    niu_____
mnib__NimrodCE-Bold                     MT  1330    niub____
mnibi__NimrodCE-BoldItalic               MT  1330    niubi___
mniri__NimrodCE-Italic                   MT  1330    niui____
mbbrn8a   BemboMT-Schlbk                    MT  1331    bsc_____
mbbbn8a   BemboMT-SchlbkBold                MT  1331    bscb____
mbbbin8a  BemboMT-SchlbkBoldItalic          MT  1331    bscbi___
mbbrin8a  BemboMT-SchlbkItalic              MT  1331    bsci____
mntrz     TimesNewRomanCyr                  MT  1332    tpc_____
mntbz     TimesNewRomanCyr-Bold             MT  1332    tpcb____
mntbiz    TimesNewRomanCyr-BoldIncl         MT  1332    tpcbi___
mntriz    TimesNewRomanCyr-Inclined         MT  1332    tpci____
mbbrd8a   BemboTitlingMT                    MT  1333    bet_____
mbbrid8a  BemboTitlingMT-Italic             MT  1333    beti____
mcard8a   CaslonTitlingMT                   MT  1333    cat_____
mpnrd8a   PlantinTitlingMT                  MT  1333    plt_____
mpnbd8ac  PlantinHeadlineMT-BoldCond        MT  1334    plhbc___
mpnld8ac  PlantinHeadlineMT-LightCond       MT  1334    plhlc___
mpnmd8ac  PlantinHeadlineMT-MedCond         MT  1334    plhmc___
zms7rn    SassoonInfant                     MT  1335    szi_____

zms7rna   SassoonInfant-Alt                 MT  1335    szia____
zms7bn    SassoonInfant-Bold                MT  1335    szib____
zms7bna   SassoonInfant-BoldAlt             MT  1335    sziba___
zms7rp    SassoonPrimary                    MT  1336    szp_____
zms7rpa   SassoonPrimary-Alt                MT  1336    szpa____
zms7bp    SassoonPrimary-Bold               MT  1336    szpb____
zms7bpa   SassoonPrimary-BoldAlt            MT  1336    szpba___
zms7rs    SassoonSans                       MT  1337    szs_____
zms7rsa   SassoonSans-Alt                   MT  1337    szsa____
zms7bs    SassoonSans-Bold                  MT  1337    szsb____
zms7bsa   SassoonSans-BoldAlt               MT  1337    szsba___
zms7ms    SassoonSans-Medium                MT  1337    szsm____
zms7msa   SassoonSans-MediumAlt             MT  1337    szsma___
zms7rsi   SassoonSanSlope                   MT  1338    szz_____
zms7rsia  SassoonSanSlope-Alt               MT  1338    szza____
zms7bsi   SassoonSanSlope-Bold              MT  1338    szzb____
zms7bsia  SassoonSanSlope-BoldAlt           MT  1338    szzba___
zms7msi   SassoonSanSlope-Medium            MT  1338    szzm____
zms7msia  SassoonSanSlope-MediumAlt         MT  1338    szzma___
mgsr__GillSansCE                        MT  1339    gsc_____
mgsb__GillSansCE-Bold                   MT  1339    gscb____
mgsbi__GillSansCE-BoldItalic             MT  1339    gscbi___
mgsri__GillSansCE-Italic                 MT  1339    gsci____
mgsl__GillSansCE-Light                  MT  1339    gscl____
mgsli__GillSansCE-LightItalic            MT  1339    gscli___
mgsr__GillSansTur                       MT  1339    gst_____
mgsb__GillSansTur-Bold                  MT  1339    gstb____
mgsbi__GillSansTur-BoldItalic            MT  1339    gstbi___
mgsri__GillSansTur-Italic                MT  1339    gsti____
mgsl__GillSansTur-Light                 MT  1339    gstl____
mgsli__GillSansTur-LightItalic           MT  1339    gstli___
myhr8a    StrayhornMT-Regular               MT  1340    str_____
myhb8a    StrayhornMT-Bold                  MT  1340    strb____
myhbi8a   StrayhornMT-BoldItalic            MT  1340    strbi___
myhri8a   StrayhornMT-Italic                MT  1340    stri____
myhl8a    StrayhornMT-Light                 MT  1340    strl____
myhli8a   StrayhornMT-LightItalic           MT  1340    strli___
myhx8a    StrayhornMT-ExtraBold             MT  1340    strx____
myhxi8a   StrayhornMT-ExtraBoldItalic       MT  1340    strxi___
mgsrn8a   GillSansMT-Schlbk                 MT  1341    gsk_____
mgsbn8a   GillSansMT-SchlbkBold             MT  1341    gskb____
mgsbin8a  GillSansMT-SchlbkBoldItalic       MT  1341    gskbi___
mgsrin8a  GillSansMT-SchlbkItalic           MT  1341    gski____
myhbij8a  StrayhornMT-BoldItalicOsF         MT  1342    stbis___
myhbj8a   StrayhornMT-BoldOsF               MT  1342    stbs____
myhrij8a  StrayhornMT-ItalicOsF             MT  1342    stis____
myhlij8a  StrayhornMT-LightItalicOsF        MT  1342    stlis___

myhlc8a   StrayhornMT-LightSC               MT  1342    stls____
myhrc8a   StrayhornMT-RegularSC             MT  1342    sts_____
myhxij8a  StrayhornMT-ExtraBoldItaOsF       MT  1342    stxis___
myhxj8a   StrayhornMT-ExtraBoldOsF          MT  1342    stxs____
mpi205    DiamondBodoni                     MT  1343    dibod___
mpi206    DiamondPositive                   MT  1343    dipos___
mpi207    DiamondNegative                   MT  1343    dineg___
mf2r8ac   FacadeMT-Condensed                MT  1343    fac_____
mfsra8a   FalstaffFestivalMT                MT  1343    falfe___
mg4r8a    GoudyModernMT-Bold                MT  1344    gomb____
mg4ri8a   GoudyModernMT-BoldItalic          MT  1344    gombi___
mgyrp     GoudySortsMT                      MT  1344    gous____
mpi303    BonApetit                         MT  1345    bonap___
mpi306    Holidays                          MT  1345    holid___
mpi308    KeystrokesMT                      MT  1345    keya____
mpi307    KeystrokesExtrasMT                MT  1345    keyb____
mpi323    KeystrokesShadowMT                MT  1345    keyc____
mpi322    KeystrokesShadowExtrasMT          MT  1345    keyd____
mpi309    Melodies                          MT  1345    melod___
mpi312    MonkeyBusiness                    MT  1345    monke___
mpi315    Parties                           MT  1345    parti___
mpi320    SignsMT                           MT  1345    sig_____
mpi108    SportsThreeMT                     MT  1345    sport___
mpi302    AquaLifeMT                        MT  1346    aqua____
mpi304    DinosaursMT                       MT  1346    dino____
mpi305    FrontierMT                        MT  1346    frnt____
mpi310    MilestonesOneMT                   MT  1346    milea___
mpi311    MilestonesTwoMT                   MT  1346    mileb___
mpi314    NatureMT                          MT  1346    nat_____
mpi208    OvalFrameMT                       MT  1346    ovaa____
mpi209    OvalFrameNegativeMT               MT  1346    ovab____
mpi319    PlaytimeMT                        MT  1346    play____
mpi109    SportsFourMT                      MT  1346    sporfm__
mpi321    SouthwestOrnamentsMT              MT  1346    sworna__
ma1rt8a   ArialMonospacedMT                 MT  1347    arm_____
ma1bt8a   ArialMonospacedMT-Bold            MT  1347    armb____
ma1bot8a  ArialMonospacedMT-BoldObliq       MT  1347    armbo___
ma1rot8a  ArialMonospacedMT-Oblique         MT  1347    armo____
mpwr8ac   PerrywoodMT-Condensed             MT  1348    pwc_____
mpwb8ac   PerrywoodMT-BoldCondensed         MT  1348    pwcb____
mpwbi8ac  PerrywoodMT-BoldItalicCon         MT  1348    pwcbi___
mpwri8ac  PerrywoodMT-ItalicCondensed       MT  1348    pwci____
mpwx8ac   PerrywoodMT-ExtraBoldCon          MT  1349    pwce____
mpwxi8ac  PerrywoodMT-ExtraBoldItaCon       MT  1349    pwcei___
mpwl8ac   PerrywoodMT-LightCondensed        MT  1349    pwcl____
mpwli8ac  PerrywoodMT-LightItalicCon        MT  1349    pwcli___
mpws8ac   PerrywoodMT-SemiBoldCon           MT  1349    pwcs____

mpwsi8ac  PerrywoodMT-SemiBoldItaCon        MT  1349    pwcsi___
mpwr8ax   PerrywoodMT-Expanded              MT  1350    pwe_____
mpwb8ax   PerrywoodMT-BoldExpanded          MT  1350    pweb____
mpwbi8ax  PerrywoodMT-BoldItalicExp         MT  1350    pwebi___
mpwri8ax  PerrywoodMT-ItalicExpanded        MT  1350    pwei____
mpwx8ax   PerrywoodMT-ExtraBoldExp          MT  1351    pwee____
mpwxi8ax  PerrywoodMT-ExtraBoldItaExp       MT  1351    pweei___
mpwl8ax   PerrywoodMT-LightExpanded         MT  1351    pwel____
mpwli8ax  PerrywoodMT-LightItalicExp        MT  1351    pweli___
mpws8ax   PerrywoodMT-SemiBoldExp           MT  1351    pwes____
mpwsi8ax  PerrywoodMT-SemiBoldItaExp        MT  1351    pwesi___
mgsr8a    GillSansCyrMT                     MT  1356    gcy_____
mgsb8a    GillSansCyrMT-Bold                MT  1356    gcyb____
mgsbi8a   GillSansCyrMT-BoldInclined        MT  1356    gcybi___
mgsri8a   GillSansCyrMT-Inclined            MT  1356    gcyi____
mgsl8a    GillSansCyrMT-Light               MT  1356    gcyl____
mgsli8a   GillSansCyrMT-LightIncl           MT  1356    gcyli___
mhob8a    HorleyOldStyleMT-Bold             MT  1357    hsb_____
mhobi8a   HorleyOldStyleMT-BoldItalic       MT  1357    hsbi____
mhori8a   HorleyOldStyleMT-Italic           MT  1357    hsi_____
mhol8a    HorleyOldStyleMT-Light            MT  1357    hsl_____
mholi8a   HorleyOldStyleMT-LightItalic      MT  1357    hsli____
mhor8a    HorleyOldStyleMT                  MT  1357    hsrg____
mhosi8a   HorleyOldStyleMT-SbItalic         MT  1357    hssb____
mhos8a    HorleyOldStyleMT-SemiBold         MT  1357    hssbi___
m_____GillFloriatedCapsMT               MT  1363    gilf____
m_____GillFloriatedCapsMT-Alt           MT  1363    gilfa___
m_____MonotypeGallia                    MT  1364    gall____
mbwr8a    MonotypeBroadway                  MT  1365    broa____
mbwre8a   MonotypeBroadway-Engraved         MT  1365    broae___
m_____CurlzMT-Regular                   MT  1367    crl_____
m_____KidprintMT-Regular                MT  1368    kidp____
m_____KidprintMT-Bold                   MT  1368    kidpb___
mgsbij8a  GillSans-BoldItalicOsF            MT  1369    gnbio___
mgsbc8a   GillSans-BoldSC                   MT  1369    gnbsc___
mgsrij8a  GillSans-ItalicOsF                MT  1369    gnio____
mgslij8a  GillSans-LightItalicOsF           MT  1369    gnlio___
mgslc8a   GillSans-LightSC                  MT  1369    gnlsc___
mgsrc8a   GillSans-SC                       MT  1369    gnsc____
mcrb8a    CourierLD-Bold                    MT  1371    cldb____
mcrbi8a   CourierLD-BoldItalic              MT  1371    cldbi___
mcrri8a   CourierLD-Italic                  MT  1371    cldi____
mcrr8a    CourierLD-Regular                 MT  1371    cldr____
mgshij8a  GillSans-HeavyItalicOsF           MT  1372    gnhio___
mgshc8a   GillSans-HeavySC                  MT  1372    gnhsc___
mgskij8a  GillSans-BookItalicOsF            MT  1372    gnwio___
mgskc8a   GillSans-BookSC                   MT  1372    gnwsc___

m_____GillFaciaMT-Italic                MT  1373    gfai____
m_____GillFaciaMT-ItalicAlt             MT  1373    gfaia___
m_____GillFaciaMT-ItalicDisplay         MT  1373    gfaid___
m_____GillFaciaMT-Regular               MT  1373    gfar____
m_____GillFaciaMT-RegularAlt            MT  1373    gfara___
m_____GillFaciaMT-RegularDisplay        MT  1373    gfard___
m_____GillFaciaMT-ItDisplayAlt          MT  1373    gfida___
m_____GillFaciaMT-RegDisplayAlt         MT  1373    gfrda___
mngr8a    NewsGothicMT                      MT  1374    nwg_____
mngb8a    NewsGothicMT-Bold                 MT  1374    nwgb____
mngb8a    NewsGothicMT-BoldItalic           MT  1374    nwgbi___
mngri8a   NewsGothicMT-Italic               MT  1374    nwgi____
m_____AlbertinaMT-Bold                  MT  1378    abb_____
m_____AlbertinaMT-BoldItalic            MT  1378    abbi____
m_____AlbertinaMT-BoldItalicOsF         MT  1378    abbio___
m_____AlbertinaMT-BoldItalicSC          MT  1378    abbis___
m_____AlbertinaMT-BoldOsF               MT  1378    abbos___
m_____AlbertinaMT-BoldSC                MT  1378    abbsc___
m_____AlbertinaMT-Italic                MT  1378    abi_____
m_____AlbertinaMT-ItalicOsF             MT  1378    abios___
m_____AlbertinaMT-ItalicSC              MT  1378    abisc___
m_____AlbertinaMT-Medium                MT  1378    abm_____
m_____AlbertinaMT-MediumItalic          MT  1378    abmi____
m_____AlbertinaMT-MediumItalicOsF       MT  1378    abmio___
m_____AlbertinaMT-MediumItalicSC        MT  1378    abmis___
m_____AlbertinaMT-MediumOsF             MT  1378    abmos___
m_____AlbertinaMT-MediumSC              MT  1378    abmsc___
m_____AlbertinaMT-Regular               MT  1378    abr_____
m_____AlbertinaMT-RegularOsF            MT  1378    abros___
m_____AlbertinaMT-RegularSC             MT  1378    abrsc___
mngr__NewsGothicCE                      MT  1379    ngu_____
mngb__NewsGothicCE-Bold                 MT  1379    ngub____
mngbi__NewsGothicCE-BoldItalic           MT  1379    ngubi___
mngri__NewsGothicCE-Italic               MT  1379    ngui____
mngrk     NewsGothicGre                     MT  1380    ngg_____
mngbk     NewsGothicGre-Bold                MT  1380    nggb____
mngbik    NewsGothicGre-BoldInclined        MT  1380    nggbi___
mngrik    NewsGothicGre-Inclined            MT  1380    nggi____
mngrz     NewsGothicCyr                     MT  1381    ngc_____
mngbz     NewsGothicCyr-Bold                MT  1381    ngcb____
mngbiz    NewsGothicCyr-BoldInclined        MT  1381    ngcbi___
mngriz    NewsGothicCyr-Inclined            MT  1381    ngci____
m_____BuffaloGalMT                      MT  1382    bufg____
mrwr__RockwellTur                       MT  1383    rkt_____
mrwb__RockwellTur-Bold                  MT  1383    rktb____
mrwbi__RockwellTur-BoldItalic            MT  1383    rktbi___
mrwri__RockwellTur-Italic                MT  1383    rkti____

mrwl__RockwellTur-Light                 MT  1383    rktl____
mrwli__RockwellTur-LightItalic           MT  1383    rktli___
mrwr__RockwellCE                        MT  1383    rku_____
mrwb__RockwellCE-Bold                   MT  1383    rkub____
mrwbi__RockwellCE-BoldItalic             MT  1383    rkubi___
mrwri__RockwellCE-Italic                 MT  1383    rkui____
mrwl__RockwellCE-Light                  MT  1383    rkul____
mrwli__RockwellCE-LightItalic            MT  1383    rkuli___
mo4bij8a  OceanSansMT-BoldItalicOsF         MT  1397    oebio___
mo4bic8a  OceanSansMT-BoldItalicSC          MT  1397    oebis___
mo4bj8a   OceanSansMT-BoldOsF               MT  1397    oebo____
mo4bc8a   OceanSansMT-BoldSC                MT  1397    oebs____
mo4kij8a  OceanSansMT-BookItalicOsF         MT  1397    oekio___
mo4kic8a  OceanSansMT-BookItalicSC          MT  1397    oekis___
mo4kj8a   OceanSansMT-BookOsF               MT  1397    oeko____
mo4kc8a   OceanSansMT-BookSC                MT  1397    oeks____
mo4lij8a  OceanSansMT-LightItalicOsF        MT  1398    oelio___
mo4lic8a  OceanSansMT-LightItalicSC         MT  1398    oelis___
mo4lj8a   OceanSansMT-LightOsF              MT  1398    oelo____
mo4lc8a   OceanSansMT-LightSC               MT  1398    oels____
mo4sij8a  OceanSansMT-SemiBoldItaOsF        MT  1398    oesio___
mo4sic8a  OceanSansMT-SemiBoldItaSC         MT  1398    oesis___
mo4sj8a   OceanSansMT-SemiBoldOsF           MT  1398    oeso____
mo4sc8a   OceanSansMT-SemiBoldSC            MT  1398    oess____
mo4xij8a  OceanSansMT-ExtraBoldItaOsF       MT  1398    oexio___
mo4xic8a  OceanSansMT-ExtraBoldItaSC        MT  1398    oexis___
mo4xj8a   OceanSansMT-ExtraBoldOsF          MT  1398    oexo____
mo4xc8a   OceanSansMT-ExtraBoldSC           MT  1398    oexs____
mgmrz     GaramondCyrMT-Regular             MT  1401    gac_____
mgmbz     GaramondCyrMT-Bold                MT  1401    gacb____
mgmriz    GaramondCyrMT-Inclined            MT  1401    gaci____
mgmrk     GaramondDualGreekMT-Regular       MT  1402    gad_____
mgmbk     GaramondDualGreekMT-Bold          MT  1402    gadb____
mgmrik    GaramondDualGreekMT-Incl          MT  1402    gadi____
mgmr__GaramondCEMT-Regular              MT  1403    gae_____
mgmb__GaramondCEMT-Bold                 MT  1403    gaeb____
mgmri__GaramondCEMT-Italic               MT  1403    gaei____
mgmr__GaramondTurMT-Regular             MT  1403    gat_____
mgmb__GaramondTurMT-Bold                MT  1403    gatb____
mgmri__GaramondTurMT-Italic              MT  1403    gati____
mngr8ac   NewsGothicMT-Condensed            MT  1451    ngoc____
mngb8ac   NewsGothicMT-CondensedBold        MT  1451    ngocb___
m_____MahsuriSansMT-Bold                MT  1454    msb_____
m_____MahsuriSansMT-BoldItalic          MT  1454    msbi____
m_____MahsuriSansMT-BoldItalicOsF       MT  1454    msbio___
m_____MahsuriSansMT-BoldOsF             MT  1454    msbo____
m_____MahsuriSansMT-Italic              MT  1454    msi_____

m_____MahsuriSansMT-ItalicOsF           MT  1454    msio____
m_____MahsuriSansMT-Light               MT  1454    msl_____
m_____MahsuriSansMT-LightItalic         MT  1454    msli____
m_____MahsuriSansMT-LightItalicOsF      MT  1454    mslio___
m_____MahsuriSansMT-LightOsF            MT  1454    mslo____
m_____MahsuriSansMT-Regular             MT  1454    msr_____
m_____MahsuriSansMT-RegularOsF          MT  1454    msro____
m_____MahsuriSansMT-ExtraBold           MT  1454    msx_____
m_____MahsuriSansMT-ExtraBoldItalic     MT  1454    msxi____
m_____MahsuriSansMT-ExtBoldItaOsF       MT  1454    msxio___
m_____MahsuriSansMT-ExtraBoldOsF        MT  1454    msxo____
mu0r8a    ColumbusMT                        MT  1492    cu______
mu0b8a    ColumbusMTBold                    MT  1492    cub_____
mu0bi8a   ColumbusMTBoldItalic              MT  1492    cubi____
mu0ri8a   ColumbusMTItalic                  MT  1492    cui_____
mu0s8a    ColumbusMTSemiBold                MT  1492    cus_____
mu0si8a   ColumbusMTSemiBoldItalic          MT  1492    cusi____
mpi001    ArabesqueOrnamentsOneMT  ornaments
mpi002    ArabesqueOrnamentsTwoMT  ditto
mpi003    ArabesqueOrnamentsThreeMT  ditto
mpi004    RococoOrnamentsOneMT  ornaments
mpi005    RococoOrnamentsTwoMT  ditto
mpi006    RococoOrnamentsThreeMT  ditto
mpi007    ChristmasOrnamentsMT-One  ornaments
mpi008    ChristmasOrnamentsMT-Two  ditto
mpi009    ChristmasOrnamentsMT-Three  ditto
mpi010    ChristmasOrnamentsMT-Four  ditto
mpi011    ChristmasOrnamentsMT-Five  ditto
mpi012    ChristmasOrnamentsMT-Six  ditto
mpi013    ContemporaryOrnMT-One  ornaments
mpi014    ContemporaryOrnMT-Two  ditto
mpi015    ContemporaryOrnMT-Three  ditto
mpi016    ContemporaryOrnMT-Four  ditto
mpi017    ContemporaryOrnMT-Five  ditto
mpi101    AlmanacMT  zodiac
mpi102    BotanicalMT  flowers
mpi103    DirectionsMT  arrows & pointers
mpi104    SportsOneMT  sports
mpi105    SportsTwoMT  ditto
mpi106    TransportMT  signs
mpi107    VacationMT  signs
mpi108    SportsThreeMT  sports
mpi109    SportsFourMT  ditto
mpi201    CircleFrameMT  circled numerals
mpi202    CircleFrameNegMT  ditto
mpi203    SquareFrameMT  squared numerals
mpi204    SquareFrameNegMT  ditto

mpi205    DiamondBodoni  bullets & markers
mpi206    DiamondPositive  ditto
mpi207    DiamondNegative  ditto
mpi208    OvalFrameMT  circled numerals
mpi209    OvalFrameNegativeMT  ditto
mpi300    Giddyup-ThangsMT  cowboys & indians
mpi301    PetrucchioMT  ???
mpi302    AquaLifeMT  fish
mpi303    BonApetit  food
mpi304    DinosaursMT  dinosaurs
mpi305    FrontierMT  cowboys & indians
mpi306    Holidays  greeting cards
mpi307    KeystrokesExtrasMT  keyboards
mpi308    KeystrokesMT  ditto

mpi309 Melodies musical instruments

mpi310    MilestonesOneMT  clip art
mpi311    MilestonesTwoMT  ditto
mpi312    MonkeyBusiness  animals
mpi313    MonotypeSorts  dingbats
mpi314    NatureMT  animals
mpi315    Parties  clip art
mpi316    BingoMT  clip art
mpi317    BongoMT  clip art
mpi318    PieMT  food & drinks
mpi319    PlaytimeMT  toys
mpi320    SignsMT  signs
mpi321    SouthwestOrnamentsMT  geometric ornaments
mpi322    KeystrokesShadowExtrasMT  keyboards
mpi323    KeystrokesShadowMT  ditto

A.9 URW fonts

Information graciously supplied through URW, via Vincent Ertner: Each URW font has an assigned font number. The last character describes the size or purpose of the font, as the design of the font is optimized accordingly:

X000000E Extreme: 8 pt and smaller

X000000T Text: between 8 and 18 pt

X000000D Display: 18 pt and larger

X000000P Poster: 48 pt and larger

X000000M Monospaced / Typewriter

X000000L LaserWriter / PostScript / Adobe compatible

X000000I Intellifont PCL (5) compatible An ‘F’ identifier is not known.

The ‘urw.map’ file has abbreviations for URW fonts.

uagd8a    URWGothicL-Demi                               a010015l
uagdo8a   URWGothicL-DemiObli                           a010035l
uagk8a    URWGothicL-Book                               a010013l
uagko8a   URWGothicL-BookObli                           a010033l
ubkd8a    URWBookmanL-DemiBold                          b018015l
ubkdi8a   URWBookmanL-DemiBoldItal                      b018035l
ubkl8a    URWBookmanL-Ligh                              b018012l
ubkli8a   URWBookmanL-LighItal                          b018032l
ucrb8a    NimbusMonL-Bold                               n022004l
ucrbo8a   NimbusMonL-BoldObli                           n022024l
ucrr8a    NimbusMonL-Regu                               n022003l
ucrro8a   NimbusMonL-ReguObli                           n022023l
uhvb8a    NimbusSanL-Bold                               n019004l
uhvb8ac   NimbusSanL-BoldCond                           n019044l
uhvbo8a   NimbusSanL-BoldItal                           n019024l
uhvbo8ac  NimbusSanL-BoldCondItal                       n019064l
uhvr8a    NimbusSanL-Regu                               n019003l
uhvr8ac   NimbusSanL-ReguCond                           n019043l
uhvro8a   NimbusSanL-ReguItal                           n019023l
uhvro8ac  NimbusSanL-ReguCondItal                       n019063l
uncb8a    CenturySchL-Bold                              c059016l
uncbi8a   CenturySchL-BoldItal                          c059036l
uncr8a    CenturySchL-Roma                              c059013l
uncri8a   CenturySchL-Ital                              c059033l
uplb8a    URWPalladioL-Bold                             p052004l
uplbi8a   URWPalladioL-BoldItal                         p052024l
uplr8a    URWPalladioL-Roma                             p052003l
uplri8a   URWPalladioL-Ital                             p052023l
usyr      StandardSymL                                  s050000l
utmb8a    NimbusRomNo9L-Medi                            n021004l
utmbi8a   NimbusRomNo9L-MediItal                        n021024l
utmr8a    NimbusRomNo9L-Regu                            n021003l
utmri8a   NimbusRomNo9L-ReguItal                        n021023l
uzcmi8a   URWChanceryL-MediItal                         z003034l
uzdr      Dingbats                                      d050000l

Appendix B Encodings

The following sections define various encodings as PostScript encoding vectors. These files can be installed in the Dvips header directory and then used in ‘psfonts.map’ to change font encodings of downloadable or builtin fonts (see Section “psfonts.map” in Dvips).

You can also specify these ‘.enc’ files to Afm2tfm when building a font (see Section “Changing font encodings” in Dvips).

For a discussion of TEX font encodings in general, and defining new ones for use with

fontinst in particular, see http://www.ctan.org/tex-archive/fonts/utilities/fontinst/doc/encspecs.

A tutorial on character code issues in general is available at http://www.cs.tut.fi/~jkorpela/chars.html. Naming specific characters is a tangled and difficult area. Here are some links to charac-

ter tables for Unicode, Adobe, and WGL4, for your researching pleasure (thanks to Boguslaw Jackowski for the list):

  • http://www.unicode.org/Public/UNIDATA/UnicodeData.txt
  • http://www.unicode.org/Public/UNIDATA/NamesList.txt
  • http://partners.adobe.com/asn/tech/type/corporateuse-old.txt
  • http://partners.adobe.com/asn/tech/type/aglfn13.txt
  • http://partners.adobe.com/asn/tech/type/unicodegn.jsp
  • http://www.microsoft.com/OpenType/otspec/WGL4.htm

B.1 ‘6w.enc’: Cyrillic CP-1251

% @psencodingfile{
%   author = "Adobe",
%   version = "1.1",
%   date = "12jun05",
%   filename = "8a.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "Adobe StandardEncoding as an afm2tfm/dvips .enc file."
% }
% 12jun05 - comment changes only.
/StandardEncoding [
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/space
/exclam
/quotedbl
/numbersign
/dollar
/percent
/ampersand
/quoteright
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
/zero
/one
/two
/three
/four
/five
/six
/seven
/eight
/nine

/colon
/semicolon
/less
/equal
/greater
/question
/at
/A
/B
/C
/D
/E
/F
/G
/H
/I
/J
/K
/L
/M
/N
/O
/P
/Q
/R
/S
/T
/U
/V
/W
/X
/Y
/Z
/bracketleft
/backslash
/bracketright
/asciicircum
/underscore
/quoteleft
/a
/b
/c
/d
/e
/f
/g
/h

/i
/j
/k
/l
/m
/n
/o
/p
/q
/r
/s
/t
/u
/v
/w
/x
/y
/z
/braceleft
/bar
/braceright
/asciitilde
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/exclamdown
/cent
/sterling
/fraction
/yen
/florin
/section
/currency
/quotesingle
/quotedblleft
/guillemotleft
/guilsinglleft
/guilsinglright
/fi
/fl
/.notdef
/endash
/dagger
/daggerdbl
/periodcentered
/.notdef
/paragraph
/bullet
/quotesinglbase
/quotedblbase
/quotedblright
/guillemotright
/ellipsis
/perthousand
/.notdef
/questiondown
/.notdef
/grave
/acute
/circumflex
/tilde
/macron
/breve

/dotaccent
/dieresis
/.notdef
/ring
/cedilla
/.notdef
/hungarumlaut
/ogonek
/caron
/emdash
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/AE
/.notdef
/ordfeminine
/.notdef
/.notdef
/.notdef
/.notdef
/Lslash
/Oslash
/OE
/ordmasculine
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/ae
/.notdef
/.notdef
/.notdef
/dotlessi

/.notdef
/.notdef
/lslash
/oslash
/oe
/germandbls
/.notdef
/.notdef
/.notdef
/.notdef
] def

B.2 ‘8a.enc’: Adobe standard encoding

% @psencodingfile{
%   author = "Adobe",
%   version = "1.1",
%   date = "12jun05",
%   filename = "8a.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "Adobe StandardEncoding as an afm2tfm/dvips .enc file."
% }
% 12jun05 - comment changes only.
/StandardEncoding [
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/space
/exclam
/quotedbl
/numbersign
/dollar
/percent
/ampersand
/quoteright
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
/zero
/one
/two
/three
/four
/five
/six
/seven
/eight
/nine
/colon
/semicolon
/less
/equal
/greater
/question
/at
/A
/B

/D
/E
/F
/G
/H
/I
/J
/K
/L
/M
/N
/O
/P
/Q
/R
/S
/T
/U
/V
/W
/X
/Y
/Z
/bracketleft
/backslash
/bracketright
/asciicircum
/underscore
/quoteleft
/a
/b
/c
/d
/e
/f
/g
/h
/i
/j
/k
/l
/m
/n
/o
/p
/q
/r

/s
/t
/u
/v
/w
/x
/y
/z
/braceleft
/bar
/braceright
/asciitilde
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/exclamdown

/cent
/sterling
/fraction
/yen
/florin
/section
/currency
/quotesingle
/quotedblleft
/guillemotleft
/guilsinglleft
/guilsinglright
/fi
/fl
/.notdef
/endash
/dagger
/daggerdbl
/periodcentered
/.notdef
/paragraph
/bullet
/quotesinglbase
/quotedblbase
/quotedblright
/guillemotright
/ellipsis
/perthousand
/.notdef
/questiondown
/.notdef
/grave
/acute
/circumflex
/tilde
/macron
/breve
/dotaccent
/dieresis
/.notdef
/ring
/cedilla
/.notdef
/hungarumlaut
/ogonek
/caron
/emdash

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/AE
/.notdef
/ordfeminine
/.notdef
/.notdef
/.notdef
/.notdef
/Lslash
/Oslash
/OE
/ordmasculine
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/ae
/.notdef
/.notdef
/.notdef
/dotlessi
/.notdef
/.notdef
/lslash
/oslash
/oe
/germandbls
/.notdef
/.notdef
/.notdef
/.notdef

] def

B.3 ‘8r.enc’: TEX base encoding

% File 8r.enc  TeX Base 1 Encoding  Revision 2.0  2002-10-30
%
% @@psencodingfile@{
%   author    = "S. Rahtz, P. MacKay, Alan Jeffrey, B. Horn, K. Berry,
%                W. Schmidt, P. Lehman",
%   version   = "2.0",
%   date      = "27nov06",
%   filename  = "8r.enc",
%   email     = "tex-fonts@@tug.org",
%   docstring = "This is the encoding vector for Type1 and TrueType
%                fonts to be used with TeX.  This file is part of the
%                PSNFSS bundle, version 9"
% @}
%
% The idea is to have all the characters normally included in Type 1 fonts
% available for typesetting.  This is effectively the characters in Adobe
% Standard encoding, ISO Latin 1, Windows ANSI including the euro symbol,
% MacRoman, and some extra characters from Lucida.
%
% Character code assignments were made as follows:
%
% (1) the Windows ANSI characters are almost all in their Windows ANSI
% positions, because some Windows users cannot easily reencode the
% fonts, and it makes no difference on other systems.  The only Windows
% ANSI characters not available are those that make no sense for
% typesetting -- rubout (127 decimal), nobreakspace (160), softhyphen
% (173).  quotesingle and grave are moved just because it's such an
% irritation not having them in TeX positions.
%
% (2) Remaining characters are assigned arbitrarily to the lower part
% of the range, avoiding 0, 10 and 13 in case we meet dumb software.
%
% (3) Y&Y Lucida Bright includes some extra text characters; in the
% hopes that other PostScript fonts, perhaps created for public
% consumption, will include them, they are included starting at 0x12.
% These are /dotlessj /ff /ffi /ffl.
%
% (4) hyphen appears twice for compatibility with both ASCII and Windows.
%
% (5) /Euro was assigned to 128, as in Windows ANSI
%
% (6) Missing characters from MacRoman encoding incorporated as follows:
%

%     PostScript      MacRoman        TeXBase1
%     --------------  --------------  --------------
%     /notequal       173             0x16
%     /infinity       176             0x17
%     /lessequal      178             0x18
%     /greaterequal   179             0x19
%     /partialdiff    182             0x1A
%     /summation      183             0x1B
%     /product        184             0x1C
%     /pi             185             0x1D
%     /integral       186             0x81
%     /Omega          189             0x8D
%     /radical        195             0x8E
%     /approxequal    197             0x8F
%     /Delta          198             0x9D
%     /lozenge        215             0x9E
%
/TeXBase1Encoding [
% 0x00
 /.notdef /dotaccent /fi /fl
 /fraction /hungarumlaut /Lslash /lslash
 /ogonek /ring /.notdef /breve
 /minus /.notdef /Zcaron /zcaron
% 0x10
 /caron /dotlessi /dotlessj /ff
 /ffi /ffl /notequal /infinity
 /lessequal /greaterequal /partialdiff /summation
 /product /pi /grave /quotesingle
% 0x20
 /space /exclam /quotedbl /numbersign
 /dollar /percent /ampersand /quoteright
 /parenleft /parenright /asterisk /plus
 /comma /hyphen /period /slash
% 0x30
 /zero /one /two /three
 /four /five /six /seven
 /eight /nine /colon /semicolon
 /less /equal /greater /question
% 0x40
 /at /A /B /C
 /D /E /F /G
 /H /I /J /K
 /L /M /N /O
% 0x50
 /P /Q /R /S
 /T /U /V /W
 /X /Y /Z /bracketleft

/backslash /bracketright /asciicircum /underscore
% 0x60
/quoteleft /a /b /c
/d /e /f /g
/h /i /j /k
/l /m /n /o
% 0x70
/p /q /r /s
/t /u /v /w
/x /y /z /braceleft
/bar /braceright /asciitilde /.notdef
% 0x80
/Euro /integral /quotesinglbase /florin
/quotedblbase /ellipsis /dagger /daggerdbl
/circumflex /perthousand /Scaron /guilsinglleft
/OE /Omega /radical /approxequal
% 0x90
/.notdef /.notdef /.notdef /quotedblleft
/quotedblright /bullet /endash /emdash
/tilde /trademark /scaron /guilsinglright
/oe /Delta /lozenge /Ydieresis
% 0xA0
/.notdef /exclamdown /cent /sterling
/currency /yen /brokenbar /section
/dieresis /copyright /ordfeminine /guillemotleft
/logicalnot /hyphen /registered /macron
% 0xB0
/degree /plusminus /twosuperior /threesuperior
/acute /mu /paragraph /periodcentered
/cedilla /onesuperior /ordmasculine /guillemotright
/onequarter /onehalf /threequarters /questiondown
% 0xC0
/Agrave /Aacute /Acircumflex /Atilde
/Adieresis /Aring /AE /Ccedilla
/Egrave /Eacute /Ecircumflex /Edieresis
/Igrave /Iacute /Icircumflex /Idieresis
% 0xD0
/Eth /Ntilde /Ograve /Oacute
/Ocircumflex /Otilde /Odieresis /multiply
/Oslash /Ugrave /Uacute /Ucircumflex
/Udieresis /Yacute /Thorn /germandbls
% 0xE0
/agrave /aacute /acircumflex /atilde
/adieresis /aring /ae /ccedilla
/egrave /eacute /ecircumflex /edieresis
/igrave /iacute /icircumflex /idieresis
% 0xF0

 /eth /ntilde /ograve /oacute
 /ocircumflex /otilde /odieresis /divide
 /oslash /ugrave /uacute /ucircumflex
 /udieresis /yacute /thorn /ydieresis
] def

B.4 ‘dvips.enc’: Dvips encoding

% @psencodingfile{
%   author = "T. Rokicki, K. Berry",
%   version = "2.1",
%   date = "12jun05",
%   filename = "dvips.enc",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "An encoding for PostScript fonts to be used with TeX.
%
% This is almost the default encoding produced by the afm2tfm program
% that comes with the dvips distribution, when it is applied to a
% StandardEncoding PostScript font.    The positions of the accented
% characters (which afm2tfm assigned more or less randomly in the
% previous set of fonts) have been held fixed.
%
% It is almost upward-compatible with TeX text.    (The primary exceptions
% are dotaccent and hungarumlaut.)    Some of the ligatures are also in
% different positions, but this is not important, since ligature
% positions are defined in the TFM file, not by macros.
%
% The character positions were assigned in this order:
%   1) non-composite characters in the afm2tfm encoding (except dotlessj
%      was included in its plain TeX position);
%   2) available characters in the Cork (EC) encoding;
%   3) remaining composite characters in the base encoding (starting at
%      position 130 decimal).    We don't include all the other base
%      characters because that changes the height/depth/width tables in
%      the TFM files (because more characters mean more distinct values).
%
% We leave the initial positions open for the Greek letters in Times-Roman.
% "
% }
% 12jun05 - don't use extra @'s for Texinfo any more.
%
/DvipsEncoding [
  /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
  /.notdef /.notdef /.notdef /.notdef
    /.notdef /quotesingle /exclamdown /questiondown

/dotlessi /dotlessj /grave /acute /caron /breve /macron /ring
/cedilla /germandbls /ae /oe /oslash /AE /OE /Oslash
/space /exclam /quotedbl /numbersign /dollar /percent /ampersand /quoteright
/parenleft /parenright /asterisk /plus /comma /hyphen /period /slash
/zero /one /two /three /four /five /six /seven
/eight /nine /colon /semicolon /less /equal /greater /question
/at /A /B /C /D /E /F /G
/H /I /J /K /L /M /N /O
/P /Q /R /S /T /U /V /W
/X /Y /Z /bracketleft /backslash /bracketright /circumflex /underscore
/quoteleft /a /b /c /d /e /f /g
/h /i /j /k /l /m /n /o
/p /q /r /s /t /u /v /w
/x /y /z /braceleft /bar /braceright /tilde /dieresis
/asciicircum /asciitilde /Ccedilla /Iacute
    /Icircumflex /atilde /edieresis /egrave
/scaron /zcaron /Eth /ff /ffi /ffl /.notdef /.notdef
/.notdef /.notdef /Scaron /.notdef /.notdef /.notdef /.notdef /.notdef
/Ydieresis /.notdef /Zcaron  /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /cent /sterling /fraction /yen /florin /section
/currency /copyright /quotedblleft /guillemotleft
    /guilsinglleft /guilsinglright /fi /fl
/degree /endash /dagger /daggerdbl /periodcentered /.notdef /paragraph /bullet
/quotesinglbase /quotedblbase /quotedblright /guillemotright
    /ellipsis /perthousand /.notdef /.notdef
/Agrave /Aacute /Acircumflex /Atilde /Adieresis /Aring /.notdef /dotaccent
/Egrave /Eacute /Ecircumflex /Edieresis
    /Igrave /hungarumlaut /ogonek /Idieresis
/emdash /Ntilde /Ograve /Oacute /Ocircumflex /Otilde /Odieresis /.notdef
/.notdef /Ugrave /Uacute /Ucircumflex /Udieresis /Yacute /Thorn /.notdef
/agrave /aacute /acircumflex /ordfeminine
    /adieresis /aring /.notdef /ccedilla
/Lslash /eacute /ecircumflex /ordmasculine
    /igrave /iacute /icircumflex /idieresis
/.notdef /ntilde /ograve /oacute /ocircumflex /otilde /odieresis /.notdef
/lslash /ugrave /uacute /ucircumflex /udieresis /yacute /thorn /ydieresis
] def

B.5 ‘ec.enc’: Cork encoding

% @@psencodingfile@{
%   date      = "24feb10",
%   filename  = "ec.enc",
%   email     = "tex-fonts@@tug.org",
%   docstring = "This is the EC (aka Cork aka T1) encoding vector
%                for 8-bit fonts to be used with TeX."
% @}

%
% The original paper can be found on
%    https://www.tug.org/TUGboat/Articles/tb11-4/tb30ferguson.pdf
%
% Names of the glyphs try to comply with:
%    Adobe Glyph List For New Fonts, ver.  1.6, 30 Januaury 2006
%    http://partners.adobe.com/public/developer/en/opentype/aglfn13.txt
% but some exceptions are listed below.
%
% The following characters are not present in Unicode:
% - /perthousandzero = used to compose /perthousand, /pertenthousand, ...  out of /percent + this character
% -  Germandbls      = now there is a related character in Unicode,
%                      U+1E9E - LATIN CAPITAL LETTER SHARP S
%                      but probably /S_S ligature name should be used instead
%
% The following two characters have different Unicode mappings than on the AGL:
% - /tcedilla     = U+0163 - LATIN SMALL LETTER T WITH CEDILLA
%                   according to AGL it should be /tcommaaccent
% - /Tcedilla     = U+0162 - LATIN CAPITAL LETTER T WITH CEDILLA
%                   according to AGL it should be /Tcommaaccent
% But having /scedilla and /tcommaaccent mixed is even more ugly.
%
% The following characters are not present on the AGL:
% - /dotlessj     = U+0237 - LATIN SMALL LETTER DOTLESS J
% - /cwm          = U+200B - ZERO WIDTH SPACE
% - /visiblespace = U+2423 - OPEN BOX
% and should probably better be labeled with uniXXXX notation
% (though hardly any font has those glyphs anyway).
%
% The following ligatures are deprecated by recent AGL:
% - /ff, /fi, /fl, /ffi, /ffl = /f_i and similar should have been used instead
%
% Hyphen character:
% There are two character named "hyphen".  The original article explains
% the background of this decision (if one wanted to use a different
% hyphenation character and still be able to use the dash).  So the first
% character should serve as "hyphen char" and the other one as "dash" in
% compound words.    The character 0x2D is supposed to be the hyphen char,
% but since there used to be two characters of the same name in this
% file for some time, the tools like afmtotfm have been confused and
% remapped the hyphen char to 0x7F automatically.    After the second
% character's name was changed, hyphenation characters in existing
% tfm fonts (generated with the old ec.enc) have disappeared, so the
% second character has been renamed back to "hyphen".  Anyone generating
% tfm fonts is strongly encouraged to make a copy of this file, rename
% it and change glyph names as required for their font.
%

% Other remarks:
% - /Eth = U+00D0 - LATIN CAPITAL LETTER ETH
%          must also be used as /Dcroat
%          U+0110 - LATIN CAPITAL LETTER D WITH STROKE
%
% Misleading names and shapes of letters (that lead to different names in past)
% - /lcaron = U+013E - LATIN SMALL LETTER L WITH CARON
%             /lquoteright
% - /Lcaron = U+013D - LATIN CAPITAL LETTER L WITH CARON
%             /Lquoteright
% - /dcaron = U+010F - LATIN SMALL LETTER D WITH CARON
%             /dquoteright
% - /tcaron = U+0165 - LATIN SMALL LETTER T WITH CARON
%             /tquoteright
% - /dcroat = U+0111 - LATIN SMALL LETTER D WITH STROKE
%             /dbar, /dmacron
%
% - /cwm        - /bom, /zerowidthnobreakspace
% - /underscore - /underline
% - /hyphen.alt - /sfthyphen, /hyphen, /dash
%
%
% Turkish & Romanian
% --------------------
% EC encoding has been defined to include /tcedilla and /scedilla.
% For Turkish one needs /scedilla, while for Romanian one would need
% "t with comma below" (/tcommaaccent) and "s with comma below" (/scommaaccent).
% "t with cedilla" in not used in any language except in inferior
% Romanian typography that can use cedillas instead of commas below.
% AGL has a strange mapping as well.
%
% Without adding an additional character, it is impossible to support both
% languages at the same time, so EC is not really best choice for Romanians,
% even though it can be "misused" if one replaces both /scedilla and /tcedilla
% with /scommaaccent and /tcommaaccent.
%
% Characters that have been renamed in July 2008:
%    0x18        /zeroinferior -> /perthousandzero
%    0x20 U+2423 /visualspace  -> /visiblespace (if font doesn't have it, use /space)
%    0x8D U+014A /Ng           -> /Eng
%    0x9E U+0111 /dbar         -> /dcroat
%    0xAD U+014B /ng           -> /eng
%    0xB4 U+0165 /tquoteright  -> /tcaron
%
% Although these ligatures and kerns are somewhat questionable to have
% in an encoding file, they've been here forever, and they are
% traditional in the TeX world, so we preserve them.

%
% LIGKERN space l =:  lslash ; space L =:  Lslash ;
% LIGKERN question quoteleft =:  questiondown ; exclam quoteleft =:  exclamdown ;
% LIGKERN hyphen hyphen =:  endash ; endash hyphen =:  emdash ;
% LIGKERN quoteleft quoteleft =:  quotedblleft ;
% LIGKERN quoteright quoteright =:  quotedblright ;
%
%   We blow away kerns to and from spaces (TeX doesn't have a
%   space) and also remove any kerns from the numbers (although
%   the only kern pair that mentions a number in Times-Roman.afm
%   is one one.)
%
% LIGKERN space {} * ; * {} space ; zero {} * ; * {} zero ;
% LIGKERN one {} * ; * {} one ; two {} * ; * {} two ;
% LIGKERN three {} * ; * {} three ; four {} * ; * {} four ;
% LIGKERN five {} * ; * {} five ; six {} * ; * {} six ;
% LIGKERN seven {} * ; * {} seven ; eight {} * ; * {} eight ;
% LIGKERN nine {} * ; * {} nine ;
% LIGKERN comma comma =:  quotedblbase ; less less =:  guillemotleft ;
% LIGKERN greater greater =:  guillemotright ;
%
/ECEncoding [    % now 256 chars follow
/grave           % 0x00 U+0060
/acute           % 0x01 U+00B4
/circumflex      % 0x02 U+02C6
/tilde           % 0x03 U+02DC
/dieresis        % 0x04 U+00A8
/hungarumlaut    % 0x05 U+02DD
/ring            % 0x06 U+02DA
/caron           % 0x07 U+02C7
/breve           % 0x08 U+02D8
/macron          % 0x09 U+00AF
/dotaccent       % 0x0A U+02D9
/cedilla         % 0x0B U+00B8
/ogonek          % 0x0C U+02DB
/quotesinglbase  % 0x0D U+201A
/guilsinglleft   % 0x0E U+2039
/guilsinglright  % 0x0F U+203A
/quotedblleft    % 0x10 U+201C
/quotedblright   % 0x11 U+201D
/quotedblbase    % 0x12 U+201E
/guillemotleft   % 0x13 U+00AB
/guillemotright  % 0x14 U+00BB
/endash          % 0x15 U+2013
/emdash          % 0x16 U+2014
/cwm             % 0x17 U+200B
/perthousandzero % 0x18 ______

/dotlessi        % 0x19 U+0131
/dotlessj        % 0x1A U+0237
/ff              % 0x1B U+FB00
/fi              % 0x1C U+FB01
/fl              % 0x1D U+FB02
/ffi             % 0x1E U+FB03
/ffl             % 0x1F U+FB04
/visiblespace    % 0x20 U+2423
/exclam          % 0x21
/quotedbl        % 0x22
/numbersign      % 0x23
/dollar          % 0x24
/percent         % 0x25
/ampersand       % 0x26
/quoteright      % 0x27 U+2019
/parenleft       % 0x28
/parenright      % 0x29
/asterisk        % 0x2A
/plus            % 0x2B
/comma           % 0x2C
/hyphen          % 0x2D
/period          % 0x2E
/slash           % 0x2F
/zero            % 0x30
/one             % 0x31
/two             % 0x32
/three           % 0x33
/four            % 0x34
/five            % 0x35
/six             % 0x36
/seven           % 0x37
/eight           % 0x38
/nine            % 0x39
/colon           % 0x3A
/semicolon       % 0x3B
/less            % 0x3C
/equal           % 0x3D
/greater         % 0x3E
/question        % 0x3F
/at              % 0x40
/A               % 0x41
/B               % 0x42
/C               % 0x43
/D               % 0x44
/E               % 0x45
/F               % 0x46
/G               % 0x47

/H               % 0x48
/I               % 0x49
/J               % 0x4A
/K               % 0x4B
/L               % 0x4C
/M               % 0x4D
/N               % 0x4E
/O               % 0x4F
/P               % 0x50
/Q               % 0x51
/R               % 0x52
/S               % 0x53
/T               % 0x54
/U               % 0x55
/V               % 0x56
/W               % 0x57
/X               % 0x58
/Y               % 0x59
/Z               % 0x5A
/bracketleft     % 0x5B
/backslash       % 0x5C
/bracketright    % 0x5D
/asciicircum     % 0x5E
/underscore      % 0x5F
/quoteleft       % 0x60 U+2018
/a               % 0x61
/b               % 0x62
/c               % 0x63
/d               % 0x64
/e               % 0x65
/f               % 0x66
/g               % 0x67
/h               % 0x68
/i               % 0x69
/j               % 0x6A
/k               % 0x6B
/l               % 0x6C
/m               % 0x6D
/n               % 0x6E
/o               % 0x6F
/p               % 0x70
/q               % 0x71
/r               % 0x72
/s               % 0x73
/t               % 0x74
/u               % 0x75
/v               % 0x76

/w               % 0x77
/x               % 0x78
/y               % 0x79
/z               % 0x7A
/braceleft       % 0x7B
/bar             % 0x7C
/braceright      % 0x7D
/asciitilde      % 0x7E
/hyphen          % 0x7F U+002D (explicit "hyphen char")
/Abreve          % 0x80 U+0102
/Aogonek         % 0x81 U+0104
/Cacute          % 0x82 U+0106
/Ccaron          % 0x83 U+010C
/Dcaron          % 0x84 U+010E
/Ecaron          % 0x85 U+011A
/Eogonek         % 0x86 U+0118
/Gbreve          % 0x87 U+011E
/Lacute          % 0x88 U+0139
/Lcaron          % 0x89 U+013D
/Lslash          % 0x8A U+0141
/Nacute          % 0x8B U+0143
/Ncaron          % 0x8C U+0147
/Eng             % 0x8D U+014A
/Ohungarumlaut   % 0x8E U+0150
/Racute          % 0x8F U+0154
/Rcaron          % 0x90 U+0158
/Sacute          % 0x91 U+015A
/Scaron          % 0x92 U+0160
/Scedilla        % 0x93 U+015E
/Tcaron          % 0x94 U+0164
/Tcedilla        % 0x95 U+0162
/Uhungarumlaut   % 0x96 U+0170
/Uring           % 0x97 U+016E
/Ydieresis       % 0x98 U+0178
/Zacute          % 0x99 U+0179
/Zcaron          % 0x9A U+017D
/Zdotaccent      % 0x9B U+017B
/IJ              % 0x9C U+0132
/Idotaccent      % 0x9D U+0130
/dcroat          % 0x9E U+0111
/section         % 0x9F U+00A7
/abreve          % 0xA0 U+0103
/aogonek         % 0xA1 U+0105
/cacute          % 0xA2 U+0107
/ccaron          % 0xA3 U+010D
/dcaron          % 0xA4 U+010F
/ecaron          % 0xA5 U+011B

/eogonek         % 0xA6 U+0119
/gbreve          % 0xA7 U+011F
/lacute          % 0xA8 U+013A
/lcaron          % 0xA9 U+013E
/lslash          % 0xAA U+0142
/nacute          % 0xAB U+0144
/ncaron          % 0xAC U+0148
/eng             % 0xAD U+014B
/ohungarumlaut   % 0xAE U+0151
/racute          % 0xAF U+0155
/rcaron          % 0xB0 U+0159
/sacute          % 0xB1 U+015B
/scaron          % 0xB2 U+0161
/scedilla        % 0xB3 U+015F
/tcaron          % 0xB4 U+0165
/tcedilla        % 0xB5 U+0163
/uhungarumlaut   % 0xB6 U+0171
/uring           % 0xB7 U+016F
/ydieresis       % 0xB8 U+00FF
/zacute          % 0xB9 U+017A
/zcaron          % 0xBA U+017E
/zdotaccent      % 0xBB U+017C
/ij              % 0xBC U+0133
/exclamdown      % 0xBD U+00A1
/questiondown    % 0xBE U+00BF
/sterling        % 0xBF U+00A3
/Agrave          % 0xC0
/Aacute          % 0xC1
/Acircumflex     % 0xC2
/Atilde          % 0xC3
/Adieresis       % 0xC4
/Aring           % 0xC5
/AE              % 0xC6
/Ccedilla        % 0xC7
/Egrave          % 0xC8
/Eacute          % 0xC9
/Ecircumflex     % 0xCA
/Edieresis       % 0xCB
/Igrave          % 0xCC
/Iacute          % 0xCD
/Icircumflex     % 0xCE
/Idieresis       % 0xCF
/Eth             % 0xD0
/Ntilde          % 0xD1
/Ograve          % 0xD2
/Oacute          % 0xD3
/Ocircumflex     % 0xD4

/Otilde          % 0xD5
/Odieresis       % 0xD6
/OE              % 0xD7 U+0152
/Oslash          % 0xD8
/Ugrave          % 0xD9
/Uacute          % 0xDA
/Ucircumflex     % 0xDB
/Udieresis       % 0xDC
/Yacute          % 0xDD
/Thorn           % 0xDE
/Germandbls      % 0xDF (U+1E9E, /S_S)
/agrave          % 0xE0
/aacute          % 0xE1
/acircumflex     % 0xE2
/atilde          % 0xE3
/adieresis       % 0xE4
/aring           % 0xE5
/ae              % 0xE6
/ccedilla        % 0xE7
/egrave          % 0xE8
/eacute          % 0xE9
/ecircumflex     % 0xEA
/edieresis       % 0xEB
/igrave          % 0xEC
/iacute          % 0xED
/icircumflex     % 0xEE
/idieresis       % 0xEF
/eth             % 0xF0
/ntilde          % 0xF1
/ograve          % 0xF2
/oacute          % 0xF3
/ocircumflex     % 0xF4
/otilde          % 0xF5
/odieresis       % 0xF6
/oe              % 0xF7 U+0153
/oslash          % 0xF8
/ugrave          % 0xF9
/uacute          % 0xFA
/ucircumflex     % 0xFB
/udieresis       % 0xFC
/yacute          % 0xFD
/thorn           % 0xFE
/germandbls      % 0xFF U+00DF
] def

B.6 ‘groff.enc’: groff encoding

%!PS-Adobe-3.0 Resource-Encoding
%
% @psencodingfile{
%   author = "Werner Lemberg, Michail Vidiassov",
%   version = "1.0",
%   date = "2006-Feb-03",
%   filename = "groff.enc",
%   license = "public domain",
%   email = "groff@gnu.org",
%   docstring = "Groff default text encoding.    Used with MetaPost."
% }
%%BeginResource:  encoding groffEncoding
/groffEncoding [
% 0x00 | 0
/asciicircum     /asciitilde        /Scaron            /Zcaron
/scaron          /zcaron            /Ydieresis         /trademark
/quotesingle     /Euro              /.notdef           /.notdef
/.notdef         /.notdef           /.notdef           /.notdef
% 0x10 | 16
/.notdef         /.notdef           /.notdef           /.notdef
/.notdef         /.notdef           /.notdef           /.notdef
/.notdef         /.notdef           /.notdef           /.notdef
/.notdef         /.notdef           /.notdef           /.notdef
% 0x20 | 32
/space           /exclam            /quotedbl          /numbersign
/dollar          /percent           /ampersand         /quoteright
/parenleft       /parenright        /asterisk          /plus
/comma           /hyphen            /period            /slash
% 0x30 | 48
/zero            /one               /two               /three
/four            /five              /six               /seven
/eight           /nine              /colon             /semicolon
/less            /equal             /greater           /question
% 0x40 | 64
/at              /A                 /B                 /C
/D               /E                 /F                 /G
/H               /I                 /J                 /K
/L               /M                 /N                 /O
% 0x50 | 80
/P               /Q                 /R                 /S
/T               /U                 /V                 /W
/X               /Y                 /Z                 /bracketleft
/backslash       /bracketright      /circumflex        /underscore
% 0x60 | 96

/quoteleft       /a                 /b                 /c
/d               /e                 /f                 /g
/h               /i                 /j                 /k
/l               /m                 /n                 /o
% 0x70 | 112
/p               /q                 /r                 /s
/t               /u                 /v                 /w
/x               /y                 /z                 /braceleft
/bar             /braceright        /tilde             /.notdef
% 0x80 | 128
/quotesinglbase  /guillemotleft     /guillemotright    /bullet
/florin          /fraction          /perthousand       /dagger
/daggerdbl       /endash            /emdash            /ff
/fi              /fl                /ffi               /ffl
% 0x90 | 144
/dotlessi        /dotlessj          /grave             /hungarumlaut
/dotaccent       /breve             /caron             /ring
/ogonek          /quotedblleft      /quotedblright     /oe
/lslash          /quotedblbase      /OE                /Lslash
% 0xA0 | 160
/.notdef         /exclamdown        /cent              /sterling
/currency        /yen               /brokenbar         /section
/dieresis        /copyright         /ordfeminine       /guilsinglleft
/logicalnot      /minus             /registered        /macron
% 0xB0 | 176
/degree          /plusminus         /twosuperior       /threesuperior
/acute           /mu                /paragraph         /periodcentered
/cedilla         /onesuperior       /ordmasculine      /guilsinglright
/onequarter      /onehalf           /threequarters     /questiondown
% 0xC0 | 192
/Agrave          /Aacute            /Acircumflex       /Atilde
/Adieresis       /Aring             /AE                /Ccedilla
/Egrave          /Eacute            /Ecircumflex       /Edieresis
/Igrave          /Iacute            /Icircumflex       /Idieresis
% 0xD0 | 208
/Eth             /Ntilde            /Ograve            /Oacute
/Ocircumflex     /Otilde            /Odieresis         /multiply
/Oslash          /Ugrave            /Uacute            /Ucircumflex
/Udieresis       /Yacute            /Thorn             /germandbls
% 0xE0 | 224
/agrave          /aacute            /acircumflex       /atilde
/adieresis       /aring             /ae                /ccedilla
/egrave          /eacute            /ecircumflex       /edieresis
/igrave          /iacute            /icircumflex       /idieresis
% 0xF0 | 240
/eth             /ntilde            /ograve            /oacute
/ocircumflex     /otilde            /odieresis         /divide

  /oslash          /ugrave            /uacute            /ucircumflex
  /udieresis       /yacute            /thorn             /ydieresis
] def
%%EndResource
% eof

B.7 ‘t5.enc’: T5 encoding

% Copyright 2000-2005 Werner Lemberg <wl@gnu.org>.
% This file is part of vntex.    License:  LPPL, version 1.3 or newer,
% according to http://www.latex-project.org/lppl.txt
%
%
% t5.enc
%
%
% This is LaTeX T5 encoding for Vietnamese
%
% history:
%   version 1.0 (28-Jan-2000):
%     First version.
%
%   version 1.1 (29-Apr-2002):
%     Rewritten to make it AGL compliant.
%
%   version 1.2 (01-May-2002):
%     Use uni2423 instead uni2420.
%
%   version 1.3 (27-Sep-2002):
%     .  Changed the following glyph names to Unicode entities:
%         0xB3  uni1EBD  etilde
%         0x93  uni1EBC  Etilde
%     .  Ytilde (uni1EF8) was used in slot for ytilde (uni1EF9).
%     .  Using .notdef in slot for compound word mark (on position 0x17)
%       caused some problems so it was changed to cwm (for compatibilty
%       with ec.enc).
%     .  Use sfthyphen in slot 127 instead of hyphen.
%
%   version 1.4 (20-Apr-2005):
%     Add copyright.
%     Minor documentation updates.
%
% LIGKERN hyphen hyphen =:  endash ;
% LIGKERN endash hyphen =:  emdash ;
% LIGKERN quoteleft quoteleft =:  quotedblleft ;
% LIGKERN quoteright quoteright =:  quotedblright ;

% LIGKERN comma comma =:  quotedblbase ;
%
%   We blow away kerns to and from spaces (TeX doesn't have a
%   space) and also remove any kerns from the numbers.
%
% LIGKERN space {} * ; * {} space ;
% LIGKERN zero {} * ; * {} zero ;
% LIGKERN one {} * ; * {} one ;
% LIGKERN two {} * ; * {} two ;
% LIGKERN three {} * ; * {} three ;
% LIGKERN four {} * ; * {} four ;
% LIGKERN five {} * ; * {} five ;
% LIGKERN six {} * ; * {} six ;
% LIGKERN seven {} * ; * {} seven ;
% LIGKERN eight {} * ; * {} eight ;
% LIGKERN nine {} * ; * {} nine ;
%
%
% Some notes on the glyph names:
%
% .  We follow the Adobe Glyph List (AGL), also using its associated
%   algorithm.
%
% .  The glyphs `spacing dot below' (on position 0x05), `spacing hook above'
%   (on position 0x0C), and `compound word mark' (on position 0x17) are not
%   in Unicode.    We call them `dotbelow', `hookabove', and `cwm',
%   respectively.
%
% .  The following glyphs have no special name in the AGL (or names which
%   should not be used in general) and are thus represented as a Unicode
%   entity:
%
%     0x1A  uni1EF6  Y hook above
%     0x1B  uni1EF7  y hook above
%     0x1C  uni1EF4  Y dot below
%     0x1D  uni1EF5  y dot below
%
%     0x20  uni2423  graphic symbol for space
%
%     0x83  uni1EA2  A hook above
%     0x84  uni1EA0  A dot below
%     0x86  uni1EA6  A circumflex grave
%     0x87  uni1EA4  A circumflex acute
%     0x88  uni1EAA  A circumflex tilde
%     0x89  uni1EA8  A circumflex hook above
%     0x8A  uni1EAC  A circumflex dot below
%     0x8C  uni1EB0  A breve grave

%     0x8D  uni1EAE  A breve acute
%     0x8E  uni1EB4  A breve tilde
%     0x8F  uni1EB2  A breve hook above
%     0x90  uni1EB6  A breve dot below
%
%     0x93  uni1EBC  E tilde
%     0x94  uni1EBA  E hook above
%     0x95  uni1EB8  E dot below
%     0x97  uni1EC0  E circumflex grave
%     0x98  uni1EBE  E circumflex acute
%     0x99  uni1EC4  E circumflex tilde
%     0x9A  uni1EC2  E circumflex hook above
%     0x9B  uni1EC6  E circumflex dot below
%
%     0x9F  uni1EC8  I hook above
%
%     0xA3  uni1EA3  a hook above
%     0xA4  uni1EA1  a dot below
%     0xA6  uni1EA7  a circumflex grave
%     0xA7  uni1EA5  a circumflex acute
%     0xA8  uni1EAB  a circumflex tilde
%     0xA9  uni1EA9  a circumflex hook above
%     0xAA  uni1EAD  a circumflex dot below
%     0xAC  uni1EB1  a breve grave
%     0xAD  uni1EAF  a breve acute
%     0xAE  uni1EB5  a breve tilde
%     0xAF  uni1EB3  a breve hook above
%     0xB0  uni1EB7  a breve dot below
%
%     0xB3  uni1EBD  e tilde
%     0xB4  uni1EBB  e hook above
%     0xB5  uni1EB9  e dot below
%     0xB7  uni1EC1  e circumflex grave
%     0xB8  uni1EBF  e circumflex acute
%     0xB9  uni1EC5  e circumflex tilde
%     0xBA  uni1EC3  e circumflex hook above
%     0xBB  uni1EC7  e circumflex dot below
%
%     0xBF  uni1EC9  i hook above
%
%     0xC0  uni1ECA  I dot below
%
%     0xC4  uni1ECE  O hook above
%     0xC5  uni1ECC  O dot below
%     0xC7  uni1ED2  O circumflex grave
%     0xC8  uni1ED0  O circumflex acute
%     0xC9  uni1ED6  O circumflex tilde

%     0xCA  uni1ED4  O circumflex hook above
%     0xCB  uni1ED8  O circumflex dot below
%     0xCD  uni1EDC  O horn grave
%     0xCE  uni1EDA  O horn acute
%     0xCF  uni1EE0  O horn tilde
%     0xD0  uni1EDE  O horn hook above
%     0xD1  uni1EE2  O horn dot below
%
%     0xD5  uni1EE6  U hook above
%     0xD6  uni1EE4  U dot below
%     0xD8  uni1EEA  U horn grave
%     0xD9  uni1EE8  U horn acute
%     0xDA  uni1EEE  U horn tilde
%     0xDB  uni1EEC  U horn hook above
%     0xDC  uni1EF0  U horn dot below
%
%     0xDF  uni1EF8  Y tilde
%
%     0xE0  uni1ECB  i dot below
%
%     0xE4  uni1ECF  o hook above
%     0xE5  uni1ECD  o dot below
%     0xE7  uni1ED3  o circumflex grave
%     0xE8  uni1ED1  o circumflex acute
%     0xE9  uni1ED7  o circumflex tilde
%     0xEA  uni1ED5  o circumflex hook above
%     0xEB  uni1ED9  o circumflex dot below
%     0xED  uni1EDD  o horn grave
%     0xEE  uni1EDB  o horn acute
%     0xEF  uni1EE1  o horn tilde
%     0xF0  uni1EDF  o horn hook above
%     0xF1  uni1EE3  o horn dot below
%
%     0xF5  uni1EE7  u hook above
%     0xF6  uni1EE5  u dot below
%     0xF8  uni1EEB  u horn grave
%     0xF9  uni1EE9  u horn acute
%     0xFA  uni1EEF  u horn tilde
%     0xFB  uni1EED  u horn hook above
%     0xFC  uni1EF1  u horn dot below
%
%     0xFF  uni1EF9  y tilde
/T5Encoding [
% 0x00
  /grave /acute /circumflex /tilde
  /dieresis /dotbelow /ring /caron
  /breve /macron /dotaccent /cedilla

/hookabove /quotesinglbase /guilsinglleft /guilsinglright
% 0x10
/quotedblleft /quotedblright /quotedblbase /guillemotleft
/guillemotright /endash /emdash /cwm
/zeroinferior /dotlessi /uni1EF6 /uni1EF7
/uni1EF4 /uni1EF5 /Dcroat /dcroat
% 0x20
/uni2423 /exclam /quotedbl /numbersign
/dollar /percent /ampersand /quoteright
/parenleft /parenright /asterisk /plus
/comma /hyphen /period /slash
% 0x30
/zero /one /two /three
/four /five /six /seven
/eight /nine /colon /semicolon
/less /equal /greater /question
% 0x40
/at /A /B /C
/D /E /F /G
/H /I /J /K
/L /M /N /O
% 0x50
/P /Q /R /S
/T /U /V /W
/X /Y /Z /bracketleft
/backslash /bracketright /asciicircum /underscore
% 0x60
/quoteleft /a /b /c
/d /e /f /g
/h /i /j /k
/l /m /n /o
% 0x70
/p /q /r /s
/t /u /v /w
/x /y /z /braceleft
/bar /braceright /asciitilde /sfthyphen
% 0x80
/Agrave /Aacute /Atilde /uni1EA2
/uni1EA0 /Acircumflex /uni1EA6 /uni1EA4
/uni1EAA /uni1EA8 /uni1EAC /Abreve
/uni1EB0 /uni1EAE /uni1EB4 /uni1EB2
% 0x90
/uni1EB6 /Egrave /Eacute /uni1EBC
/uni1EBA /uni1EB8 /Ecircumflex /uni1EC0
/uni1EBE /uni1EC4 /uni1EC2 /uni1EC6
/Igrave /Iacute /Itilde /uni1EC8
% 0xA0

/agrave /aacute /atilde /uni1EA3
/uni1EA1 /acircumflex /uni1EA7 /uni1EA5
/uni1EAB /uni1EA9 /uni1EAD /abreve
/uni1EB1 /uni1EAF /uni1EB5 /uni1EB3
% 0xB0
/uni1EB7 /egrave /eacute /uni1EBD
/uni1EBB /uni1EB9 /ecircumflex /uni1EC1
/uni1EBF /uni1EC5 /uni1EC3 /uni1EC7
/igrave /iacute /itilde /uni1EC9
% 0xC0
/uni1ECA /Ograve /Oacute /Otilde
/uni1ECE /uni1ECC /Ocircumflex /uni1ED2
/uni1ED0 /uni1ED6 /uni1ED4 /uni1ED8
/Ohorn /uni1EDC /uni1EDA /uni1EE0
% 0xD0
/uni1EDE /uni1EE2 /Ugrave /Uacute
/Utilde /uni1EE6 /uni1EE4 /Uhorn
/uni1EEA /uni1EE8 /uni1EEE /uni1EEC
/uni1EF0 /Ygrave /Yacute /uni1EF8
% 0xE0
/uni1ECB /ograve /oacute /otilde
/uni1ECF /uni1ECD /ocircumflex /uni1ED3
/uni1ED1 /uni1ED7 /uni1ED5 /uni1ED9
/ohorn /uni1EDD /uni1EDB /uni1EE1
% 0xF0
/uni1EDF /uni1EE3 /ugrave /uacute
/utilde /uni1EE7 /uni1EE5 /uhorn
/uni1EEB /uni1EE9 /uni1EEF /uni1EED
/uni1EF1 /ygrave /yacute /uni1EF9
] def
% end of t5.enc
B.8 ‘qx.enc’

From GUST.

% =========================================================================
                       QX LAYOUT ver.  0.06
% =========================================================================
Release date:  24 August 2003
Original URL: ftp://bop.eps.gda.pl/pub/qx/
E-mail:  JNSteam at gust.org.pl
QX Layout is, actually, a pair of encodings:  one for TeX and one for GUI
systems.  QX layout has been worked out by the Polish TeX Users Group GUST as
a difficult compromise between needs and abilities.  The character set of the
QX layout is expected to suffice for the typesetting of texts (using TeX) in

most of European languages.
QX Encoding for TeX is based on EC (Cork Encoding) with a few characters
exchanged.  E.g., gbreve, Gbreve, uring, and Uring, which can be obtained by
superimposing of an accent, are replaced by Lithuanian iogonek, Iogonek,
uogonek, and Uogonek, respectively, which should not be assembled form
separate elements; QX Encoding for GUI systems is a variant of the Code Page
1250 (and also includes Lithuanian characters with ogonek).
After a long debate, we decided (in ver.  0.04 of the QX layout) to replace
omnipresent but actually nonexistent [tT]cedilla by Romanian [tT]commaaccent;
censequently, Turkish [sS]cedilla was replaced by [sS]commaaccent (still we
would recommend to keep the glyph of a [sS]cedilla character in a font).
Under the pressure from euro currency funs, we included moreover an
Euro symbol (in ver.  0.05 of the QX layout).  We assigned the slot
128 for it complying with Microsoft proposal for non-cyrillic fonts
(http://www.microsoft.com/typography/faq/faq12.htm):
  The symbol has been added to the following codepages at position
  '0x80'; 1250 Eastern European, 1252 Western, 1253 Greek, 1254 Turkish,
  1257 Baltic, 1255 Hebrew, 1256 Arabic, 1258 Vietnamese, 874 Thai.
  In 1251 Cyrillic the symbol will be added at position '0x88'.
Therefore, we had to do something with a Lithuanian Iogonek character that
previously had occupied this slot.  We decided to remove Thorn (222, Tcedilla
in CP 1250) and thorn (254, tcedilla in CP 1250) characters from the Windows
QX layout (we recommend, as with [sS]cedilla characters, to keep the glyphs
in a font) and use the freed slots for Iogonek (previously 128) and iogonek
(previously 144, now free), respectively.  The TeX QX layout, fortunately, had
had the slot 128 already free, hence there was no need for rearrangement.
Finally, we decided to change the name of a (non-standard) ligature
fk to f_k, complying with Adobe recommendations contained in:
http://partners.adobe.com/asn/developer/typeforum/unicodegn.html
Most recent change (24.08.2003, ver.  0.06) was the extension of
the character set by ligatures ij, IJ (available only in the TeX
encoding) and a punctuation mark threequartersemdash.
Below you'll find the comparison of basic layouts for TeX and GUI systems
with QX ver.  0.06.
% ===========================================================================
%                          COMPARISON OF TeX LAYOUTS
% ===========================================================================
% CODE             CM                EC                PL        QX ver.  0.06

% ===========================================================================
 0  00         Gamma             grave             Gamma               alpha
 1  01         Delta             acute             Delta               Delta
 2  02         Theta        circumflex             Theta                beta
 3  03        Lambda             tilde            Lambda               delta
 4  04            Xi          dieresis                Xi                  pi
 5  05            Pi      hungarumlaut                Pi                  Pi
 6  06         Sigma              ring             Sigma               Sigma
 7  07       Upsilon             caron           Upsilon                  mu
 8  08           Phi             breve               Phi            ellipsis
 9  09           Psi            macron               Psi                 f_k
10  0A         Omega         dotaccent             Omega               Omega
11  0B            ff           cedilla                ff                  ff
12  0C            fi            ogonek                fi                  fi
13  0D            fl    quotesinglbase                fl                  fl
14  0E           ffi     guilsinglleft               ffi                 ffi
15  0F           ffl    guilsinglright               ffl                 ffl
16  10      dotlessi      quotedblleft          dotlessi            dotlessi
17  11      dotlessj     quotedblright          dotlessj            dotlessj
18  12         grave      quotedblbase             grave               grave
19  13         acute     guillemotleft             acute               acute
20  14         caron    guillemotright             caron               caron
21  15         breve            endash             breve               breve
22  16        macron            emdash            macron              macron
23  17          ring  compoundwordmark              ring                ring
24  18       cedilla       perthousand           cedilla             cedilla
25  19    germandbls          dotlessi        germandbls          germandbls
26  1A            ae          dotlessj                ae                  ae
27  1B            oe                ff                oe                  oe
28  1C        oslash                fi            oslash              oslash
29  1D            AE                fl                AE                  AE
30  1E            OE               ffi                OE                  OE
31  1F        Oslash               ffl            Oslash              Oslash
32  20      suppress       visualspace            ogonek               space
33  21        exclam            exclam            exclam              exclam
34  22 quotedblright          quotedbl     quotedblright       quotedblright
35  23    numbersign        numbersign        numbersign          numbersign
36  24        dollar            dollar            dollar              dollar
37  25       percent           percent           percent             percent
38  26     ampersand         ampersand         ampersand           ampersand
39  27    quoteright        quoteright        quoteright          quoteright
40  28     parenleft         parenleft         parenleft           parenleft
41  29    parenright        parenright        parenright          parenright
42  2A      asterisk          asterisk          asterisk            asterisk
43  2B          plus              plus              plus                plus
44  2C         comma             comma             comma               comma
45  2D        hyphen            hyphen            hyphen              hyphen

46  2E        period            period            period              period
47  2F         slash             slash             slash               slash
48  30          zero              zero              zero                zero
49  31           one               one               one                 one
50  32           two               two               two                 two
51  33         three             three             three               three
52  34          four              four              four                four
53  35          five              five              five                five
54  36           six               six               six                 six
55  37         seven             seven             seven               seven
56  38         eight             eight             eight               eight
57  39          nine              nine              nine                nine
58  3A         colon             colon             colon               colon
59  3B     semicolon         semicolon         semicolon           semicolon
60  3C    exclamdown              less        exclamdown          exclamdown
61  3D         equal             equal             equal               equal
62  3E  questiondown           greater      questiondown        questiondown
63  3F      question          question          question            question
64  40            at                at                at                  at
65  41             A                 A                 A                   A
66  42             B                 B                 B                   B
67  43             C                 C                 C                   C
68  44             D                 D                 D                   D
69  45             E                 E                 E                   E
70  46             F                 F                 F                   F
71  47             G                 G                 G                   G
72  48             H                 H                 H                   H
73  49             I                 I                 I                   I
74  4A             J                 J                 J                   J
75  4B             K                 K                 K                   K
76  4C             L                 L                 L                   L
77  4D             M                 M                 M                   M
78  4E             N                 N                 N                   N
79  4F             O                 O                 O                   O
80  50             P                 P                 P                   P
81  51             Q                 Q                 Q                   Q
82  52             R                 R                 R                   R
83  53             S                 S                 S                   S
84  54             T                 T                 T                   T
85  55             U                 U                 U                   U
86  56             V                 V                 V                   V
87  57             W                 W                 W                   W
88  58             X                 X                 X                   X
89  59             Y                 Y                 Y                   Y
90  5A             Z                 Z                 Z                   Z
91  5B   bracketleft       bracketleft       bracketleft         bracketleft
92  5C  quotedblleft         backslash      quotedblleft        quotedblleft

 93  5D  bracketright      bracketright      bracketright        bracketright
 94  5E    circumflex       asciicircum        circumflex          circumflex
 95  5F     dotaccent        underscore         dotaccent           dotaccent
 96  60     quoteleft         quoteleft         quoteleft           quoteleft
 97  61             a                 a                 a                   a
 98  62             b                 b                 b                   b
 99  63             c                 c                 c                   c
100  64             d                 d                 d                   d
101  65             e                 e                 e                   e
102  66             f                 f                 f                   f
103  67             g                 g                 g                   g
104  68             h                 h                 h                   h
105  69             i                 i                 i                   i
106  6A             j                 j                 j                   j
107  6B             k                 k                 k                   k
108  6C             l                 l                 l                   l
109  6D             m                 m                 m                   m
110  6E             n                 n                 n                   n
111  6F             o                 o                 o                   o
112  70             p                 p                 p                   p
113  71             q                 q                 q                   q
114  72             r                 r                 r                   r
115  73             s                 s                 s                   s
116  74             t                 t                 t                   t
117  75             u                 u                 u                   u
118  76             v                 v                 v                   v
119  77             w                 w                 w                   w
120  78             x                 x                 x                   x
121  79             y                 y                 y                   y
122  7A             z                 z                 z                   z
123  7B        endash         braceleft            endash              endash
124  7C        emdash               bar            emdash              emdash
125  7D  hungarumlaut        braceright      hungarumlaut        hungarumlaut
126  7E         tilde        asciitilde             tilde               tilde
127  7F      dieresis            hyphen          dieresis            dieresis
128  80             @            Abreve                 @                Euro
129  81             @           Aogonek           Aogonek             Aogonek
130  82             @            Cacute            Cacute              Cacute
131  83             @            Ccaron                 @             greater
132  84             @            Dcaron                 @        greaterequal
133  85             @            Ecaron                 @         approxequal
134  86             @           Eogonek           Eogonek             Eogonek
135  87             @            Gbreve                 @             Iogonek
136  88             @            Lacute                 @                less
137  89             @       Lquoteright                 @           lessequal
138  8A             @            Lslash            Lslash              Lslash
139  8B             @            Nacute            Nacute              Nacute

140  8C             @            Ncaron                 @          asciitilde
141  8D             @               Eng                 @         asciicircum
142  8E             @     Ohungarumlaut                 @                 ell
143  8F             @            Racute                 @              dagger
144  90             @            Rcaron                 @           daggerdbl
145  91             @            Sacute            Sacute              Sacute
146  92             @            Scaron                 @              Scaron
147  93             @          Scedilla                 @        Scommaaccent
148  94             @            Tcaron                 @              degree
149  95             @          Tcedilla                 @        Tcommaaccent
150  96             @     Uhungarumlaut                 @              ogonek
151  97             @             Uring                 @             Uogonek
152  98             @         Ydieresis                 @           Ydieresis
153  99             @            Zacute            Zacute              Zacute
154  9A             @            Zcaron                 @              Zcaron
155  9B             @        Zdotaccent        Zdotaccent          Zdotaccent
156  9C             @                IJ                 @                  IJ
157  9D             @        Idotaccent                 @           braceleft
158  9E             @              dbar                 @          braceright
159  9F             @           section                 @             section
160  A0             @            abreve                 @                   @
161  A1             @           aogonek           aogonek             aogonek
162  A2             @            cacute            cacute              cacute
163  A3             @            ccaron                 @          registered
164  A4             @       dquoteright                 @           copyright
165  A5             @            ecaron                 @              divide
166  A6             @           eogonek           eogonek             eogonek
167  A7             @            gbreve                 @             iogonek
168  A8             @            lacute                 @               minus
169  A9             @       lquoteright                 @            multiply
170  AA             @            lslash            lslash              lslash
171  AB             @            nacute            nacute              nacute
172  AC             @            ncaron                 @           plusminus
173  AD             @               eng                 @            infinity
174  AE             @     ohungarumlaut     guillemotleft       guillemotleft
175  AF             @            racute    guillemotright      guillemotright
176  B0             @            rcaron                 @           paragraph
177  B1             @            sacute            sacute              sacute
178  B2             @            scaron                 @              scaron
179  B3             @          scedilla                 @        scommaaccent
180  B4             @       tquoteright                 @              bullet
181  B5             @          tcedilla                 @        tcommaaccent
182  B6             @     uhungarumlaut                 @ threequartersemdash
183  B7             @             uring                 @             uogonek
184  B8             @         ydieresis                 @           ydieresis
185  B9             @            zacute            zacute              zacute
186  BA             @            zcaron                 @              zcaron

187  BB             @        zdotaccent        zdotaccent          zdotaccent
188  BC             @                ij                 @                  ij
189  BD             @        exclamdown                 @      periodcentered
190  BE             @      questiondown                 @            quotedbl
191  BF             @          sterling                 @         quotesingle
192  C0             @            Agrave                 @              Agrave
193  C1             @            Aacute                 @              Aacute
194  C2             @       Acircumflex                 @         Acircumflex
195  C3             @            Atilde                 @              Atilde
196  C4             @         Adieresis                 @           Adieresis
197  C5             @             Aring                 @               Aring
198  C6             @                AE                 @           backslash
199  C7             @          Ccedilla                 @            Ccedilla
200  C8             @            Egrave                 @              Egrave
201  C9             @            Eacute                 @              Eacute
202  CA             @       Ecircumflex                 @         Ecircumflex
203  CB             @         Edieresis                 @           Edieresis
204  CC             @            Igrave                 @              Igrave
205  CD             @            Iacute                 @              Iacute
206  CE             @       Icircumflex                 @         Icircumflex
207  CF             @         Idieresis                 @           Idieresis
208  D0             @               Eth                 @                 Eth
209  D1             @            Ntilde                 @              Ntilde
210  D2             @            Ograve                 @              Ograve
211  D3             @            Oacute            Oacute              Oacute
212  D4             @       Ocircumflex                 @         Ocircumflex
213  D5             @            Otilde                 @              Otilde
214  D6             @         Odieresis                 @           Odieresis
215  D7             @                OE                 @            currency
216  D8             @            Oslash                 @         perthousand
217  D9             @            Ugrave                 @              Ugrave
218  DA             @            Uacute                 @              Uacute
219  DB             @       Ucircumflex                 @         Ucircumflex
220  DC             @         Udieresis                 @           Udieresis
221  DD             @            Yacute                 @              Yacute
222  DE             @             Thorn                 @               Thorn
223  DF             @        Germandbls                 @                 bar
224  E0             @            agrave                 @              agrave
225  E1             @            aacute                 @              aacute
226  E2             @       acircumflex                 @         acircumflex
227  E3             @            atilde                 @              atilde
228  E4             @         adieresis                 @           adieresis
229  E5             @             aring                 @               aring
230  E6             @                ae                 @          underscore
231  E7             @          ccedilla                 @            ccedilla
232  E8             @            egrave                 @              egrave
233  E9             @            eacute                 @              eacute

234  EA             @       ecircumflex                 @         ecircumflex
235  EB             @         edieresis                 @           edieresis
236  EC             @            igrave                 @              igrave
237  ED             @            iacute                 @              iacute
238  EE             @       icircumflex                 @         icircumflex
239  EF             @         idieresis                 @           idieresis
240  F0             @               eth                 @                 eth
241  F1             @            ntilde                 @              ntilde
242  F2             @            ograve                 @              ograve
243  F3             @            oacute            oacute              oacute
244  F4             @       ocircumflex                 @         ocircumflex
245  F5             @            otilde                 @              otilde
246  F6             @         odieresis                 @           odieresis
247  F7             @                oe                 @            anglearc
248  F8             @            oslash                 @            diameter
249  F9             @            ugrave                 @              ugrave
250  FA             @            uacute                 @              uacute
251  FB             @       ucircumflex                 @         ucircumflex
252  FC             @         udieresis                 @           udieresis
253  FD             @            yacute                 @              yacute
254  FE             @             thorn                 @               thorn
255  FF             @        germandbls      quotedblbase        quotedblbase
% ===========================================================================

% ==========================================================
%          COMPARISON OF LAYOUTS FOR GUI SYSTEMS
% ==========================================================
% CODE        NEW ANSI (0--127)    Macintosh    QX ver.  0.06
%             +  CODE PAGE 1250     PLEURO
% ==========================================================
  0  00              @                 @                   @
  1  01              @                 @               acute
  2  02              @                 @          circumflex
  3  03              @                 @             cedilla
  4  04              @                 @            dieresis
  5  05              @                 @               breve
  6  06              @                 @           dotaccent
  7  07              @                 @              macron
  8  08              @                 @                ring
  9  09              @                 @                   @
 10  0A              @                 @        hungarumlaut
 11  0B              @                 @              ogonek
 12  0C              @                 @               caron
 13  0D              @                 @                   @
 14  0E              @                 @            dotlessi
 15  0F              @                 @            dotlessj

16  10              @                 @               tilde
17  11              @                 @                  ff
18  12              @                 @                 ffi
19  13              @                 @                 ffl
20  14              @                 @                 f_k
21  15              @                 @         approxequal
22  16              @                 @                 ell
23  17              @                 @            infinity
24  18              @                 @               alpha
25  19              @                 @                beta
26  1A              @                 @               delta
27  1B              @                 @                  pi
28  1C              @                 @               Delta
29  1D              @                 @                  Pi
30  1E              @                 @               Sigma
31  1F              @                 @               Omega
32  20          space             space               space
33  21         exclam            exclam              exclam
34  22       quotedbl          quotedbl            quotedbl
35  23     numbersign        numbersign          numbersign
36  24         dollar            dollar              dollar
37  25        percent           percent             percent
38  26      ampersand         ampersand           ampersand
39  27    quotesingle       quotesingle         quotesingle
40  28      parenleft         parenleft           parenleft
41  29     parenright        parenright          parenright
42  2A       asterisk          asterisk            asterisk
43  2B           plus              plus                plus
44  2C          comma             comma               comma
45  2D         hyphen            hyphen              hyphen
46  2E         period            period              period
47  2F          slash             slash               slash
48  30           zero              zero                zero
49  31            one               one                 one
50  32            two               two                 two
51  33          three             three               three
52  34           four              four                four
53  35           five              five                five
54  36            six               six                 six
55  37          seven             seven               seven
56  38          eight             eight               eight
57  39           nine              nine                nine
58  3A          colon             colon               colon
59  3B      semicolon         semicolon           semicolon
60  3C           less              less                less
61  3D          equal             equal               equal
62  3E        greater           greater             greater

63  3F       question          question            question
64  40             at                at                  at
65  41              A                 A                   A
66  42              B                 B                   B
67  43              C                 C                   C
68  44              D                 D                   D
69  45              E                 E                   E
70  46              F                 F                   F
71  47              G                 G                   G
72  48              H                 H                   H
73  49              I                 I                   I
74  4A              J                 J                   J
75  4B              K                 K                   K
76  4C              L                 L                   L
77  4D              M                 M                   M
78  4E              N                 N                   N
79  4F              O                 O                   O
80  50              P                 P                   P
81  51              Q                 Q                   Q
82  52              R                 R                   R
83  53              S                 S                   S
84  54              T                 T                   T
85  55              U                 U                   U
86  56              V                 V                   V
87  57              W                 W                   W
88  58              X                 X                   X
89  59              Y                 Y                   Y
90  5A              Z                 Z                   Z
91  5B    bracketleft       bracketleft         bracketleft
92  5C      backslash         backslash           backslash
93  5D   bracketright      bracketright        bracketright
94  5E    asciicircum       asciicircum         asciicircum
95  5F     underscore        underscore          underscore
96  60          grave             grave               grave
97  61              a                 a                   a
98  62              b                 b                   b
99  63              c                 c                   c
100  64              d                 d                   d
101  65              e                 e                   e
102  66              f                 f                   f
103  67              g                 g                   g
104  68              h                 h                   h
105  69              i                 i                   i
106  6A              j                 j                   j
107  6B              k                 k                   k
108  6C              l                 l                   l
109  6D              m                 m                   m

110  6E              n                 n                   n
111  6F              o                 o                   o
112  70              p                 p                   p
113  71              q                 q                   q
114  72              r                 r                   r
115  73              s                 s                   s
116  74              t                 t                   t
117  75              u                 u                   u
118  76              v                 v                   v
119  77              w                 w                   w
120  78              x                 x                   x
121  79              y                 y                   y
122  7A              z                 z                   z
123  7B      braceleft         braceleft           braceleft
124  7C            bar               bar                 bar
125  7D     braceright        braceright          braceright
126  7E     asciitilde        asciitilde          asciitilde
127  7F              @                 @                   @
128  80           Euro                 @                Euro
129  81              @                 @                  fi
130  82 quotesinglbase    quotesinglbase                  fl
131  83              @         Ydieresis           Ydieresis
132  84   quotedblbase      quotedblbase        quotedblbase
133  85       ellipsis          ellipsis            ellipsis
134  86         dagger            dagger              dagger
135  87      daggerdbl         daggerdbl           daggerdbl
136  88              @       Ecircumflex         Ecircumflex
137  89    perthousand       perthousand         perthousand
138  8A         Scaron            Scaron              Scaron
139  8B  guilsinglleft     guilsinglleft        Scommaaccent
140  8C         Sacute            Sacute              Sacute
141  8D         Tcaron                OE                  OE
142  8E         Zcaron            Zcaron              Zcaron
143  8F         Zacute            Zacute              Zacute
144  90              @                 @ threequartersemdash
145  91      quoteleft         quoteleft           quoteleft
146  92     quoteright        quoteright          quoteright
147  93   quotedblleft      quotedblleft        quotedblleft
148  94  quotedblright     quotedblright       quotedblright
149  95         bullet            bullet              bullet
150  96         endash            endash              endash
151  97         emdash            emdash              emdash
152  98              @       ecircumflex         ecircumflex
153  99      trademark         trademark            anglearc
154  9A         scaron            scaron              scaron
155  9B guilsinglright    guilsinglright        scommaaccent
156  9C         sacute            sacute              sacute

157  9D    tquoteright                oe                  oe
158  9E         zcaron            zcaron              zcaron
159  9F         zacute            zacute              zacute
160  A0              @           nbspace             nbspace
161  A1          caron        exclamdown          exclamdown
162  A2          breve      questiondown        questiondown
163  A3         Lslash            Lslash              Lslash
164  A4       currency          currency            currency
165  A5        Aogonek           Aogonek             Aogonek
166  A6      brokenbar         brokenbar            diameter
167  A7        section           section             section
168  A8       dieresis          dieresis             Uogonek
169  A9      copyright         copyright           copyright
170  AA       Scedilla                AE                  AE
171  AB  guillemotleft     guillemotleft       guillemotleft
172  AC     logicalnot            Ntilde              Ntilde
173  AD         hyphen             minus               minus
174  AE     registered        registered          registered
175  AF     Zdotaccent        Zdotaccent          Zdotaccent
176  B0         degree            degree              degree
177  B1      plusminus         plusminus           plusminus
178  B2         ogonek       twosuperior           lessequal
179  B3         lslash            lslash              lslash
180  B4          acute             acute        greaterequal
181  B5             mu                mu                  mu
182  B6      paragraph         paragraph           paragraph
183  B7 periodcentered    periodcentered      periodcentered
184  B8        cedilla                 @             uogonek
185  B9        aogonek           aogonek             aogonek
186  BA       scedilla                ae                  ae
187  BB guillemotright    guillemotright      guillemotright
188  BC    Lquoteright            ntilde              ntilde
189  BD   hungarumlaut           onehalf        Tcommaaccent
190  BE    lquoteright     threequarters        tcommaaccent
191  BF     zdotaccent        zdotaccent          zdotaccent
192  C0         Racute            Agrave              Agrave
193  C1         Aacute            Aacute              Aacute
194  C2    Acircumflex       Acircumflex         Acircumflex
195  C3         Abreve            Atilde              Atilde
196  C4      Adieresis         Adieresis           Adieresis
197  C5         Lacute             Aring               Aring
198  C6         Cacute            Cacute              Cacute
199  C7       Ccedilla          Ccedilla            Ccedilla
200  C8         Ccaron            Egrave              Egrave
201  C9         Eacute            Eacute              Eacute
202  CA        Eogonek           Eogonek             Eogonek
203  CB      Edieresis         Edieresis           Edieresis

204  CC         Ecaron            Igrave              Igrave
205  CD         Iacute            Iacute              Iacute
206  CE    Icircumflex       Icircumflex         Icircumflex
207  CF         Dcaron         Idieresis           Idieresis
208  D0            Eth               Eth                 Eth
209  D1         Nacute            Nacute              Nacute
210  D2         Ncaron            Ograve              Ograve
211  D3         Oacute            Oacute              Oacute
212  D4    Ocircumflex       Ocircumflex         Ocircumflex
213  D5  Ohungarumlaut            Otilde              Otilde
214  D6      Odieresis         Odieresis           Odieresis
215  D7       multiply          multiply            multiply
216  D8         Rcaron            Oslash              Oslash
217  D9          Uring            Ugrave              Ugrave
218  DA         Uacute            Uacute              Uacute
219  DB  Uhungarumlaut       Ucircumflex         Ucircumflex
220  DC      Udieresis         Udieresis           Udieresis
221  DD         Yacute            Yacute              Yacute
222  DE       Tcedilla             Thorn             Iogonek
223  DF     germandbls        germandbls          germandbls
224  E0         racute            agrave              agrave
225  E1         aacute            aacute              aacute
226  E2    acircumflex       acircumflex         acircumflex
227  E3         abreve            atilde              atilde
228  E4      adieresis         adieresis           adieresis
229  E5         lacute             aring               aring
230  E6         cacute            cacute              cacute
231  E7       ccedilla          ccedilla            ccedilla
232  E8         ccaron            egrave              egrave
233  E9         eacute            eacute              eacute
234  EA        eogonek           eogonek             eogonek
235  EB      edieresis         edieresis           edieresis
236  EC         ecaron            igrave              igrave
237  ED         iacute            iacute              iacute
238  EE    icircumflex       icircumflex         icircumflex
239  EF    dquoteright         idieresis           idieresis
240  F0           dbar               eth                 eth
241  F1         nacute            nacute              nacute
242  F2         ncaron            ograve              ograve
243  F3         oacute            oacute              oacute
244  F4    ocircumflex       ocircumflex         ocircumflex
245  F5  ohungarumlaut            otilde              otilde
246  F6      odieresis         odieresis           odieresis
247  F7         divide            divide              divide
248  F8         rcaron            oslash              oslash
249  F9          uring            ugrave              ugrave
250  FA         uacute            uacute              uacute

251  FB  uhungarumlaut       ucircumflex         ucircumflex
252  FC      udieresis         udieresis           udieresis
253  FD         yacute            yacute              yacute
254  FE       tcedilla             thorn             iogonek
255  FF      dotaccent         ydieresis           ydieresis
% ==========================================================

/encqxoosix[
/alpha
/Delta
/beta
/delta
/pi
/Pi
/Sigma
/mu
/ellipsis
/f_k
/Omega
/ff
/fi
/fl
/ffi
/ffl
/dotlessi
/dotlessj
/grave
/acute
/caron
/breve
/macron
/ring
/cedilla
/germandbls
/ae
/oe
/oslash
/AE
/OE
/Oslash
/space
/exclam
/quotedblright
/numbersign
/dollar
/percent

/ampersand
/quoteright
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
/zero
/one
/two
/three
/four
/five
/six
/seven
/eight
/nine
/colon
/semicolon
/exclamdown
/equal
/questiondown
/question
/at
/A
/B
/C
/D
/E
/F
/G
/H
/I
/J
/K
/L
/M
/N
/O
/P
/Q
/R
/S
/T

/U
/V
/W
/X
/Y
/Z
/bracketleft
/quotedblleft
/bracketright
/circumflex
/dotaccent
/quoteleft
/a
/b
/c
/d
/e
/f
/g
/h
/i
/j
/k
/l
/m
/n
/o
/p
/q
/r
/s
/t
/u
/v
/w
/x
/y
/z
/endash
/emdash
/hungarumlaut
/tilde
/dieresis
/Euro
/Aogonek
/Cacute
/greater

/greaterequal
/approxequal
/Eogonek
/Iogonek
/less
/lessequal
/Lslash
/Nacute
/asciitilde
/asciicircum
/ell
/dagger
/daggerdbl
/Sacute
/Scaron
/Scommaaccent
/degree
/Tcommaaccent
/ogonek
/Uogonek
/Ydieresis
/Zacute
/Zcaron
/Zdotaccent
/IJ
/braceleft
/braceright
/section
/.notdef
/aogonek
/cacute
/registered
/copyright
/divide
/eogonek
/iogonek
/minus
/multiply
/lslash
/nacute
/plusminus
/infinity
/guillemotleft
/guillemotright
/paragraph
/sacute
/scaron

/scommaaccent
/bullet
/tcommaaccent
/threequartersemdash
/uogonek
/ydieresis
/zacute
/zcaron
/zdotaccent
/ij
/periodcentered
/quotedbl
/quotesingle
/Agrave
/Aacute
/Acircumflex
/Atilde
/Adieresis
/Aring
/backslash
/Ccedilla
/Egrave
/Eacute
/Ecircumflex
/Edieresis
/Igrave
/Iacute
/Icircumflex
/Idieresis
/Eth
/Ntilde
/Ograve
/Oacute
/Ocircumflex
/Otilde
/Odieresis
/currency
/perthousand
/Ugrave
/Uacute
/Ucircumflex
/Udieresis
/Yacute
/Thorn
/bar
/agrave
/aacute

/acircumflex
/atilde
/adieresis
/aring
/underscore
/ccedilla
/egrave
/eacute
/ecircumflex
/edieresis
/igrave
/iacute
/icircumflex
/idieresis
/eth
/ntilde
/ograve
/oacute
/ocircumflex
/otilde
/odieresis
/anglearc
/diameter
/ugrave
/uacute
/ucircumflex
/udieresis
/yacute
/thorn
/quotedblbase
] def

B.9 ‘texmext.enc’: TEX math extension encoding

% @psencodingfile{
%   author = "Donald Knuth",
%   version = "1.0",
%   date = "12jun05",
%   filename = "texmext.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "TeX math extension encoding."
% }
%
% 12jun05 - add header.

%
% No ligatures or kerns.
%
/TeXMathExtensionEncoding [          % now 256 chars follow
% 0x0
/parenleftbig
/parenrightbig
/bracketleftbig
/bracketrightbig
/floorleftbig
/floorrightbig
/ceilingleftbig
/ceilingrightbig
% 0x8
/braceleftbig
/bracerightbig
/angbracketleftbig
/angbracketrightbig
/vextendsingle
/vextenddouble
/slashbig
/backslashbig
% 0x10
/parenleftBig
/parenrightBig
/parenleftbigg
/parenrightbigg
/bracketleftbigg
/bracketrightbigg
/floorleftbigg
/floorrightbigg
% 0x18
/ceilingleftbigg
/ceilingrightbigg
/braceleftbigg
/bracerightbigg
/angbracketleftbigg
/angbracketrightbigg
/slashbigg
/backslashbigg
% 0x20
/parenleftBigg

/parenrightBigg
/bracketleftBigg
/bracketrightBigg
/floorleftBigg
/floorrightBigg
/ceilingleftBigg
/ceilingrightBigg
% 0x28
/braceleftBigg
/bracerightBigg
/angbracketleftBigg
/angbracketrightBigg
/slashBigg
/backslashBigg
/slashBig
/backslashBig
% 0x30
/parenlefttp
/parenrighttp
/bracketlefttp
/bracketrighttp
/bracketleftbt
/bracketrightbt
/bracketleftex
/bracketrightex
% 0x38
/bracelefttp
/bracerighttp
/braceleftbt
/bracerightbt
/braceleftmid
/bracerightmid
/braceex
/arrowvertex
% 0x40
/parenleftbt
/parenrightbt
/parenleftex
/parenrightex
/angbracketleftBig
/angbracketrightBig
/unionsqtext
/unionsqdisplay

% 0x48
/contintegraltext
/contintegraldisplay
/circledottext
/circledotdisplay
/circleplustext
/circleplusdisplay
/circlemultiplytext
/circlemultiplydisplay
% 0x50
/summationtext
/producttext
/integraltext
/uniontext
/intersectiontext
/unionmultitext
/logicalandtext
/logicalortext
% 0x58
/summationdisplay
/productdisplay
/integraldisplay
/uniondisplay
/intersectiondisplay
/unionmultidisplay
/logicalanddisplay
/logicalordisplay
% 0x60
/coproducttext
/coproductdisplay
/hatwide
/hatwider
/hatwidest
/tildewide
/tildewider
/tildewidest
% 0x68
/bracketleftBig
/bracketrightBig
/floorleftBig
/floorrightBig
/ceilingleftBig

/ceilingrightBig
/braceleftBig
/bracerightBig
% 0x70
/radicalbig
/radicalBig
/radicalbigg
/radicalBigg
/radicalbt
/radicalvertex
/radicaltp
/arrowvertexdbl
% 0x78
/arrowtp
/arrowbt
/bracehtipdownleft
/bracehtipdownright
/bracehtipupleft
/bracehtipupright
/arrowdbltp
/arrowdblbt
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def

B.10 ‘texmsym.enc’: TEX math symbol encoding

% @psencodingfile{

%   author = "Donald Knuth",
%   version = "1.0",
%   date = "12jun05",
%   filename = "texmsym.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "TeX math symbols encoding."
% }
% 12jun05 - add header.
%
% No ligatures or kerns.
%
/TeXMathSymbolEncoding [          % now 256 chars follow
% 0x0
/minus
/periodcentered
/multiply
/asteriskmath
/divide
/diamondmath
/plusminus
/minusplus
% 0x8
/circleplus
/circleminus
/circlemultiply
/circledivide
/circledot
/circlecopyrt
/openbullet
/bullet
% 0x10
/equivasymptotic
/equivalence
/reflexsubset
/reflexsuperset
/lessequal
/greaterequal
/precedesequal
/followsequal
% 0x18
/similar
/approxequal

/propersubset
/propersuperset
/lessmuch
/greatermuch
/precedes
/follows
% 0x20
/arrowleft
/arrowright
/arrowup
/arrowdown
/arrowboth
/arrownortheast
/arrowsoutheast
/similarequal
% 0x28
/arrowdblleft
/arrowdblright
/arrowdblup
/arrowdbldown
/arrowdblboth
/arrownorthwest
/arrowsouthwest
/proportional
% 0x30
/prime
/infinity
/element
/owner
/triangle
/triangleinv
/negationslash
/mapsto
% 0x38
/universal
/existential
/logicalnot
/emptyset
/Rfractur
/Ifractur
/latticetop
/perpendicular

% 0x40
/aleph
/A
/B
/C
/D
/E
/F
/G
% 0x48
/H
/I
/J
/K
/L
/M
/N
/O
% 0x50
/P
/Q
/R
/S
/T
/U
/V
/W
% 0x58
/X
/Y
/Z
/union
/intersection
/unionmulti
/logicaland
/logicalor
% 0x60
/turnstileleft
/turnstileright
/floorleft
/floorright
/ceilingleft
/ceilingright

/braceleft
/braceright
% 0x68
/angbracketleft
/angbracketright
/bar
/bardbl
/arrowbothv
/arrowdblbothv
/backslash
/wreathproduct
% 0x70
/radical
/coproduct
/nabla
/integral
/unionsq
/intersectionsq
/subsetsqequal
/supersetsqequal
% 0x78
/section
/dagger
/daggerdbl
/paragraph
/club
/diamond
/heart
/spade
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef

/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def

B.11 ‘texmital.enc’: TEX math italic encoding

% @psencodingfile{
%   author = "Donald Knuth",
%   version = "1.0",
%   date = "12jun05",
%   filename = "texmital.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "TeX math italic encoding."
% }
%
% 12jun05 - add header.
%
% No ligatures or kerns.
%
/TeXMathItalicEncoding [          % now 256 chars follow
% 0x0
/Gamma
/Delta
/Theta
/Lambda
/Xi
/Pi
/Sigma
/Upsilon
% 0x8
/Phi
/Psi
/Omega
/alpha
/beta
/gamma
/delta
/epsilon1
% 0x10
/zeta
/eta

/theta
/iota
/kappa
/lambda
/mu
/nu
% 0x18
/xi
/pi
/rho
/sigma
/tau
/upsilon
/phi
/chi
% 0x20
/psi
/omega
/epsilon
/theta1
/pi1
/rho1
/sigma1
/phi1
% 0x28
/arrowlefttophalf
/arrowleftbothalf
/arrowrighttophalf
/arrowrightbothalf
/arrowhookleft
/arrowhookright
/triangleright
/triangleleft
% 0x30
/zerooldstyle
/oneoldstyle
/twooldstyle
/threeoldstyle
/fouroldstyle
/fiveoldstyle
/sixoldstyle
/sevenoldstyle

% 0x38
/eightoldstyle
/nineoldstyle
/period
/comma
/less
/slash
/greater
/star
% 0x40
/partialdiff
/A
/B
/C
/D
/E
/F
/G
% 0x48
/H
/I
/J
/K
/L
/M
/N
/O
% 0x50
/P
/Q
/R
/S
/T
/U
/V
/W
% 0x58
/X
/Y
/Z
/flat
/natural
/sharp

/slurbelow
/slurabove
% 0x60
/lscript
/a
/b
/c
/d
/e
/f
/g
% 0x68
/h
/i
/j
/k
/l
/m
/n
/o
% 0x70
/p
/q
/r
/s
/t
/u
/v
/w
% 0x78
/x
/y
/z
/dotlessi
/dotlessj
/weierstrass
/vector
/tie
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef

/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def
B.12 ‘texnansi.enc’

From Y&Y.

% @psencodingfile{
%   date = "22apr09",
%   filename = "texnansi.enc",
%   email = "tex-fonts@tug.org",
%   author = "Y&Y, Inc.",
%   version = "1.1",
%   codetable = "ISO/ASCII",
%   docstring = "texnansi (aka LY1) encoding vector for Type 1 fonts to
%                be used with TeX, following ASCII and Windows ANSI."
% }
%
% The idea is to have all 228 characters normally included in Type 1 text
% fonts (plus a few more) available for typesetting.    This is effectively
% the character set in Adobe Standard Encoding, ISO Latin 1, plus a few more.
%
% Character code assignments were made as follows:
%
% (1) The character layout largely matches `ASCII' in the 32 -- 126 range,
% except for `circumflex' in 94 and `tilde' in 126, to match `TeX text'
% (`asciicircumflex' and `asciitilde' appear in 158 and 142 instead).
%
% (2) The character layout matches `Windows ANSI' in almost all places,
% except for `quoteright' in 39 and `quoteleft' in 96 to match ASCII
% (`quotesingle' and `grave' appear in 129 and 18 instead).
%
% (3) The character layout matches `TeX typewriter' used by CM text fonts
% in most places (except for discordant positions such as hungarumlaut
% (instead of braceright), dotaccent (instead of underscore) etc.

%
% (4) Remaining characters are assigned arbitrarily to the `control character'
% range (0 -- 31), avoiding 0, 9, 10 and 13 in case we meet dumb software
% - similarly one should really avoid 127 and 128 if possible.
% In addition, the 8 open slots in Windows ANSI between 128 and 159 are used.
%
% (5) Y&Y Lucida Bright includes some extra ligatures and such; ff, ffi, ffl,
% and `dotlessj,' these are included 11 -- 15, and 17.
%
% (6) Hyphen appears both at 45 and 173 for compatibility with both ASCII
% and Windows ANSI.
%
% (7) It doesn't really matter where ligatures appear (both real, such as ffi,
% and pseudo such as ---) since these should not be accessed directly, only
% via ligature information in the TFM file.
%
% SAMPLE USAGE (in `psfonts.map' file for DVIPS):
%
% lbr LucidaBright "TeXnANSIEncoding ReEncodeFont" <texnansi.enc <lbr.pfb
%
% This tells DVIPS that the font called `lbr' in TeX has PostScript
% FontName `LucidaBright.'    It also asks DVIPS to expand the file `lbr.pfb'
% into PFA form, to include the attached `texnansi.enc' encoding vector,
% and to then actually reencode the font based on that encoding vector.
%
% Revised 2009 April 20 with doc changes only (Uni20AC -> uni20ac).
% Revised 2009 March 9 with doc changes only.
% Revised 1996 June 1 by adding second position for `fl' to avoid Acrobat bug.
% Revised 1996 June 1 by adding second position for `fraction' for same reason.
% Revised 1997 Oct 1 by adding cwm (used in boundary char TFM code).
% Revised 1998 Mar 1 by adding Unicode for Euro character.
% emacs-page
%
/TeXnANSIEncoding [     % now 256 chars follow
/.notdef                % 0x00 0
/Euro                   % 0x01 1 /uni20AC
/.notdef                % 0x02 2
/.notdef                % 0x03 3
/fraction               % 0x04 4
/dotaccent              % 0x05 5
/hungarumlaut           % 0x06 6
/ogonek                 % 0x07 7
/fl                     % 0x08 8
/.notdef                % 0x09 9 /fraction at 4, backward compatibility only
/cwm                    % 0x0A 10 not used, except boundary char internally
/ff                     % 0x0B 11
/fi                     % 0x0C 12

/.notdef                % 0x0D 13 /fl at 8, backward compatibility only
/ffi                    % 0x0E 14
/ffl                    % 0x0F 15
/dotlessi               % 0x10 16
/dotlessj               % 0x11 17
/grave                  % 0x12 18
/acute                  % 0x13 19
/caron                  % 0x14 20
/breve                  % 0x15 21
/macron                 % 0x16 22
/ring                   % 0x17 23
/cedilla                % 0x18 24
/germandbls             % 0x19 25
/ae                     % 0x1A 26
/oe                     % 0x1B 27
/oslash                 % 0x1C 28
/AE                     % 0x1D 29
/OE                     % 0x1E 30
/Oslash                 % 0x1F 31
/space                  % 0x20 32 /suppress in TeX text
/exclam                 % 0x21 33
/quotedbl               % 0x22 34 /quotedblright in TeX text
/numbersign             % 0x23 35
/dollar                 % 0x24 36
/percent                % 0x25 37
/ampersand              % 0x26 38
/quoteright             % 0x27 39 /quotesingle in ANSI
/parenleft              % 0x28 40
/parenright             % 0x29 41
/asterisk               % 0x2A 42
/plus                   % 0x2B 43
/comma                  % 0x2C 44
/hyphen                 % 0x2D 45
/period                 % 0x2E 46
/slash                  % 0x2F 47
/zero                   % 0x30 48
/one                    % 0x31 49
/two                    % 0x32 50
/three                  % 0x33 51
/four                   % 0x34 52
/five                   % 0x35 53
/six                    % 0x36 54
/seven                  % 0x37 55
/eight                  % 0x38 56
/nine                   % 0x39 57
/colon                  % 0x3A 58
/semicolon              % 0x3B 59

/less                   % 0x3C 60 /exclamdown in Tex text
/equal                  % 0x3D 61
/greater                % 0x3E 62 /questiondown in TeX text
/question               % 0x3F 63
/at                     % 0x40 64
/A                      % 0x41 65
/B                      % 0x42 66
/C                      % 0x43 67
/D                      % 0x44 68
/E                      % 0x45 69
/F                      % 0x46 70
/G                      % 0x47 71
/H                      % 0x48 72
/I                      % 0x49 73
/J                      % 0x4A 74
/K                      % 0x4B 75
/L                      % 0x4C 76
/M                      % 0x4D 77
/N                      % 0x4E 78
/O                      % 0x4F 79
/P                      % 0x50 80
/Q                      % 0x51 81
/R                      % 0x52 82
/S                      % 0x53 83
/T                      % 0x54 84
/U                      % 0x55 85
/V                      % 0x56 86
/W                      % 0x57 87
/X                      % 0x58 88
/Y                      % 0x59 89
/Z                      % 0x5A 90
/bracketleft            % 0x5B 91
/backslash              % 0x5C 92 /quotedblleft in TeX text
/bracketright           % 0x5D 93
/circumflex             % 0x5E 94 /asciicircum in ASCII
/underscore             % 0x5F 95 /dotaccent in TeX text
/quoteleft              % 0x60 96 /grave accent in ANSI
/a                      % 0x61 97
/b                      % 0x62 98
/c                      % 0x63 99
/d                      % 0x64 100
/e                      % 0x65 101
/f                      % 0x66 102
/g                      % 0x67 103
/h                      % 0x68 104
/i                      % 0x69 105
/j                      % 0x6A 106

/k                      % 0x6B 107
/l                      % 0x6C 108
/m                      % 0x6D 109
/n                      % 0x6E 110
/o                      % 0x6F 111
/p                      % 0x70 112
/q                      % 0x71 113
/r                      % 0x72 114
/s                      % 0x73 115
/t                      % 0x74 116
/u                      % 0x75 117
/v                      % 0x76 118
/w                      % 0x77 119
/x                      % 0x78 120
/y                      % 0x79 121
/z                      % 0x7A 122
/braceleft              % 0x7B 123 /endash in TeX text
/bar                    % 0x7C 124 /emdash in TeX test
/braceright             % 0x7D 125 /hungarumlaut in TeX text
/tilde                  % 0x7E 126 /asciitilde in ASCII
/dieresis               % 0x7F 127 also at 168
/Lslash                 % 0x80 128 position is unfortunate, but too late to fix
/quotesingle            % 0x81 129
/quotesinglbase         % 0x82 130
/florin                 % 0x83 131
/quotedblbase           % 0x84 132
/ellipsis               % 0x85 133
/dagger                 % 0x86 134
/daggerdbl              % 0x87 135
/circumflex             % 0x88 136
/perthousand            % 0x89 137
/Scaron                 % 0x8A 138
/guilsinglleft          % 0x8B 139
/OE                     % 0x8C 140
/Zcaron                 % 0x8D 141
/asciicircum            % 0x8E 142
/minus                  % 0x8F 143
/lslash                 % 0x90 144
/quoteleft              % 0x91 145
/quoteright             % 0x92 146
/quotedblleft           % 0x93 147
/quotedblright          % 0x94 148
/bullet                 % 0x95 149
/endash                 % 0x96 150
/emdash                 % 0x97 151
/tilde                  % 0x98 152
/trademark              % 0x99 153

/scaron                 % 0x9A 154
/guilsinglright         % 0x9B 155
/oe                     % 0x9C 156
/zcaron                 % 0x9D 157
/asciitilde             % 0x9E 158
/Ydieresis              % 0x9F 159
/nbspace                % 0xA0 160 no break space
/exclamdown             % 0xA1 161
/cent                   % 0xA2 162
/sterling               % 0xA3 163
/currency               % 0xA4 164
/yen                    % 0xA5 165
/brokenbar              % 0xA6 166
/section                % 0xA7 167
/dieresis               % 0xA8 168
/copyright              % 0xA9 169
/ordfeminine            % 0xAA 170
/guillemotleft          % 0xAB 171
/logicalnot             % 0xAC 172
/sfthyphen              % 0xAD 173 /hyphen (hanging hyphen)
/registered             % 0xAE 174
/macron                 % 0xAF 175
/degree                 % 0xB0 176
/plusminus              % 0xB1 177
/twosuperior            % 0xB2 178
/threesuperior          % 0xB3 179
/acute                  % 0xB4 180
/mu                     % 0xB5 181
/paragraph              % 0xB6 182
/periodcentered         % 0xB7 183
/cedilla                % 0xB8 184
/onesuperior            % 0xB9 185
/ordmasculine           % 0xBA 186
/guillemotright         % 0xBB 187
/onequarter             % 0xBC 188
/onehalf                % 0xBD 189
/threequarters          % 0xBE 190
/questiondown           % 0xBF 191
/Agrave                 % 0xC0 192
/Aacute                 % 0xC1 193
/Acircumflex            % 0xC2 194
/Atilde                 % 0xC3 195
/Adieresis              % 0xC4 196
/Aring                  % 0xC5 197
/AE                     % 0xC6 198
/Ccedilla               % 0xC7 199
/Egrave                 % 0xC8 200

/Eacute                 % 0xC9 201
/Ecircumflex            % 0xCA 202
/Edieresis              % 0xCB 203
/Igrave                 % 0xCC 204
/Iacute                 % 0xCD 205
/Icircumflex            % 0xCE 206
/Idieresis              % 0xCF 207
/Eth                    % 0xD0 208
/Ntilde                 % 0xD1 209
/Ograve                 % 0xD2 210
/Oacute                 % 0xD3 211
/Ocircumflex            % 0xD4 212
/Otilde                 % 0xD5 213
/Odieresis              % 0xD6 214
/multiply               % 0xD7 215 OE in T1
/Oslash                 % 0xD8 216
/Ugrave                 % 0xD9 217
/Uacute                 % 0xDA 218
/Ucircumflex            % 0xDB 219
/Udieresis              % 0xDC 220
/Yacute                 % 0xDD 221
/Thorn                  % 0xDE 222
/germandbls             % 0xDF 223
/agrave                 % 0xE0 224
/aacute                 % 0xE1 225
/acircumflex            % 0xE2 226
/atilde                 % 0xE3 227
/adieresis              % 0xE4 228
/aring                  % 0xE5 229
/ae                     % 0xE6 230
/ccedilla               % 0xE7 231
/egrave                 % 0xE8 232
/eacute                 % 0xE9 233
/ecircumflex            % 0xEA 234
/edieresis              % 0xEB 235
/igrave                 % 0xEC 236
/iacute                 % 0xED 237
/icircumflex            % 0xEE 238
/idieresis              % 0xEF 239
/eth                    % 0xF0 240
/ntilde                 % 0xF1 241
/ograve                 % 0xF2 242
/oacute                 % 0xF3 243
/ocircumflex            % 0xF4 244
/otilde                 % 0xF5 245
/odieresis              % 0xF6 246
/divide                 % 0xF7 247 oe in T1

/oslash                 % 0xF8 248
/ugrave                 % 0xF9 249
/uacute                 % 0xFA 250
/ucircumflex            % 0xFB 251
/udieresis              % 0xFC 252
/yacute                 % 0xFD 253
/thorn                  % 0xFE 254
/ydieresis              % 0xFF 255 germandbls in T1
] def
B.13 ‘texnansx.enc’

From Y&Y. ‘texnansi’ without repeats.

% @psencodingfile{
%   author = "Y&Y, Inc.",
%   version = "1.3",
%   date = "22apr09",
%   filename = "texnansx.enc",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "Encoding for fonts in Adobe Type 1 format for use with TeX."
% }
% MODIFIED FOR AFM2TFM to avoid repeated encodings, namely these
% positions have become .notdef:
%
% 137/circumflex 141/OE     146/quoteleft 147/quoteright 153/tilde 157/oe
% 169/dieresis   176/macron 181/acute     185/cedilla    199/AE    217/Oslash
% 224/germandbls 231/ae     249/oslash
%
% 22apr09 document the precise differences from texnansi.
% 17aug08 change /Uni20AC (wrong capitalization of /Uni) to /Euro for
% consistency with other encodings.    tex-live mail from Vasile Gaburici,
% 16 Aug 2008 20:44:08.
%
/TeXnANSIEncoding [
/.notdef % 0
/Euro % /uni20AC 1
/.notdef % 2
/.notdef % 3

/fraction %     4
/dotaccent %    5
/hungarumlaut % 6
/ogonek %       7
/fl     %       8
/.notdef % /fraction %  9       not used (see 4), backward compatability only
/cwm    %       10      not used, except boundary char internally maybe
/ff    %        11

/fi    %        12
/.notdef % /fl    %     13      not used (see 8), backward compatability only
/ffi   %        14
/ffl   %        15
/dotlessi %     16
/dotlessj %     17
/grave %        18
/acute %        19
/caron %        20
/breve %        21
/macron %       22
/ring  %        23
/cedilla %      24
/germandbls %   25
/ae    %        26
/oe    %        27
/oslash %       28
/AE    %        29
/OE    %        30
/Oslash %       31
/space %        32      % /suppress in TeX text
/exclam %       33
/quotedbl %     34      % /quotedblright in TeX text
/numbersign %   35
/dollar %       36
/percent %      37
/ampersand %    38
/quoteright %   39      % /quotesingle in ANSI
/parenleft %    40
/parenright %   41
/asterisk %     42
/plus  %        43
/comma %        44
/hyphen %       45
/period %       46
/slash %        47
/zero  %        48
/one   %        49
/two   %        50
/three %        51
/four  %        52
/five  %        53
/six   %        54
/seven %        55
/eight %        56
/nine  %        57
/colon %        58

/semicolon %    59
/less  %        60      % /exclamdown in Tex text
/equal %        61
/greater %      62      % /questiondown in TeX text
/question %     63
/at %   64
/A %    65
/B %    66
/C %    67
/D %    68
/E %    69
/F %    70
/G %    71
/H %    72
/I %    73
/J %    74
/K %    75
/L %    76
/M %    77
/N %    78
/O %    79
/P %    80
/Q %    81
/R %    82
/S %    83
/T %    84
/U %    85
/V %    86
/W %    87
/X %    88
/Y %    89
/Z %    90
/bracketleft %  91
/backslash %    92      % /quotedblleft in TeX text
/bracketright % 93
/circumflex %   94      % /asciicircum in ASCII
/underscore %   95      % /dotaccent in TeX text
/quoteleft %    96      % /grave accent in ANSI
/a %    97
/b %    98
/c %    99
/d %    100
/e %    101
/f %    102
/g %    103
/h %    104
/i %    105

/j %    106
/k %    107
/l %    108
/m %    109
/n %    110
/o %    111
/p %    112
/q %    113
/r %    114
/s %    115
/t %    116
/u %    117
/v %    118
/w %    119
/x %    120
/y %    121
/z %    122
/braceleft %    123     % /endash in TeX text
/bar   %        124     % /emdash in TeX test
/braceright %   125     % /hungarumlaut in TeX text
/tilde %        126     % /asciitilde in ASCII
/dieresis %     127     not used (see 168), use higher up instead
/Lslash %       128     this position is unfortunate, but now too late to fix
/quotesingle %  129
/quotesinglbase %       130
/florin %       131
/quotedblbase % 132
/ellipsis %     133
/dagger %       134
/daggerdbl %    135
/.notdef % /circumflex %        136
/perthousand %  137
/Scaron %       138
/guilsinglleft %        139
/.notdef % /OE    %     140
/Zcaron %       141
/asciicircum %  142
/minus %        143
/lslash %       144
/.notdef % /quoteleft % 145
/.notdef % /quoteright %        146
/quotedblleft % 147
/quotedblright %        148
/bullet %       149
/endash %       150
/emdash %       151
/.notdef % /tilde %     152

/trademark %    153
/scaron %       154
/guilsinglright %       155
/.notdef % /oe    %     156
/zcaron %       157
/asciitilde %   158
/Ydieresis %    159
/nbspace %      160     % /space (no break space)
/exclamdown %   161
/cent  %        162
/sterling %     163
/currency %     164
/yen   %        165
/brokenbar %    166
/section %      167
/.notdef % /dieresis %  168
/copyright %    169
/ordfeminine %  170
/guillemotleft %        171
/logicalnot %   172
/sfthyphen %    173 % /hyphen (hanging hyphen)
/registered %   174
/.notdef % /macron %    175
/degree %       176
/plusminus %    177
/twosuperior %  178
/threesuperior %        179
/.notdef % /acute %     180
/mu    %        181
/paragraph %    182
/periodcentered %       183
/.notdef % /cedilla %   184
/onesuperior %  185
/ordmasculine % 186
/guillemotright %       187
/onequarter %   188
/onehalf %      189
/threequarters %        190
/questiondown % 191
/Agrave %       192
/Aacute %       193
/Acircumflex %  194
/Atilde %       195
/Adieresis %    196
/Aring %        197
/.notdef % /AE    %     198
/Ccedilla %     199

/Egrave %       200
/Eacute %       201
/Ecircumflex %  202
/Edieresis %    203
/Igrave %       204
/Iacute %       205
/Icircumflex %  206
/Idieresis %    207
/Eth   %        208
/Ntilde %       209
/Ograve %       210
/Oacute %       211
/Ocircumflex %  212
/Otilde %       213
/Odieresis %    214
/multiply %     215     % OE in T1
/.notdef % /Oslash %    216
/Ugrave %       217
/Uacute %       218
/Ucircumflex %  219
/Udieresis %    220
/Yacute %       221
/Thorn %        222
/.notdef % /germandbls %        223
/agrave %       224
/aacute %       225
/acircumflex %  226
/atilde %       227
/adieresis %    228
/aring %        229
/.notdef % /ae    %     230
/ccedilla %     231
/egrave %       232
/eacute %       233
/ecircumflex %  234
/edieresis %    235
/igrave %       236
/iacute %       237
/icircumflex %  238
/idieresis %    239
/eth   %        240
/ntilde %       241
/ograve %       242
/oacute %       243
/ocircumflex %  244
/otilde %       245
/odieresis %    246

/divide %       247     % oe in T1
/.notdef % /oslash %    248
/ugrave %       249
/uacute %       250
/ucircumflex %  251
/udieresis %    252
/yacute %       253
/thorn %        254
/ydieresis %    255     % germandbls in T1

] def

B.14 ‘xl2.enc’: OT1 + ISO Latin 2 extended

% @psencodingfile{
%   author = "Petr Olsak, Zdenek Wagner",
%   date = "27aug05",
%   filename = "xl2.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "
% some of our (CSTUG- czech TeX Users Group) users want to support
% 8bit font coding such that:
% -- lower 7bit  is exactly OT1 (but with differences imposed
%   by DEK -- e.g.  layout of cmr is different from cmtt)
% -- upper part is taken from ISO-Latin 2 (iso 8859-2),
%    but some of empty positions are filled with useful characters
%    usually available in type-1 font (permill sign etc.)
% "
% }
%
% 27aug05 - record Petr Olsak and Zdenek Wagner as the authors.
% 12jun05 - remove Texinfo @'s from LIGKERN's.
%
/XL2encoding [
%%%%%   '000 = "00
   /Gamma
   /Delta
   /Theta
   /Lambda
   /Xi
   /Pi
   /Sigma
   /Upsilon
%%%%%   '010 = "08
   /Phi
   /Psi

/Omega
/ff
/fi
/fl
/ffi
/ffl
%%%%%   '020 = "10
/dotlessi
/dotlessj
/grave
/acute
/caron
/breve
/macron
/ring
%%%%%   '030 = "18
/cedilla
/germandbls
/ae
/oe
/oslash
/AE
/OE
/Oslash
%%%%%   '040 = "20
/.notdef
/exclam
/quotedblright
/numbersign
/dollar
/percent
/ampersand
/quoteright
%%%%%   '050 = "28
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
%%%%%   '060 = "30
/zero
/one
/two
/three

/four
/five
/six
/seven
%%%%%   '070 = "38
/eight
/nine
/colon
/semicolon
/exclamdown
/equal
/questiondown
/question
%%%%%   '100 = "40
/at
/A
/B
/C
/D
/E
/F
/G
%%%%%   '110 = "48
/H
/I
/J
/K
/L
/M
/N
/O
%%%%%   '120 = "50
/P
/Q
/R
/S
/T
/U
/V
/W
%%%%%   '130 = "58
/X
/Y
/Z
/bracketleft
/quotedblleft
/bracketright

/circumflex
/dotaccent
%%%%%   '140 = "60
/quoteleft
/a
/b
/c
/d
/e
/f
/g
%%%%%   '150 = "68
/h
/i
/j
/k
/l
/m
/n
/o
%%%%%   '160 = "70
/p
/q
/r
/s
/t
/u
/v
/w
%%%%%   '170 = "78
/x
/y
/z
/endash
/emdash
/hungarumlaut
/tilde
/dieresis
%%%%%   '200 = "80
/ellipsis
/dagger
/daggerdbl
/bullet
/sterling
/paragraph
/.notdef
/.notdef

%%%%%   '210 = "88
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/perthousand
/.notdef
/.notdef
%%%%%   '220 = "90
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
%%%%%   '230 = "98
/Agrave
/.notdef
/.notdef
/.notdef
/hyphen
/ogonek
/guillemotleft
/guillemotright
%%%%%   '240 = "A0
/.notdef
/Aogonek
/breve
/Lslash
/currency
/Lcaron
/Sacute
/section
%%%%%   '250 = "A8
/dieresis
/Scaron
/Scedilla
/Tcaron
/Zacute
/.notdef
/Zcaron
/Zdotaccent
%%%%%   '260 = "B0
/ring

/aogonek
/cedilla
/lslash
/acute
/lcaron
/sacute
/caron
%%%%%   '270 = "B8
/agrave
/scaron
/scedilla
/tcaron
/zacute
/hungarumlaut
/zcaron
/zdotaccent
%%%%%   '300 = "C0
/Racute
/Aacute
/Acircumflex
/Abreve
/Adieresis
/Lacute
/Cacute
/Ccedilla
%%%%%   '310 = "C8
/Ccaron
/Eacute
/Eogonek
/Edieresis
/Ecaron
/Iacute
/Icircumflex
/Dcaron
%%%%%   '320 = "D0
/Eth
/Nacute
/Ncaron
/Oacute
/Ocircumflex
/Ohungarumlaut
/Odieresis
/multiply
%%%%%   '330 = "D8
/Rcaron
/Uring
/Uacute

/Uhungarumlaut
/Udieresis
/Yacute
/Togonek
/germandbls
%%%%%   '340 = "E0
/racute
/aacute
/acircumflex
/abreve
/adieresis
/lacute
/cacute
/ccedilla
%%%%%   '350 = "E8
/ccaron
/eacute
/eogonek
/edieresis
/ecaron
/iacute
/icircumflex
/dcaron
%%%%%   '360 = "F0
/eth
/nacute
/ncaron
/oacute
/ocircumflex
/ohungarumlaut
/odieresis
/divide
%%%%%   '370 = "F8
/rcaron
/uring
/uacute
/uhungarumlaut
/udieresis
/yacute
/quotedblbase
/quotedblleft
] def
% LIGKERN hyphen hyphen =:  endash ; endash hyphen =:  emdash ;
% LIGKERN quoteleft quoteleft =:  quotedblleft ;
% LIGKERN quoteright quoteright =:  quotedblright ;
% LIGKERN exclamdown exclamdown =:  guillemotleft ;

% frenchdblquotes
% LIGKERN questiondown questiondown =:  guillemotright ;
% csquoteleft
% LIGKERN comma comma =:  quotedblbase ;
% LIGKERN space {} * ; * {} space ; zero {} * ; * {} zero ;
% LIGKERN one {} * ; * {} one ; two {} * ; * {} two ;
% LIGKERN three {} * ; * {} three ; four {} * ; * {} four ;
% LIGKERN five {} * ; * {} five ; six {} * ; * {} six ;
% LIGKERN seven {} * ; * {} seven ; eight {} * ; * {} eight ;
% LIGKERN nine {} * ; * {} nine ;
% LIGKERN question {} quoteleft ; exclam {} quoteleft ;

B.15 ‘xt2.enc’: typewriter OT1 + ISO Latin 2 extended

% @psencodingfile{
%   author = "Petr Olsak, Zdenek Wagner",
%   date = "27aug05",
%   filename = "xt2.enc",
%   license = "public domain",
%   email = "tex-fonts@tug.org",
%   codetable = "ISO/ASCII",
%   docstring = "This is the typewriter version of xl2."
% }
%
% 27aug05 - record Petr Olsak and Zdenek Wagner as the authors.
% 12jun05 - remove Texinfo @'s from LIGKERN's.
%
/XT2encoding [
%%%%%   '000 = "00
   /Gamma
   /Delta
   /Theta
   /Lambda
   /Xi
   /Pi
   /Sigma
   /Upsilon
%%%%%   '010 = "08
   /Phi
   /Psi
   /Omega
   /arrowup
   /arrowdown
   /quotesingle
   /exclamdown
   /questiondown
%%%%%   '020 = "10

/dotlessi
/dotlessj
/grave
/acute
/caron
/breve
/macron
/ring
%%%%%   '030 = "18
/cedilla
/germandbls
/ae
/oe
/oslash
/AE
/OE
/Oslash
%%%%%   '040 = "20
/.notdef
/exclam
/quotedblright
/numbersign
/dollar
/percent
/ampersand
/quoteright
%%%%%   '050 = "28
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
%%%%%   '060 = "30
/zero
/one
/two
/three
/four
/five
/six
/seven
%%%%%   '070 = "38
/eight
/nine

/colon
/semicolon
/less
/equal
/greater
/question
%%%%%   '100 = "40
/at
/A
/B
/C
/D
/E
/F
/G
%%%%%   '110 = "48
/H
/I
/J
/K
/L
/M
/N
/O
%%%%%   '120 = "50
/P
/Q
/R
/S
/T
/U
/V
/W
%%%%%   '130 = "58
/X
/Y
/Z
/bracketleft
/backslash
/bracketright
/circumflex
/underscore
%%%%%   '140 = "60
/quoteleft
/a
/b

/d
/e
/f
/g
%%%%%   '150 = "68
/h
/i
/j
/k
/l
/m
/n
/o
%%%%%   '160 = "70
/p
/q
/r
/s
/t
/u
/v
/w
%%%%%   '170 = "78
/x
/y
/z
/braceleft
/bar
/braceright
/tilde
/dieresis
%%%%%   '200 = "80
/ellipsis
/dagger
/daggerdbl
/bullet
/sterling
/paragraph
/.notdef
/.notdef
%%%%%   '210 = "88
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/perthousand

/.notdef
/.notdef
%%%%%   '220 = "90
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
%%%%%   '230 = "98
/Agrave
/.notdef
/.notdef
/.notdef
/hyphen
/ogonek
/guillemotleft
/guillemotright
%%%%%   '240 = "A0
/.notdef
/Aogonek
/breve
/Lslash
/currency
/Lcaron
/Sacute
/section
%%%%%   '250 = "A8
/dieresis
/Scaron
/Scedilla
/Tcaron
/Zacute
/.notdef
/Zcaron
/Zdotaccent
%%%%%   '260 = "B0
/ring
/aogonek
/cedilla
/lslash
/acute
/lcaron
/sacute
/caron

%%%%%   '270 = "B8
/agrave
/scaron
/scedilla
/tcaron
/zacute
/hungarumlaut
/zcaron
/zdotaccent
%%%%%   '300 = "C0
/Racute
/Aacute
/Acircumflex
/Abreve
/Adieresis
/Lacute
/Cacute
/Ccedilla
%%%%%   '310 = "C8
/Ccaron
/Eacute
/Eogonek
/Edieresis
/Ecaron
/Iacute
/Icircumflex
/Dcaron
%%%%%   '320 = "D0
/Eth
/Nacute
/Ncaron
/Oacute
/Ocircumflex
/Ohungarumlaut
/Odieresis
/multiply
%%%%%   '330 = "D8
/Rcaron
/Uring
/Uacute
/Uhungarumlaut
/Udieresis
/Yacute
/Togonek
/germandbls
%%%%%   '340 = "E0
/racute

/aacute
/acircumflex
/abreve
/adieresis
/lacute
/cacute
/ccedilla
%%%%%   '350 = "E8
/ccaron
/eacute
/eogonek
/edieresis
/ecaron
/iacute
/icircumflex
/dcaron
%%%%%   '360 = "F0
/eth
/nacute
/ncaron
/oacute
/ocircumflex
/ohungarumlaut
/odieresis
/divide
%%%%%   '370 = "F8
/rcaron
/uring
/uacute
/uhungarumlaut
/udieresis
/yacute
/quotedblbase
/quotedblleft
] def
% LIGKERN quoteleft quoteleft =:  quotedblleft ;
% LIGKERN quoteright quoteright =:  quotedblright ;
% LIGKERN exclamdown exclamdown =:  guillemotleft ;
% frenchdblquotes
% LIGKERN questiondown questiondown =:  guillemotright ;
% csquoteleft
% LIGKERN comma comma =:  quotedblbase ;
% LIGKERN space {} * ; * {} space ; zero {} * ; * {} zero ;
% LIGKERN one {} * ; * {} one ; two {} * ; * {} two ;
% LIGKERN three {} * ; * {} three ; four {} * ; * {} four ;
% LIGKERN five {} * ; * {} five ; six {} * ; * {} six ;

% LIGKERN seven {} * ; * {} seven ; eight {} * ; * {} eight ;
% LIGKERN nine {} * ; * {} nine ;
% LIGKERN question {} quoteleft ; exclam {} quoteleft ;
% LIGKERN f {} f ; f {} i ; f {} l ;

General index 331

General index

3

35 standard PostScript fonts . . . . . . . . . . . . . . . . . . . 27

6

62.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

7

7-bit encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

8

8-bit encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 8a.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 8r.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

9

9s variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

A

Adobe standard encoding . . . . . . . . . . . . . . . . . . . . . 246 adobe.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Afm2tfm encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 apple.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 attribution of fonts, missing. . . . . . . . . . . . . . . . . . . . . 3 automatic expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Avant Garde . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

B

base encoding, for TEX . . . . . . . . . . . . . . . . . . . . . . . 252 Beebe, Nelson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Beeton, Barbara . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Bernstein, Rocky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 bitstrea.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 bizarre font names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Bouche, Thierry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 bright variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

C

category codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 character scaling, automatic . . . . . . . . . . . . . . . . . . . 24 compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 contributors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Cork encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 CP-1251 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Cugley, Damian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Cyrillic encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Cyrillic encodings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Cyrillic fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

D

design size specification . . . . . . . . . . . . . . . . . . . . . . . . . 3 distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 dtc.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Dviljk translator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Dvips encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 Dvips translator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 dvips.enc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255

E

EC encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 ec.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 encoding files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 encoding vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 encoding, Dvips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 encoding, Groff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 encoding, T5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 encodings, Adobe standard . . . . . . . . . . . . . . . . . . . 246 encodings, Cork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 encodings, Cyrillic CP-1251. . . . . . . . . . . . . . . . . . . 240 encodings, EC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 encodings, qx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 encodings, TEX base . . . . . . . . . . . . . . . . . . . . . . . . . . 252 encodings, TEX Latin 1 . . . . . . . . . . . . . . . . . . . . . . . 256 encodings, TEX math italic . . . . . . . . . . . . . . . . . . . 299 encodings, TEX math symbol . . . . . . . . . . . . 290, 294 encodings, texnansi . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 encodings, texnansx . . . . . . . . . . . . . . . . . . . . . . . . . . 310 encodings, unspecified . . . . . . . . . . . . . . . . . . . . . . . . . 21 Ertner, Vincent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 escape for nonstandard names . . . . . . . . . . . . . . . . . . 4 expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 expert encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 expertised encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 expertised font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 extension, TEX math encoding . . . . . . . . . . . . . . . . 290

F

fax variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 filenames, basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 font scaling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 font_coding_scheme . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Fontinst distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 FontName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 fontname mapping file . . . . . . . . . . . . . . . . . . . . . . . . . 25 Fontname version 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 foundries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

Garamond, versions of . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Greek fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Groff encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 groff.enc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265

H

hexadecimal design size . . . . . . . . . . . . . . . . . . . . . . . . . 3 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Horn, Berthol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

I

individual font designers . . . . . . . . . . . . . . . . . . . . . . . . 3 informal variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ISO Latin 2 extended and typewriter, TEX

encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 ISO Latin 2 extended, TEX encoding. . . . . . . . . . 316 italic, TEX math encoding . . . . . . . . . . . . . . . . . . . . 299 itc.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

J

Jeffrey, Alan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

K

Knuth, Donald E. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

L

Lang, Russell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 LaserJet 4 fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 letterspacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ligatures and encodings . . . . . . . . . . . . . . . . . . . . . . . . 20 linotype.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 long naming scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 lslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Lslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Lubalin, Herb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Lucida Bright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

M

MacKay, Pierre . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 mapping file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 mapping files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 math extension, TEX encoding . . . . . . . . . . . . . . . . 290 math italic, TEX encoding . . . . . . . . . . . . . . . . . . . . 299 math symbol, TEX encoding . . . . . . . . . . . . . . . . . . 294 math variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Mittelbach, Frank . . . . . . . . . . . . . . . . . . . . . . . . 1, 3, 21 modes.mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 monotype.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 multiple variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

N

nonstandard names, escape for . . . . . . . . . . . . . . . . . . 4

P

phonetic encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 pi fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 PostScript encoding vectors. . . . . . . . . . . . . . . . . . . 240 PostScript fonts, standard . . . . . . . . . . . . . . . . . . . . . 27

Q

qx encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 qx.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272

R

Rahtz, Sebastian. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 raw fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 references. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 related files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Rivlin, Jean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Rokicki, Tom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

S

sans variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Schoepf, Rainer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 schoolbook variant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 semisans variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 semiserif variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 small caps fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 small foundries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Southall, Richard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 standard PostScript fonts . . . . . . . . . . . . . . . . . . . . . . 27 Stone Informal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 SuperFont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 supplier.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 suppliers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 symbol, TEX math encoding . . . . . . . . . . . . . . . . . . 294

T

T1.etx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 T5 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 t5.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 TDS standard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 tex-fonts@tug.org. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 TEX base encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 TEX Directory Structure standard. . . . . . . . . . . . . . . 1 TEX Latin 1 encoding . . . . . . . . . . . . . . . . . . . . . . . . 256 TEX math extension encoding . . . . . . . . . . . . . . . . 290 TEX math italic encoding . . . . . . . . . . . . . . . . . . . . . 299 TEX math symbol encoding . . . . . . . . . . . . . . . . . . . 294 texfonts.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 texmext.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 texmital.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 texmsym.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 texnansi encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 texnansi.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 texnansx encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 texnansx.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 type foundries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 typeface.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 typewriter variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

U

unattributed fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 URW fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 urw.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

V

variant.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Vietnamese encoding . . . . . . . . . . . . . . . . . . . . . . . . . 267

W

weight. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 weight.map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 width.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 widths. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Windows ANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

X

X Window System font names . . . . . . . . . . . . . . . . . 25 xl2.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 xt2.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323